### v1

# Anderson localized states for the quasi-periodic nonlinear Schrödinger equation on $\mathbb{Z}^{d}$

Abstract.
We establish large sets of Anderson localized states for the quasi-periodic nonlinear Schrödinger equation on $\mathbb{Z}^{d}$, thus extending Anderson localization from the linear (cf. Bourgain [GAFA 17(3), 682–706, 2007]) to a nonlinear setting, and the random (cf. Bourgain-Wang [JEMS 10(1), 1–45, 2008]) to a deterministic setting. Among the main ingredients are a new Diophantine estimate of quasi-periodic functions in arbitrarily dimensional phase space, and the application of Bourgain’s geometric lemma in [GAFA 17(3), 682–706, 2007].
Key words and phrases:
Anderson localization, discrete nonlinear Schrödinger equation, quasi-periodic potential, Diophantine estimates, semi-algebraic sets, Lyapunov-Schmidt decomposition, Green’s function estimates, Newton iteration
1. Introduction and Main Result
Let
$\bm{\alpha}$ $=(\alpha_{1},\cdots,\alpha_{d})\in[0,1]^{d},$
$\bm{\theta}$ $=(\theta_{1},\cdots,\theta_{d})\in[0,1]^{d},$
and consider the nonlinear Schrödinger equation with a quasi-periodic potential on $\mathbb{Z}^{d}$:
$iu_{t}+(\varepsilon\Delta+V(\bm{\theta}+\bm{n}\bm{\alpha})\delta_{\bm{n},\bm{n}^{\prime}})u+\delta|u|^{2p}u=0,$ (1.1)
where $\varepsilon$ and $\delta$ are parameters in $[0,1]$, $\bm{n}\in\mathbb{Z}^{d}$, $p\in\mathbb{N}$ and $\Delta(\bm{n},\bm{n}^{\prime})=\delta_{|\bm{n}-\bm{n}^{\prime}|_{1},1}$ denotes the discrete Laplacian with $|\cdot|_{1}$ the $\ell^{1}$ norm and
$\bm{n}\bm{\alpha}=(n_{1}\alpha_{1},\cdots,n_{d}\alpha_{d})\in\mathbb{R}^{d};$
the potential $V$ is a trigonometric polynomial:
$V(\bm{\theta})=\sum_{\bm{\ell}\in\Gamma_{K}}v_{\bm{\ell}}\cos 2\pi(\bm{\ell}\cdot\bm{\theta})\ {\rm with}\ \bm{\ell}\cdot\bm{\theta}=\sum_{k=1}^{d}\ell_{k}\theta_{k},$
where $\Gamma_{K}\subset[-K,K]^{d}\setminus\{\bm{0}\},\ K\geq 1$ and
$\bm{v}=(v_{\bm{\ell}})_{\bm{\ell}\in\Gamma_{K}}\in\mathbb{R}^{\#\Gamma_{K}}\setminus\{\bm{0}\},$
where #()#\#($\cdot$)# ( ⋅ ) denotes the cardinality of a set. We further assume that the set $\Gamma_{K}$ satisfies:
$\text{if }\bm{\ell}\in\Gamma_{K},{\rm then},\ell_{k}\neq 0,\forall\ k=1,2,\cdots,d,$ (1.2)
$\text{and if }\ \bm{\ell},\bm{\ell}^{\prime}\in\Gamma_{K},\ {\rm then}\ \bm{\ell}+\bm{\ell}^{\prime}\neq\bm{0}.$
We assume that $V$ is non-degenerate: Let $\ell\in\{1,2,\cdots,d\}$ and fix $\theta_{i}$, $\forall\ i\neq\ell$. Then $V$ is not a constant function in $\theta_{\ell}$ for all fixed $\theta_{i}$, $i\neq\ell$, for all $\ell$. Note that the first condition in (1.2) then says that, in fact, each term in the polynomial is non-degenerate.
Remark 1.1.
We may assume, in addition that
$\min_{\bm{\theta}\in[0,1]^{d}}V(\bm{\theta})\geq 1.$ (1.3)
Indeed, if we replace $V$ in (1.1) with $V+m$ for some constant $m>\|V\|_{\infty}+1$ and have constructed its corresponding solution $u(t,\bm{n})$, then $e^{-imt}u(t,\bm{n})$ must be a solution to the original equation (1.1). Since our goal is to establish the existence of time quasi-periodic solutions to (1.1), such a phase shift is largely immaterial. The bound (1.3) is convenient for imposing certain lower bounds in the analysis later.
When $\delta=0$, it is known from the breakthrough paper [Bou07] (cf. also the recent refinement [JLS20]) that the linear operator
$\mathcal{H}=\varepsilon\Delta+V(\bm{n}\bm{\alpha}+\bm{\theta})\delta_{\bm{n},\bm{n}^{\prime}},\,\text{on }\mathbb{Z}^{d},$ (1.4)
exhibits Anderson localization, namely pure point spectrum with exponentially decaying eigenfunctions, for small $\varepsilon$, on a large set in $(\bm{\alpha},\bm{\theta})$. It follows that when $0<\varepsilon\ll 1$, the linear Schrödinger equation
$iu_{t}+(\varepsilon\Delta+V(\bm{n}\bm{\alpha}+\bm{\theta})\delta_{\bm{n},\bm{n}^{\prime}})u=0,$ (1.5)
has only Anderson localized states, i.e., roughly speaking, wave packets localized about the origin remain localized for all time.
The present paper addresses the persistence question when $\delta\neq 0$, namely the existence of Anderson localized type solutions for the nonlinear Schrödinger equation (1.1) when $0<\varepsilon,\delta\ll 1$, under appropriate conditions on $\bm{\alpha}$ and $\bm{\theta}$. Since both $\varepsilon$ and $\delta$ are small, we start from the (decoupled) equation
$iu_{t}+V(\bm{n}\bm{\alpha}+\bm{\theta})\delta_{\bm{n},\bm{n}^{\prime}}u=0.$ (1.6)
It has solutions of the form
$u^{(0)}(t,\bm{n})=\sum_{\bm{n}\in\mathbb{Z}^{d}}a_{\bm{n}}e^{i\mu_{\bm{n}}t},$ (1.7)
where
$\mu_{\bm{n}}={V(\bm{n}\bm{\alpha}+\bm{\theta})}\geq 1,$
and $a_{\bm{n}}$ decays rapidly. The solutions (1.7), in general, have infinite number of frequencies and are almost periodic in time.
We study the persistence of the above type of solutions with finite (but arbitrary) number of frequencies in time, i.e., the quasi-periodic in time solutions. Denoting the frequency by $\bm{\omega}$, $\bm{\omega}=(\omega_{1},\omega_{2},\cdots,\omega_{b})$, as an Ansatz, we seek solutions to (1.1) in the form of a convergent series:
$u(t,\bm{n})=\sum_{(\bm{k},\bm{n})\in\mathbb{Z}^{b}\times\mathbb{Z}^{d}}\hat{u}(\bm{k},\bm{n})e^{i\bm{k}\cdot\bm{\omega}t}.$ (1.8)
Note that solutions of the above form are consistent with the nonlinearity in (1.1).
Denote by ${\rm meas}(\cdot)$ the Lebesgue measure of a set. We have
Theorem 1.1.
Fix (any) $\{\bm{n}_{l}\}_{l=1}^{b}\subset\mathbb{Z}^{d}$ and let
$u^{(0)}(t,\bm{n})=\sum_{l=1}^{b}a_{l}e^{i\omega_{l}^{(0)}t}\delta_{\bm{n}_{l},\bm{n}},$ (1.9)
be a solution to (1.6), where
$\bm{a}=(a_{l})_{l=1}^{b}\in[1,2]^{b},$
and
$\bm{\omega}^{(0)}=(\omega_{l}^{(0)})_{l=1}^{b}=\left({V(\bm{n}_{l}\bm{\alpha}+\bm{\theta})}\right)_{l=1}^{b}\in[-\|V\|_{\infty},\|V\|_{\infty}]^{b}.$
Then for $0<\varepsilon\lesssim\delta\ll 1$, there is a set $\mathcal{W}\subset[0,1]^{2d}$ in $(\bm{\alpha},\bm{\theta})$, satisfying ${\rm meas}([0,1]^{2d}\setminus\mathcal{W})\leq\log^{-c}\frac{1}{\varepsilon+\delta}$ $(c=c(b,d,K)>0)$, such that for any $(\bm{\alpha},\bm{\theta})\in\mathcal{W}$, there is a set in $\bm{a}$, $\mathcal{R}=\mathcal{R}_{\bm{\alpha},\bm{\theta}}\subset[1,2]^{b}$ of ${\rm meas}([1,2]^{b}\setminus\mathcal{R})\leq(\varepsilon+\delta)^{c}$, so that the following holds true: If $\bm{a}\in\mathcal{R}$, then there is $\bm{\omega}=\bm{\omega}(\bm{a})$ satisfying $|\bm{\omega}-\bm{\omega}^{(0)}|\lesssim\delta$ so that
$u(t,\bm{n})=\sum_{\bm{k}\in\mathbb{Z}^{b}}\hat{u}(\bm{k},\bm{n})e^{i\bm{k}\cdot\bm{\omega}t}$
is a solution to (1.1). Moreover, one has
[…]
$\sum_{(\bm{k},\bm{n})\notin\mathcal{S}_{+}}|\hat{u}(\bm{k},\bm{n})|e^{\rho(|\bm{k}|+|\bm{n}|)}$ $<\sqrt{\varepsilon+\delta},\ \rho>0,$
where $\mathcal{S}_{+}=\{(\bm{e}_{l},\bm{n}_{l})_{l=1}^{b}\}$ with $\bm{e}_{l}$ the standard basis vectors (i.e., $e_{l}(j)=\delta_{l,j},j=1,\cdots,b$) for $\mathbb{Z}^{b}.$
Remark 1.2.
Theorem 1.1 readily generalizes to potentials $V$, which are given by finite Fourier series with both cosine and sine terms.
Remark 1.3.
Note that we do not require the linear equation (1.5) to satisfy Anderson localization. However, linear Anderson localization does hold on a large subset of $\mathcal{W}$.
1.1. Ideas of the proof
The multi-dimensional phase space, namely $\bm{\theta}$ being a vector instead of a scalar as in e.g., [SW23] poses a major challenge. Diophantine properties of the eigenvalues of the linear Schrödinger operator (1.4) plays an essential role in the proof. To $O(\delta)$, this could be replaced by Diophantine properties of the potential $V(\bm{n}\bm{\alpha}+\bm{\theta})$ at different lattice sites $\bm{n}\in\mathbb{Z}^{d}$. We need to show, that the values of the function at these chosen sites are, in some sense, linearly independent. This is, however, a priori, not obvious, since $V$ is a given function on $\mathbb{R}^{d}$. To prove linear independence, we use a generalized Wronskian and bound the determinant away from zero by carefully removing some $(\bm{\alpha},\bm{\theta})$. This method is applicable to any trigonometric polynomials, generalizing the approach in [SW23] for the cosine function on $\mathbb{R}$. It is independent of the main body of the proof and could be of independent interest. Such arguments may be closely related to Diophantine approximations on manifold developed by Kleinbock and Margulis [KM98].
The other main ingredient of the proof is Bourgain’s approach to Anderson localization of linear quasi-periodic Schrödinger operators in $d$-dimensions [Bou07], such as the $\mathcal{H}$ in (1.4). This seems to be, so far, the only method available to deal with the $d$-dimensional phase space, i.e., $\bm{\theta}\in[0,1]^{d}$, $d\geq 1$. Among other innovations, it uses semi-algebraic geometry arguments (e.g., Yomdin-Gromov algebraic lemma) to control the resonances. The difficulty of $d$-dimensional phase space manifests as well in our paper, in that for the linear analysis, semi-algebraic geometry already seems indispensable to control the resonances; while in previous works, see [BW08, Wan16, LW22, GSW23, SW23], this role is filled by Diophantine properties. To control the resonances in time, we also need a weak second Melnikov-type estimates. This could be readily obtained, similar to [BW08, LW22, GSW23, SW23]. The proof is then accomplished by doing multi-scale analysis as in the aforementioned papers.
We mention the work of Yuan [Yua02] in which a delicate KAM scheme was first developed to prove the existence of quasi-periodic solutions for some nonlinear discrete equations without external parameters (i.e., regarding the initial states as parameters). Later, Geng-You-Zhao [GYZ14] also applied the KAM method to obtain the existence of quasi-periodic solutions for the nonlinear quasi-periodic Schrödinger equation on $\mathbb{Z}$.
1.2. The nonlinear random Schrödinger equation
When the potential $V$ with $V(\bm{n})=V(\bm{n}\bm{\alpha}+\bm{\theta})$ is replaced by a random potential, for example, with $V(\bm{n})$, a family of independently identically distributed random variables, it is known from [BW08, GSW23] that the nonlinear random Schrödinger equation:
$iu_{t}+\varepsilon\Delta u+Vu+\delta|u|^{2p}u=0,\,\text{on }\mathbb{Z}^{d},$
has large sets of Anderson localized states for small $\varepsilon$ and $\delta$. When $d=1$, it is shown further in [LW22] that large sets of Anderson localized states persist for all $\varepsilon\neq 0$. (Similar results could be proven for the nonlinear wave equation.) Recall that Anderson localization for the linear random Schrödinger equation,
$iu_{t}+\varepsilon\Delta u+Vu=0,\,\text{on }\mathbb{Z}^{d},$
for small $\varepsilon$ was first proven in [FS83], see also [AM93].
It is generally believed that the linear random Schrödinger equation and the linear quasi-periodic Schrödinger equation should share common features, despite the quasi-periodic problems being more delicate and the results more difficult to obtain. The use of semi-algebraic geometry and Cartan’s Lemma [BGS02, Bou07, JLS20], for example, certainly renders the quasi-periodic issue more involved than that of the random, and this seems to continue to the nonlinear case, as mentioned earlier. Theorem 1.1 generalizes this circle of ideas to the nonlinear setting by providing a concrete example where the quasi-periodic and the random have indeed similar behavior.
1.3. Structure of the paper
In sect. 2, we briefly present the general scheme. Sect. 3 makes linear estimates, in the form of large deviation theorems. The paper concludes by constructing the quasi-periodic in time solutions in sect. 4. The proof of the Diophantine-type estimates is presented in the Appendix.
2. The general scheme
We work directly with the nonlinear equation (1.1). Let
$\mu_{\bm{n}}=V(\bm{n}\bm{\alpha}+\bm{\theta})\geq 1.$
Then
[…]
Using the Ansatz, we obtain the nonlinear lattice equation
${\rm diag}_{(\bm{k},\bm{n})}(-\bm{k}\cdot\bm{\omega}+\mu_{\bm{n}})\hat{u}+\varepsilon\Delta\hat{u}+\delta(\hat{u}*\hat{v})^{*p}*\hat{u}=0,$ (2.1)
where $\hat{v}(\bm{k},\bm{n})=\overline{\hat{u}}(-\bm{k},\bm{n})$, the convolution is only in the $\bm{k}$-variables:
$\hat{u}*\hat{v}(\bm{k},\bm{n})=\sum_{\bm{k}_{1}+\bm{k}_{2}=\bm{k}}\hat{u}(\bm{k}_{1},\bm{n})\hat{v}(\bm{k}_{2},\bm{n})$
and
$(\hat{u}*\hat{v})^{*p}=\underbrace{(\hat{u}*\hat{v})*\cdots*(\hat{u}*\hat{v})}_{p\ {\rm times}};$
while the Laplacian, $\Delta$ acts only on the $\bm{n}$-variables. To solve for $\hat{u},\hat{v}$, we also need a conjugate equation of (2.1)
${\rm diag}_{(\bm{k},\bm{n})}(\bm{k}\cdot\bm{\omega}+\mu_{\bm{n}})\hat{v}+\varepsilon\Delta\hat{v}+\delta(\hat{u}*\hat{v})^{*p}*\hat{v}=0.$
As a result, we obtain the following vector-valued nonlinear lattice equation
$F(\vec{u})=0,\ \vec{u}=\binom{\hat{u}}{\hat{v}},$ (2.2)
where
$F(\vec{u})=H_{0}\vec{u}+\delta P(\vec{u})$
with
$H_{0}={\rm diag}_{(\bm{k},\bm{n})\in\mathbb{Z}^{b+d}}$ $\left(\begin{array}[]{cc}{-\bm{k}\cdot\bm{\omega}+\mu_{\bm{n}}+\varepsilon\Delta}&{0}\\ {0}&{\bm{k}\cdot\bm{\omega}+\mu_{\bm{n}}+\varepsilon\Delta}\\ \end{array}\right)$ (2.5)
and
$P(\vec{u})=\binom{(\hat{u}*\hat{v})^{*p}*\hat{u}}{(\hat{u}*\hat{v})^{*p}*\hat{v}}.$
The linearized operator of $P$ at $\vec{u}$ is then given by
${T}_{\vec{u}}=\left(\begin{array}[]{cc}{(p+1)(\hat{u}*\hat{v})^{*p}*}&{p(\hat{u}*\hat{v})^{*(p-1)}*\hat{u}*\hat{u}*}\\ {p(\hat{u}*\hat{v})^{*(p-1)}*\hat{v}*\hat{v}*}&{(p+1)(\hat{u}*\hat{v})^{*p}*}\\ \end{array}\right).$ (2.8)
2.1. The Lyapunov-Schmidt decomposition
We look for quasi-periodic solutions to (1.1) near $u^{(0)}$, using the Lyapunov-Schmidt decomposition. In view of (2.2), we define the lattice:
$\mathbb{Z}^{b+d}_{{\rm pm}}=\mathbb{Z}^{b+d}\times\{+,-\}.$
Recall from Theorem 1.1, the Fourier support of $u^{(0)}$, the set $\mathcal{S}_{+}=\{(\bm{e}_{l},\bm{n}_{l})_{l=1}^{b}\}\subset\mathbb{Z}^{b+d}$. We consequently identify $\mathcal{S}_{+}$ with $\{(\bm{e}_{l},\bm{n}_{l})_{l=1}^{b}\}\times\{+\}\subset\mathbb{Z}^{b+d}_{{\rm pm}}$. Similarly, let $\mathcal{S}_{-}=\{(-\bm{e}_{l},\bm{n}_{l})_{l=1}^{b}\}\subset\mathbb{Z}^{b+d}$ be the Fourier support of ${\bar{u}}^{(0)}$, which is then identified with $\{(-\bm{e}_{l},\bm{n}_{l})_{l=1}^{b}\}\times\{-\}\subset\mathbb{Z}^{b+d}_{{\rm pm}}$.
Let
$\mathcal{S}=\mathcal{S}_{+}\cup\mathcal{S}_{-},$
and write
$\mathbb{Z}^{b+d}_{{\rm pm},*}=\mathbb{Z}^{b+d}_{{\rm pm}}\setminus\mathcal{S}:=\mathcal{S}^{c}.$
We divide the equation (2.2) into the $P$-equations
${F}\big{|}_{\mathbb{Z}^{b+d}_{{\rm pm},*}}\left(\vec{u}\right)=0,$ (2.9)
and the $Q$-equations
${F}\big{|}_{\mathcal{S}}(\vec{u})=0.$ (2.10)
We use (2.9) to solve for $\vec{u}|_{\mathbb{Z}^{b+d}_{{\rm pm},*}}$ using a Newton scheme. On $\mathcal{S}$, $\vec{u}$ is held fixed, $\vec{u}_{\mathcal{S}}=\vec{u}_{\mathcal{S}}^{(0)}$, the $Q$-equations (2.10) are viewed instead as equations for $\bm{\omega}$, and will be solved using the implicit function theorem. Solving the $P$-equations requires analyzing the invertibility of the linearized operator $H=H_{0}+\delta T_{\vec{u}}$ restricted to $\mathbb{Z}^{b+d}_{{\rm pm},*}$, which we do in the next section.
3. Linear analysis
In this section, we investigate properties of linearized operators. Of particular importance is the large deviation theorem (LDT) for the corresponding Green’s functions.
[…]
In the following, we study on $\mathbb{Z}^{b+d}_{{\rm pm},*}$ the operator
${H}(\sigma)={D}(\sigma)+\varepsilon(\Delta\oplus\Delta)+\delta S,\ \sigma\in\mathbb{R},$
where
${D}(\sigma)$ $={\rm diag}_{(\bm{k},\bm{n})\in\mathbb{Z}^{b+d}}\left(\begin{array}[]{cc}{-\sigma-\bm{k}\cdot\bm{\omega}+\mu_{\bm{n}}}&{0}\\ {0}&{\sigma+\bm{k}\cdot\bm{\omega}+\mu_{\bm{n}}}\\ \end{array}\right),$
and the operator $S$ refers typically to the linearized operator of the nonlinear perturbation $P(\vec{u})$ in the previous section. So we may let $S$ satisfy the following:
  • First
    $S:\ \ell^{2}(\mathbb{Z}^{b+d}_{{\rm pm}})\to\ell^{2}(\mathbb{Z}^{b+d}_{{\rm pm}})$
    is a bounded self-adjoint operator.
  • For all $\bm{k},\bm{k}^{\prime},\bm{k}^{\prime\prime}\in\mathbb{Z}^{b}$, $\bm{n},\bm{n}^{\prime}\in\mathbb{Z}^{d}$ and $\xi,\xi^{\prime}\in\{+,-\}$, we have the Töplitz property in the $\bm{k}$-variables:
    $S((\bm{k}^{\prime}+\bm{k},\bm{n},\xi);(\bm{k}^{\prime\prime}+\bm{k},\bm{n}^{\prime},\xi^{\prime}))=S((\bm{k}^{\prime},\bm{n},\xi);(\bm{k}^{\prime\prime},\bm{n}^{\prime},\xi^{\prime})).$
  • We have for some $C>0,\gamma>0,$
    $|S((\bm{k},\bm{n},\xi);(\bm{k}^{\prime},\bm{n}^{\prime},\xi^{\prime}))|\leq Ce^{-\gamma|\bm{k}-\bm{k}^{\prime}|-\gamma|\bm{n}|}\delta_{\bm{n},\bm{n}^{\prime}}.$
The frequency $\bm{\omega}$,
$\bm{\omega}\in\Omega=\bm{\omega}^{(0)}+[-C\delta,C\delta]^{b},$ (3.1)
for some $C>0.$
Denote by $\Lambda_{N}$,
$\Lambda_{N}=\{(\bm{k},\bm{n})\in\mathbb{Z}^{b+d}:\ |(\bm{k},\bm{n})|\leq N\},$
the cube of radius $N$ and center $\bm{0}$.
We introduce the elementary regions on $\mathbb{Z}^{b+d}$. These sets were first introduced in [BGS02] in the $2$-dimensional case, which can be easily extended to higher dimensions. For $\bm{i}\in\mathbb{Z}^{b+d}$ and $\bm{w}\in\mathbb{R}_{+}^{b+d}$, let $R_{\bm{w}}(\bm{i})=\{\bm{j}\in\mathbb{Z}^{b+d}:\ |j_{l}-i_{l}|\leq w_{l},\,l=1,2,\cdots,b+d\}$ denote the rectangle. An elementary region is defined to be a set $\Lambda$ of the form
$\Lambda=R_{\bm{w}}(\bm{i})\setminus(R_{\bm{w}}(\bm{i})+\bm{z}),$
where $\bm{z}\in\mathbb{Z}^{b+d}$ is arbitrary. The size of $\Lambda$, denoted by diam $\text{diam }\Lambda$, is simply its diameter. The set of all elementary regions of size $M$ will be denoted by $\mathcal{ER}(M)$. Elements of $\mathcal{ER}(M)$ are also referred to as $M$-regions. Define
$\mathcal{ER}_{\bm{0}}(M)=\left\{\Lambda\in\mathcal{ER}(M):\ \Lambda=R_{\bm{w}}(\bm{0})\setminus(R_{\bm{w}}(\bm{0})+\bm{z}),\text{ diam }\Lambda=M\right\}.$
We now define Green’s functions on subsets of $\mathbb{Z}_{{\rm pm},*}^{b+d}$. Let $\Lambda\subset\mathbb{Z}_{{\rm pm},*}^{b+d}$ be a non-empty set, $R_{\Lambda}$ the restriction operator to $\Lambda$, and define the $|\Lambda|\times|\Lambda|$ matrix:
$(R_{\Lambda}{H}(\sigma)R_{\Lambda})((\bm{j},\xi);(\bm{j}^{\prime},\xi^{\prime}))={H}(\sigma)((\bm{j},\xi);(\bm{j}^{\prime},\xi^{\prime})),\ (\bm{j},\xi),(\bm{j}^{\prime},\xi^{\prime})\in\Lambda.$ (3.2)
Define the inverse, the Green’s function, if it exists,
$G_{\Lambda}(\sigma)=(R_{\Lambda}{H}(\sigma)R_{\Lambda})^{-1}.$
Denote by $\Pi_{b,d}\Lambda$ the projection of $\Lambda\subset\mathbb{Z}_{{\rm pm},*}^{b+d}$ onto $\mathbb{Z}^{b+d}$.
Before addressing the LDT for Green’s functions, let us first give the definition of large deviation estimates (LDE):
Definition 3.1.
Let $10<\rho_{1}<\rho_{2}<\rho_{3}<$ 1, $0<\gamma^{\prime}<\gamma$ and $M\in\mathbb{N}$. We say ${H}(\sigma)$ satisfies the $(\rho_{1},\rho_{2},\rho_{3},\gamma^{\prime},M)$-LDE if there exists a set $\Sigma_{M}\subset\mathbb{R}$ with
${\rm meas}(\Sigma_{M})\leq e^{-M^{\rho_{1}}},$
so that for $\sigma\notin\Sigma_{M}$ the following estimates hold: If $\Lambda\subset\mathbb{Z}_{{\rm pm},*}^{b+d}$ is such that $\Pi_{b,d}\Lambda\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(M)$ satisfies $|\bm{n}|\leq 2M$, then
$\|G_{\Lambda}(\sigma)\|$ $\leq e^{M^{\rho_{2}}},$
$|G_{\Lambda}(\sigma)((\bm{j},\xi);(\bm{j}^{\prime},\xi^{\prime})))|$ $\leq e^{-\gamma^{\prime}|\bm{j}-\bm{j}^{\prime}|}\ {\rm for}\ |\bm{j}-\bm{j}^{\prime}|\geq M^{\rho_{3}}.$
where $\bm{j}=(\bm{k},\bm{n}),\bm{j}^{\prime}=(\bm{k}^{\prime},\bm{n}^{\prime})\in{\Pi_{b,d}\Lambda}$, and $\xi,{\xi^{\prime}}\in\{+,-\}$.
We will show in this section that the $(\rho_{1},\rho_{2},\rho_{3},\gamma^{\prime},N)$-LDE hold for all $N\gg 1$, under certain non-resonant conditions on $\bm{\alpha},\bm{\theta}$ and $\bm{\omega}$. This leads to the LDT, Theorem 3.6 below. Since we view $\bm{\omega}$ as an $O(\delta)$ perturbation of $\bm{\omega}^{(0)}$, the properties of $\bm{\omega}^{(0)}$ become essential in the proof of LDT. So, in the following, we first establish non-resonant properties (called Diophantine estimates) of $\bm{\omega}^{(0)}$. Then the proof of LDT will follow from a multi-scale analysis scheme in the spirit of [BGS02, Bou07].
3.1. Diophantine estimates
Recall that
$\bm{\omega}^{(0)}$ $=\bm{\omega}^{(0)}(\bm{\alpha},\bm{\theta})=(V(\bm{n}_{l}\bm{\alpha}+\bm{\theta}))_{l=1}^{b},$
$\mu_{\bm{n}}$ $=\mu_{\bm{n}}(\bm{\alpha},\bm{\theta})=V(\bm{n}\bm{\alpha}+\bm{\theta}),\ \bm{n}\in\mathbb{Z}^{d}.$
The main purpose of this section is to obtain lower bounds on
$|\bm{k}\cdot\bm{\omega}^{(0)}|\ {\rm for}\ \bm{k}\in\mathbb{Z}^{b}\setminus\{0\},$
$|\bm{k}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}}|\ {\rm for}\ (\bm{k},\bm{n})\in\mathbb{Z}^{b+d}\backslash{\mathcal{S}_{-}},$
$|\bm{k}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}}-\mu_{\bm{n}^{\prime}}|\ {\rm for}\ \bm{n}\neq\bm{n}^{\prime}\in\mathbb{Z}^{d},$
under certain restrictions on $(\bm{\alpha},\bm{\theta}).$ As mentioned in the Introduction, due to the $d$-dimensional phase space, such Diophantine-type estimates are highly non-trivial. We use a generalized Wronskian approach and work out the details in the Appendix. These estimates can be read independently, and are key to the existence of quasi-periodic solutions to the nonlinear equation.
Fix
$\bm{n}^{\prime}_{1},\cdots,\bm{n}^{\prime}_{m}\in\mathbb{Z}^{d}$
to be $m$, $m\geq 2$, distinct lattice points, and write
$\tilde{\bm{\omega}}=(V(\bm{n}^{\prime}_{l}\bm{\alpha}+\bm{\theta}))_{l=1}^{m}\in\mathbb{R}^{m}.$
In this section, all constants $C,c>0$ are assumed to depend only on
$\tau_{m}=m\cdot(\#\Gamma_{K})=mK^{d}2^{d-1},$ (3.3)
$\max_{\ell}|\bm{n}^{\prime}_{\ell}|$ and $|\bm{v}|$. For convenience, we do not make this dependence explicit. We have
Theorem 3.2.
Let $0<\iota\ll 1$, and define
$\ \ \ {\rm DC}_{\rm nls}(\iota)=\left\{(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}:\ |\bm{k}\cdot\tilde{\bm{\omega}}|\geq\frac{\iota}{|\bm{k}|^{4\tau_{m}^{2}}}\ {\rm for}\ \forall\ \bm{k}\in\mathbb{Z}^{m}\setminus\{\bm{0}\}\right\}.$
Then
${\rm meas}({\rm DC}_{\rm nls}(\iota))\geq 1-C\iota^{\frac{1}{10d\tau_{m}^{2}}}.$
Proof.
This is the same as Theorem A.1 in the Appendix, see the proof there. These estimates are closely related to Diophantine approximations on manifold in the celebrated work of Kleinbock and Margulis [KM98]. ∎
Remark 3.1.
To our knowledge, using transversality properties to make Diophantine approximations on manifolds was first developed by Pyartli [Pya69]. This method was later introduced into the study of KAM theory [XYQ97, Eli02, Bam03]. It turns out that this idea also plays a key role in the Craig-Wayne-Bourgain type argument used in the present paper. In addition, we can handle Diophantine estimates on general quasi-periodic polynomials on $\mathbb{T}^{d}$. The main difference between [XYQ97, Eli02, Bam03] and the present work is that we have no prior transversality estimate, which requires much additional effort.
In the following application, we may set $m=2,b,b+1,b+2$ in the above theorem. Fix any $L>\max_{l=1,\cdots,b}|\bm{n}_{l}|$. We shall consider finite scale Diophantine conditions. Let $\mathcal{J}_{1}$ be the set of all $(\bm{k},\bm{n},\bm{n}^{\prime})\in\mathbb{Z}^{b}\times\mathbb{Z}^{d}\times\mathbb{Z}^{d}$ satisfying
$|\bm{k}|\leq 2L,\ \bm{n}\neq\bm{n}^{\prime},\ |(\bm{n},\bm{n}^{\prime})|\leq L,\ (\bm{n},\bm{n}^{\prime})\notin\{{\bm{n}_{l}}\}_{l=1}^{b}\times\{\bm{n}_{l}\}_{l=1}^{b};$
$\mathcal{J}_{2}$ is defined to be the set of $(\bm{k},\bm{n},\bm{n}^{\prime})\in\mathbb{Z}^{b}\times\mathbb{Z}^{d}\times\mathbb{Z}^{d}$ with
[…]
For a given scale $L$, let $\Lambda_{L}=[-L,L]^{b+d}$, we define
Definition 3.3.
$\mathcal{M}_{1}=\bigcap_{|(\bm{n},\bm{n}^{\prime})|\leq L,\bm{n}\neq\bm{n}^{\prime}}\left\{(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}:\ |\mu_{\bm{n}}-\mu_{\bm{n}^{\prime}}|>\iota\right\},$
[…]
$\mathcal{M}_{3}=\bigcap_{(\bm{k},\bm{n})\in\Lambda_{L}\setminus\mathcal{S_{-}}}\left\{(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}:\ |\bm{k}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}}|>\frac{\iota}{(2L)^{4\tau_{b+1}^{2}}}\right\},$
$\mathcal{M}_{4}=\bigcap_{(\bm{k},\bm{n},\bm{n}^{\prime})\in(\mathcal{J}_{1}\cup\mathcal{J}_{2})}\left\{(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}:\ |\bm{k}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}}-\mu_{\bm{n}^{\prime}}|>\frac{\iota}{(2L)^{4\tau_{b+2}^{2}}}\right\},$
and
$\mathcal{M}=\bigcap_{1\leq i\leq 4}\mathcal{M}_{i}.$
In anticipation of the upcoming multi-scale analysis, we take
$L=L_{\varepsilon,\delta}=(\varepsilon+\delta)^{-\frac{1}{10^{3}b^{2}d^{2}\tau_{b+2}^{2}}},$ (3.4)
and
$\iota=(\varepsilon+\delta)^{\frac{1}{10b}}.$
Remark 3.2.
We should note at this point that the exponent in (3.4) is less than $1$, which is, roughly speaking, the scales where one could replace $\bm{\omega}$ by $\bm{\omega}^{(0)}$ in the analysis. The precise exponent, however, is rather technical, and will only become clear later in Theorem 3.4 and the proofs in sect. 3.3. The factor $(\varepsilon+\delta)$ appears because we perturb from the decoupled equation (1.6).
Applying Theorem 3.2 using (3.4) then leads to
Theorem 3.4.
For $0<\varepsilon+\delta\ll 1,$ there is a set $\mathcal{M}=\mathcal{M}_{\varepsilon,\delta}\subset[0,1]^{2d}$ with
${\rm meas}([0,1]^{2d}\setminus\mathcal{M})\leq(\varepsilon+\delta)^{\frac{1}{10^{3}bd\tau_{b+2}^{2}}}$ (3.5)
so that for $(\bm{\alpha},\bm{\theta})\in\mathcal{M},$ the following holds true.
  • (1)
    For all $(\bm{k},\bm{n})\in\Lambda_{L_{\varepsilon,\delta}}\setminus\mathcal{S_{-}}$, we have
    $|\bm{k}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}}|>(\varepsilon+\delta)^{\frac{1}{8b}}.$ (3.6)
  • (2)
    We have
    $\sup_{\sigma\in\mathbb{R}}\#\left\{(\bm{k},\bm{n})\in\Lambda_{L_{\varepsilon,\delta}}:\ \min_{\xi=\pm 1}|\xi(\sigma+\bm{k}\cdot\bm{\omega}^{(0)})+\mu_{\bm{n}}|<\frac{(\varepsilon+\delta)^{\frac{1}{8b}}}{4}\right\}\leq 2b.$ (3.7)
    where #()#\#($\cdot$)# ( ⋅ ) denotes the cardinality of a set, as before.
Proof.
The proof is similar to that in [SW23]. First, the measure estimate (3.5) is a consequence of Theorem 3.2 and the definition of $\mathcal{M}.$ Next, the inequality (3.6) in (1) follows directly from the definition of $\mathcal{M}_{3}.$ So, it suffices to establish (3.7) of (2). We prove it by contradiction. Without loss of generality, we consider the case $\xi=+1$. Assume that there are $b+1$ distinct $\{(\bm{k}_{l},\bm{n}_{l}^{\prime})\}_{1\leq l\leq b+1}$ satisfying
$|\sigma+\bm{k}_{l}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}_{l}^{\prime}}|<\frac{(\varepsilon+\delta)^{\frac{1}{8b}}}{4}.$ (3.8)
We claim that all $\bm{k}_{l}$ ($11\leq l\leq b+$ 1) are distinct. Actually, if there are $\bm{k}_{l_{1}}=\bm{k}_{l_{2}}$ for some $11\leq l_{1}\neq l_{2}\leq b+$ 1, then it must be that $\bm{n}_{l_{1}}^{\prime}\neq\bm{n}_{l_{2}}^{\prime}$. As a result, we obtain using (3.8) that
$|\mu_{\bm{n}_{l_{1}}^{\prime}}-\mu_{\bm{n}_{l_{2}}^{\prime}}|$ $\leq|\sigma+\bm{k}_{l_{1}}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}_{l_{1}}^{\prime}}|$
$\ \ \ +|\sigma+\bm{k}_{l_{2}}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}_{l_{2}}^{\prime}}|$
$\leq\frac{(\varepsilon+\delta)^{\frac{1}{8b}}}{2},$
which contradicts the definition of $\mathcal{M}_{1}$. Next, we claim that $\{\bm{n}_{l}^{\prime}\}_{1\leq l\leq b+1}\subset\{\bm{n}_{l}\}_{1\leq l\leq b}.$ In fact, if $\bm{n}_{l_{1}}^{\prime}\notin\{\bm{n}_{l}\}_{1\leq l\leq b}$ for some $11\leq l_{1}\leq b+$ 1, we can assume that there is $l_{2}\neq l_{1}$ so that $\bm{n}_{l_{2}}^{\prime}\neq\bm{n}_{l_{1}}^{\prime}$. Otherwise, we may have
$|(\bm{k}_{l_{1}}-\bm{k}_{l_{2}})\cdot\bm{\omega}^{(0)}|\leq\frac{(\varepsilon+\delta)^{\frac{1}{8b}}}{2},$
which contradicts the definition of $\mathcal{M}_{2}$, since $\bm{k}_{l_{1}}\neq\bm{k}_{l_{2}}$ as claimed above. So we obtain for some $l_{2}\neq l_{1}$ and $11\leq l_{2}\leq b=$ 1 with $\bm{n}_{l_{2}}^{\prime}\neq\bm{n}_{l_{1}}^{\prime}$ that
$|(\bm{k}_{l_{1}}-\bm{k}_{l_{2}})\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}_{l_{1}}^{\prime}}-\mu_{\bm{n}_{l_{2}}^{\prime}}|\leq\frac{(\varepsilon+\delta)^{\frac{1}{8b}}}{2},$
which contradicts the definition of $\mathcal{M}_{4}$. So it suffices to assume $\{\bm{n}_{l}^{\prime}\}_{1\leq l\leq b+1}\subset\{\bm{n}_{l}\}_{1\leq l\leq b}.$ However, in this case, we have by the pigeonhole principle that there are $11\leq l_{1}\neq l_{2}\leq b+$ 1 so that $\bm{n}_{l_{1}}^{\prime}=\bm{n}_{l_{2}}^{\prime}$. Then we get
$|(\bm{k}_{l_{1}}-\bm{k}_{l_{2}})\cdot\bm{\omega}^{(0)}|\leq\frac{(\varepsilon+\delta)^{\frac{1}{8b}}}{2},$
which contradicts the definition of $\mathcal{M}_{2}$ as shown above. This proves (3.7). ∎
3.2. Bourgain’s geometric lemma
To proceed with the analysis, we will also need Anderson localization-type properties for the linear quasi-periodic Schrödinger operator $\mathcal{H}$ in (1.4). This was proven by Bourgain in the remarkable work [Bou07], where he established Anderson localization for general quasi-periodic Schrödinger operators on $\mathbb{Z}^{d}$ with analytic potentials on $\mathbb{T}^{d}$. This result was later significantly extended by Jitomirskaya-Liu-Shi [JLS20] (cf. also [Liu22]) to potentials on $\mathbb{T}^{b}$ for $b\geq d.$ In the present, due to the $d$-dimensional phase space, we need to apply Bourgain-type arguments. Using short-range property of $S$ in the $\bm{n}$-variables, Bourgain’s geometric lemma is made available to handle LDE for Green’s functions at large scales. This is a main new aspect not present in [SW23].
Denote by $\mathcal{ER}_{\mathbb{Z}^{d}}(M)$ the set of all elementary $M$-regions in $\mathbb{Z}^{d}$. We have
Lemma 3.5 (cf. Claim (p. 694) in [Bou07] and Theorem 2.7 in [JLS20]).
For $0<\varepsilon<1,$ there is $N_{0}\in\mathbb{N}$, so that for every $N>N_{0}$, $N\in\mathbb{N}$, there is $\mathcal{A}_{N}\subset[0,1]^{d}$ satisfying
${\rm meas}(\cap_{N\geq N_{0}}\mathcal{A}_{N})=1-O(\log^{-c}\frac{1}{\varepsilon})\ {\rm as}\ \varepsilon\to 0,c>0,$
so that the following holds true for $\bm{\alpha}\in\mathcal{A}_{N}:$ There exist $10<\kappa_{1}<\kappa_{2}<$ 1 (depending only on $d$) so that, for all $E\in\mathbb{R}$, all $\bm{\theta}\in[0,1]^{d}$, there is $L\in[N^{\kappa_{1}},N^{\kappa_{2}}]$ so that, if $L_{1}\sim(\log N)^{C},\ C=C(d)>1$, then for all $\Lambda_{1}\in\mathcal{ER}_{\mathbb{Z}^{d}}(L_{1})$ satisfying
$\Lambda_{1}\subset[-L,L]^{d}\setminus[-L^{\frac{1}{10d}},L^{\frac{1}{10d}}]^{d},$
one has
$\|\mathcal{H}^{-1}_{\Lambda_{1}}(E;\bm{\theta})\|$ $\leq e^{\sqrt{L_{1}}},$
$|\mathcal{H}^{-1}_{\Lambda_{1}}(E;\bm{\theta})(\bm{n};\bm{n}^{\prime})|$ $\leq e^{-\frac{1}{2}|\log\varepsilon|\cdot|\bm{n}-\bm{n}^{\prime}|}\ {\rm for}\ |\bm{n}-\bm{n}^{\prime}|\geq\frac{L_{1}}{10},$
where
$\mathcal{H}_{\Lambda_{1}}(E;\bm{\theta})$ $=R_{\Lambda_{1}}\left(\varepsilon\Delta+\mu_{\bm{n}}\delta_{\bm{n},\bm{n}^{\prime}}\right)R_{\Lambda_{1}}-E$
$:=\mathcal{L}_{\Lambda_{1}}(\bm{\theta})-E.$
Denote
$\mathcal{A}=\cap_{N\geq N_{0}}\mathcal{A}_{N}.$ (3.9)
Remark 3.3.
The proof of this lemma is based on analysis of semi-algebraic sets (cf. [Bou07]), and applies to general non-degenerate analytic potentials.
Remark 3.4.
We mention that the set $\mathcal{W}$ on which Theorem 1.1 holds satisfies
$\mathcal{W}=\{\mathcal{A}\times[0,1]^{d}\}\cap\mathcal{M},$
where $\mathcal{M}$ is the set on which the Diophantine estimates in Theorem 3.4 holds.
3.3. Large deviation theorem
In this subsection, we establish that the LDE, see Definition 3.1, hold for all sufficiently large scales by a multi-scale analysis. From (3.1), however, the variation of the frequency, $\bm{\omega}$ is only $O(\delta)$ and not $O(1)$, hence as we will see shortly, the proof of LDT will be accomplished in three steps instead of the usual two:
  • The first step deals with the small scales, for which we establish LDE for all $\bm{\omega}\in\Omega$ and scales
    $N_{0}\leq N\lesssim\log^{\frac{1}{\rho_{1}}}\frac{1}{\varepsilon+\delta},$
    for some $\rho_{1}>0.$ In this step, we only use the Neumann series argument, which leads to the above scales (cf. Lemma 3.7).
  • Next, we establish intermediate scales LDE, i.e., scales in the range
    $\log^{\frac{1}{\rho_{1}}}\frac{1}{\varepsilon+\delta}\lesssim N\leq(\varepsilon+\delta)^{-\frac{1}{10^{3}b^{2}d^{2}\tau_{b+2}^{2}}},$
    for all $\bm{\omega}\in\Omega$. Note that Theorem 3.4 is available in the above interval of scales. The proof of such intermediate scales LDE is based on preparation type theorem together with clustering properties of the spectrum of the diagonal matrix $D(0)$ (cf. (3.7), Theorem 3.4).
    Remark 3.5.
    The intermediate scales are needed here because for a fixed $(\bm{\alpha},\bm{\theta})$, hence fixed $\bm{\omega}^{(0)}$, $\bm{\omega}$ only varies in an interval of size $O(\delta)$, see (3.1), as mentioned above. Consequently, there is a gap in the range of scales before Bourgain-type analysis becomes applicable. Fortunately, the Diophantine estimates for $(\bm{\alpha},\bm{\theta})\in\mathcal{M}$ can control the resonances directly as in (3.7), leading to LDE for the range of scales in the gap, which we term the intermediate scales.
  • For the large scales (i.e., $N>(\varepsilon+\delta)^{-\frac{1}{10^{3}b^{2}d^{2}\tau_{b+2}^{2}}}$) LDE, we will apply matrix-valued Cartan’s lemma and semi-algebraic sets theory of [BGS02, Bou07]. Of particular importance is Bourgain’s geometric lemma (cf. Lemma 3.5). This step requires both the Diophantine condition and the so called weak second Melnikov’s condition on $\bm{\omega}.$
    To be coherent, since the variation in $\bm{\omega}$ is only $O(\delta)$, the excisied measure in $\Omega$ needs to be less than $\delta^{b}$. So we make a particular choice of the exponent and impose the following Diophantine condition:
    $\ \ \ {\rm DC}_{\bm{\omega}}(N)$
[…]
    Then we have for $\varepsilon\lesssim\delta^{\frac{1}{2b}}$,
    ${\rm meas}\left(\Omega\setminus\left(\bigcap_{N>(\varepsilon+\delta)^{-\frac{1}{10^{3}b^{2}d^{2}\tau_{b+2}^{2}}}}{\rm DC}_{\bm{\omega}}(N)\right)\right)\ll\delta^{b}.$ (3.11)
    The second Melnikov’s condition is imposed in order to restrict the time-$\bm{k}$ directions where resonances could possibly occur. This greatly reduces the number of resonances, and paves the way for the application of Lemma 3.13.
    Remark 3.6.
    The condition on the size of $\varepsilon$ relative to $\delta$ stems again from the modulation in $\bm{\omega}$ being $O(\delta)$; while the excised measure depends on both $\varepsilon$ and $\delta$, as we will see in the proof.
The central theorem of this section is
Theorem 3.6 (LDT).
Fix $\bm{\alpha}\in\mathcal{A}$ and $(\bm{\alpha},\bm{\theta})\in\mathcal{M}$ (cf. (3.9) and Theorem 3.4). Let $10<\rho_{1}<\rho_{2}<\rho_{3}<$ 1 satisfy $\rho_{1}<\frac{\kappa_{1}\rho_{2}^{2}}{30d}$ and $\rho_{2}>\frac{2}{3}$ ($\kappa_{1}$ is given in Lemma 3.5). Then for every $N\geq N_{0}(b,d,\rho_{1},\rho_{2},\rho_{3},\kappa_{1},\max_{1\leq l\leq b}|\bm{n}_{l}|)\gg 1$, there exists a semi-algebraic set $\widetilde{\Omega}_{N}\subset\Omega$, independent of $\bm{a}$, of ${\rm meas}(\Omega\setminus\widetilde{\Omega}_{N})\leq e^{-N^{\zeta}}$ and $\deg\widetilde{\Omega}_{N}\leq N^{C}$ for some $\zeta=\zeta(b,d,\kappa_{1},\rho_{2})\in(0,1)$ and $C=C(b,d)>1,$ so that if we define $\Omega_{N}$ to be
$\Omega_{N}$ =$=\left\{\begin{array}[]{ll}\Omega&{\rm if}\ N_{0}\leq N\leq(\varepsilon+\delta)^{-\frac{1}{10^{3}b^{2}d^{2}\tau_{b+2}^{2}}},\\ {\rm DC}_{\bm{\omega}}(N)\cap\widetilde{\Omega}_{N}&{\rm if}\ N>(\varepsilon+\delta)^{-\frac{1}{10^{3}b^{2}d^{2}\tau_{b+2}^{2}}},\end{array}\right.$
then for $\bm{\omega}\in\Omega_{N}$, ${H}(\sigma)$ satisfies the $(\rho_{1},\rho_{2},\rho_{3},\gamma_{N},N)$-LDE with $\gamma_{N}=\gamma-N^{-\kappa}$ for some $\kappa=\kappa(b,d,\rho_{1},\rho_{2},\rho_{3})>0$.
Remark 3.7.
Note that the sets $\Omega_{N}$ are independent of $\bm{a}$. This will be important for later nonlinear applications.
Remark 3.8.
One may observe that this LDT is for a fixed operator; while the linearized operator is varying in a Newton scheme. This issue, however, can be remedied by using that the Green’s function estimates in LDE (at scale $N$) are stable under perturbations of order $e^{-N^{2}}$. The sup-exponentially decaying error (cf. (4.8)) in the Newton scheme can ensure this. For more details, see sect. V, F. Step 3 and the proof of Proposition 3.2 in [KLW24].
This theorem will be proved in detail in the following three sections.
3.3.1. The small scales LDE
This section is devoted to the proof of LDE for small scales, i.e.,
$N_{0}\leq N\leq c\log^{\frac{1}{\rho_{1}}}\frac{1}{\varepsilon+\delta},$
where $N_{0}\gg 1$ depends on $b,d,\rho_{1}$ and $c\in(0,1)$. In this case, we only use the standard Neumann series. As a result, we establish LDE for all $\bm{\omega}\in\Omega.$ We have
Lemma 3.7.
Let $10<\rho_{1}<$ 1 and
$\ \ \ \Sigma_{N}$
$=\bigcup_{\Pi_{b,d}\Lambda\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N),\ |\bm{n}|\leq 2N}\left\{\sigma\in\mathbb{R}:\ \min_{(\bm{k},\bm{n},\xi)\in\Lambda}|\xi(\sigma+\bm{k}\cdot\bm{\omega})+\mu_{\bm{n}}|\leq e^{-2N^{\rho_{1}}}\right\}.$
Then for $N_{0}\leq N\leq c\log^{\frac{1}{\rho_{1}}}\frac{1}{\varepsilon+\delta}$,
${\rm meas}(\Sigma_{N})\leq e^{-N^{\rho_{1}}},$ (3.12)
where $N_{0}>0$ only depends on $b,d,\rho_{1}.$ Moreover, if $\sigma\notin\Sigma_{N}$, we have for all $\Lambda\subset\mathbb{Z}_{{\rm pm},*}^{b+d}$ satisfying $\Pi_{b,d}\Lambda\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N)$ with $|\bm{n}|\leq 2N$,
$\|G_{\Lambda}(\sigma)\|$ $\leq 2e^{2N^{\rho_{1}}},$ (3.13)
$|G_{\Lambda}(\sigma)((\bm{j},\xi);(\bm{j}^{\prime},\xi^{\prime}))|$ $\leq 2e^{2N^{\rho_{1}}}e^{-\gamma|\bm{j}-\bm{j}^{\prime}|}\ {\rm for}\ \bm{j}\neq\bm{j}^{\prime}.$ (3.14)
Remark 3.9.
In this lemma, we make no restrictions on $\bm{\omega}.$
Proof.
We first establish the measure bound (3.12). For $N\geq N_{0}$, we obtain
${\rm meas}(\mathbb{R}\setminus\Sigma_{N})$ $\leq C(b,d)N^{C(b,d)}e^{-2N^{\rho_{1}}}$
$\leq e^{-N^{\rho_{1}}},$
where in the last inequality we use $N\geq N_{0}(b,d,\rho_{1})\gg 1.$ Next, if $\sigma\notin\Sigma_{N}$, then (3.13) and (3.14) follow directly from the standard Neumann series (cf. the proof of Theorem 4.3 in [JLS20]) since $e^{-2N^{\rho_{1}}}\gg\varepsilon+\delta$ if $N\leq c\log^{\frac{1}{\rho_{1}}}\frac{1}{\varepsilon+\delta}$.
3.3.2. The intermediate scales LDE
This section and the next deal with intermediate scales LDE. We will show that the LDE hold true for all $\bm{\omega}\in\Omega$ and scales in the range
$\log^{\frac{1}{\rho_{1}}}\frac{1}{\varepsilon+\delta}\lesssim N\leq(\varepsilon+\delta)^{-\frac{1}{10^{3}b^{2}d^{2}\tau_{b+2}^{2}}}.$
In this case, the analysis in section 3.1 becomes essential.
Lemma 3.8.
Assume $(\bm{\alpha},\bm{\theta})\in\mathcal{M}$ (cf. Theorem 3.4). Fix $01>\rho_{3}>\rho_{2}>2\rho_{1}>$ 0 and
$\log^{\frac{1}{\rho_{1}}}\frac{1}{\varepsilon+\delta}\lesssim N\leq(\varepsilon+\delta)^{-\frac{1}{10^{3}b^{2}d^{2}\tau_{b+2}^{2}}}.$
Then for each $\bm{\omega}\in\Omega$, there exists a set $\Sigma_{N}\subset\mathbb{R}$ with
${\rm meas}(\Sigma_{N})\leq(\varepsilon+\delta)^{-\frac{9}{10}}e^{-\frac{1}{2b}N^{\frac{\rho_{2}}{2}}},$
so that if $\sigma\notin\Sigma_{N}$, then for all $\Lambda\subset\mathbb{Z}_{{\rm pm},*}^{b+d}$ satisfying $\Pi_{b,d}\Lambda\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N)$ with $|\bm{n}|\leq 2N$,
$\|G_{\Lambda}(\sigma)\|$ $\leq e^{N^{\rho_{2}}},$
$|G_{\Lambda}(\sigma)((\bm{j},\xi);(\bm{j}^{\prime},\xi^{\prime}))|$ $\leq e^{-\gamma_{N}|\bm{j}-\bm{j}^{\prime}|}\ {\rm for}\ |\bm{j}-\bm{j}^{\prime}|\geq N^{\rho_{3}},$
where
$\gamma_{N}=\gamma-N^{-\kappa}$
with $\kappa=\kappa(b,d,\rho_{2},\rho_{3})>0.$
Remark 3.10.
Note that this lemma also holds for all $\bm{\omega}\in\Omega$.
Proof.
The proof is based on clustering property (cf. (2) of Theorem 3.4), Schur’s completement argument and Rouché theorem. It can be decomposed into two steps.
Step 1. We define
$\Sigma_{N}=\bigcup_{\sqrt{N}\leq N^{\prime}\leq N,\ \Pi_{b,d}\Lambda\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N^{\prime}),\ |\bm{n}|\leq 2N^{\prime}}\Sigma_{\Lambda},$ (3.15)
where
$\Sigma_{\Lambda}=\left\{\sigma\in\mathbb{R}:\ \|G_{\Lambda}(\sigma)\|\geq e^{({\rm diam}\ \Lambda)^{\rho_{2}}}\right\}.$
We first show that
$\Sigma_{\Lambda}\subset\bigcup_{\xi=\pm 1,\bm{i}=(\bm{k},\bm{n})\in\Pi_{b,d}\Lambda}I_{\xi,\bm{i}},$ (3.16)
where
$I_{\xi,\bm{i}}=\left\{\sigma\in\mathbb{R}:\ |\xi(\sigma+\bm{k}\cdot\bm{\omega}^{(0)})+\mu_{\bm{n}}|\leq\frac{1}{100}(\varepsilon+\delta)^{\frac{1}{8b}}\right\}.$
In fact, if $\sigma\notin\bigcup\limits_{(\xi,\bm{i})\in\Lambda}I_{\xi,\bm{i}},$ then for each $\bm{\omega}\in\Omega,$
$\ \ \ \ \inf_{(\xi,\bm{i})\in\Lambda}|\xi(\sigma+\bm{k}\cdot\bm{\omega})+\mu_{\bm{n}}|$
$\geq\inf_{(\xi,\bm{i})\in\Lambda}|\xi(\sigma+\bm{k}\cdot\bm{\omega}^{(0)})+\mu_{\bm{n}}|-\sup_{\bm{k}\in\Lambda}|\bm{k}\cdot(\bm{\omega}-\bm{\omega}^{(0)})|$
$\geq\frac{1}{200}(\varepsilon+\delta)^{\frac{1}{8b}}.$
Since $\rho_{2}>2\rho_{1}$ and $N\geq c\log^{\frac{1}{\rho_{1}}}\frac{1}{\varepsilon+\delta}$, using Neumann series argument implies that
$\|G_{\Lambda}(\sigma)\|\leq C(\varepsilon+\delta)^{-\frac{1}{4b}}<e^{({\rm diam}\ \Lambda)^{\rho_{2}}},$
when $\sigma\notin\bigcup\limits_{(\xi,\bm{i})\in\Lambda}I_{\xi,\bm{i}}.$ This proves (3.16) and
$\Sigma_{\Lambda}=\bigcup_{(\xi,\bm{i})\in\Lambda}(\Sigma_{\Lambda}\cap I_{\xi,\bm{i}}).$
In the following, we estimate ${\rm meas}(\Sigma_{\Lambda}\cap I_{\xi,\bm{i}}).$ For this purpose, we fix any $\xi^{*}\in\{+,-\}$, $\bm{i}^{*}=(\bm{k}^{*},\bm{n}^{*})\in\Pi_{b,d}\Lambda$ and denote
$I^{*}=I_{\xi*,\bm{i}^{*}}=\left[\sigma^{*}-\frac{1}{100}(\varepsilon+\delta)^{\frac{1}{8b}},\sigma^{*}+\frac{1}{100}(\varepsilon+\delta)^{\frac{1}{8b}}\right],$
where
$\sigma^{*}=-\bm{k}^{*}\cdot\bm{\omega}^{(0)}-\frac{1}{\xi^{*}}\mu_{\bm{n}^{*}}.$ (3.17)
We will study $G_{\Lambda}(\sigma)$ for $\sigma\in I^{*}.$ From (2) of Theorem 3.4, we have for every $\sigma^{*}\in\mathbb{R}$, there exists $B_{*}=B_{*}^{+}\cup B_{*}^{-}\subset\Lambda$ with #$\#B_{*}^{\pm}\leq b$ so that
$\inf_{(\xi,\bm{i})\in\Lambda\setminus B_{*}}|\xi(\sigma^{*}+\bm{k}\cdot\bm{\omega}^{(0)})+\mu_{\bm{n}}|\geq\frac{(\varepsilon+\delta)^{\frac{1}{8b}}}{4}.$
Similar to the proof of (3.16), if
$z\in\mathbb{D}_{(\varepsilon+\delta)^{\frac{1}{8b}}/10}(\sigma^{*})=\{z\in\mathbb{C}:\ |z-\sigma^{*}|\leq(\varepsilon+\delta)^{\frac{1}{8b}}/10\},$
then
$\inf_{(\xi,\bm{i})\in\Lambda\setminus B_{*}}|\xi(z+\bm{k}\cdot\bm{\omega})+\mu_{\bm{n}}|\geq\frac{(\varepsilon+\delta)^{\frac{1}{8b}}}{8}.$
So we have an analytic extension of ${H}_{\Lambda}(z)$ for $z\in\mathbb{D}_{(\varepsilon+\delta)^{\frac{1}{8b}}/10}(\sigma^{*})$ and
$\sup_{\bm{j},\bm{j}^{\prime}\in\Pi_{b,d}\Lambda,\ z\in\mathbb{D}_{(\varepsilon+\delta)^{\frac{1}{8b}}/10}(\sigma^{*})}|{D}_{\Lambda}(z)((\bm{j},\xi);(\bm{j}^{\prime},\xi^{\prime}))|\leq C(b)(\varepsilon+\delta)^{-\frac{1}{10^{3}b^{2}d^{2}\tau_{b+2}^{2}}}.$
More importantly, if we denote $\Lambda^{c}=\Lambda\setminus B_{*}$, then by Neumann series argument, we have for all $z\in\mathbb{D}_{(\varepsilon+\delta)^{\frac{1}{8b}}/10}(\sigma^{*}),$
$\|G_{\Lambda^{c}}(z)\|\leq(\varepsilon+\delta)^{-\frac{1}{4b}}.$
In the following, we estimate $G_{\Lambda}(z)$ using the Schur complement reduction. We define
$S(z)=R_{B_{*}}{H}(z)R_{B_{*}}-R_{B_{*}}{H}(z)R_{\Lambda^{c}}G_{\Lambda^{c}}(z)R_{\Lambda^{c}}{H}(z)R_{B_{*}}.$
Obviously, we have
$R_{B_{*}}{H}(z)R_{B_{*}}$ $=R_{B_{*}}{D}(z)R_{B_{*}}+O(\varepsilon+\delta),$ (3.18)
and
$R_{B_{*}}{H}(z)R_{\Lambda^{c}}G_{\Lambda^{c}}(z)R_{\Lambda^{c}}{H}(z)R_{B_{*}}=O((\varepsilon+\delta)^{2-\frac{1}{4b}}).$ (3.19)
Now we consider $\det S(z)$. Combining (3.18) and (3.19) yields
$\ \ \ \det S(z)$
$=\prod_{\bm{i}=(\bm{k},\bm{n})\in B_{*}^{+}}(z+\bm{k}\cdot\bm{\omega}+\mu_{\bm{n}})\prod_{\bm{i}=(\bm{k},\bm{n})\in B_{*}^{-}}(-z-\bm{k}\cdot\bm{\omega}+\mu_{\bm{n}})$
$\ \ \ +O((\varepsilon+\delta)^{3/4})$
$\sim\prod_{l=1}^{\#B_{*}}(z-\sigma_{l})+O((\varepsilon+\delta)^{3/4}).$
Recalling the definition of $B_{*}$, we must have
$|z-\sigma_{l}|\leq\frac{4}{5}(\varepsilon+\delta)^{\frac{1}{8b}}.$
From (3.17), there exists at least one $\sigma_{l}$ so that $\sigma_{l}=\sigma^{*}.$ At this stage, we need a useful result.
Lemma 3.9.
There is $1\leq l_{*}\leq 5b$ so that
$\left(\mathbb{D}_{l_{*}}(\sigma^{*})\setminus\mathbb{D}_{l_{*}-1}(\sigma^{*})\right)\cap\{\sigma_{l}\}_{1\leq l\leq\#B_{*}}=\emptyset.$
where for $0\leq l\leq 5b$,
$\mathbb{D}_{l}(\sigma^{*})=\mathbb{D}_{(\frac{1}{10}+\frac{l}{50b})(\varepsilon+\delta)^{\frac{1}{8b}}}(\sigma^{*}).$
Proof of Lemma 3.9.
It suffices to note that
$\mathbb{D}_{5b}(\sigma^{*})\setminus\mathbb{D}_{0}(\sigma^{*})=\bigcup_{l=1}^{5b}\left(\mathbb{D}_{l}(\sigma^{*})\setminus\mathbb{D}_{l-1}(\sigma^{*})\right).$
The proof then follows from the pigeonhole principle since #$\#B_{*}\leq 2b$. ∎
From Lemma 3.9, we can restrict our considerations on $\mathbb{D}_{1}$ with
$\mathbb{D}_{1}=\mathbb{D}_{(\frac{1}{10}+\frac{1}{100b}+\frac{l_{*}-1}{50b})(\varepsilon+\delta)^{\frac{1}{8b}}}(\sigma^{*})$
Then
$\mathbb{D}_{\frac{1}{10}(\varepsilon+\delta)^{\frac{1}{8b}}}(\sigma^{*})\subset\mathbb{D}_{l_{*}-1}(\sigma^{*})\subset\mathbb{D}_{1}\subset\mathbb{D}_{\frac{1}{5}(\varepsilon+\delta)^{\frac{1}{8b}}}(\sigma^{*}).$
Let
$K=\{1\leq l\leq\#B_{*}:\ \sigma_{l}\in\mathbb{D}_{l_{*}-1}(\sigma^{*})\}.$
We have $K\neq\emptyset$ and for $l\not\in K$,
$\sigma_{l}\notin\mathbb{D}_{1},\ {\rm dist}(\sigma_{l},\partial\mathbb{D}_{1})\geq\frac{1}{100b}(\varepsilon+\delta)^{\frac{1}{8b}}.$
As a result, we can write for $z\in\mathbb{D}_{1},$
$\frac{\det S(z)}{\prod_{l\notin K}(z-\sigma_{l})}\sim{\prod_{l\in K}(z-\sigma_{l})}+O((\varepsilon+\delta)^{\frac{1}{2}}).$ (3.20)
It is easy to see that
${\prod_{l\in K}|z-\sigma_{l}|}\big{|}_{\partial\mathbb{D}_{1}}\geq(\frac{1}{100b})^{\#K}(\varepsilon+\delta)^{\frac{\#K}{8b}}\geq c(\varepsilon+\delta)^{1/4}\gg C(\varepsilon+\delta)^{\frac{1}{2}}.$
Applying Rouché theorem shows that the function $\frac{\det S(z)}{\prod_{l\notin K}(z-\sigma_{l})}$ has exactly $K$ zeros (denoted by $z_{l},\ l\in K$) in $\mathbb{D}_{1}.$ By (3.20), we also get
$|\sigma_{l}-z_{l}|=O((\varepsilon+\delta)^{\frac{1}{2}}).$ (3.21)
So on $\mathbb{D}_{1}$, we have
$\frac{\det S(z)}{\prod_{l\notin K}(z-\sigma_{l})}=g(z)\prod_{l\in K}(z-z_{l}),$ (3.22)
where $g$ is analytic on $\mathbb{D}_{1}$ with $\inf\limits_{z\in\mathbb{D}_{1}}|g(z)|>0.$ Using (3.21) and (3.22), we have
$g\big{|}_{\partial\mathbb{D}_{1}}\sim 1+O((\varepsilon+\delta)^{\frac{1}{4}}).$
The maximum principle then leads to
$g\big{|}_{\mathbb{D}_{1}}\sim 1+O((\varepsilon+\delta)^{\frac{1}{4}}).$
We have established on $\mathbb{D}_{1},$
$\det S(z)\sim\prod_{l\notin K}(z-\sigma_{l})\prod_{l\in K}(z-z_{l}).$
In particular, one has for all $z\in\mathbb{D}_{1},$
$|\det S(z)|$ $\geq C(b)(\varepsilon+\delta)^{\frac{\#B_{*}-\#K}{8b}}\prod_{l\in K}|z-z_{l}|$
$\geq(\varepsilon+\delta)^{\frac{1}{4}}\prod_{l\in K}|z-z_{l}|,$
assuming $0<\varepsilon+\delta\leq c(b)\ll 1$ Thus combining Hadamard’s inequality and Cramer’s rule implies for $z\in\mathbb{D}_{1},$
$\|S^{-1}(z)\|$ $\leq C(b)(\varepsilon+\delta)^{-\frac{2b}{10^{3}b^{2}d^{2}\tau_{b+2}^{2}}}(\varepsilon+\delta)^{-\frac{1}{4}}\prod_{l\in K}|z-z_{l}|^{-1}$
$\leq(\varepsilon+\delta)^{-\frac{1}{3}}\prod_{l\in K}|z-z_{l}|^{-1}.$
Applying Schur’s complement reduction shows that for $z\in\mathbb{D}_{1},$
$\|G_{\Lambda}(z)\|$ $\leq 4(1+\|G_{\Lambda^{c}}(z)\|)^{2}(1+\|S^{-1}(z)\|)$
$\leq(\varepsilon+\delta)^{-\frac{7}{8}}\prod_{l\in K}|z-z_{l}|^{-1}.$
Recall that $I^{*}\subset\mathbb{D}_{1}$. So for $\sigma\in I^{*},$ we have
$\|G_{\Lambda}(\sigma)\|$ $\leq(\varepsilon+\delta)^{-\frac{7}{8}}\prod_{l\in K}|\sigma-z_{l}|^{-1},$
and
$\Sigma_{\Lambda}\cap I^{*}\subset\bigcup_{l\in K}\left\{\sigma\in I^{*}:\ |\sigma-z_{l}|\leq(\varepsilon+\delta)^{-\frac{7}{8}}e^{-\frac{1}{2b}({\rm diam}\ \Lambda)^{\rho_{2}}}\right\}.$
This implies
${\rm meas}(\Sigma_{\Lambda}\cap I^{*})\leq C(\varepsilon+\delta)^{-\frac{7}{8}}e^{-\frac{1}{2b}({\rm diam}\ \Lambda)^{\rho_{2}}},$
and subsequently
${\rm meas}(\Sigma_{\Lambda})\leq(\varepsilon+\delta)^{-\frac{8}{9}}e^{-\frac{1}{2b}({\rm diam}\ \Lambda)^{\rho_{2}}},$
which combined with (3.15) shows
${\rm meas}(\Sigma_{N})\leq(\varepsilon+\delta)^{-\frac{9}{10}}e^{-\frac{1}{2b}N^{\frac{\rho_{2}}{2}}}.$
Step 2. In this step, we prove the exponential off-diagonal decay of $G_{\Lambda}(\sigma)((\bm{j},\xi);(\bm{j}^{\prime},\xi^{\prime}))$ assuming $\sigma\not\in\Sigma_{N}$ and $\Pi_{b,d}\Lambda\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N)$ with $|\bm{n}|\leq 2N.$ By Theorem 3.4, we know that there exists $B\subset\Lambda$ so that
$\inf_{(\xi,\bm{k},\bm{n})\in\Lambda\setminus B}|\xi(\sigma+\bm{k}\cdot\bm{\omega})+\mu_{\bm{n}}|\geq\frac{(\varepsilon+\delta)^{\frac{1}{8b}}}{8}.$
Neumann series yields that for any $\Lambda^{\prime}\subset\Lambda$ with $\Lambda^{\prime}\cap B=\emptyset,$
$\|G_{\Lambda^{\prime}}(\sigma)\|$ $\leq 50(\varepsilon+\delta)^{-\frac{1}{4b}},$ (3.23)
$|G_{\Lambda^{\prime}}(\sigma)((\bm{j},\xi);(\bm{j}^{\prime},\xi^{\prime}))|$ $\leq 50(\varepsilon+\delta)^{-\frac{1}{4b}}e^{-\gamma|\bm{j}-\bm{j}^{\prime}|}\ {\rm for}\ \bm{j}\neq\bm{j}^{\prime}.$ (3.24)
We say that $\mathcal{ER}(\sqrt{N})\ni\Pi_{b,d}Q,\ Q\subset\Lambda$ is $\sqrt{N}$-regular if both (3.23) and (3.24) hold with $\Lambda^{\prime}=Q.$ Otherwise, $Q$ is called $\sqrt{N}$-singular. Let $\mathcal{F}$ be any family of pairwise disjoint $\sqrt{N}$-singular sets contained in $\Lambda$. We obtain
#$\#\mathcal{F}\leq 2b.$
We have established the sublinear bound. Then applying the coupling lemma of [HSSY20] (cf. Lemma 4.2) shows
$|G_{\Lambda}(\sigma)((\bm{j},\xi);(\bm{j}^{\prime},\xi^{\prime}))|\leq e^{-\gamma_{N}|\bm{j}-\bm{j}^{\prime}|}\ {\rm for}\ |\bm{j}-\bm{j}^{\prime}|>N^{\rho_{3}},$
where
$\gamma_{N}=\gamma-N^{-\kappa},\ \kappa=\kappa(b,d,\rho_{2},\rho_{3})>0.$
This proves the exponential off-diagonal decay estimates.
We have completed the proof of Lemma 3.8.
3.3.3. The large scales LDE
In this section, we will prove the LDE for
$(\varepsilon+\delta)^{-\frac{1}{10^{3}b^{2}d^{2}\tau_{b+2}^{2}}}<N<\infty.$
We have
Lemma 3.10.
Let $(\bm{\alpha},\bm{\theta})\in\mathcal{M}$ (cf. Theorem 3.4), $\bm{\alpha}\in\mathcal{A}$ (cf. Remark 3.3) and $N$ satisfy
$(\varepsilon+\delta)^{-\frac{1}{10^{3}b^{2}d^{2}\tau_{b+2}^{2}}}<N<\infty.$
Let $10<\rho_{1}<\rho_{2}<\rho_{3}<$ 1 satisfy $\rho_{2}>\frac{2}{3}$ and $\rho_{1}<\frac{\kappa_{1}\rho_{2}^{2}}{30d}$ ($\kappa_{1}$ is given in Lemma 3.5). Then there is a semi-algebraic set $\widetilde{\Omega}_{N}\subset\Omega$ with ${\rm meas}(\Omega\setminus\widetilde{\Omega}_{N})\leq e^{-N^{\zeta}}$ and $\deg(\widetilde{\Omega}_{N})\leq N^{C}$ for some $\zeta=\zeta(b,d,\rho_{2},\kappa_{1})\in(0,1),C=C(b,d)>1$ so that the following holds for $\bm{\omega}\in{\rm DC}_{\bm{\omega}}(N)\cap\widetilde{\Omega}_{N}$: There is a set $\Sigma_{N}\subset\mathbb{R}$ with ${\rm meas}(\Sigma_{N})\leq e^{-N^{\rho_{1}}}$, so that if $\sigma\notin\Sigma_{N}$, then for all $\Lambda\subset\mathbb{Z}_{{\rm pm},*}^{b+d}$ satisfying $\Pi_{b,d}\Lambda\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N)$ with $|\bm{n}|\leq 2N,$
$\|G_{\Lambda}(\sigma)\|$ $\leq e^{N^{\rho_{2}}},$
$|G_{\Lambda}(\sigma)((\bm{j},\xi);(\bm{j}^{\prime},\xi^{\prime}))|$ $\leq e^{-\gamma_{N}|\bm{j}-\bm{j}^{\prime}|}\ {\rm for}\ |\bm{j}-\bm{j}^{\prime}|\geq N^{\rho_{3}},$
where $\gamma_{N}=\gamma-N^{-\kappa}$ for some $\kappa=\kappa(b,d,\rho_{1},\rho_{2},\rho_{3})>0.$
To perform the multi-scale scheme, we need two small scales $N_{1}<N_{2}<N$ with
$N_{2}=N_{1}^{\widetilde{C}},\ N=N_{2}^{\widetilde{C}},$
where $\widetilde{C}>1$ will be specified below. For $\Lambda\subset\mathbb{Z}_{{\rm pm},*}^{b+d}$ with $\Pi_{b,d}\Lambda\in\mathcal{ER}(K)$, we say $\Lambda$ is $\sigma$-good if
$\|G_{\Lambda}(\sigma)\|$ $\leq e^{K^{\rho_{2}}},$
$|G_{\Lambda}(\sigma)(\bm{j},\xi);(\bm{j}^{\prime},\xi^{\prime}))|$ $\leq e^{-\gamma_{K}|\bm{j}-\bm{j}^{\prime}|}\ {\rm for}\ |\bm{j}-\bm{j}^{\prime}|\geq K^{\rho_{3}},$
where $\gamma_{K}=\gamma-K^{-\kappa}$, and $\sigma$-bad, if one of the inequalities is violated.
The proof of Lemma 3.10 is based on matrix-valued Cartan’s lemma (cf. Lemma 3.13 below) and iterations of resolvent identities. To apply Cartan’s lemma, one has to show the sublinear bound on $\sigma$-bad $\Lambda$ with $\Pi_{b,d}\Lambda\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ contained in $\Lambda_{N}$: If $|\bm{n}|\leq 2N_{1}$, the sublinear bound can be established via the Töplitz property of $H$ in the $\bm{k}$-direction together with the semi-algebraic sets theory (cf. Lemma 3.11); for $|\bm{n}|>2N_{1}$, we will use the short-range property and Bourgain’s geometric lemma to prove the sublinear bound. Once the sublinear bound is derived, one can use the Cartan’s lemma and resolvent identities to finish the proof of large scales LDE.
We need the following ingredients to prove Lemma 3.10.
Lemma 3.11.
Assume that LDE holds true at scale $N_{1}.$ There is $C>1$ depending only on $b,d$ so that if $\widetilde{C}^{2}\geq\frac{C}{\rho_{2}}$, then the number of $\sigma$-bad $\Lambda$ with $\Pi_{b,d}\Lambda\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ satisfying $|\bm{k}|\leq N$ and $|\bm{n}|\leq 2N_{1}$ is at most ${N^{\rho_{2}/4}}/2.$
Proof.
First, we define $\widetilde{\Sigma}_{N_{1}}$ to be the set of $\sigma\in\mathbb{R}$ so that, there is $\Lambda$ with $\Pi_{b,d}\Lambda\in\mathcal{ER}_{\bm{0}}(N_{1})$ which is $\sigma$-bad. We then obtain using LDE at scale $N_{1},$
${\rm meas}(\widetilde{\Sigma}_{N_{1}})\leq CN_{1}^{C}e^{-N_{1}^{\rho_{1}}}<e^{-\frac{1}{2}N_{1}^{\rho_{1}}}.$
From the standard reduction using Hilbert-Schmidt norm and Cramer’s rule, we can regard $\widetilde{\Sigma}_{N_{1}}$ as a semi-algebraic set of degree $\deg\widetilde{\Sigma}_{N_{1}}\leq N^{C}$ with $C=C(b,d)>0.$ Then using Basu-Pollack-Roy Theorem [BPR96] on Betti numbers of a semi-algebraic set (cf. also Proposition 9.2 of [Bou05]), we have a decomposition of
$\widetilde{\Sigma}_{N_{1}}=\bigcup_{1\leq l\leq N_{1}^{C}}I_{l},$
where each $I_{l}$ is an interval of length $|I_{l}|\leq e^{-\frac{1}{2}N_{1}^{\rho_{1}}}.$ Now assume that $\Pi_{b,d}\Lambda\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ and $\Pi_{b,d}\Lambda^{\prime}\in(\bm{k}^{\prime},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ and $\Lambda,\Lambda^{\prime}$ are all $\sigma$-bad for $|\bm{n}|\leq 2N_{1},\bm{k}\neq\bm{k}^{\prime}$ and $|\bm{k}|,|\bm{k}^{\prime}|\leq N$. Then from Töplitz property of ${H}(\sigma)$ in the $\bm{k}$-direction, we obtain
$\sigma+\bm{k}\cdot\bm{\omega},\ \sigma+\bm{k}^{\prime}\cdot\bm{\omega}\in\widetilde{\Sigma}_{N_{1}}.$
As a result, by $\bm{k}\neq\bm{k}^{\prime}$ with $|\bm{k}|,|\bm{k}^{\prime}|\leq N$ and $\bm{\omega}\in{\rm DC}_{\bm{\omega}}(N)$, we have
${e^{-\frac{1}{2}N_{1}^{\rho_{1}}}}\ll N^{-10^{4}b^{4}d^{2}\tau_{b+2}^{2}}\leq|(\bm{k}-\bm{k}^{\prime})\cdot\bm{\omega}|,$
which implies $\bm{k}$ and $\bm{k}^{\prime}$ can not stay in the same interval $I_{l}$ for $1\leq l\leq N_{1}^{C}.$ Thus we have shown that the number of $\sigma$-bad $\Lambda$ with $\Pi_{b,d}\Lambda\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ with $|\bm{n}|\leq 2N_{1}$ and $|\bm{k}|\leq N$ is at most
$CN_{1}^{C}N_{1}^{d}\leq N_{1}^{C}\sim N^{\frac{C}{\widetilde{C}^{2}}}\leq N^{\rho_{2}/4}/2,$
assuming $\widetilde{C}^{2}\geq\frac{C}{\rho_{2}}.$ This proves Lemma 3.11. ∎
Next, we will deal with $|\bm{n}|>2N_{1}$ by combining short-range property and Bourgain’s geometric lemma (cf. Lemma 3.5). For $\Lambda\subset\mathbb{Z}_{{\rm pm},*}^{b+d}$, denote
$\Lambda_{*}=\Pi_{b,d}\Lambda\subset\mathbb{Z}^{b+d}.$
Write further $\Lambda_{*}=\Pi_{b}\Lambda_{*}\times\Pi_{d}\Lambda_{*}$ with $\Pi_{b}\Lambda_{*}$ (resp. $\Pi_{d}\Lambda_{*}$) denoting the projection of $\Lambda_{*}$ on $\mathbb{Z}^{b}$ (resp. $\mathbb{Z}^{d}$). For each $\bm{k}\in\Pi_{b}\Lambda_{*},$ define
$\Pi_{d}\Lambda_{*}(\bm{k})=\{\bm{n}\in\mathbb{Z}^{d}:\ (\bm{k},\bm{n})\in\Lambda_{*}\}\subset\mathbb{Z}^{d}.$
We have
Lemma 3.12.
Assume that $\bm{\alpha}\in\mathcal{A}$. Then
  • (1)
    Fix $\Lambda_{*}\in(\bm{k}^{\prime},\bm{n}^{\prime})+\mathcal{ER}_{\bm{0}}(N_{1})$ with $|\bm{n}^{\prime}|>2N_{1}$. There is a collection of $\{\lambda_{l}\geq\frac{1}{2}\}_{1\leq l\leq N_{1}^{2d}}$ depending only on $\Lambda,\bm{\alpha},\bm{\theta},V$ so that, if
    $\min_{\bm{k}\in\Pi_{b}\Lambda_{*},1\leq l\leq N_{1}^{2d},\xi=\pm 1}|\sigma+\bm{k}\cdot\bm{\omega}+\xi{\lambda_{l}}|>e^{-\frac{1}{4}N_{1}^{\kappa_{1}\rho_{2}}},$
    then $\Lambda$ is $\sigma$-good, where $\kappa_{1}$ is given in Lemma 3.5.
  • (2)
    There exists some $\Sigma_{N_{2}}\subset\mathbb{R}$ with ${\rm meas}(\Sigma_{N_{2}})\leq e^{-\frac{1}{5}N_{2}^{\kappa_{1}\rho_{2}}}$ so that if $\sigma\notin\Sigma_{N_{2}}$, then each $\Lambda$ satisfying $\Lambda_{*}\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{2})$ with $|(\bm{k},\bm{n})|\leq 10N$ and $|\bm{n}|>2N_{2}$ is $\sigma$-good.
Proof.
(1) Fix $\bm{k}\in\Pi_{b}\Lambda_{*}$. For any $Q\subset\Lambda_{\bm{k},*}:=\Pi_{d}\Lambda_{*}(\bm{k})\in\mathcal{ER}_{\mathbb{Z}^{d}}(N_{1}),$ write
$A_{\bm{k}}$ $=R_{Q}({D}(\sigma)+\varepsilon(\Delta\oplus\Delta))R_{Q}$
$=(-\sigma-\bm{k}\cdot\bm{\omega}+\mathcal{L}_{Q}(\bm{\theta}))\oplus(\sigma+\bm{k}\cdot\bm{\omega}+\mathcal{L}_{Q}(\bm{\theta})),$
where $\mathcal{L}_{Q}(\bm{\theta})$ is defined in Lemma 3.5. Then the set of all eigenvalues of $A_{\bm{k}}$ is given by
{±$\{\pm(\sigma+\bm{k}\cdot\bm{\omega})+\lambda_{l}(\bm{k})\}_{l=1}^{\#Q},$
where $\{\lambda_{l}(\bm{k})\}_{l=1}^{\#Q}$ denotes the spectrum of $\mathcal{L}_{Q}(\bm{\theta}).$ So we have
$\|A_{\bm{k}}^{-1}\|$ $\leq\max_{\xi=\pm 1,1\leq l\leq\#Q}|\xi(\sigma+\bm{k}\cdot\bm{\omega})+\lambda_{l}(\bm{k})|^{-1}.$
We now use Lemma 3.5 and resolvent identities. Since $\bm{\alpha}\in\mathcal{A}\subset\mathcal{A}_{N_{1}}$, applying Lemma 3.5 with $E_{\bm{k}}=\sigma+\bm{k}\cdot\bm{\omega}$ similar to [JLS20] (cf. p. 465, the proof of Theorem 2.7) gives the following: For each $\bm{m}\in\Lambda_{\bm{k},*}$, there exist $\frac{1}{4}N_{1}^{\kappa_{1}}\leq\tilde{N}_{1}\leq N_{1}^{\kappa_{2}}$, $Q_{\bm{m}}\in\mathcal{ER}_{\mathbb{Z}^{d}}(\tilde{N}_{1})$ and $\tilde{Q}_{\bm{m}}$, such that
$Q_{\bm{m}}\subset\bm{m}+[-\tilde{N}_{1},\tilde{N}_{1}]^{d},\ \bm{m}+[-\tilde{N}_{1}^{\frac{1}{10d}},\tilde{N}_{1}^{\frac{1}{10d}}]^{d}\subset\tilde{Q}_{\bm{m}},$
$\bm{m}\in Q_{\bm{m}}\subset\Lambda_{\bm{k},*},{\rm dist}(\bm{m},\Lambda_{\bm{k},*}\setminus Q_{\bm{m}})\geq\frac{\tilde{N}_{1}}{2},\ {\rm diam}(\tilde{Q}_{\bm{m}})\leq 4\tilde{N}_{1}^{\frac{1}{10d}}.$
Also for any $\bm{n}\in Q_{\bm{m}}\setminus\tilde{Q}_{\bm{m}}$, there exist $L_{1}\sim(\log N_{1})^{C}$ and some $\mathcal{ER}_{\mathbb{Z}^{d}}(L_{1})\ni W\subset Q_{\bm{m}}\setminus\tilde{Q}_{\bm{m}}$ such that
$\bm{n}\in W,\ {\rm dist}(\bm{n},(Q_{\bm{m}}\setminus\tilde{Q}_{\bm{m}})\setminus W)\geq\frac{L_{1}}{2},$
and
$\|\mathcal{H}^{-1}_{W}(E_{\bm{k}};\bm{\theta})\|$ $\leq e^{\sqrt{L_{1}}},$
$|\mathcal{H}^{-1}_{W}(E_{\bm{k}};\bm{\theta})(\bm{n};\bm{n}^{\prime})|$ $\leq e^{-\frac{1}{2}|\log\varepsilon|\cdot|\bm{n}-\bm{n}^{\prime}|}\ {\rm for}\ |\bm{n}-\bm{n}^{\prime}|\geq\frac{L_{1}}{10}.$
Now assume that (set $Q=Q_{\bm{m}}$)
$\min_{\bm{m}\in\Lambda_{\bm{k},*},\xi=\pm 1,1\leq l\leq\#Q_{\bm{m}}}|\sigma+\bm{k}\cdot\bm{\omega}+\xi{\lambda_{l}(\bm{k})}|>e^{-\frac{1}{4}\tilde{N}_{1}^{\rho_{2}}}.$
Then applying the resolvent identities in Lemma 3.2 and Theorem 3.3 of [JLS20] yields
$\|A^{-1}_{\bm{k}}\|$ $\leq e^{\frac{1}{2}N_{1}^{\rho_{2}}},$
$\|A_{\bm{k}}^{-1}(\bm{n};\bm{n}^{\prime})\|$ $\leq e^{-\frac{1}{4}|\log\varepsilon|\cdot|\bm{n}-\bm{n}^{\prime}|}\ {\rm for}\ |\bm{n}-\bm{n}^{\prime}|\geq{N_{1}}^{\rho_{3}}.$
Note that
$\tilde{G}_{\Lambda_{*}}(\sigma)=\bigoplus_{\bm{k}\in\Pi_{b}\Lambda_{*}}A_{\bm{k}}^{-1},$
where
$\tilde{G}_{\Lambda_{*}}(\sigma)=\left(R_{\Lambda_{*}}({D}(\sigma)+\varepsilon(\Delta\oplus\Delta))R_{\Lambda_{*}}\right)^{-1}.$
By taking account of all $\bm{k}\in\Pi_{b}\Lambda_{*},$ we obtain that if
$\min_{\bm{k}\in\Pi_{b}\Lambda_{*},\bm{m}\in\Lambda_{\bm{k},*},\xi=\pm 1,1\leq l\leq\#Q_{\bm{m}}}|\sigma+\bm{k}\cdot\bm{\omega}+\xi{\lambda_{l}(\bm{k})}|>e^{-\frac{1}{4}N_{1}^{\kappa_{1}\rho_{2}}},$ (3.25)
then
$\|\tilde{G}_{\Lambda_{*}}(\sigma)\|$ $\leq e^{\frac{1}{2}N_{1}^{\rho_{2}}},$
$|\tilde{G}_{\Lambda_{*}}(\sigma)((\bm{k},\bm{n},\xi);(\bm{k}^{\prime},\bm{n}^{\prime},\xi^{\prime}))|$ $\leq\delta_{\bm{k},\bm{k}^{\prime}}e^{-\frac{1}{4}|\log\varepsilon|\cdot|\bm{n}-\bm{n}^{\prime}|}\ {\rm for}\ |\bm{n}-\bm{n}^{\prime}|\geq N_{1}^{\rho_{3}}.$
Finally, since $|\bm{n}|>2N_{1}$, we know that $S$ has the decay estimate
$|S((\bm{k},\bm{n},\xi);(\bm{k}^{\prime},\bm{n}^{\prime},\xi^{\prime}))|\leq Ce^{-2\gamma N_{1}}.$
Then applying the perturbation lemma of [Shi22] (cf. Lemma A.1) implies that $\Lambda$ is $\sigma$-good assuming (3.25) holds true. This proves (1) of Lemma 3.12.
Remark 3.11.
We would also like to remark that if we take into account of all these $\Lambda_{*}\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ with $|(\bm{k},\bm{n})|\leq 10N$ and $|\bm{n}|>2N_{1}$, then we get a sequence $\{{\lambda_{l}}\}_{1\leq l\leq N^{C}}$. As a result, there is a set $\Sigma_{N_{1}}\subset\mathbb{R}$ satisfying ${\rm meas}(\Sigma_{N_{1}})\leq N^{C}e^{-\frac{1}{4}N_{1}^{\kappa_{1}\rho_{2}}}\leq e^{-\frac{1}{5}N_{1}^{\kappa_{1}\rho_{2}}}$ so that, for $\sigma\not\in\Sigma_{N_{1}}$, all $\Lambda$ satisfying $\Lambda_{*}\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ with $|(\bm{k},\bm{n})|\leq 10N$ and $|\bm{n}|>2N_{1}$ are $\sigma$-good.
(2) The conclusion is just a corollary of (1) by replacing $N_{1}$ with $N_{2}$ in Remark 3.11.
This completes the proof of Lemma 3.12.∎
By combining Lemma 3.11 and Lemma 3.12, we are ready to prove LDE at the scale $N.$ We first recall an important lemma.
Lemma 3.13 (Matrix-valued Cartan’s lemma, Proposition 14.1 in [Bou05]).
Let $T(\sigma)$ be a self-adjoint $N\times N$ matrix-valued function of a parameter $\sigma\in[-\beta,\beta]$ satisfying the following conditions:
  • (i)
    $T(\sigma)$ is real analytic in $\sigma$ and has a holomorphic extension to
    $\mathbb{D}=\left\{z\in\mathbb{C}:\ |\Re z|\leq\beta,\ |\Im z|\leq\beta_{1}\right\}$
    satisfying $\sup\limits_{z\in\mathbb{D}}\|T(z)\|\leq K_{1},\ K_{1}\geq 1.$
  • (ii)
    For each $\sigma\in[-\beta,\beta]$, there is a subset $Y\subset[1,N]$ with $|Y|\leq M$ such that
    $\|(R_{[1,N]\setminus Y}T(\sigma)R_{[1,N]\setminus Y})^{-1}\|\leq K_{2},\ K_{2}\geq 1.$
  • (iii)
    Assume
    $\mathrm{meas}\left(\{\sigma\in[-{\beta},{\beta}]:\ \|T^{-1}(\sigma)\|\geq K_{3}\}\right)\leq 10^{-3}\beta_{1}(1+K_{1})^{-1}(1+K_{2})^{-1}.$
Let $0<\epsilon\leq(1+K_{1}+K_{2})^{-10M}.$ Then we have
$\mathrm{meas}\left(\left\{\sigma\in\left[-{\beta}/{2},{\beta}/{2}\right]:\ \|T^{-1}(\sigma)\|\geq\frac{1}{\epsilon}\right\}\right)\leq Ce^{\frac{-c\log\frac{1}{\epsilon}}{M\log(M+K_{1}+K_{2}+K_{3})}},$ (3.26)
where $C,c>0$ are some absolute constants.
Proof of Lemma 3.10.
Recall that
$N_{2}=N_{1}^{\widetilde{C}},\ N=N_{2}^{\widetilde{C}}.$
We assume that for $K=N_{1},N_{2}$ and $\Lambda\subset\mathbb{Z}_{{\rm pm},*}^{b+d}$ satisfying $\Lambda_{*}\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(K)$ with $|\bm{n}|\leq 2K$, $G_{\Lambda}(\sigma)$ satisfies the LDE.
Without loss of generality, we consider $\Lambda_{{\rm pm},N}=\{(\bm{j},\xi)\in\mathbb{Z}_{{\rm pm},*}^{b+d}:\ |\bm{j}|\leq N\}=\Lambda_{N}\times\{+,-\}\cap\mathbb{Z}_{{\rm pm},*}^{b+d}$, the other cases can be dealt with similarly. We will cover the $\Lambda_{{\rm pm},N}$ with two subregions, $\Lambda_{{\rm pm},N}=\mathcal{R}_{1}\cup\mathcal{R}_{2}$ satisfying
$\mathcal{R}_{1}=\{(\bm{k},\bm{n},\xi)\in\Lambda_{{\rm pm},N}:\ |\bm{n}|\leq N^{\frac{\rho_{2}}{10d}}\},$
$\mathcal{R}_{2}=\{(\bm{k},\bm{n},\xi)\in\Lambda_{{\rm pm},N}:\ |\bm{n}|\geq N^{\frac{\rho_{2}}{20d}}\}.$
We will apply matrix-valued Cartan’s lemma in $\mathcal{R}_{1}$, and Bourgain’s geometric lemma in $\mathcal{R}_{2}$ to remove $\sigma$. Finally, we apply the resolvent identity Theorem 3.3 of [JLS20] to prove $\Lambda_{{\rm pm},N}$ is $\sigma$-good if desired $\sigma$ are removed.
Before we enter the details, let us first explain why we need to divide $\Lambda_{{\rm pm},N}$ into the above two regions. This is quite different from that in [SW23]. Based on Lemma 3.12, we can impose the so called weak second Melnikov’s condition on $\bm{\omega}$ so that the $\sigma$-bad $\Lambda\subset\mathbb{Z}_{{\rm pm},*}^{b+d}$ with $\Lambda_{*}\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ satisfies for some $\bm{k}_{*},$
$|\bm{k}-\bm{k}_{*}|\leq 2N_{1}.$
This gives the desired control on $\sigma$-bad $\Lambda$ in the $\bm{k}$-direction. To establish the sublinear bound, it remains to handle the $\bm{n}$-direction. In [SW23], the $\bm{n}$-direction estimates can be obtained directly using the fact that $\theta\in\mathbb{R}$, Diophantine $\bm{\omega}$ and LDT for the corresponding quasi-periodic Schrödinger operators. Such arguments inevitably fail for $\bm{\theta}\in\mathbb{R}^{d}$: There is no interval decomposition of $d$-dimensional semi-algebraic sets, and the Diophantine property of $\bm{\omega}$ can not ensure the sublinear estimate in the $\bm{n}$-directions. This motivates us to restrict $|\bm{n}|\leq N^{\frac{\rho_{2}}{10d}}$ (i.e., the region $\mathcal{R}_{1}$). We can use the Cartan’s lemma and resolvent identities to control Green’s function on $\mathcal{R}_{1}$ similar to that in [SW23]. For $\mathcal{R}_{2}$, we take advantage of Bourgain’s geometric lemma and the short-range property of $H$ similar to the proof of Lemma 3.12. Finally, it suffices to cover $\Lambda_{{\rm pm},N}$ with overlapped regions $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ and apply the resolvent identities again.
Analysis in $\mathcal{R}_{1}$
We first prove the sublinear bound on the number of $\sigma$-bad elementary $N_{1}$-regions with centers in $[-N,N]^{b}\times([-N^{\frac{\rho_{2}}{10d}},N^{\frac{\rho_{2}}{10d}}]^{d}\setminus[-2N_{1},2N_{1}]^{d})$ for all $\sigma\in\mathbb{R}.$ For this purpose, an additional restriction on $\bm{\omega}$ (may be called the weak second Melnikov’s conditions) is required. Recalling (1) of Lemma 3.12 and taking into account of all $\Lambda_{*}\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ with $|(\bm{k},\bm{n})|\leq 10N$ and $|\bm{n}|>2N_{1}$ yields a sequence $\{{\lambda_{l}}\}_{\leq l\leq N^{C}}\subset\mathbb{R}$ with $C=C(b,d)>0$ and each $\lambda_{l}$ depending only on $\bm{\alpha},\bm{\theta}$ (but not on $\sigma,\bm{\omega}$), so that if $\Lambda_{*}=(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ satisfies $|(\bm{k},\bm{n})|\leq 10N$, $|\bm{n}|>2N_{1}$ and
$\min_{\bm{k}\in\Pi_{b}\Lambda_{*},\xi=\pm 1,1\leq l\leq N^{C}}|\sigma+\bm{k}\cdot\bm{\omega}+\xi{\lambda_{l}}|>e^{-\frac{1}{4}N_{1}^{\kappa_{1}\rho_{2}}},$
then $\Lambda$ is $\sigma$-good. So we can define the set
$\widetilde{\Omega}_{N}=\bigcap_{1\leq l,l^{\prime}\leq N^{C},\xi=\pm 1,\xi^{\prime}=\pm 1,0<|\bm{k}|\leq 2N}\left\{\bm{\omega}\in\Omega:\ |\bm{k}\cdot\bm{\omega}+\xi{\lambda_{l}}-\xi^{\prime}{\lambda_{l^{\prime}}}|>2e^{-\frac{1}{4}N_{1}^{\kappa_{1}\rho_{2}}}\right\}.$
Then we have ${\rm meas}(\Omega\setminus\widetilde{\Omega}_{N})\leq e^{-\frac{1}{5}N_{1}^{\kappa_{1}\rho_{2}}}.$ In the following, we always assume $\bm{\omega}\in\widetilde{\Omega}_{N}.$ Suppose now $\Lambda^{\prime}$ with $\Lambda^{\prime}_{*}\in(\bm{k}^{\prime},\bm{n}^{\prime})+\mathcal{ER}_{\bm{0}}(N_{1})$ satisfying $|\bm{k}^{\prime}|\leq N$ and $2N_{1}<|\bm{n}^{\prime}|\leq 10N$ is $\sigma$-bad. Then there are some $\bm{k}_{*}\in\Pi_{b}\Lambda_{*}^{\prime}\subset[-N,N]^{b}$, $\xi_{*}=\pm 1$ and some $l_{*}\in[1,N^{C}]$ so that
$|\sigma+\bm{k}_{*}\cdot\bm{\omega}+\xi_{*}{\lambda_{l_{*}}}|\leq e^{-\frac{1}{4}N_{1}^{\kappa_{1}\rho_{2}}}.$
Now let $\Lambda^{\prime\prime}$ (satisfying $\Lambda_{*}^{\prime\prime}\in(\bm{k}^{\prime\prime},\bm{n}^{\prime\prime})+\mathcal{ER}_{\bm{0}}(N_{1})$ with $|\bm{k}^{\prime\prime}|\leq N$ and $2N_{1}<|\bm{n}^{\prime\prime}|\leq 10N$) be another $\sigma$-bad region. From $\bm{\omega}\in\widetilde{\Omega}_{N},$ we must have
$\bm{k}_{*}\in\Pi_{b}\Lambda^{\prime\prime}_{*}.$ (3.27)
In other words, we have established all $\sigma$-bad elementary $N_{1}$-regions $\Lambda$ with centers $(\bm{k},\bm{n})$ satisfying $|\bm{k}|\leq N$ and $2N_{1}<|\bm{n}|\leq 10N$ must admit
$\Pi_{b}\Lambda_{*}\subset[\bm{k}_{*}-2N_{1},\bm{k}_{*}+2N_{1}]^{b}.$ (3.28)
Next, we estimate the number of $\sigma$-bad elementary $N_{1}$-regions satisfying (3.28) contained in $\mathcal{R}_{1}$. This needs to control the $\bm{n}$-directions and it is ensured by $|\bm{n}|\leq N^{\frac{\rho_{2}}{10d}}$. In conclusion, assuming $\bm{\omega}\in\widetilde{\Omega}_{N}$, we have proved that for all $\sigma\in\mathbb{R}$, the number of $\sigma$-bad elementary $N_{1}$-regions of centers $(\bm{k},\bm{n})$ satisfying $|\bm{n}|\leq N^{\frac{\rho_{2}}{10d}}$ and $|\bm{k}|>2N_{1}$ is at most
$CN_{1}^{2b}N^{\rho_{2}/10}\leq N^{\rho_{2}/8}.$
Combining with Lemma 3.11, we get actually the sublinear bound $N^{\rho_{2}/4}$ of all $\sigma$-bad elementary $N_{1}$-regions with centers belonging to $\mathcal{R}_{1}$.
We apply Lemma 3.13 with
$T(\sigma)={{H}}_{\mathcal{R}_{1}}(\sigma),\ \beta=\beta_{1}=e^{-10N_{1}^{\rho_{2}}}.$
It remains to verify the assumptions of Lemma 3.13. Obviously, one has
$K_{1}=O(N).$
We let $Y$ be the union of all $\sigma$-bad $\Lambda$ satisfying $\Lambda_{*}\in\mathcal{ER}(N_{1})$ with centers belonging to $\mathcal{R}_{1}$. Then by the above sublinear bound conclusion and applying the resolvent identity, we have
$M=\#Y\leq N^{\rho_{2}/4},\ K_{2}=e^{2{N_{1}}^{\rho_{2}}}.$ (3.29)
To apply the matrix-valued Cartan’s lemma, we need the scale
$N_{2}\sim N_{1}^{\widetilde{C}}.$
Recall that the LDE holds at the scale $N_{2}$. Applying the resolvent identity and using (2) of Lemma 3.12 yields
$\|T^{-1}(\sigma)\|\leq CN_{2}^{C}e^{{N_{2}}^{\rho_{2}}}\leq e^{2{N_{2}}^{\rho_{2}}}=K_{3},$
for $\sigma$ away from a set of measure at most (since $\rho_{1}<\kappa_{1}\rho_{2}$)
$C\tilde{N}^{C}e^{-{N_{2}^{\rho_{1}}}}+e^{-\frac{1}{5}{N_{2}^{\kappa_{1}\rho_{2}}}}\leq e^{-{N_{2}^{\rho_{1}}}/{2}}.$
It follows from assuming
$\widetilde{C}\rho_{1}>\rho_{2}$
that
$10^{-3}\beta_{1}(1+K_{1})^{-1}(1+K_{2})^{-1}>e^{-{N_{2}^{\rho_{1}}}/{2}}.$
This verifies (iii) of Lemma 3.13. If $\epsilon=e^{-N^{\rho_{2}}}$, then one has $\epsilon<(1+K_{1}+K_{2})^{-10M}.$ Cover $[-10N,10N]$ with disjoint intervals of length $e^{-10N_{1}^{\rho_{2}}}$. Define
$\Sigma_{N,1}=\{\sigma\in\mathbb{R}:\ \|G_{\mathcal{R}_{1}}(\sigma)\|\geq e^{N^{\rho_{2}}}\}.$
Then by (3.26) of Lemma 3.13, one obtains since $\rho_{1}<\rho_{2}/5$ that
$\mathrm{meas}(\Sigma_{N,1})\leq CN^{C}e^{10N_{1}^{\rho_{2}}}e^{-\frac{c{{N}^{\rho_{2}}}}{CN_{1}N_{2}N^{\rho_{2}/4}\log N}}\leq e^{-N^{\rho_{2}/5}}\leq e^{-N^{\rho_{1}}}/2.$
Next, for $\sigma\notin\Sigma_{N,1}$, it suffices to prove the exponential off-diagonal decay of $G_{\mathcal{R}_{1}}(\sigma)$. This will be completed by using the coupling lemma of [HSSY20] (cf. Lemma 4.2).
Collecting all the restrictions on $\widetilde{C},\rho_{1},\rho_{2}$ yields
[…]
Analysis in $\mathcal{R}_{2}$
In this region, we will use Bourgain’s geometric lemma (cf. Lemma 3.5). Choose $L\sim N^{\frac{\rho_{2}}{30d}},L\in\mathbb{N}.$ Since $\bm{\alpha}\in\mathcal{A}\subset\mathcal{A}_{L}$, similar to the proof of Lemma 3.12, we can show for $\sigma$ outside a set (denoted by $\Sigma_{N,2}$) of measure at most
$\mathrm{meas}(\Sigma_{N,2})\leq e^{-\frac{1}{5}L^{\kappa_{1}\rho_{2}}}\leq e^{-cN^{\frac{\kappa_{1}\rho_{2}^{2}}{30d}}}<e^{-N^{\rho_{1}}}/2,$
all $\Lambda_{*}\in\mathcal{ER}(L)$ with $\Lambda\subset\mathcal{R}_{2}$ are $\sigma$-good, where we have used the fact
$\rho_{1}<\frac{\kappa_{1}\rho_{2}^{2}}{30d}.$
More precisely, fix $\Lambda_{*}\in\mathcal{ER}(L)$ and $\bm{k}\in\Pi_{b}\Lambda_{*}$. For any $Q\subset\Lambda_{\bm{k},*}:=\Pi_{d}\Lambda_{*}(\bm{k})\in\mathcal{ER}_{\mathbb{Z}^{d}}(L),$ write
$A_{\bm{k}}$ $=R_{Q}({D}(\sigma)+\varepsilon(\Delta\oplus\Delta))R_{Q}$
$=(-\sigma-\bm{k}\cdot\bm{\omega}+\mathcal{L}_{Q}(\bm{\theta}))\oplus(\sigma+\bm{k}\cdot\bm{\omega}+\mathcal{L}_{Q}(\bm{\theta})),$
where $\mathcal{L}_{Q}(\bm{\theta})$ is defined in Lemma 3.5. Then the set of all eigenvalues of $A_{\bm{k}}$ is given by
{±$\{\pm(\sigma+\bm{k}\cdot\bm{\omega})+\lambda_{l}(\bm{k})\}_{l=1}^{\#Q},$
where $\{\lambda_{l}(\bm{k})\}_{l=1}^{\#Q}$ denotes the spectrum of $\mathcal{L}_{Q}(\bm{\theta}).$ So we have
$\|A_{\bm{k}}^{-1}\|$ $\leq\max_{\xi=\pm 1,1\leq l\leq\#Q}|\xi(\sigma+\bm{k}\cdot\bm{\omega})+\lambda_{l}(\bm{k})|^{-1}.$
Similar to the proof of Lemma 3.12 by using Lemma 3.5 and resolvent identities, namely Lemma 3.2 and Theorem 3.3 in [JLS20] and Lemma A.1 in [Shi22], we can show for each $\bm{m}\in\Lambda_{\bm{k},*}$, there exist $\frac{1}{4}L^{\kappa_{1}}\leq\tilde{L}\leq L^{\kappa_{2}}$, $Q_{\bm{m}}\in\mathcal{ER}_{\mathbb{Z}^{d}}(\tilde{L})$ so that, if
$\min_{\bm{m}\in\Lambda_{\bm{k},*},\xi=\pm 1,1\leq l\leq\#Q_{\bm{m}}}|\sigma+\bm{k}\cdot\bm{\omega}+\xi{\lambda_{l}(\bm{k})}|>e^{-\frac{1}{4}\tilde{L}^{\rho_{2}}},$
then
$\|A^{-1}_{\bm{k}}\|$ $\leq e^{\frac{1}{2}L^{\rho_{2}}},$
$\|A_{\bm{k}}^{-1}(\bm{n};\bm{n}^{\prime})\|$ $\leq e^{-\frac{1}{4}|\log\varepsilon|\cdot|\bm{n}-\bm{n}^{\prime}|}\ {\rm for}\ |\bm{n}-\bm{n}^{\prime}|\geq{L}^{\rho_{3}}.$
Note that
$\tilde{G}_{\Lambda_{*}}(\sigma)=\bigoplus_{\bm{k}\in\Pi_{b}\Lambda_{*}}A_{\bm{k}}^{-1},$
where
$\tilde{G}_{\Lambda_{*}}(\sigma)=\left(R_{\Lambda_{*}}({D}(\sigma)+\varepsilon(\Delta\oplus\Delta))R_{\Lambda_{*}}\right)^{-1}.$
By taking account of all $\bm{k}\in\Pi_{b}\Lambda_{*},$ we obtain that if
$\min_{\bm{k}\in\Pi_{b}\Lambda_{*},\bm{m}\in\Lambda_{\bm{k},*},\xi=\pm 1,1\leq l\leq\#Q_{\bm{m}}}|\sigma+\bm{k}\cdot\bm{\omega}+\xi{\lambda_{l}(\bm{k})}|>e^{-\frac{1}{4}L^{\kappa_{1}\rho_{2}}},$ (3.30)
then
$\|\tilde{G}_{\Lambda_{*}}(\sigma)\|$ $\leq e^{\frac{1}{2}L^{\rho_{2}}},$
$|\tilde{G}_{\Lambda_{*}}(\sigma)((\bm{k},\bm{n},\xi);(\bm{k}^{\prime},\bm{n}^{\prime},\xi^{\prime}))|$ $\leq\delta_{\bm{k},\bm{k}^{\prime}}e^{-\frac{1}{4}|\log\varepsilon|\cdot|\bm{n}-\bm{n}^{\prime}|}\ {\rm for}\ |\bm{n}-\bm{n}^{\prime}|\geq L^{\rho_{3}}.$
Finally, since $|\bm{n}|>N^{\frac{\rho_{2}}{20d}}\gg 10L$, we know that $S$ has the decay estimate
$|S((\bm{k},\bm{n},\xi);(\bm{k}^{\prime},\bm{n}^{\prime},\xi^{\prime}))|\leq Ce^{-10\gamma L}.$
Then applying the perturbation lemma of [Shi22] (cf. Lemma A.1) implies that $\Lambda$ is $\sigma$-good assuming (3.30) holds true. Taking account of all $\Lambda\subset\mathcal{R}_{2}$ leads to the desired measure estimates on $\Sigma_{N,2}$.
Application of the resolvent identity
Finally, we cover points $(\bm{k},\bm{n},\xi)\in\Lambda_{{\rm pm},N}$ satisfying $|\bm{n}|\leq N^{\frac{\rho_{2}}{20d}}$ with $\mathcal{R}_{1}$, points satisfying $|\bm{n}|>N^{\frac{\rho_{2}}{20d}}$ with $\sigma$-good $L$-regions. Applying Theorem 3.3 of [JLS20] yields $\Lambda_{{\rm pm},N}$ is $\sigma$-good for $\sigma$ outside a set $\Sigma_{N}=\Sigma_{N,1}\cup\Sigma_{N,2}$ of measure at most $e^{-N^{\rho_{1}}}.$
This completes the proof of Lemma 3.10. ∎
Proof of Theorem 3.6.
The proof is just a combination of Lemmas 3.73.8  and 3.10. ∎
4. Nonlinear Analysis
We prove the existence of nonlinear Anderson localized states using the Lyapunov-Schmidt decomposition as written in sect. 2.1, which divides the nonlinear matrix equation in (2.2),
$F(\vec{u})=0,\ \vec{u}=\binom{\hat{u}}{\hat{v}},$
into the $P$-equations in (2.9),
${F}\big{|}_{\mathbb{Z}^{b+d}_{{\rm pm},*}}\left(\vec{u}\right)=0,\ \mathbb{Z}^{b+d}_{{\rm pm},*}=\mathbb{Z}^{b+d}\times\{+,-\}\setminus\mathcal{S},$
and the $Q$-equations in (2.10),
${F}\big{|}_{\mathcal{S}}(\vec{u})=0.$
We iteratively solve the $Q$ and then the $P$-equations. On $\mathcal{S}$, $\vec{u}$ is held fixed at ${\vec{u}}^{(0)}$; the $Q$-equations are used to relate the frequencies $\bm{\omega}$ to the amplitudes $\bm{a}$, permitting amplitude-frequency modulation to solve the $P$-equations. The resolution of the $P$-equations uses a Newton scheme and the analysis in section 3. The large deviation estimates in $\sigma$ in Theorem 3.6 will be converted into estimates in the amplitudes $\bm{a}$ by using a semi-algebraic projection lemma of Bourgain (cf. Lemma 4.4 below), and amplitude-frequency modulation, $\bm{\omega}=\bm{\omega}(\bm{a})$. This is an established scheme, which has recently been extensively elaborated on with detailed proofs in sects. IV-VI, [KLW24]. It will serve as the basic reference point for the present section.
4.1. The Newton scheme
Recall first the formal Newton scheme. Let $\vec{w}$ be an approximate solution to the $P$-equations (2.9). For the next approximation ${\vec{w}}^{\prime}$, write
${\vec{w}}^{\prime}=\vec{w}+\vec{y}.$
Then ${\vec{y}}$ is set to be
$\vec{y}=-[H(\vec{w})]^{-1}F(\vec{w}),$
where the linearized operator
$H={\rm diag}_{(\bm{k},\bm{n})\in\mathbb{Z}^{b+d}}$ $\left(\begin{array}[]{cc}{-\bm{k}\cdot\bm{\omega}+\mu_{\bm{n}}+\varepsilon\Delta}&{0}\\ {0}&{\bm{k}\cdot\bm{\omega}+\mu_{\bm{n}}+\varepsilon\Delta}\\ \end{array}\right)$
$+\delta\left(\begin{array}[]{cc}{(p+1)(\hat{w}*\hat{\mathfrak{w}})^{*p}*}&{p(\hat{w}*\hat{\mathfrak{w}})^{*(p-1)}*\hat{w}*\hat{w}*}\\ {p(\hat{w}*\hat{\mathfrak{w}})^{*(p-1)}*\hat{\mathfrak{w}}*\hat{\mathfrak{w}}*}&{(p+1)(\hat{w}*\hat{\mathfrak{w}})^{*p}*}\\ \end{array}\right),$
as in (2.5) and (2.8), with $\hat{w}$ in lieu of $\hat{u}$ and $\hat{\mathfrak{w}}$ in lieu of $\hat{v}$. This is, however, only indicative and assumes that $H$ is invertible. Due to the small divisor difficulties in the inversion process, we need to regularize the linearized operator $H$ and do a multi-scale Newton iteration instead.
Toward that purpose, let $M$ be a large integer and consider the geometric sequence of scales $M^{\ell}$, $\ell=1,2,\cdots$. Denote by $\vec{u}^{(r)}$, the $r$-th approximate solution to the $P$-equations (2.9). For the $(r+1)$th approximation, write
$\vec{u}^{(r+1)}=\vec{u}^{(r)}+\Delta_{r+1}\vec{u}.$
We define (if it exsits)
$\Delta_{r+1}\vec{u}=-[R_{\Lambda_{M^{\tilde{r}+1}}}{{H(\vec{u}^{(r)})}}R_{\Lambda_{M^{\tilde{r}+1}}}{{}}]^{-1}F(\vec{u}^{(r)}),$
where $R_{\Lambda_{M^{\tilde{r}+1}}}$ is the restriction operator to the cube
$[-M^{\tilde{r}+1},M^{\tilde{r}+1}]^{b+d}\times\{+,-\}\backslash\mathcal{S},$
cf. (3.2), and $\tilde{r}=r+r_{0}$, for some $r_{0}>1$. The reason to consider a shifted sequence $\tilde{r}$ is technical, see sect. V, E. Step 2-F. Step 4, [KLW24]. This leads to estimate the inverse (the Green’s function),
${{G}}_{r+1}=\left(R_{\Lambda_{M^{\tilde{r}+1}}}{{H(\vec{u}^{(r)})}}R_{\Lambda_{M^{\tilde{r}+1}}}{{}}\right)^{-1},\,r=0,1,2,\cdots.$
The following estimates are crucial for the convergence of the approximate solutions with exponential decay:
$\|{{G}}_{r+1}(\vec{u}^{(r)})\|$ $\leq M^{({\tilde{r}}+1)^{C}},$
$|{{G}}_{r+1}(\vec{u}^{(r)})((\bm{j},\xi);(\bm{j}^{\prime},\xi^{\prime}))|$ $\leq e^{-c|\bm{j}-\bm{j}^{\prime}|}\ {\rm for}\ |\bm{j}-\bm{j}^{\prime}|\geq({\tilde{r}}+1)^{C},$
for $\bm{j}=(\bm{k},\bm{n})$ and some $c,C>0$, see sect. V, [KLW24]. The main purpose of this section is thus to establish the above estimates assuming $0<\varepsilon+\delta\ll 1$ and additional restrictions on $\bm{\omega}$ (and then $\bm{a}$).
4.2. Extraction of parameters
Write $u^{(0)}$ (cf. (1.9)) as
$u^{(0)}(t,\bm{n})=\sum_{(\bm{k},\bm{n})\in\mathcal{S}_{+}}\hat{u}^{(0)}(\bm{k},\bm{n})e^{i\bm{k}\cdot\bm{\omega}^{(0)}},$ (4.1)
where $\mathcal{S}_{+}=\{(\bm{e}_{l},\bm{n}_{l})\}_{l=1}^{b}$ is given by Theorem 1.1 and
[…]
The solutions $\vec{u}=(\hat{u},\hat{v})$ on $\mathcal{S}$ are held fixed: $\vec{u}=(\hat{u},\hat{v})=(\hat{u}^{(0)},\hat{v}^{(0)})=\vec{u}^{(0)}$ on $\mathcal{S}$, the $Q$-equations are used instead to solve for the frequencies. Due to symmetry, the $Q$-equations for $\hat{v}$ are the same as that for $\hat{u}$. So we only need to consider that for the $\hat{u}$. Solving the equations at the initial step
${F}\big{|}_{\mathcal{S_{+}}}\left(\vec{u}^{(0)}\right)=0$
leads to, for $1\leq l\leq b$, the initial modulated frequencies
$\omega_{l}^{(1)}$ $=\omega_{l}^{(0)}+\delta a_{l}^{2p}$ (4.2)
satisfying
$\det\left(\frac{\partial\bm{\omega}^{(1)}}{\partial\bm{a}}\right)=\delta^{b}\prod_{l=1}^{b}(2pa_{l}^{2p-1}).$
Along the way, we will show that $\bm{\omega}^{(r)}=\bm{\omega}^{(0)}+O(\delta)$, for all $r=0,1,2,\cdots$. So $\bm{\omega}^{(r)}\in\Omega$, the frequency set in Theorem 3.6, hence Theorem 3.6 is at our disposal to solve the $P$-equations. To construct approximate solutions to the $P$-equations, it is convenient to view $\bm{\omega}\in\Omega$ as an independent parameter, and work in the $(\bm{\omega},\bm{a})$ variables, $(\bm{\omega},\bm{a})\in\Omega\times[1,2]^{b}=\{\bm{\omega}^{(0)}+[-C\delta,C\delta]^{b}\}\times[1,2]^{b}$. The approximate solutions to the nonlinear matrix equation (2.1) will then be obtained by taking into account the solutions to the $Q$-equations as well, which restrict $(\bm{\omega},\bm{a})$ to the $b$-dimensional hyper-surface $\bm{\omega}=\bm{\omega}^{(r)}(\bm{a})$, at the $r$-th iteration.
4.3. The initial $r_{0}$ iteration steps
We may choose $r_{0}$ such that
$M^{2r_{0}}\leq(\varepsilon+\delta)^{-\frac{1}{10^{3}b^{2}d^{2}\tau_{b+2}^{2}}}<M^{2r_{0}+1},$ (4.3)
cf., sect. V, E. Step 2.-F. Step 4, [KLW24]. Let $\vec{u}^{(0)}(\bm{\omega},\bm{a})=\vec{u}^{(0)}(\bm{\omega}^{(0)},\bm{a})$. We start by constructing $\vec{u}^{(1)}(\bm{\omega},\bm{a})$. Obviously,
${F}(\vec{u}^{(0)})=O(\varepsilon+\delta)$
and ${\rm supp}\ {F}(\vec{u}^{(0)})\subset\Lambda_{C}$ for some $C=C(p,\mathcal{S})>0.$ It suffices to estimate
$H_{M^{r_{0}+1}}^{-1}=(R_{\Lambda_{M^{r_{0}+1}}}({D}(0)+\varepsilon(\Delta\oplus\Delta)+\delta{T}_{\vec{u}^{(0)}})R_{\Lambda_{M^{r_{0}+1}}})^{-1}.$
It follows from (3.6) and the Neumann series that
$\|H_{M^{r_{0}+1}}^{-1}\|$ $\leq 2(\varepsilon+\delta)^{-\frac{1}{4b}},$
$|H_{M^{r_{0}+1}}^{-1}((\bm{j},\xi);(\bm{j}^{\prime},\xi^{\prime}))|$ $\leq 2(\varepsilon+\delta)^{-\frac{1}{4b}}e^{-c|\bm{j}-\bm{j}^{\prime}|}\ {\rm for}\ \bm{j}\neq\bm{j}^{\prime}\in\Lambda_{M},$
since $M^{r_{0}+1}\ll(\varepsilon+\delta)^{-\frac{1}{10^{3}b^{2}d^{2}\tau_{b+2}^{2}}}$ is fixed. Since
$\Delta_{1}\vec{u}=-H_{M^{r_{0}+1}}^{-1}{F}(\vec{u}^{(0)}),$
and consequently
$\vec{u}^{(1)}=\vec{u}^{(0)}+\Delta_{1}\vec{u}=\vec{u}^{(0)}-H_{M^{r_{0}+1}}^{-1}{F}(\vec{u}^{(0)}).$
Substituting $\vec{u}^{(1)}$ into the $Q$-equations using
$\hat{u}^{(1)}\big{|}_{\mathcal{S}+}=\hat{v}^{(1)}\big{|}_{\mathcal{S}-}=\bm{a},$
we obtain
$\omega_{l}$ $=\omega_{l}^{(0)}+\varepsilon[\frac{1}{a_{l}}(\Delta\hat{u}^{(1)})(\bm{e}_{l},\bm{n}_{l})](\bm{\omega},\bm{a})$
$\ \ +\delta[\frac{1}{a_{l}}(\hat{u}^{(1)}*\hat{v}^{(1)})^{*p}*\hat{u}^{(1)}(\bm{e}_{l},\bm{n}_{l})](\bm{\omega},\bm{a}),\,\lambda=1,2,\cdots,b.$
Since the second and the third terms are smooth in $\bm{\omega}$ and $\bm{a}$, implicit function theorem yields $\bm{\omega}^{(2)}=\bm{\omega}^{(2)}(\bm{a})=\{\omega_{l}^{(2)}(\bm{a})\}_{l=1}^{b}$, written in the form,
$\omega_{l}^{(2)}(\bm{a})=\omega_{l}^{(0)}+(\varepsilon+\delta)\varphi_{l}^{(2)}(\bm{a}),\,\lambda=1,2,\cdots,b,$
with a smooth function $\varphi_{l}^{(2)}$. (Compare with (4.2).) The above constructions can be performed for $r_{0}$ steps, with $r_{0}$ satisfying (4.3). So we have obtained $\vec{u}^{(l)}=\vec{u}^{(l)}(\bm{\omega},\bm{a})$ ($1\leq l\leq r_{0}$) for all $(\bm{\omega},\bm{a})\in\Omega\times[1,2]^{b}$.
4.4. The Inductive Theorem
We introduce the Inductive Theorem.
Theorem 4.1.
For $r\geq r_{0},$ we have
  • (Hi)
    ${\rm supp}\ \vec{u}^{(r)}\subset\Lambda_{M^{\tilde{r}}}$, $\tilde{r}=r+r_{0}$
  • (Hii)
    $\|\Delta_{r}\vec{u}\|<\delta_{r},\ \|\partial\Delta_{r}\vec{u}\|<{\bar{\delta}}_{r}$, where $\partial$ refers to derivation in $\bm{\omega}$ or $\bm{a}$,
    $\vec{u}^{(r)}=\vec{u}^{(r-1)}+\Delta_{r}\vec{u},$
    and $\|\cdot\|=\sup_{\bm{\omega},\bm{a}}\|\cdot\|_{\ell^{2}(\mathbb{Z}_{\rm pm}^{b+d})}.$
    Remark 4.1.
    The size of $\delta_{r},\bar{\delta}_{r}$ will satisfy $\log\log\frac{1}{\delta_{r}+\bar{\delta}_{r}}\sim r.$ Precisely, we have
    $\delta_{r}<\sqrt{\varepsilon+\delta}e^{-(\frac{4}{3})^{r}},\ \bar{\delta}_{r}<\sqrt{\varepsilon+\delta}e^{-\frac{1}{2}(\frac{4}{3})^{r}}.$
  • (Hiii)
    $|\vec{u}^{(r)}(\bm{k},\bm{n})|_{2}\leq e^{-c(|\bm{k}|+|\bm{n}|)}$ for some $c>0.$
    Remark 4.2.
    The constant $c>0$ will decrease slightly along the iterations but remain bounded away from 00. This will become clear in the proof, cf. sect. V, G. Step 4, [KLW24]. We also remark that $\vec{u}^{(r)}$ can be defined as a $C^{1}$ function on the entire parameter space $(\bm{\omega},\bm{a})\in\Omega\times[1,2]^{b}$ by using an extension argument, cf., [Bou98, BW08] and sect. V, L. Step 9, [KLW24]. So one can apply the implicit function theorem to solve the $Q$-equations leading to
    $\omega_{l}^{(r)}(\bm{a})=\omega_{l}^{(0)}+(\varepsilon+\delta)\varphi_{l}^{(r)}(\bm{a})\ (1\leq l\leq b),$ (4.4)
    where $\|\partial\bm{\varphi}^{(r)}\|=\sup_{1\leq l\leq b}\|\partial\varphi_{l}^{(r)}\|\lesssim 1$ and $\bm{\varphi}^{(r)}=(\varphi^{(r)}_{l})_{l=1}^{b}.$ By (Hii), we have
    $|\bm{\varphi}^{(r)}-\bm{\varphi}^{(r-1)}|\lesssim(\varepsilon+\delta)\|\vec{u}^{(r)}-\vec{u}^{(r-1)}\|\lesssim{(\varepsilon+\delta)}\delta_{r}.$
    Denote by $\Gamma_{r}$ the graph of $\bm{\omega}^{(r)}=\bm{\omega}^{(r)}(\bm{a})$. We have $\|\Gamma_{r}-\Gamma_{r-1}\|\lesssim{(\varepsilon+\delta)}\delta_{r}.$ Recall that $\bm{\omega}$ is given by (4.2) and $\bm{\varphi}^{(0)}=\bm{0}$. Thus we have a diffeomorphism from $\bm{\omega}^{(r)}\in\Omega$ to $\bm{a}\in[1,2]^{b}$.
  • (Hiv)
    There is a collection $\mathcal{I}_{r}$ of intervals $I\subset\Omega\times[1,2]^{b}$ of size $M^{-{\tilde{r}}^{C}},$ so that
    • (a)
      On each $I\in\mathcal{I}_{r},$ both $\hat{u}^{(r)}(\bm{\omega},\bm{a})$ and $\hat{v}^{(r)}(\bm{\omega},\bm{a})$ are given by rational functions in $(\bm{\omega},\bm{a})$ of degree at most $M^{{\tilde{r}}^{3}}$.
    • (b)
      For $(\bm{\omega},\bm{a})\in\bigcup_{I\in\mathcal{I}_{r}}I$,
      $\|F(\vec{u}^{(r)})\|\leq\kappa_{r},\ \|\partial F(\vec{u}^{(r)})\|\leq\bar{\kappa}_{r},$
      where $\partial$ refers to derivation in $\bm{\omega}$ or $\bm{a}$, and $\log\log\frac{1}{\kappa_{r}+\bar{\kappa}_{r}}\sim r$. More precisely, we have
      $\kappa_{r}<\sqrt{\varepsilon+\delta}e^{-(\frac{4}{3})^{r+2}},\ \bar{\kappa}_{r}<\sqrt{\varepsilon+\delta}e^{-\frac{1}{2}(\frac{4}{3})^{r+2}}.$
    • (c)
      For $(\bm{\omega},\bm{a})\in\bigcup_{I\in\mathcal{I}_{r}}I$ and $H=H(\vec{u}^{(r-1)})$, one has
      $\|H_{M^{\tilde{r}}}^{-1}\|$ $\leq M^{{\tilde{r}}^{C}},$ (4.5)
      $|H_{M^{\tilde{r}}}^{-1}((\bm{j},\xi);(\bm{j}^{\prime},\xi^{\prime}))|$ $\leq e^{-c|\bm{j}-\bm{j}^{\prime}|}\ {\rm for}\ |\bm{j}-\bm{j}^{\prime}|>{\tilde{r}}^{C},$ (4.6)
      where $\bm{j}=(\bm{k},\bm{n}),\bm{j}^{\prime}=(\bm{k}^{\prime},\bm{n}^{\prime})$ and $H_{M^{\tilde{r}}}$ refers to the restriction of $H$ on $\Lambda_{M^{\tilde{r}}}$.
    • (d)
      Each $I\in\mathcal{I}_{r}$ is contained in some $I^{\prime}\in\mathcal{I}_{r-1}$ and
      ${\rm meas}\left(\Pi_{\bm{a}}(\Gamma_{r-1}\cap(\bigcup_{I^{\prime}\in\mathcal{I}_{r-1}}I^{\prime}\setminus\bigcup_{I\in\mathcal{I}_{r}}I))\right)\leq M^{-\frac{\tilde{r}}{C(b)}},$
      where $C(b)>0$ depends only on $b$, and $\Pi_{\bm{a}}$ denotes the projection of the set on the $\bm{a}$-variables.
4.5. Proof of the Inductive Theorem
We assume that the Inductive Theorem holds for $r_{0}\leq l\leq r.$ We will prove it for $l=r+1.$ It suffices to establish (Hiv, c) and (Hiv, d), and the other inductive assumptions can be verified in the usual way, cf., [BW08] and sect. V, [KLW24]. For this purpose, we set $N=M^{\tilde{r}+1}$ and a smaller scale,
$N_{1}=(\log N)^{C^{\prime}},$
where $C^{\prime}>1$ will be specified below.
To establish (Hiv, c), we first make approximations on $T_{\vec{u}^{(l)}}$ for different $\vec{u}^{(l)}$. It follows from
$\|T_{\vec{u}^{(r)}}-T_{\vec{u}^{(r-1)}}\|\lesssim\|\Delta_{r}\vec{u}\|<\delta_{r},\ \log\log\frac{1}{\delta_{r}}\sim r,$ (4.7)
and assumptions on $H_{\Lambda_{M^{\tilde{r}}}}({\vec{u}^{(r-1)}})$ that $H_{\Lambda_{M^{\tilde{r}}}}({\vec{u}^{(r)}})$ also satisfies essentially the estimates (4.5) and (4.6). For the set
$W=\left(([-N,N]^{b+d}\setminus[-N/2,N/2]^{b+d})\times\{+,-\}\right)\cap\mathbb{Z}_{{\rm pm},*}^{b+d},$
we use $N_{1}$-regions to do estimates, in order to lower the degree of the associated semi-algebraic set, which will be essential for the upcoming semi-algebraic projection arguments. We set
$r_{1}=2\left[\frac{\log N_{1}}{\log\frac{4}{3}}\right]+1.$
Then
$\|T_{\vec{u}^{(r)}}-T_{\vec{u}^{(r_{1})}}\|\lesssim\delta_{r_{1}}\lesssim\sqrt{\varepsilon+\delta}e^{-(\frac{4}{3})^{r_{1}}}<e^{-N_{1}^{2}}.$ (4.8)
So we can estimate $H_{W}^{-1}(\vec{u}^{(r)})$ using $H_{Q}^{-1}(\vec{u}^{(r_{1})})$ for $Q\subset W$. More precisely, we want to show that the following estimates hold for every $Q\subset W$ with $\Pi_{b,d}Q\in\mathcal{ER}(N_{1})$:
$\|H_{Q}^{-1}(\vec{u}^{(r_{1})})\|$ $\leq e^{N_{1}^{\rho_{2}}},$ (4.9)
$|H_{Q}^{-1}(\vec{u}^{(r_{1})})((\bm{j},\xi);(\bm{j}^{\prime},\xi^{\prime}))|$ $\leq e^{-c|\bm{j}-\bm{j}^{\prime}|}\ {\rm for}\ |\bm{j}-\bm{j}^{\prime}|>N_{1}^{\rho_{3}}.$ (4.10)
This requires additional restrictions on $(\bm{\omega},\bm{a}).$ For this, we divide into the following cases:
Case 1. Assume $\Pi_{b,d}Q\in(\bm{k},\bm{n})+\mathcal{ER}_{0}(N_{1})$ with $|\bm{n}|>2N_{1}$ and $Q\subset W.$ Denote by $\mathcal{C}_{1}$ the set of all these $Q$. To estimate $H_{Q}^{-1}(\vec{u}^{(r_{1})})$, we can impose as in the proof of Lemma 3.12 the following condition
$\min_{\xi=\pm 1,1\leq l\leq N_{1}^{C}}|\bm{k}\cdot\bm{\omega}+\xi{\lambda_{l}}|>e^{-\frac{1}{5}N_{1}^{\kappa_{1}\rho_{2}}},$
where $\lambda_{l}>1/2$ is given by Lemma 3.12. For $\bm{k}=\bm{0}$, the above condition holds trivially, since $\lambda_{l}>1/2.$ For $\bm{k}\neq\bm{0}$, we can remove $\bm{\omega}$ directly. Assume that $C^{\prime}\kappa_{1}\rho_{2}>3.$ So by taking account of all $Q\in\mathcal{C}_{1}$ and by Fubini’s theorem, we can find $I_{1}\subset\Omega\times[1,2]^{b}$ with
${\rm meas}(I_{1})\leq N^{C}e^{-\frac{1}{5}(\log N)^{C^{\prime}\kappa_{1}\rho_{2}}}\leq e^{-(\log N)^{2}}\ll(\varepsilon+\delta)^{b}M^{-{\tilde{r}}/2},$
so that, for all $(\bm{\omega},\bm{a})\notin I_{1}$ and all $Q\in\mathcal{C}_{1}$, $H_{Q}^{-1}(\vec{u}^{(r_{1})})$ satisfies (4.9) and (4.10).
Case 2. Assume $\Pi_{b,d}Q\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ with $|\bm{n}|\leq 2N_{1}$ and $Q\subset W.$ Denote by $\mathcal{C}_{2}$ the set of all these $Q$. In this case, it must be that $|\bm{k}|\geq N/2.$ We will use the projection lemma of Bourgain and LDT to remove $\bm{\omega}.$ For any $Q\in\mathcal{C}_{2}$, we can write $\Pi_{b,d}Q=(\bm{k},\bm{0})+\Pi_{b,d}Q_{0}$ with $\Pi_{b,d}Q_{0}\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ for $|\bm{n}|\leq 2N_{1},$ and $N/2\leq|\bm{k}|\leq N.$ This motivates us to consider
$H(\sigma)=H(\vec{u}^{(r_{1})};\sigma)=D(\sigma)+\varepsilon(\Delta\oplus\Delta)+\delta T_{\vec{u}^{(r_{1})}},$
which has been investigated in section 3. Recall that $G_{Q}(\sigma)$ denotes the Green’s function of $H(\sigma)$ restricted to $Q$. Then the Toeplitz property in the $\bm{k}$-direction of $H(\sigma)$ implies
$H^{-1}_{Q}(\vec{u}^{(r_{1})})=G_{Q}(0)=G_{Q_{0}}(\bm{k}\cdot\bm{\omega}).$
Note that $\Omega_{N_{1}}$ given by Theorem 3.6 is a semi-algebraic set, independent of $\bm{a}$, and of degree at most $N_{1}^{C}$ (for $N_{1}>(\varepsilon+\delta)^{-\frac{1}{10^{3}b^{2}d^{2}\tau_{b+2}^{2}}}$) and measure at least $\delta^{b}-O(N_{1}^{-C})$. We may therefore assume $\Pi_{\bm{\omega}}I\subset\Omega_{N_{1}}$ for each $I\in\mathcal{I}_{r_{1}}$, and apply the LDE at the scale $N_{1}$ on each $I\in\mathcal{I}_{r_{1}}$: for all $\Pi_{b,d}Q\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ with $|\bm{n}|\leq 2N_{1},$ we have for $\sigma$ outside a set of measure at most $e^{-N_{1}^{\rho_{1}}}$, the following bounds,
$\|G_{Q}(\sigma)\|$ $\leq e^{N_{1}^{\rho_{2}}},$ (4.11)
$|G_{Q}(\sigma)((\bm{j},\xi);(\bm{j}^{\prime},\xi^{\prime}))|$ $\leq e^{-c|\bm{j}-\bm{j}^{\prime}|}\ {\rm for}\ |\bm{j}-\bm{j}^{\prime}|>N_{1}^{\rho_{3}}.$ (4.12)
For the usage of projection lemma of Bourgain in the present setting (i.e., ${\rm meas}(\Omega)\sim\delta^{b}$), we need to make more precise descriptions of the admitted $\sigma$ first. We say $\sigma$ is $Q$-bad if either (4.11) or (4.12) fails.
Lemma 4.2.
Let $C^{\prime}\rho_{1}>3$. Fix $\Pi_{b,d}Q\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ with $|\bm{n}|\leq 2N_{1}.$ Denote by $\Sigma_{Q}$ the set of $Q$-bad $\sigma\in\mathbb{R}$. Then
$\Sigma_{Q}\subset\bigcup_{1\leq l\leq N_{1}^{C}}J_{l},$
where each $J_{l}$ is an interval of length $\sim\varepsilon+\delta.$
Proof.
Note that $N=M^{\tilde{r}+1}\geq M^{2r_{0}+1}>(\varepsilon+\delta)^{-\frac{1}{10^{3}b^{2}d^{2}\tau_{b+2}^{2}}}$ from our assumptions. So we have
$N_{1}\sim(\log N)^{C^{\prime}}>c\log^{C^{\prime}}|\varepsilon+\delta|>\log^{2/\rho_{1}}|\varepsilon+\delta|$
assuming $C^{\prime}\rho_{1}>3,$ which implies
$(\varepsilon+\delta)^{-1}<e^{N_{1}^{\rho_{1}/2}}.$ (4.13)
So we can define for each $\bm{j}=(\bm{k},\bm{n})\in\Pi_{b,d}Q$ and $\xi=\pm 1$ the set
$J_{\bm{j},\xi}=\{\sigma\in\mathbb{R}:\ |\sigma+\bm{k}\cdot\bm{\omega}+\xi\mu_{\bm{n}}|\leq C(\varepsilon+\delta)\},$
where $C>1$ depends only on $\|\Delta\|,\|T_{\vec{u}^{(r_{1})}}\|.$ From (4.13) and the Neumann series, we must have
$\Sigma_{Q}\subset\bigcup_{\bm{j}\in\Pi_{b,d}Q,\xi=\pm 1}J_{\bm{j},\xi}.$
This proves Lemma 4.2.
Now fix $I_{0}\in\mathcal{I}_{r_{1}}.$ Solving the $Q$-equation at $r=r_{1}$ leads to the graph $\Gamma_{r_{1}}$. Then $\tilde{I}=\Pi_{\bm{\omega}}(\Gamma_{r_{1}}\cap I_{0})$ is an interval of size $\sim\varepsilon+\delta.$ For $\bm{\omega}\in\tilde{I}$, let $\tilde{\Sigma}$ be the set of $\sigma\in\mathbb{R}$ so that either (4.11) or (4.12) fails for some $\Pi_{b,d}Q\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ with $|\bm{n}|\leq 2N_{1}$. Then by Lemma 4.2, $\tilde{\Sigma}$ can be covered by $N_{1}^{C}$ intervals of size $\sim\varepsilon+\delta.$ Pick $J$ to be one of such intervals and consider $\mathcal{K}:=\tilde{I}\times(\tilde{\Sigma}\cap J)\subset I_{1}\times J\subset\mathbb{R}^{b}\times\mathbb{R}.$ We will show that $\mathcal{K}$ is a semi-algebraic set of degree $\deg\mathcal{K}\leq N_{1}^{C}M^{C{\tilde{r}}_{1}^{3}}$ and measure ${\rm meas}(\mathcal{K})\leq C(\delta+\varepsilon)^{b}e^{-N_{1}^{\rho_{1}}}$. This measure bound follows directly from the LDT at scale $N_{1}$ and the Fubini’s theorem. For the semi-algebraic description of $\mathcal{K}$, we need the following lemma.
Lemma 4.3 (Tarski-Seidenberg Principle, [Bou07]).
Denote by $(\bm{x},\bm{y})\in\mathbb{R}^{d_{1}+d_{2}}$ the product variable. If $A\subset\mathbb{R}^{d_{1}+d_{2}}$ is semi-algebraic of degree $B$, then its projections $\Pi_{\bm{x}}A\subset\mathbb{R}^{d_{1}}$ and $\Pi_{\bm{y}}A\subset\mathbb{R}^{d_{2}}$ are semi-algebraic of degree at most $B^{C}$, where $C=C(d_{1},d_{2})>0$.
We define $X\subset I_{0}\times J$ to be the set of all $(\bm{\omega},\bm{a},\sigma)$ so that either (4.11) or (4.12) fails for some $\Pi_{b,d}Q\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ with $|\bm{n}|\leq 2N_{1}$. By standard arguments, we can regard $X$ as a semi-algebraic set of degree $N_{1}^{C}$. From (Hiv) and solving the $Q$-equation for $r=r_{1}$, we know that $\Gamma_{r_{1}}\cap I_{0}$ is given by an algebraic equation in $\bm{\omega},\bm{a}$ (cf. (4.4)) of degree $M^{C{\tilde{r}}_{1}^{3}}$. Note also that
$\mathcal{K}=\Pi_{\bm{\omega},\sigma}\left(X\cap((\Gamma_{r_{1}}\cap I_{0})\times\mathbb{R})\right)$
which together with Lemma 4.3 implies $\deg\mathcal{K}\leq N_{1}^{C}M^{C{\tilde{r}}_{1}^{3}}.$ Our aim is to estimate
${\rm meas}_{b}(\bigcup_{N/2\leq|\bm{k}|\leq N}\{\bm{\omega}:\ (\bm{\omega},\bm{k}\cdot\bm{\omega})\in\mathcal{K}\}).$ (4.14)
At this stage, we need a projection lemma of Bourgain.
Lemma 4.4 ([Bou07]).
Let $A\subset[0,1]^{d=d_{1}+d_{2}}$ be a semi-algebraic set of degree $\deg(A)=B$ and $\mathrm{meas}_{d}(A)\leq\eta$, where $\log B\ll\log\frac{1}{\eta}.$ Denote by $(\bm{x},\bm{y})\in[0,1]^{d_{1}}\times[0,1]^{d_{2}}$ the product variable. Suppose $\eta^{\frac{1}{d}}\leq\epsilon.$ Then there is a decomposition of $A$ as
$A=A_{1}\cup A_{2}$
with the following properties: The projection of $A_{1}$ on $[0,1]^{d_{1}}$ has small measure
$\mathrm{meas}_{d_{1}}(\Pi_{\bm{x}}A_{1})\leq{B}^{C(d)}\epsilon,$
and $A_{2}$ has the transversality property
$\mathrm{meas}_{d_{2}}(\mathcal{L}\cap A_{2})\leq B^{C(d)}\epsilon^{-1}\eta^{\frac{1}{d}},$
where $\mathcal{L}$ is any $d_{2}$-dimensional hyperplane in $\mathbb{R}^{d}$ s.t., $\max\limits_{1\leq j\leq d_{1}}|\Pi_{\mathcal{L}}(\bm{e}_{j})|<{\epsilon},$ where we denote by $\bm{e}_{1},\cdots,{\bm{e}}_{d_{1}}$ the $\bm{x}$-coordinate vectors.
Remark 4.3.
This lemma permits scaling. It is based on the Yomdin-Gromov triangulation theorem [Gro87]; for a complete proof, see [BN19].
Taking $\eta=C(\varepsilon+\delta)^{b}e^{-N_{1}^{\rho_{1}}},\epsilon=10/N$, we obtain by $C^{\prime}\rho_{1}>3$ that
$C(\varepsilon+\delta)^{b}e^{-\frac{1}{b+1}N_{1}^{\rho_{1}}}\leq C(\varepsilon+\delta)^{b}e^{-\frac{1}{b+1}(\log N)^{C^{\prime}\rho_{1}}}\ll\epsilon.$
We have a decomposition
$\mathcal{K}=\mathcal{K}_{1}\cup\mathcal{K}_{2},$
where ${\rm meas}\ (\Pi_{\bm{\omega}}\mathcal{K}_{1})\leq C(\varepsilon+\delta)^{b+1}N_{1}^{C}M^{C{\tilde{r}}_{1}^{3}}M^{-{\tilde{r}}}\leq(\varepsilon+\delta)^{b+1}M^{-\frac{\tilde{r}}{2}}$ and the factor $(\varepsilon+\delta)^{b+1}$ comes from scaling when applying Lemma 4.4. Note however that $|\bm{k}|\geq N/2$ can not ensure $\min_{1\leq l\leq b}|k_{l}|\geq N/2$. So the hyperplane $\{(\bm{\omega},\bm{k}\cdot\bm{\omega})\}$ does not satisfy the steepness condition. To address this issue, we will need to take into account of all possible directions and apply Lemma 4.4 $b$ times, as done by Bourgain (cf. (3.26) of [Bou07] and also (135) of [LW22]). This then leads to an upper bound $\ C(\varepsilon+\delta)^{b+1}M^{-\frac{{\tilde{r}}+1}{C(b)}}$ on (4.14), where $C(b)>0$ only depends on $b$.
Taking into account of all $J$, we have shown the existence of $\tilde{I}_{1}\subset\tilde{I}$ with ${\rm meas}_{b}(\tilde{I}_{1})\leq C(\varepsilon+\delta)^{b+1}N_{1}^{C}M^{-\frac{{\tilde{r}}+1}{C(b)}}\leq C(\varepsilon+\delta)^{b+1}M^{-\frac{{\tilde{r}}+1}{C(b)}}$ so that for $\bm{\omega}\in\tilde{I}\setminus\tilde{I}_{1}$, the estimates (4.9) and (4.10) hold for all $Q\in\mathcal{C}_{2}$. Let $I_{0}$ range over $\mathcal{I}_{r_{1}}$. The total measure removed from $\Pi_{\bm{\omega}}\Gamma_{r_{1}}$ is at most $C(\varepsilon+\delta)^{b+1}M^{{\tilde{r}}_{1}^{C}}M^{-\frac{{\tilde{r}}+1}{C(b)}}\leq(\varepsilon+\delta)^{b+1}M^{-\frac{{\tilde{r}}+1}{C(b)}}$. Since (4.9) and (4.10) allow $O(e^{-N^{2}_{1}})$ perturbation of $(\bm{\omega},\bm{a})$, and $\|\Gamma_{r}-\Gamma_{r_{1}}\|\lesssim\delta_{r_{1}}\ll e^{-N^{2}_{1}}$, we obtain a subset $\Gamma_{r}^{\prime}\subset\Gamma_{r}$ with $(\varepsilon+\delta)^{b+1}M^{-\frac{{\tilde{r}}+1}{C(b)}}$ so that (4.9) and (4.10) hold on
$\bigcup_{I\in\mathcal{I}_{r_{1}}}(I\cap(\Gamma_{r}\setminus\Gamma_{r}^{\prime})),$
and hence on
$\bigcup_{I\in\mathcal{I}_{r}}(I\cap(\Gamma_{r}\setminus\Gamma_{r}^{\prime})).$
Since $M^{{\tilde{r}}^{C}}+e^{N_{1}^{\rho_{2}}}\ll M^{({\tilde{r}}+1)^{C}}$, the estimates (4.5) and (4.9) allow $O(M^{-({\tilde{r}}+1)^{C}})$ perturbation of $(\bm{\omega},\bm{a})$. Combining conclusions in the above two cases and the resolvent identity of [HSSY20] (cf. Lemma 3.6) gives a collection $\mathcal{I}_{r+1}$ of intervals of size $M^{-({\tilde{r}}+1)^{C}}$ so that for $I\in\mathcal{I}_{r+1}$,
$\|H_{N}^{-1}(\vec{u}^{(r)})\|$ $\leq M^{-({\tilde{r}}+1)^{C}},$
$|H_{N}^{-1}(\vec{u}^{(r)})((\bm{j},\xi);(\bm{j}^{\prime},\xi^{\prime}))|$ $\leq e^{-c_{r+1}|\bm{j}-\bm{j}^{\prime}|}\ {\rm for}\ |\bm{j}-\bm{j}^{\prime}|>({\tilde{r}}+1)^{C},$
which concludes (Hiv, c) at the scale $r+1.$ We remark that in the above exponential off-diagonal decay estimates, $c_{r+1}\geq c_{r}-(\log M)^{8}({\tilde{r}}+1)^{-8}$, which implies $\inf_{r}c_{r}>0$. This explains why we need the sublinear distant off-diagonal decay (i.e., $\rho_{3}\in(0,1)$) in LDT.
Finally, we have
${\rm meas}_{b}\left(\Gamma_{r}\cap(\bigcup_{I^{\prime}\in\mathcal{I}_{r}}I^{\prime}\setminus\bigcup_{I\in\mathcal{I}_{r+1}}I)\right)\leq{\rm meas}(\Gamma_{r}^{\prime})\leq(\varepsilon+\delta)^{b+1}M^{-\frac{{\tilde{r}}+1}{C(b)}}.$
Recalling that $\bm{\omega}\to\bm{a}$ is a $C^{1}$ diffeomorphism and $\det(\frac{\partial\bm{\omega}}{\partial\bm{a}})\sim\delta^{b}$, we obtain since $\varepsilon\lesssim\delta$
${\rm meas}\left(\Pi_{\bm{a}}(\Gamma_{r}\cap(\bigcup_{I^{\prime}\in\mathcal{I}_{r}}I^{\prime}\setminus\bigcup_{I\in\mathcal{I}_{r+1}}I)))\right)\leq M^{-\frac{{\tilde{r}}+1}{C(b)}},$
which yields (Hiv, d) at the step $r+1.$square
Proof of Theorem 1.1.
The proof of Theorem 1.1 is a direct corollary of the inductive theorem. We refer to [BW08] and sects. V and VI, [KLW24] for further details. ∎
Acknowledgements
Y. Shi was supported by the NSFC (No. 12271380). W.-M. Wang acknowledges support from the CY Initiative of Excellence, “Investissements d’Avenir” Grant No. ANR-16-IDEX-0008.
Appendix A Diophantine properties of quasi-periodic functions
A.1. Introduction and main results
Consider a periodic function $f$ on $\mathbb{R}^{d}$ with period $1$ in each variable,
$f(\bm{\theta})=f(\theta_{1},\cdots,\theta_{d})=f(\theta_{1}+1,\cdots,\theta_{d}+1).$
We identify $f$ with a function defined on the torus $\mathbb{T}^{d}:=[0,1)^{d}$. Let ${\bm{\alpha}}=(\alpha_{1},\cdots,\alpha_{d})\in(0,1]^{d}$. Given ${\bm{n}}\in\mathbb{Z}^{d}$, set ${\bm{n}}{\bm{\alpha}}=(n_{1}\alpha_{1},\cdots,n_{d}\alpha_{d})$, and we call $g$,
$g({\bm{\theta}})=f({\bm{n}}{\bm{\alpha}}+{\bm{\theta}})=f(n_{1}\alpha_{1}+\theta_{1},\cdots,n_{d}\alpha_{d}+\theta_{d}),$
a shift of $f$. The main purpose of this Appendix is to study Diophantine properties of finite families of such shifts and that of their square roots (assuming $f>0$), when $f$ is a trigonometric polynomial. We note that when $\bm{\alpha}$ is an irrational vector, considered as a function on $\mathbb{Z}^{d}$, $g$ is quasi-periodic. As we shall explain below, questions about Diophantine properties of shifts arise naturally from the analysis of the quasi-periodic linear Schrödinger equations and the quasi-periodic linear wave equations on $\mathbb{Z}^{d}$.
The linear Schrödinger equation is the following equation:
$iu_{t}+\varepsilon\Delta u+Vu=0,\,\text{on }\mathbb{Z}^{d},$ (A.1)
where $\varepsilon$ is a parameters in $[0,1]$, $\Delta$ is the usual discrete Laplacian, $\Delta{(\bm{n},\bm{n}^{\prime})}=\delta_{|{\bm{n}}-{\bm{n}^{\prime}}|_{1},1}$ with $|\bm{n}|_{1}$ the $\ell^{1}$ norm, and $V$ is a smooth, say analytic function, on $\mathbb{T}^{d}$, acting as a diagonal operator, with diagonals
$V_{\bm{n}}=V({\bm{\alpha}\bm{n}+\bm{\theta}})=V(n_{1}\alpha_{1}+\theta_{1},\cdots,n_{d}\alpha_{d}+\theta_{d});$
while the linear wave equation:
$u_{tt}+\varepsilon\Delta u+Vu+mu=0,\,\text{on }\mathbb{Z}^{d},$ (A.2)
where $m>\|V\|_{\infty}$.
Using separation of variables, the analysis of the time evolution of (A.1) can be reduced to the spectral analysis of the Schrödinger operator
$H=\varepsilon\Delta+V,\,\text{on }\ell^{2}(\mathbb{Z}^{d});$ (A.3)
while for (A.2), the wave operator
$W=\sqrt{\varepsilon\Delta+V+m},\,\text{on }\ell^{2}(\mathbb{Z}^{d}),$ (A.4)
where $m>\|V\|_{\infty}+2d$. The eigenvalues of $H$ and $W$ serve as characteristic frequencies for the time evolution.
When $\varepsilon$ is small, to zeroth order approximation, $V_{\bm{n}}=V({\bm{\alpha}\bm{n}+\bm{\theta}})$ are the eigenvalues of $H$; while $\sqrt{V_{\bm{n}}+m}=\sqrt{V({\bm{\alpha}\bm{n}+\bm{\theta}})+m}$ are that of $W$. This is why we are interested in shifts and square roots of shifts of $V$. The linear combinations of shifts enter in the analysis of the nonlinear Schrödinger equation (NLS):
$iu_{t}+\varepsilon\Delta u+Vu+\delta|u|^{2p}u=0,\,p\in\mathbb{N},\text{on }\mathbb{Z}^{d};$ (A.5)
while the square roots of shifts enter in that of the nonlinear wave equation (NLW):
$u_{tt}+\varepsilon\Delta u+Vu+mu+\delta u^{2p+1}=0,\,p\in\mathbb{N},\text{on }\mathbb{Z}^{d}.$ (A.6)
In particular, Diophantine properties of shifts and square roots of shifts are essential when seeking quasi-periodic in time solutions (cf. [SW23]), or when proving long time stability of more general solutions.
Below we provide such Diophantine estimates when $V$ are (arbitrary) trigonometric polynomials. These estimates hold on a large set in $(\bm{\alpha},\bm{\theta})$. The main idea is to utilize an appropriate generalized Wronskian. Given the relation between analytic functions (hence trigonometric polynomials) and generalized Wronskians, this approach also seems natural in hindsight.
A.1.1. The Diophantine estimates
For $\bm{x},\bm{y}\in\mathbb{R}^{d}$, denotel
$\bm{x}\bm{y}$ $=(x_{1}y_{1},\cdots,x_{d}y_{d})\in\mathbb{R}^{d},$
$\bm{x}\cdot\bm{y}$ $=x_{1}y_{1}+\cdots+x_{d}y_{d}\in\mathbb{R},$
and consider
$P(\bm{\theta})=\sum_{\bm{\ell}\in\Gamma_{K}}p_{\bm{\ell}}\cos 2\pi(\bm{\ell}\cdot\bm{\theta})\ {\rm with}\ \bm{\theta}\in[0,1]^{d},$
where $\Gamma_{K}\subset[-K,K]^{d}\setminus\{0\},\ K\geq 1$ has the property that
$\text{if }\bm{\ell}\in\Gamma_{K},{\rm then},\bm{\ell}_{k}\neq 0,\forall k=1,2,...,d,$ (A.7)
$\text{and if }\ \bm{\ell},\bm{\ell}^{\prime}\in\Gamma_{K},\ {\rm then}\ \bm{\ell}+\bm{\ell}^{\prime}\neq 0;$
and
$\bm{p}=(p_{\bm{\ell}})_{\bm{\ell}\in\Gamma_{K}}\in\mathbb{R}^{\#\Gamma_{K}}\setminus\{0\}.$
Remark A.1.
Note that (A.7) is a non-degeneracy condition on each term in the polynomial $P(\bm{\theta})$, cf. the paragraph below (1.2).
For $\bm{n}\in\mathbb{Z}^{d}$, let
$V_{\bm{n}}=V_{\bm{n}}(\bm{\theta},\bm{\alpha})=P(\bm{\theta}+\bm{n}\bm{\alpha}),$
where
$\bm{\alpha}\in(0,1]^{d}.$
Fix
$\bm{n}_{1},\cdots,\bm{n}_{D}\in\mathbb{Z}^{d}$
to be $D$, $D\geq 2$, distinct lattice points. We are mainly concerned with the following.
  • NLS case: Let
    $\bm{\omega}_{{\rm NLS}}=(\omega_{1},\cdots,\omega_{D}),\ \omega_{\ell}=P({\bm{\theta}+\bm{n}_{\ell}\bm{\alpha}})\ (1\leq\ell\leq D).$
  • NLW case: Let
    $m\in[\|P\|_{\infty}+1,\|P\|_{\infty}+2]$
    with $\|P\|_{\infty}=\sup_{\bm{\theta}\in[0,1]^{d}}|P(\bm{\theta})|.$ Define
    $\bm{\omega}_{{\rm NLW}}=(\omega_{1},\cdots,\omega_{D}),\ \omega_{\ell}=\sqrt{P({\bm{\theta}+\bm{n}_{\ell}\bm{\alpha}})+m}\ (1\leq\ell\leq D).$
For ${\bm{x}}=(x_{1},\cdots,x_{d})$, let
$|\bm{x}|=\max_{1\leq i\leq d}|x_{i}|.$
Throughout this Appendix, all constants $C,c>0$ are assumed to depend only on $R=D\cdot(\#\Gamma_{K}),\max_{\ell}|\bm{n}_{\ell}|,d,|\bm{p}|$. For convenience, we omit this dependence.
Denote by ${\rm meas(\cdot)}$ the Lebesgue measure. Our main result on NLS frequencies is
Theorem A.1.
Let $0<\delta\ll 1$. We have for all $0<\gamma\leq\delta^{10dR^{2}}$ and $\tau\geq 4R^{2}$ ($R=D\cdot(\#\Gamma_{K})=D\frac{(2K+1)^{d}-1}{2}$),
${\rm meas}({\rm DC}_{\rm nls}(\gamma;\tau))\geq 1-C\delta,$
where
$\ \ \ {\rm DC}_{\rm nls}(\gamma;\tau)$
=$=\left\{(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}:\ |\bm{k}\cdot\bm{\omega}|\geq\frac{\gamma}{|\bm{k}|^{\tau}}\ {\rm for}\ \forall\ \bm{k}\in\mathbb{Z}^{D}\setminus\{0\}\right\}$
with $\bm{\omega}=\bm{\omega}_{\rm NLS}$.
Remark A.2.
In contrast, if $\bm{\omega}$ were a free parameter varying in $[0,1]^{D}$, we would have
${\rm meas}\left(\left\{\bm{\omega}\in[0,1]^{D}:\ |\bm{k}\cdot\bm{\omega}|\geq\frac{\gamma}{|\bm{k}|^{\tau}}\ {\rm for}\ \forall\ \bm{k}\in\mathbb{Z}^{D}\setminus\{0\}\right\}\right)\geq 1-O(\gamma).$
So the absence of independence of the coordinates of $\bm{\omega}_{\rm NLS}$ may lead to the corresponding measure bound (from $1-O(\gamma)$ to) $1-O(\gamma^{\frac{1}{10dR^{2}}}).$
Next, we consider the NLW case. Denote by $\{\bm{e}_{\ell}\}_{1\leq\ell\leq D}$ the canonical basis of $\mathbb{Z}^{D}.$ Define for $\gamma>0,$
${G}_{\gamma}=\left\{(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}:\ \min_{1\leq\ell<\ell^{\prime}\leq D}|(\bm{e}_{\ell}-\bm{e}_{\ell^{\prime}})\cdot\bm{\omega}_{\rm NLS}|\leq{\gamma}\right\}.$ (A.8)
For the NLW frequencies, we have
Theorem A.2.
Let $0<\delta\ll 1,0<\gamma\leq\delta^{10dR^{2}}$ and $(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}\setminus G_{\gamma}$ with $G_{\gamma}$ given by (A.8). We have for all $0<\gamma_{1}\leq\gamma^{2D(D^{2}-1)}$ and $\tau>2D^{2}$,
${\rm meas}({\rm DC}_{\rm nlw}(\gamma_{1};\tau))\geq 1-C\gamma_{1}^{\frac{1}{2D}}$
where
$\ \ \ {\rm DC}_{\rm nlw}(\gamma_{1};\tau)$
=$=\left\{m\in[\|P\|_{\infty}+1,\|P\|_{\infty}+2]:\ |\bm{k}\cdot\bm{\omega}|\geq\frac{\gamma_{1}}{|\bm{k}|^{\tau}}\ {\rm for}\ \forall\ \bm{k}\in\mathbb{Z}^{D}\setminus\{0\}\right\}$
with $\bm{\omega}=\bm{\omega}_{\rm NLW}$.
Remark A.3.
The above two theorems and their proofs can be easily generalized to finite Fourier series with both cosine and sine terms.
A.2. Some useful lemmas
In this section, we will introduce some useful lemmas.
Lemma A.3 ([KM98]).
Let $I\subset\mathbb{R}$ be an interval of finite length (i.e., $0<|I|<\infty$) and $k\geq 1.$ If $f\in C^{k}(I;\mathbb{R})$ satisfies
$\inf_{x\in I}|\frac{d^{k}}{dx^{k}}f(x)|\geq A>0,$ (A.9)
then for all $\varepsilon>0,$
${\rm meas}(\{x\in I:\ |f(x)|\leq\varepsilon\})\leq\zeta_{k}(\frac{\varepsilon}{A})^{\frac{1}{k}},$ (A.10)
where $\zeta_{k}=k(k+1)[(k+1)!]^{\frac{1}{k}}.$
Proof.
The following proof is taken from Kleinbock-Margulis [KM98].
From (A.9) and the Rolle’s theorem, the equation $\frac{d}{dx}f(x)=0$ has at most $k$ zeros on $I$. This then implies that the set $\{x\in I:\ |f(x)|\leq\varepsilon\}$ consists of at most $k+1$ intervals. Denote by $I_{1}$ one of those subintervals having maximal length. So we get
${\rm meas}(\{x\in I:\ |f(x)|\leq\varepsilon\})\leq(k+1)|I_{1}|.$
To prove (A.10), it remains to estimate $|I_{1}|>0$. We divide $I_{1}$ into $k$ equal parts by points $x_{1},\cdots,x_{k+1}.$ Let $P(x)$ be the Lagrange polynomial of $f$ on points $x_{1},\cdots,x_{k+1},$ namely,
$P(x)=\sum_{i=1}^{k+1}f(x_{i})\frac{\prod_{1\leq j\neq i\leq k+1}(x-x_{i})}{\prod_{1\leq j\neq i\leq k+1}(x_{j}-x_{i})}.$
By applying the Rolle’s theorem $k$ times and since $f(x_{i})=P(x_{i})$ ($11\leq i\leq k+$ 1), we can find $x_{0}\in I_{1}$ so that
$\frac{d^{k}}{dx^{k}}f(x_{0})=\frac{d^{k}}{dx^{k}}P(x_{0})=k!\sum_{i=1}^{k+1}\frac{f(x_{i})}{\prod_{1\leq j\neq i\leq k+1}(x_{j}-x_{i})},$
which together with (A.9) yields
$A\leq k!\sum_{i=1}^{k+1}\frac{\varepsilon}{|I_{1}|^{k}k^{-k}}.$
This implies
$|I_{1}|\leq k[(k+1)!]^{\frac{1}{k}}(\frac{\varepsilon}{A})^{\frac{1}{k}}.$
We have finished the proof. ∎
In the following we turn to the analysis of multi-variable functions. We first introduce the notations.
  • For any ${\bm{x}}=(x_{1},\cdots,x_{d})$ and any $d\geq 1$, let
    $|{\bm{x}}|_{2}=\sqrt{\sum_{i=1}^{d}|x_{i}|^{2}},\ |\bm{x}|_{1}=\sum_{i=1}^{d}|x_{i}|.$
  • Let $\bm{a}=(a_{1},\cdots,a_{d}),\bm{b}=(b_{1},\cdots,b_{d})$. Define the $d$-dimensional interval to be
    $I=I_{\bm{a},\bm{b}}=\prod_{i=1}^{d}[a_{i},b_{i}]\subset\mathbb{R}^{d}.$ (A.11)
    Denote
    $|\bm{a}\vee\bm{b}|=\sum_{i=1}^{d}\max(|a_{i}|,|b_{i}|).$
  • For ${\bm{\beta}}=(\beta_{1},\cdots,\beta_{d})\in\mathbb{R}^{d}\setminus\{0\}$, define
    $d_{\bm{\beta}}:=\sum_{i=1}^{d}\beta_{i}{\partial}_{i},\ \partial_{i}:=\frac{\partial}{\partial x_{i}}.$
    We also write for $l\geq 1,$
    $d_{\bm{\beta}}^{l}=d_{\bm{\beta}}\cdots d_{\bm{\beta}}\ (l-{\rm times}).$
    Denote
    $\nabla=(\partial_{1},\cdots,\partial_{d}).$
  • For a multi-index ${\bm{\gamma}}=(\gamma_{1},\cdots,\gamma_{d})\in\mathbb{N}^{d}$, denote
    $|\bm{\gamma}|=\sum_{i=1}^{d}\gamma_{i},\ \partial^{\bm{\gamma}}=\partial^{\gamma_{1}}_{1}\cdots\partial^{\gamma_{d}}_{d}.$
We have
Lemma A.4.
Fix $k\in\mathbb{N}$. Let $I=I_{\bm{a},\bm{b}}\subset\mathbb{R}^{d}$ be defined by (A.11). Assume that the function $f\in C^{k+1}(I;\mathbb{R})$ satisfies for some $A>0,$
$\inf_{\bm{x}\in I}\sup_{1\leq l\leq k}\left|{d_{\bm{\beta}}^{l}}f(\bm{x})\right|\geq A.$ (A.12)
Let
$\|f\|_{k+1}=\sup_{\bm{x}\in I}\sup_{1\leq|\bm{\gamma}|\leq k+1}\left|{\partial^{\bm{\gamma}}}f(\bm{x})\right|<\infty.$
Then for $0<\varepsilon<A<1,$
$\ \ \ {\rm meas}(\{\bm{x}\in I:\ |f(\bm{x})|\leq\varepsilon\})$
$\leq C(\bm{\beta},k,d)(\|f\|_{k+1}|\bm{b}-\bm{a}|_{2}+1)^{d}{|\bm{a}\vee\bm{b}|}^{d-1}A^{-(d+\frac{1}{k})}{\varepsilon}^{\frac{1}{k}},$ (A.13)
where $C=C(\bm{\beta},k,d)>0$ depends only on $\bm{\beta},k,d$ (but not on $f$).
Proof.
The proof can be divided into two steps:
Step 1. Assume that for some $1\leq k_{1}\leq k$ and interval $I_{1}=\prod_{i=1}^{d}[h_{i},m_{i}],$
$\inf_{\bm{x}\in I_{1}}\left|{d_{\bm{\beta}}^{k_{1}}}f(\bm{x})\right|\geq A.$ (A.14)
We will combine Lemma A.3 and Fubini’s theorem to deal with the measure estimate. Let $\bm{e}_{1}=\frac{\bm{\beta}}{|\bm{\beta}|_{2}}$ and choose a normalized orthogonal basis $\bm{e}_{2},\cdots,\bm{e}_{d}\in\mathbb{R}^{d}$ of
$\bm{\beta}^{\perp}=\{\bm{x}\in\mathbb{R}^{d}:\ \bm{x}\cdot\bm{\beta}=0\}.$
We can then define the coordinate transform:
$\bm{x}=\bm{\varphi}(s,\bm{y}),(s,\bm{y})\in\mathbb{R}\times\mathbb{R}^{d-1},$
via
$\bm{\varphi}(s,\bm{y})=s\bm{e}_{1}+\sum_{i=2}^{d}y_{i}\bm{e}_{i}.$
So the Jacobian is
$|\det\frac{\partial\bm{x}}{\partial(s,\bm{y})}|=1.$
Define
$S_{\varepsilon}=\{\bm{x}\in I_{1}:\ |f(\bm{x})|\leq\varepsilon\},$
and denote by $\chi_{(\cdot)}$ the indicator function. We have
[…]
We proceed to control $\bm{\varphi}^{-1}(I_{1})$. From $\bm{x}=s\bm{e}_{1}+\sum_{i=2}^{d}y_{i}\bm{e}_{i}\in I_{1}$, we obtain
$s=\bm{x}\cdot\bm{e}_{1}\in[-|\bm{h}\vee\bm{m}|,|\bm{h}\vee\bm{m}|].$
Similarly, we also have $y_{i}\in[-|\bm{h}\vee\bm{m}|,|\bm{h}\vee\bm{m}|]$ for $2\leq i\leq d,$ which yields $\bm{\varphi}^{-1}(I_{1})\subset[-|\bm{h}\vee\bm{m}|,|\bm{h}\vee\bm{m}|]^{d}$. We define
$\Pi_{1}\bm{\varphi}^{-1}(I_{1})=\{\bm{y}\in\mathbb{R}^{d-1}:\ \exists\ s\in\mathbb{R}\ s.t.,\ (s,\bm{y})\in\bm{\varphi}^{-1}(I_{1})\}$
and, define for $\bm{y}\in\Pi_{1}\bm{\varphi}^{-1}(I_{1}),$ the interval $I_{\bm{y}}=\{s\in\mathbb{R}:\ (s,\bm{y})\in\bm{\varphi}^{-1}(I_{1})\}$ of finite length. Applying the Fubini’s theorem together with (A.15) implies that
[…]
$\ \ \ ({\rm since}\ \Pi_{1}\bm{\varphi}^{-1}(I_{1})\subset[-|\bm{h}\vee\bm{m}|,|\bm{h}\vee\bm{m}|]^{d-1}).$
For fixed $\bm{y}\in\Pi_{1}\bm{\varphi}^{-1}(I_{1})$, we get
[…]
$={\rm meas}(\{s\in I_{\bm{y}}:\ |f(s\bm{e}_{1}+\sum_{i=2}^{d}y_{i}\bm{e}_{i})|\leq\varepsilon\}).$
Denote $g(s)=f(s\bm{e}_{1}+\sum_{i=2}^{d}y_{i}\bm{e}_{i})$. From (A.14), we have for all $s\in I_{\bm{y}},$
$|\frac{d^{k_{1}}}{ds^{k_{1}}}g(s)|=|\frac{1}{|\bm{\beta}|_{2}^{k_{1}}}d^{k_{1}}_{\bm{\beta}}f(s\bm{e}_{1}+\sum_{i=2}^{d}y_{i}\bm{e}_{i})|\geq\frac{A}{|\bm{\beta}|_{2}^{k_{1}}}.$
So applying Lemma A.3 yields
${\rm meas}(S_{\varepsilon})\leq\zeta_{k_{1}}2^{d-1}|\bm{h}\vee\bm{m}|^{d-1}|\bm{\beta}|_{2}(\frac{\varepsilon}{A})^{\frac{1}{k_{1}}}.$ (A.16)
Step 2. We deal with the general case in this step. So we first divide each one dimensional interval $[a_{i},b_{i}]$ into $N$ (will be specified below) equal subintervals $I_{i,j_{i}}$ ($1\leq j_{i}\leq N$). This then induces a decomposition of $I$, namely,
$I=\bigcup_{1\leq j_{1},\cdots,j_{d}\leq N}\prod_{i=1}^{d}I_{i,j_{i}}=\bigcup_{\bm{J}\in([1,N]\cap\mathbb{N})^{d}}I_{\bm{J}}.$
We will apply the argument proved in Step 1 on each $I_{\bm{J}}$. For this purpose, we fix $I_{\bm{J}}$. For any $\bm{x},\bm{y}\in I_{\bm{J}}$, we have
$|\bm{x}-\bm{y}|_{2}$ $\leq\sqrt{\sum_{i=1}^{d}\frac{|b_{i}-a_{i}|^{2}}{N^{2}}}$
$\leq\frac{|\bm{b}-\bm{a}|_{2}}{N}.$
Fixing any $\bm{x}_{0}\in I_{\bm{J}}$, we get from the assumption (A.12) that there exists $1\leq k_{1}\leq k$ with
$|d_{\bm{\beta}}^{k_{1}}f(\bm{x}_{0})|\geq A>0.$
As a result, for any $\bm{x}\in I_{\bm{J}}$, we obtain
$|d_{\bm{\beta}}^{k_{1}}f(\bm{x})|$ $\geq|d_{\bm{\beta}}^{k_{1}}f(\bm{x}_{0})|-|d_{\bm{\beta}}^{k_{1}}f(\bm{x})-d_{\bm{\beta}}^{k_{1}}f(\bm{x}_{0})|$
$\geq A-\sup_{\bm{x}\in I_{\bm{J}}}|\nabla d_{\bm{\beta}}^{k_{1}}f(\bm{x})|_{2}\cdot|\bm{x}-\bm{x}_{0}|_{2}$
$\geq A-\frac{\|f\|_{k_{1}+1}L_{\bm{\beta},d}|\bm{b}-\bm{a}|_{2}}{N}$
$\geq A-\frac{\|f\|_{k+1}L_{\bm{\beta},d}|\bm{b}-\bm{a}|_{2}}{N},$ (A.17)
where
$L_{\bm{\beta},d}=\sqrt{d}\sum_{1\leq l_{1},\cdots,l_{d}\leq d}|\beta_{l_{1}}\cdots\beta_{l_{d}}|.$
In fact, $L_{\bm{\beta},d}$ is derived from
$|\nabla d_{\bm{\beta}}^{k_{1}}f(\bm{x})|_{2}$ $=\sqrt{\sum_{j=1}^{d}|\sum_{1\leq l_{1},\cdots,l_{d}\leq d}\beta_{l_{1}}\cdots\beta_{l_{d}}\partial_{jl_{1}\cdots l_{d}}^{k_{1}+1}f(\bm{x})|^{2}}$
$\leq\|f\|_{k_{1}+1}\sqrt{\sum_{j=1}^{d}(\sum_{1\leq l_{1},\cdots,l_{d}\leq d}|\beta_{l_{1}}\cdots\beta_{l_{d}}|)^{2}}$
$=\|f\|_{k_{1}+1}L_{\bm{\beta},d}.$
From (A.17), we can set
$N=[2\|f\|_{k+1}L_{\bm{\beta},d}|\bm{b}-\bm{a}|_{2}A^{-1}]+1$
so that
$\inf_{\bm{x}\in I_{\bm{J}}}|d_{\bm{\beta}}^{k_{1}}f(\bm{x})|\geq A/2>0,$
where $[x]$ denotes the integer part of $x\in\mathbb{R}.$ Thus applying (A.16) yields (since $I_{J}\subset I,0<\varepsilon<A<1$)
${\rm meas}(\{\bm{x}\in I_{\bm{J}}:\ |f(\bm{x})|\leq\varepsilon\})$ $\leq\zeta_{k_{1}}2^{d-1}|\bm{a}\vee\bm{b}|^{d-1}2^{\frac{1}{k_{1}}}|\bm{\beta}|_{2}(\frac{\varepsilon}{A})^{\frac{1}{k_{1}}}$
$\leq C(k,d)|\bm{a}\vee\bm{b}|^{d-1}|\bm{\beta}|_{2}(\frac{\varepsilon}{A})^{\frac{1}{k}}.$
Collecting all $I_{\bm{J}}$ leads to,
$\ \ \ {\rm meas}(\{\bm{x}\in I:\ |f(\bm{x})|\leq\varepsilon\})$
$\leq C(k,d)N^{d}|\bm{a}\vee\bm{b}|^{d-1}|\bm{\beta}|_{2}(\frac{\varepsilon}{A})^{\frac{1}{k}}$
$\leq C(\bm{\beta},k,d)(\|f\|_{k+1}|\bm{b}-\bm{a}|_{2}A^{-1}+1)^{d}|\bm{a}\vee\bm{b}|^{d-1}(\frac{\varepsilon}{A})^{\frac{1}{k}}$
$\leq C(\bm{\beta},k,d)(\|f\|_{k+1}|\bm{b}-\bm{a}|_{2}+1)^{d}|\bm{a}\vee\bm{b}|^{d-1}A^{-(d+\frac{1}{k})}{\varepsilon}^{\frac{1}{k}}.$
We have proven (A.4). ∎
Finally, we introduce a key lemma (cf. Proposition in Appendix B, [BGG85] for a more precise form) on determinant estimates.
Lemma A.5.
Let $\bm{v}^{(1)},\cdots,\bm{v}^{(r)}\in\mathbb{R}^{r}$ be $r$ linearly independent vectors with $|\bm{v}^{(l)}|_{1}\leq M$ for $1\leq l\leq r.$ Then for any $\bm{w}\in\mathbb{R}^{r}$, we have
$\max_{1\leq l\leq r}|\bm{w}\cdot\bm{v}^{(l)}|\geq r^{-3/2}M^{1-r}|\bm{w}|_{2}\cdot|\det\left[\bm{v}^{(l)}\right]_{1\leq l\leq r}|.$
Proof.
For completeness, we give a proof based on Cramer’s rule and the Hadamard’s inequality. We assume $\bm{v}^{(l)}$ is the row vector in $\mathbb{R}^{r}$ ($1\leq l\leq r$). Denote by $V=[\bm{v}^{(1)},\cdots,\bm{v}^{(r)}]$ the $r\times r$ matrix and consider the equation
$V\bm{w}=\bm{\xi},\ \bm{\xi}=(\xi_{1},\cdots,\xi_{r})\ {\rm with}\ \xi_{l}=\bm{w}\cdot\bm{v}^{(l)}.$
By the Cramer’s rule, we obtain
$\bm{w}=V^{-1}\bm{\xi}=\frac{V^{*}}{\det V}\bm{\xi},$
where $V^{*}=[\tilde{v}_{ij}]_{1\leq i,j\leq r}$ denotes the adjugate matrix of $V$. As a result, we get
$|\bm{w}|_{2}\cdot|\det V|$ $=|V^{*}\bm{\xi}|_{2}$
$=\sqrt{\sum_{i=1}^{r}|\sum_{j=1}^{r}\tilde{v}_{ij}\xi_{j}|^{2}}$
$\leq\max_{1\leq j\leq r}|\xi_{j}|\sqrt{\sum_{i=1}^{r}(\sum_{j=1}^{r}|\tilde{v}_{ij}|)^{2}}$
$\leq r^{3/2}\max_{1\leq j\leq r}|\xi_{j}|\max_{1\leq i,j\leq r}|\tilde{v}_{ij}|.$
It remains to bound $\max_{1\leq i,j\leq r}|\tilde{v}_{ij}|,$ which will be completed by using the Hadamard’s inequality. In fact, we have
$|\tilde{v}_{ij}|$ $\leq\prod_{1\leq l\neq i\leq r}\sqrt{\sum_{1\leq m\neq j\leq r}|v^{(l)}_{m}|^{2}}$
$\leq\prod_{1\leq l\neq i\leq r}{\sum_{1\leq m\neq j\leq r}|v^{(l)}_{m}|}$
$\leq M^{r-1}\ ({\rm since}\ \max_{1\leq l\leq r}|\bm{v}^{(l)}|_{1}\leq M).$
This finishes the proof. ∎
A.3. Proof of Theorem A.1
In this section we prove Theorem A.1.
Proof of Theorem A.1.
Recall that for $\bm{\omega}=\bm{\omega}_{\rm NLS},$
$\bm{k}\cdot\bm{\omega}$ $=\sum_{\ell=1}^{D}k_{\ell}P(\bm{\theta}+\bm{n}_{\ell}\bm{\alpha})$
$=\sum_{\ell=1}^{D}\sum_{\bm{\ell}\in\Gamma_{K}}k_{\ell}p_{\bm{\ell}}\cos 2\pi\bm{\ell}\cdot(\bm{\theta}+\bm{n}_{\ell}\bm{\alpha})$
$=(\bm{k}\otimes\bm{p})\cdot\bm{V},$
where
$\bm{k}\otimes\bm{p}=(k_{\ell}p_{\bm{\ell}})_{1\leq\ell\leq D,\bm{\ell}\in\Gamma_{K}},$
$\bm{V}=(\cos 2\pi\bm{\ell}\cdot(\bm{\theta}+\bm{n}_{\ell}\bm{\alpha}))_{1\leq\ell\leq D,\bm{\ell}\in\Gamma_{K}}.$
Denote $\bm{\xi}=(\bm{\beta},\bm{q})$ for $\bm{\beta},\bm{q}\in(0,1]^{d}$. Write also
$d_{\bm{\xi}}=\sum_{i=1}^{d}(\beta_{i}\frac{\partial}{\partial\alpha_{i}}+q_{i}\frac{\partial}{\partial\theta_{i}})$
and accordingly $d_{\bm{\xi}}^{l}$ for $l\geq 1.$
A key observation is that for all $j\geq 1$ and $\bm{\ell}\in\Gamma_{K}$,
$\ \ \ d_{\bm{\xi}}^{2j}\cos 2\pi\bm{\ell}\cdot(\bm{\theta}+\bm{n}_{\ell}\bm{\alpha})$
$=(-(2\pi(\bm{\ell}\bm{n}_{i})\cdot\bm{\beta}+2\pi\bm{q}\cdot\bm{\ell})^{2})^{j}\cos 2\pi\bm{\ell}\cdot(\bm{\theta}+\bm{n}_{\ell}\bm{\alpha}).$
This motivates us to consider the Wronskian
$W=[d_{\bm{\xi}}^{2j}{V}_{(\ell,\bm{\ell})}]_{(\ell,\bm{\ell})}^{1\leq j\leq R}\ {\rm with}\ R=D\cdot(\#\Gamma_{K}),$
which is a $R\times R$ real matrix. Direct computations show
$\ \ \ |\det W|$
$=\prod_{\ell,\bm{\ell}}|\cos 2\pi{\bm{\ell}}\cdot(\bm{\theta}+\bm{n}_{\ell}\bm{\alpha})|\times\prod_{\ell,\bm{\ell}}(2\pi)^{2}((\bm{\ell}\bm{n}_{\ell})\cdot\bm{\beta}+\bm{q}\cdot\bm{\ell})^{2}\times|\det W_{1}|,$ (A.18)
where $W_{1}$ is a Vandermonde matrix with
$|\det W_{1}|$ $=\prod_{(\ell,\bm{\ell})\neq(\ell^{\prime},\bm{\ell}^{\prime})}(2\pi)^{2}|((\bm{\ell}\bm{n}_{\ell})\cdot\bm{\beta}+\bm{q}\cdot\bm{\ell})^{2}-((\bm{\ell}^{\prime}\bm{n}_{\ell^{\prime}})\cdot\bm{\beta}+\bm{q}\cdot\bm{\ell}^{\prime})^{2}|.$ (A.19)
Then the proof can be divided into three steps:
Step 1. Prior conditions on $(\bm{\alpha},\bm{\theta})$
We aim to get a lower bound for (A.18), which requires the prior restrictions on $\bm{\alpha},\bm{\theta}.$ Let $\|x\|_{\mathbb{T}/2}={\rm dist}(x,\mathbb{Z}/2)$. From
$4\|x-\frac{1}{4}\|_{\mathbb{T}/2}\leq|\cos 2\pi x|\leq 2\pi\|x-\frac{1}{4}\|_{\mathbb{T}/2},$
we can define for $\delta>0$ the resonant set (of $\bm{\alpha},\bm{\theta}$)
$\ \ \ S_{\delta}$
$=\bigcup_{1\leq\ell\leq D,\bm{\ell}\in\Gamma_{K}}\left\{(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}:\ \|\bm{\ell}\cdot(\bm{\theta}+\bm{n}_{\ell}\bm{\alpha})-\frac{1}{4}\|_{\mathbb{T}/2}\leq\delta\right\}.$ (A.20)
We have
Lemma A.6.
Let $10<\delta<$ 1. Then $S_{\delta}$ can be covered by $C\delta^{-2d+1}$ intervals of side length $\leq\delta.$
Proof.
Denote $L=\max_{1\leq\ell\leq D}K|\bm{n}_{\ell}|$. It suffices to cover for each $\bm{n}_{\ell},\bm{\ell}$ the set
$S_{\bm{n}_{\ell},\bm{\ell}}=\{(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}:\ \|\bm{\ell}\cdot(\bm{\theta}+\bm{n}_{\ell}\bm{\alpha})-\frac{1}{4}\|_{\mathbb{T}/2}\leq\delta\}$
[…]
$\ \ \ I_{\bm{n}_{\ell},\bm{\ell}}$
$=\bigcup_{1\leq i\leq C_{1}}\{(\bm{\alpha},\bm{\theta}_{1}^{c},\theta_{1})\in\bm{C}_{i}\times[0,1]:\ \|\ell_{1}\theta_{1}+\bm{\ell}_{1}^{c}\cdot\bm{\theta}^{c}(i)+(\bm{n}_{\ell}\bm{\ell})\cdot\bm{\alpha}_{i}-\frac{1}{4}\|_{\mathbb{T}/2}\leq 2\delta\}$
$:=\bigcup_{1\leq i\leq C_{1}}J_{i}.$
We first claim that $S_{\bm{n}_{\ell},\bm{\ell}}\subset I_{\bm{n}_{\ell},\bm{\ell}}.$ Let $(\bm{\alpha},\bm{\theta})\in S_{\bm{n}_{\ell},\bm{\ell}}$. Then by the definition (of $S_{\bm{n}_{\ell},\bm{\ell}}$), we have
$\|\ell_{1}\theta_{1}+\bm{\ell}_{1}^{c}\cdot\bm{\theta}_{1}^{c}+(\bm{\ell}\bm{n}_{\ell})\cdot\bm{\alpha}-\frac{1}{4}\|_{\mathbb{T}/2}\leq\delta.$
Since $\bm{C}_{i}$ ($1\leq i\leq C_{1}$) covers $[0,1]^{2d-1}$, there is $1\leq i_{0}\leq C_{1}$ so that $(\bm{\alpha},\bm{\theta}_{1}^{c})\in\bm{C}_{i_{0}}$, namely, $|\bm{\alpha}-\bm{\alpha}_{i_{0}}|+|\bm{\theta}_{1}^{c}-\bm{\theta}^{c}(i_{0})|\leq\frac{\delta}{2L}.$ This implies that
$\ \ \ \|\ell_{1}\theta_{1}+\bm{\ell}_{1}^{c}\cdot\bm{\theta}^{c}(i_{0})+(\bm{n}_{\ell}\bm{\ell})\cdot\bm{\alpha}_{i_{0}}-\frac{1}{4}\|_{\mathbb{T}/2}$
$\leq\|\bm{\ell}\cdot(\bm{\theta}+\bm{n}_{\ell}\bm{\alpha})-\frac{1}{4}\|_{\mathbb{T}/2}+|\bm{\ell}_{1}^{c}\cdot(\bm{\theta}^{c}(i_{0})-\bm{\theta}_{1}^{c})|$
$\ \ +|(\bm{\ell}\bm{n}_{\ell})\cdot(\bm{\alpha}-\bm{\alpha}_{i_{0}})|$
$\leq 2\delta.$
The claim is proved. So it remains to cover each $J_{i}$ with intervals of side length $\leq\delta$. This can be accomplished by the fact that the set $\{\theta_{1}\in[0,1]:\ \|\ell_{1}\theta_{1}+x\|_{\mathbb{T}/2}\leq 2\delta\}$ ($x\in\mathbb{R}$) consists of $\sim|\ell_{1}|$ one dimensional intervals of length $2\delta/|\ell_{1}|$.
We have finished the proof. ∎
Given $\delta<1/10,$ we divide $[0,1]$ into $N=[\delta^{-2}]+1$ intervals of equal length $\frac{1}{N}\sim\delta^{2}$. This then induces a partition of $[0,1]^{2d}$ into cubes of equal side length $1/N$. We denote by $\Lambda_{1/N}(\bm{x}_{i})$ ($1\leq i\leq N^{2d}$) all these cubes. We have
Lemma A.7.
Let $100<\delta<1/$ 10 and let $S_{\delta}$ be defined by (A.20). Then we have for $N=[\delta^{-2}]+1$,
#$\#\{1\leq i\leq N^{2d}:\ \Lambda_{1/N}(\bm{x}_{i})\cap S_{\delta}\neq\emptyset\}\leq C\delta^{-4d+1}.$ (A.21)
In particular, we have
${\rm meas}\left(\bigcup_{1\leq i\leq N^{2d}:\ \Lambda_{1/N}(\bm{x}_{i})\cap S_{\delta}\neq\emptyset}\Lambda_{1/N}(\bm{x}_{i})\right)\leq C\delta.$ (A.22)
Proof.
Let $\bm{C}_{i}$ be given as in the proof of Lemma A.6. Then $S_{\delta}\subset\cup_{1\leq i\leq C\delta^{-2d+1}}\bm{C}_{i}$. So it suffices to bound
$L_{1}=\#\{1\leq i\leq N^{2d}:\ (\cup_{1\leq j\leq C\delta^{-2d+1}}\bm{C}_{j})\cap\Lambda_{1/N}(\bm{x}_{i})\neq\emptyset\}.$
Note that if $\Lambda_{1/N}(\bm{x}_{i})\cap\bm{C}_{j}\neq\emptyset$, then
$\Lambda_{1/N}(\bm{x}_{i})\subset\bm{C}_{j}^{\prime},$
where $\bm{C}_{j}^{\prime}$ is also an interval containing $\bm{C}_{j}$ and satisfying
${\rm dist}(\partial\bm{C}_{j},\partial\bm{C}_{j}^{\prime})\leq 10/N,$
with $\partial$ the boundary. Regarding the disjointness (of the exterior) of $\Lambda_{1/N}(\bm{x}_{i})$ for different $i$, we must have
$L_{1}\leq C\delta^{-2d+1}(\delta+10/N)^{2d}\delta^{-4d}\leq C\delta^{-4d+1}.$
This implies the bound (A.21). Then the measure bound follows directly from (A.22). ∎
From this lemma, we get immediately that
Lemma A.8.
Let $100<\delta<1/$ 10 and define $N=[\delta^{-2}]+1$. Then there is a collection of cubes $\Lambda_{1/N}(\bm{x}_{i})\subset[0,1]^{d+d}$ ($1\leq i\leq N_{1}$) with disjoint exteriors so that
${\rm meas}\left([0,1]^{2d}\setminus\cup_{1\leq i\leq N_{1}}\Lambda_{1/N}(\bm{x}_{i})\right)\leq C\delta,$
and if $(\bm{\alpha},\bm{\theta})\in\cup_{1\leq i\leq N_{1}}\Lambda_{1/N}(\bm{x}_{i}),$ then
$\prod_{1\leq\ell\leq D,\bm{\ell}\in\Gamma_{K}}|\cos 2\pi\bm{\ell}\cdot(\bm{\theta}+\bm{n}_{\ell}\bm{\alpha})|\geq c\delta^{R},c>0.$
Step 2. Conditions on $(\bm{\beta},\bm{q})$
Next we impose conditions on $(\bm{\beta},\bm{q})$.
Lemma A.9.
Let $\epsilon\in(0,1)$ and $\bm{B}\subset(0,1]^{d}\times(0,1]^{d}$ be the set of $\bm{\xi}=(\bm{\beta},\bm{q})$ satisfying
$\min_{\ell,\bm{\ell}}((\bm{\ell}\bm{n}_{\ell})\cdot\bm{\beta}+\bm{q}\cdot\bm{\ell})^{2}\geq\epsilon>0,$
$\min_{(\ell,\bm{\ell})\neq(\ell^{\prime},\bm{\ell}^{\prime})}|(\bm{\ell}\bm{n}_{\ell}-\bm{\ell}^{\prime}\bm{n}_{\ell^{\prime}})\cdot\bm{\beta}+\bm{q}\cdot(\bm{\ell}-\bm{\ell}^{\prime})|\geq\epsilon>0,$
$\min_{(\ell,\bm{\ell})\neq(\ell^{\prime},\bm{\ell}^{\prime})}|(\bm{\ell}\bm{n}_{\ell}+\bm{\ell}^{\prime}\bm{n}_{\ell^{\prime}})\cdot\bm{\beta}+\bm{q}\cdot(\bm{\ell}+\bm{\ell}^{\prime})|\geq\epsilon>0.$
Then
${\rm meas}\left((0,1]^{2d}\setminus\bm{B}\right)\leq 3R^{2}\sqrt{\epsilon}.$
Moreover, for $\bm{\xi}\in\bm{B},$ we have
$\ \ \ \prod_{\ell,\bm{\ell}}((\bm{\ell}\bm{n}_{\ell})\cdot\bm{\beta}+\bm{q}\cdot\bm{\ell})^{2}\times|\det W_{1}|$
$=\prod_{\ell,\bm{\ell}}((\bm{\ell}\bm{n}_{\ell})\cdot\bm{\beta}+\bm{q}\cdot\bm{\ell})^{2}$
×$\ \ \times\prod_{(\ell,\bm{\ell})\neq(\ell^{\prime},\bm{\ell}^{\prime})}|((\bm{\ell}\bm{n}_{\ell})\cdot\bm{\beta}+\bm{q}\cdot\bm{\ell})^{2}-((\bm{\ell}^{\prime}\bm{n}_{\ell^{\prime}})\cdot\bm{\beta}+\bm{q}\cdot\bm{\ell}^{\prime})^{2}|$
$\geq\epsilon^{R^{3}}.$
Proof.
The proof is based on the Fubini’s theorem, (A.7) and the assumption on $\bm{n}_{\ell}$ ($1\leq\ell\leq D$). Actually, in all cases we have
$(\bm{\ell}\bm{n}_{\ell},\bm{\ell})\neq 0,$
$(\bm{\ell}\bm{n}_{\ell}-\bm{\ell}^{\prime}\bm{n}_{\ell^{\prime}},\bm{\ell}-\bm{\ell}^{\prime})\neq 0,$
$(\bm{\ell}\bm{n}_{\ell}+\bm{\ell}^{\prime}\bm{n}_{\ell^{\prime}},\bm{\ell}+\bm{\ell}^{\prime})\neq 0,$
which combined with $\ell\leq D$ and $\bm{\ell}\in\Gamma_{K}$ leads to the excision of $(\bm{\beta},\bm{q})$ in a set of measure $O(\sqrt{\epsilon})$. ∎
Step 3. Usage of the Lemma A.5
Now combing Lemma A.8 and Lemma A.9 (we set $\epsilon=10^{-2}R^{-4}$) leads to
Lemma A.10.
Let $100<\delta<1/$ 10 and define $N=[\delta^{-2}]+1$. Then there are a collection of cubes $\Lambda_{1/N}(\bm{x}_{i})\subset[0,1]^{2d}$ ($1\leq i\leq N_{1}$ for some $N_{1}>1$) with disjoint exteriors, and a set $\bm{B}\subset(0,1]^{2d}$ so that
${\rm meas}\left([0,1]^{2d}\setminus\cup_{1\leq i\leq N_{1}}\Lambda_{1/N}(\bm{x}_{i})\right)$ $\leq C\delta,$
${\rm meas}((0,1]^{2d}\setminus\bm{B})$ $<1/3.$
Moreover, if $(\bm{\alpha},\bm{\theta})\in\cup_{1\leq i\leq N_{1}}\Lambda_{1/N}(\bm{x}_{i})$ and $\bm{\xi}\in\bm{B},$ then
$|\det W|\geq c\delta^{R},\ c>0.$
We are ready to conclude the proof of Theorem A.1. Fix $\bm{\xi}\in\bm{B}$ and a $1/N\sim\delta^{2}$-cube $\Lambda_{i}=\Lambda_{1/N}(\bm{x}_{i})\subset[0,1]^{2d}$ in Lemma A.10. We will deal with $(\bm{\alpha},\bm{\theta})\in\Lambda_{i}.$ Denote for $\bm{k}\in\mathbb{Z}^{D}\setminus\{0\}$,
$f(\bm{\alpha},\bm{\theta})=\bm{k}\cdot\bm{\omega}=(\bm{k}\otimes\bm{p})\cdot\bm{V}.$
Direct computations show that for $\bm{\xi}=(\bm{\beta},\bm{q}),$
$\ \ \ \sup_{(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}}\sup_{1\leq j\leq R}|\partial^{2j}_{\bm{\xi}}\bm{V}|_{1}$
$=\sup_{(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}}\sup_{1\leq j\leq R}\sum_{1\leq\ell\leq D,\bm{\ell}\in\Gamma_{K}}|(2\pi(\bm{\ell}\bm{n}_{\ell})\cdot\bm{\beta}+2\pi\bm{q}\cdot\bm{\ell})^{2j}\cos 2\pi\bm{\ell}\cdot(\bm{\theta}+\bm{n}_{\ell}\bm{\alpha})|$
$\leq CR(\sup_{\ell}|\bm{n}_{\ell}|+dK)^{2R}:=M.$
We have by Lemma A.10 and Lemma A.5 that
$\inf_{(\bm{\alpha},\bm{\theta})\in\Lambda_{i}}\max_{1\leq j\leq R}|d^{2j}_{\bm{\xi}}f(\bm{\alpha},\bm{\theta})|\geq c\delta^{R}|\bm{k}|\cdot|\bm{p}|,$ (A.23)
where $c>0.$
It suffices to apply Lemma A.4 by letting $k=2R$. The transversality condition has been verified by (A.23). Obviously, for $f(\bm{\alpha},\bm{\theta})=\bm{k}\cdot\bm{\omega}$, we have
$\|f\|_{2R+1}=\sup_{|\bm{\gamma}|\leq 2R+1,(\bm{\alpha},\bm{\theta})\in\Lambda_{i}}|\partial^{\bm{\gamma}}_{\bm{\xi}}f(\bm{\alpha},\bm{\theta})|\leq C|\bm{k}|.$
So using Lemma A.4 by setting $\varepsilon=\frac{\gamma}{|\bm{k}|^{\tau}}$ leads to
$\ \ \ {\rm meas}\left(\left\{(\bm{\alpha},\bm{\theta})\in\Lambda_{i}:\ |\bm{k}\cdot\bm{\omega}|\leq\frac{\gamma}{|\bm{k}|^{\tau}}\right\}\right)$
$\leq C(R,\max_{\ell}|\bm{n}_{\ell}|,d,|\bm{p}|)\delta^{-R(d+\frac{1}{2R})}\gamma^{\frac{1}{2R}}|\bm{k}|^{-\frac{\tau}{2R}}$
[…]
Together with Lemma A.10 this yields,
$\ \ \ {\rm meas}([0,1]^{2d}\setminus{\rm DC}_{\rm nls}(\gamma;\tau))$
$\leq C\delta+C\delta^{-2d}\sum_{\bm{k}\in\mathbb{Z}^{D}\setminus\{0\}}\frac{\delta^{3d}}{|\bm{k}|^{2D}}$
$\leq C\delta.$
This proves Theorem A.1. ∎
A.4. Proof of Theorem A.2
In this section, we aim to prove Theorem A.2.
Proof of Theorem A.2.
Recall that $\bm{\omega}=\bm{\omega}_{\rm NLW}=(\omega_{\ell})_{1\leq\ell\leq D}$ with
$\omega_{\ell}$ $=\sqrt{P(\bm{\theta}+\bm{n}_{\ell}\bm{\alpha})+m}\ (1\leq\ell\leq D),$
$m$ $\in[\|P\|_{\infty}+1,\|P\|_{\infty}+2].$
The ideas of the proof are as follows. First, we give a lower bound on $|\omega_{\ell}-\omega_{\ell^{\prime}}|$ after an excision of ($\bm{\alpha},\bm{\theta}$) as done in the NLS case. Together with the Wronskian determinant type estimates (after taking derivatives on $m$) this then leads to the transversality estimates. The remaining issue is just to apply Lemma A.4 to remove $m.$
The proof can be divided into two steps:
Step 1. The gap estimates
We aim to estimate for $1\leq\ell<\ell^{\prime}\leq D,$
$\omega_{\ell}^{2}-\omega_{\ell^{\prime}}^{2}$ $=(\bm{e}_{\ell}-\bm{e}_{\ell^{\prime}})\cdot\bm{\omega}_{\rm NLS}.$
So recalling the proof of Theorem A.1, we have
Lemma A.11.
For $0<\delta\ll 1$ and $0<\gamma\leq\delta^{10dR^{2}}$, we have
${\rm meas}(G_{\gamma})\leq C\delta,$
where $G_{\gamma}$ is defined by (A.8).
Step 2. Restrictions on $m$
In the following we always assume $(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}\setminus G_{\gamma}$ so that
$\min_{1\leq\ell<\ell^{\prime}\leq D}|\omega_{\ell}^{2}-\omega_{\ell^{\prime}}^{2}|\geq\gamma>0.$ (A.24)
Then we can repeat the argument of [SW23]. Assume $\bm{\omega}(m)=\bm{\omega}_{\rm NLW}(m)$ is the row vector. Consider the $D\times D$ matrix $W(m)$ with the $\ell$-th row given by
${W}_{\ell}(m)=\left(\frac{d^{\ell}}{dm^{\ell}}\omega_{\ell}(m)\right)_{\ell=1}^{D}.$
We want to estimate the Wronskian $\det{W}$. A key observation is that
$\frac{d^{l}}{dm^{l}}\omega_{s}(m)$ $=\frac{1}{2}(\frac{1}{2}-1)\cdots(\frac{1}{2}-l+1)\omega_{s}^{-(2l-1)}(m)$
$:=\lambda_{l}\omega_{s}^{-(2l-1)}(m).$
Then
$|\det{W}(m)|=\prod_{s=1}^{D}|\lambda_{s}\omega_{s}^{-1}(m)|^{D}\cdot|\det{W}^{\prime}(m)|,$
where ${W}^{\prime}(m)$ is a Vandermonde matrix with the $l$-th ($1\leq l\leq D$) row given by
${W}_{l}^{\prime}(m)=\left(\omega_{s}^{-2(l-1)}(m)\right)_{s=1}^{D}.$
So we obtain
$|\det{W}(m)|=\prod_{s=1}^{D}|\lambda_{s}\omega_{s}^{-1}(m)|^{D}\prod_{1\leq l_{1}<l_{2}\leq D}|\omega_{l_{1}}^{-2}(m)-\omega_{l_{2}}^{-2}(m)|.$ (A.25)
Recall that $|\omega_{s}(m)|\sim\sqrt{\|P\|_{\infty}}$. Thus by (A.24), we obtain
$|\omega_{l_{1}}^{-2}(m)-\omega_{l_{2}}^{-2}(m)|$ $=|\omega_{l_{1}}^{-2}(m)\omega_{l_{2}}^{-2}(m)|\cdot|\omega_{l_{1}}^{2}(m)-\omega_{l_{2}}^{2}(m)|$
$\geq c(\|P\|_{\infty})\gamma,$
which combined with (A.25) implies
$|\det{W}(m)|\geq c^{\prime}\gamma^{D(D-1)},$ (A.26)
where $c^{\prime}=c^{\prime}(\|P\|_{\infty},D)>0$ depends only on $\|P\|_{\infty},D$.
Now, we fix $\bm{k}\in\mathbb{Z}^{D}\setminus\{0\}$ and $m\in[\|P\|_{\infty}+1,\|P\|_{\infty}+2]$. Then $\omega_{s}(m)\sim\sqrt{\|P\|_{\infty}}$. Observe first that for $1\leq l\leq D$,
$\sup_{1\leq l\leq D}|{W}_{l}(m)|_{1}$ $\leq\sup_{1\leq l\leq D}\sum_{s=1}^{D}|\lambda_{l}\omega_{s}^{-l}|\leq C(\|P\|_{\infty},D).$
Applying Lemma A.5 with $r=D,\bm{v}^{(l)}={V}_{l}(m)$, $M=C,\bm{w}=\bm{k}$ yields
$\inf_{m\in[\|P\|_{\infty}+1,\|P\|_{\infty}+2]}\sup_{1\leq l\leq D}\left|\frac{d^{l}}{dm^{l}}\bm{k}\cdot\bm{\omega}\right|\geq c|\bm{k}|_{2}\gamma^{D(D-1)},$ (A.27)
where $c=c(\|P\|_{\infty},D)>0$.
Finally, it suffices to apply Lemma A.4 by letting $k=D$. The transversality condition has been verified by (A.27). Obviously, for $f(m)=\bm{k}\cdot\bm{\omega}$, we have
$\|f\|_{D+1}=\sup_{1\leq l\leq D+1,\|P\|_{\infty}+1\leq m\leq\|P\|_{\infty}+2}|\frac{d^{l}}{dm^{l}}f(m)|\leq C|\bm{k}|.$
So using Lemma A.4 by setting $\varepsilon=\frac{\gamma_{1}}{|\bm{k}|^{\tau}}$ leads to
$\ \ \ {\rm meas}\left(\left\{m\in[\|P\|_{\infty}+1,\|P\|_{\infty}+2]:\ |\bm{k}\cdot\bm{\omega}|\leq\frac{\gamma_{1}}{|\bm{k}|^{\tau}}\right\}\right)$
$\leq C(R,\max_{\ell}|\bm{n}_{\ell}|,d,|\bm{p}|,\|P\|_{\infty})\gamma^{-(D^{2}-1)}\gamma_{1}^{\frac{1}{D}}|\bm{k}|^{-\frac{\tau}{D}}$
[…]
This yields
$\ \ \ {\rm meas}([\|P\|_{\infty}+1,\|P\|_{\infty}+2]\setminus{\rm DC}_{\rm nlw}(\gamma_{1};\tau))$
$\leq C\gamma_{1}^{\frac{1}{2D}}\sum_{\bm{k}\in\mathbb{Z}^{D}\setminus\{0\}}\frac{1}{|\bm{k}|^{2D}}$
$\leq C\gamma_{1}^{\frac{1}{2D}}.$
We have completed the proof of Theorem A.2. ∎