### v2

# Anderson localized states for the quasi-periodic nonlinear Schrödinger equation on $\mathbb{Z}^{d}$

Abstract.
We establish large sets of Anderson localized states for the quasi-periodic nonlinear Schrödinger equation on $\mathbb{Z}^{d}$, thus extending Anderson localization from the linear (cf. Bourgain [Geom. Funct. Anal., 17(3):682–706, 2007]) to a nonlinear setting, and the random (cf. Bourgain-Wang [J. Eur. Math. Soc., 10(1):1–45, 2008]) to a deterministic setting. Among the main ingredients are a new Diophantine estimate of quasi-periodic functions in arbitrarily dimensional phase space, and the application of Bourgain’s geometric lemma in [Geom. Funct. Anal., 17(3):682–706, 2007].
Key words and phrases:
Anderson localization, discrete nonlinear Schrödinger equation, quasi-periodic potential, Diophantine estimates, semi-algebraic sets, Lyapunov-Schmidt decomposition, Green’s function estimates, Newton iteration
1. Introduction and Main Result
Let
$\bm{\alpha}$ $=(\alpha_{1},\cdots,\alpha_{d})\in[0,1]^{d},$
$\bm{\theta}$ $=(\theta_{1},\cdots,\theta_{d})\in[0,1]^{d},$
and consider the nonlinear Schrödinger equation with a quasi-periodic potential on $\mathbb{Z}^{d}$:
$iu_{t}+(\varepsilon\Delta+V(\bm{\theta}+\bm{n}\bm{\alpha})\delta_{\bm{n},\bm{n}^{\prime}})u+\delta|u|^{2p}u=0,$ (1.1)
where $\varepsilon$ and $\delta$ are parameters in $[0,1]$, $\bm{n}\in\mathbb{Z}^{d}$, $p\in\mathbb{N}$ and $\Delta(\bm{n},\bm{n}^{\prime})=\delta_{|\bm{n}-\bm{n}^{\prime}|_{1},1}$ denotes the adjacency Laplacian with $|\bm{n}|_{1}:=\sum\limits_{\ell=1}^{d}|n_{\ell}|$ and
$\bm{n}\bm{\alpha}=(n_{1}\alpha_{1},\cdots,n_{d}\alpha_{d})\in\mathbb{R}^{d};$
the potential $V$ is a trigonometric polynomial:
$V(\bm{\theta})=\sum_{\bm{\ell}\in\Gamma_{K}}v_{\bm{\ell}}\cos 2\pi(\bm{\ell}\cdot\bm{\theta}),\ \bm{\ell}\cdot\bm{\theta}=\sum_{s=1}^{d}\ell_{s}\theta_{s},$
with $\Gamma_{K}\subset[-K,K]^{d}\setminus\{\bm{0}\},\ K\geq 1$ and
$\bm{v}=(v_{\bm{\ell}})_{\bm{\ell}\in\Gamma_{K}}\in\mathbb{R}^{\#\Gamma_{K}}\setminus\{\bm{0}\},$
where $#(\cdot)$ denotes the cardinality of a set. We further assume that the set $\Gamma_{K}$ is maximal so that:
$\text{if }\bm{\ell}\in\Gamma_{K},{\rm then},\ell_{s}\neq 0\ {\rm for}\ \forall 1\leq s\leq d;$ (1.2)
$\text{and if }\ \bm{\ell},\bm{\ell}^{\prime}\in\Gamma_{K},\ {\rm then}\ \bm{\ell}+\bm{\ell}^{\prime}\neq\bm{0}.$
We assume that $V$ is non-degenerate: For any fixed $1\leq\ell\leq d$ and $(\theta_{s})_{s\neq\ell}$, $f(\theta_{\ell}):=V((\theta_{s})_{s\neq\ell},\theta_{\ell})$ is not a constant function in $\theta_{\ell}.$ Note that the first condition in (1.2) then says that, in fact, each term in the polynomial is non-degenerate.
When $\delta=0$, it is known from the breakthrough paper [Bou07] (cf. also the recent refinement [JLS20]) that the linear operator
$\mathcal{H}=\varepsilon\Delta+V(\bm{n}\bm{\alpha}+\bm{\theta})\delta_{\bm{n},\bm{n}^{\prime}}\ {\rm on}\ \mathbb{Z}^{d},$ (1.3)
exhibits Anderson localization, namely pure point spectrum with exponentially decaying eigenfunctions, for small $\varepsilon$, on a large set in $(\bm{\alpha},\bm{\theta})$. It follows that when $0<\varepsilon\ll 1$, the linear Schrödinger equation
$iu_{t}+(\varepsilon\Delta+V(\bm{n}\bm{\alpha}+\bm{\theta})\delta_{\bm{n},\bm{n}^{\prime}})u=0$ (1.4)
has only Anderson localized states, i.e., roughly speaking, wave packets localized about the origin remain localized for all time.
The present paper addresses the persistence question when $\delta\neq 0$, namely the existence of Anderson localized type solutions for the nonlinear Schrödinger equation (1.1) when $0<\varepsilon,\delta\ll 1$, under appropriate conditions on $\bm{\alpha}$ and $\bm{\theta}$. Since both $\varepsilon$ and $\delta$ are small, we start from the (decoupled) equation
$iu_{t}+V(\bm{n}\bm{\alpha}+\bm{\theta})\delta_{\bm{n},\bm{n}^{\prime}}u=0.$ (1.5)
It has solutions of the form
$u^{(0)}(t,\bm{n})=\sum_{\bm{n}\in\mathbb{Z}^{d}}a_{\bm{n}}e^{i\mu_{\bm{n}}t},$ (1.6)
where
$\mu_{\bm{n}}={V(\bm{n}\bm{\alpha}+\bm{\theta})}$
and $a_{\bm{n}}$ decays rapidly. The solutions (1.6), in general, have infinite number of frequencies and are almost-periodic in time.
We study the persistence of the above type of solutions with finite (but arbitrary) number of frequencies in time, i.e., the quasi-periodic in time solutions. Denoting the frequency by $\bm{\omega}$, $\bm{\omega}=(\omega_{1},\omega_{2},\cdots,\omega_{b})$, as an Ansatz, we seek solutions to (1.1) in the form of a convergent series:
$u(t,\bm{n})=\sum_{(\bm{k},\bm{n})\in\mathbb{Z}^{b}\times\mathbb{Z}^{d}}\hat{u}(\bm{k},\bm{n})e^{i\bm{k}\cdot\bm{\omega}t}.$ (1.7)
Note that solutions of the above form are consistent with the nonlinearity in (1.1).
Denote by ${\rm meas}(\cdot)$ the Lebesgue measure of a set. Let $|\cdot|$ denote the supremum norm. We have
Theorem 1.1.
Fix (any) distinct $\bm{n}_{1},\cdots,\bm{n}_{b}\in\mathbb{Z}^{d}$ and let
$u^{(0)}(t,\bm{n})=\sum_{\ell=1}^{b}a_{\ell}e^{i\omega_{\ell}^{(0)}t}\delta_{\bm{n}_{\ell},\bm{n}}$ (1.8)
be a solution to (1.5), where $\bm{a}=(a_{\ell})_{\ell=1}^{b}\in[1,2]^{b}$ and
$\bm{\omega}^{(0)}=(\omega_{\ell}^{(0)})_{\ell=1}^{b}=\left({V(\bm{n}_{\ell}\bm{\alpha}+\bm{\theta})}\right)_{\ell=1}^{b}.$
Then for $0<\varepsilon\leq\delta\leq\log^{-1}\frac{1}{\varepsilon}\leq\delta_{0}(b,d,K,V,\max\limits_{1\leq\ell\leq b}|\bm{n}_{\ell}|)\ll 1$, there is a set $\mathcal{W}\subset[0,1]^{2d}$ in $(\bm{\alpha},\bm{\theta})$, satisfying ${\rm meas}([0,1]^{2d}\setminus\mathcal{W})\leq\log^{-c}\frac{1}{\varepsilon+\delta}$ $(c=c(b,d)>0)$ such that, for any $(\bm{\alpha},\bm{\theta})\in\mathcal{W}$, there exist a diffeomorphism $\bm{\omega}=\bm{\omega}(\bm{a})$ on $0,1]^{b}$ and a set in $\bm{a}$, $\mathcal{R}\subset[1,2]^{b}$ of ${\rm meas}([1,2]^{b}\setminus\mathcal{R})\leq(\varepsilon+\delta)^{c}$ so that the following holds true: If $\bm{a}\in\mathcal{R}$ and $\bm{\omega}=\bm{\omega}(\bm{a})$, then $|\bm{\omega}-\bm{\omega}^{(0)}|\lesssim\delta$ and
$u(t,\bm{n})=\sum_{\bm{k}\in\mathbb{Z}^{b}}\hat{u}(\bm{k},\bm{n})e^{i\bm{k}\cdot\bm{\omega}t}$
is a solution to (1.1). Moreover, one has
$\hat{u}(\bm{e}_{\ell},\bm{n}_{\ell})$ $=a_{\ell},\ 1\leq\ell\leq b,$
$\sum_{(\bm{k},\bm{n})\notin\mathcal{S}_{+}}|\hat{u}(\bm{k},\bm{n})|e^{|\bm{k}|+|\bm{n}|}$ $<\sqrt{\varepsilon+\delta},$
where $\mathcal{S}_{+}=\{(\bm{e}_{\ell},\bm{n}_{\ell})_{\ell=1}^{b}\}$ with $\bm{e}_{\ell}\ (1\leq\ell\leq b)$ the standard basis vectors for $\mathbb{Z}^{b}$ (i.e., $\bm{e}_{\ell}=(\delta_{\ell,j})_{j=1}^{b}$).
Remark 1.2.
Theorem 1.1 readily generalizes to potentials $V$, which are given by finite Fourier series with both cosine and sine terms.
Remark 1.3.
Note that we do not require the linear equation (1.4) to satisfy Anderson localization. However, linear Anderson localization does hold on a large subset of $\mathcal{W}$ (cf. [Bou07, JLS20]).
1.1. Ideas of the proof
The multi-dimensional phase space, namely $\bm{\theta}$ being a vector instead of a scalar as in e.g., [SW23] poses a major challenge. Diophantine properties of the eigenvalues of the linear Schrödinger operator (1.3) plays an essential role in the proof. To $O(\delta)$, this could be replaced by Diophantine properties of the potential $V(\bm{n}\bm{\alpha}+\bm{\theta})$ at different lattice sites $\bm{n}\in\mathbb{Z}^{d}$. We need to show, that the values of the function at these chosen sites are, in some sense, linearly independent. This is, however, a priori, not obvious, since $V$ is a given function on $\mathbb{R}^{d}$. To prove linear independence, we use a generalized Wronskian approach and bound the determinant away from zero by carefully removing some $(\bm{\alpha},\bm{\theta})$. This method is applicable to any trigonometric polynomials, generalizing the approach in [SW23] for the cosine function on $\mathbb{R}$. It is independent of the main body of the proof and could be of independent interest. Such arguments may be closely related to Diophantine approximations on manifold developed by Kleinbock and Margulis [KM98].
The other main ingredient of the proof is Bourgain’s approach to Anderson localization of linear quasi-periodic Schrödinger operators in $d$-dimensions [Bou07], such as the $\mathcal{H}$ in (1.3). This seems to be, so far, the only method available to deal with the $d$-dimensional phase space, i.e., $\bm{\theta}\in[0,1]^{d}$, $d\geq 3$. Among other innovations, it uses semi-algebraic geometry arguments (e.g., Yomdin-Gromov algebraic lemma) to control the resonances. The difficulty of $d$-dimensional phase space manifests as well in our paper, in that for the linear analysis, semi-algebraic geometry already seems indispensable to control the resonances; while in previous works, see [BW08, Wan16, LW24, SW23], this role is filled by Diophantine properties. To control the resonances in time, we also need a weak second Melnikov type estimates. This could be readily obtained, similar to [Bou05a, BW08, LW24, SW23]. In this paper, however, we have to overcome much more serious resonances in space (cf. Figure 1), which seems novel. The proof is then accomplished by doing multi-scale analysis with some new ideas.
The general problem discussed here is the persistency of quasi-periodic solutions of linear or integrable equations after Hamiltonian perturbation. This subject is closely related to the well-known “KAM theory” of invariant tori in smooth dynamical systems. Results along this line were first obtained by Kuksin [Kuk87] and Wayne [Way90], stimulating considerable progress, cf. e.g., [KP96, CY00, BG01, Bou05b, EK10, LY10, LY11, GT11, BBP13, PP15, EGK16, BBHM18, BKM18, CLSY18, Yua21, BHM23] to name but a few. The existence of quasi-periodic solutions can also be obtained using the direct Craig-Wayne-Bourgain method [CW93, Bou94, Bou98, Bou05a] (cf. [BW08, BB13, Wan16, HSSY20, Wan21a, Wan21b, LW24, SW23, KLW24] for more recent results), which turns out to be more robust when dealing with multi-dimensional Hamiltonian PDEs (cf. Chapters 18–20 of [Bou05a]).
Finally, we mention the work [Yua02] in which a KAM scheme was first developed to prove the existence of quasi-periodic solutions for some nonlinear discrete equations without external parameters (i.e., regarding the initial states as parameters). Later in [GYZ14], the authors also applied the KAM method to obtain the existence of quasi-periodic solutions for the nonlinear quasi-periodic Schrödinger equation on $\mathbb{Z}$.
1.2. The nonlinear random Schrödinger equation
When the potential $V$ with $V(\bm{n})=V(\bm{n}\bm{\alpha}+\bm{\theta})$ is replaced by a random potential, for example, with $V(\bm{n})$, a family of independently identically distributed random variables (with the uniform distribution), it is known from [BW08] that the nonlinear random Schrödinger equation
$iu_{t}+\varepsilon\Delta u+Vu+\delta|u|^{2p}u=0\ {\rm on}\ \mathbb{Z}^{d}$
has large sets of Anderson localized states for small $\varepsilon$ and $\delta$. When $d=1$, it is shown further in the important work [LW24] that large sets of Anderson localized states persist for all $\varepsilon\neq 0$ (similar results could be proven for the nonlinear wave equation). Recall that the first proof of the Anderson localization for the linear random Schrödinger equation
$iu_{t}+\varepsilon\Delta u+Vu=0\ {\rm on}\ \mathbb{Z}^{d}$
for small $\varepsilon$ was based on multi-scale analysis type Green’s function estimates developed in [FS83], see also [AM93].
It is generally believed that the linear random Schrödinger equation and the linear quasi-periodic Schrödinger equation should share common localization features in the perturbative regime, despite the quasi-periodic problems being more delicate and the results more difficult to obtain. The use of semi-algebraic geometry and Cartan’s Lemma [BGS02, Bou07, JLS20], for example, certainly renders the quasi-periodic issue more involved than that of the random, and this seems to continue to the nonlinear case, as mentioned earlier. Theorem 1.1 generalizes this circle of ideas to the nonlinear setting by providing a concrete example where the quasi-periodic and the random have indeed similar behavior.
1.3. Structure of the paper
In Section 2, we focus on the linear estimates, in the form of large deviation theorem. The paper concludes by constructing the quasi-periodic in time solutions in Section 3. The proof of the Diophantine type estimates and some important lemmas on resolvent identities are presented in the Appendix.
2. Linear analysis
In this section, we investigate properties of linearized operators on $\mathbb{Z}^{b+d}$. Of particular importance is the large deviation theorem (LDT) for the corresponding Green’s functions.
2.1. The operator on the lattice
For a vector $\bm{x}\in\mathbb{R}^{r}$, we denote by $|\bm{x}|$ (resp. $|\bm{x}|_{2}$) the supremum norm (resp. the Euclidean norm). For an operator (or a matrix), we denote by $\|\cdot\|$ its operator norm.
We define the lattice:
$\mathbb{Z}^{b+d}_{{\rm pm}}=\mathbb{Z}^{b+d}\times\{+,-\}.$
Let $\mathcal{S}_{+}=\{(\bm{e}_{\ell},\bm{n}_{\ell})_{\ell=1}^{b}\}\subset\mathbb{Z}^{b+d}$ and we identify $\mathcal{S}_{+}$ with $\{(\bm{e}_{\ell},\bm{n}_{\ell})_{\ell=1}^{b}\}\times\{+\}\subset\mathbb{Z}^{b+d}_{{\rm pm}}$. Similarly, let $\mathcal{S}_{-}=\{(-\bm{e}_{\ell},\bm{n}_{\ell})_{\ell=1}^{b}\}\subset\mathbb{Z}^{b+d}$, which is then identified with $\{(-\bm{e}_{\ell},\bm{n}_{\ell})_{\ell=1}^{b}\}\times\{-\}\subset\mathbb{Z}^{b+d}_{{\rm pm}}$. We let $\mathcal{S}=\mathcal{S}_{+}\cup\mathcal{S}_{-}$ and
$\mathbb{Z}^{b+d}_{{\rm pm},*}=\mathbb{Z}_{\rm pm}^{b+d}\setminus\mathcal{S}.$
In the following, we study on $\mathbb{Z}^{b+d}_{{\rm pm},*}$ the operator
${H}(\sigma)={D}(\sigma)+\varepsilon(\Delta\oplus\Delta)+\delta S,\ \sigma\in\mathbb{R},$ (2.1)
where
${D}(\sigma)$ $={\rm diag}_{(\bm{k},\bm{n})\in\mathbb{Z}^{b+d}}\left(\begin{array}[]{cc}{-\sigma-\bm{k}\cdot\bm{\omega}+\mu_{\bm{n}}}&{0}\\ {0}&{\sigma+\bm{k}\cdot\bm{\omega}+\mu_{\bm{n}}}\\ \end{array}\right),\ \mu_{\bm{n}}=V(\bm{n}\bm{\alpha}+\bm{\theta})$
and the operator $S$ refers typically to the linearized operator of the nonlinear perturbation. So we may let $S$ satisfy the following properties:
  • First
[…]
    is a bounded self-adjoint operator.
  • For all $\bm{k},\bm{k}^{\prime},\bm{k}^{\prime\prime}\in\mathbb{Z}^{b}$, $\bm{n},\bm{n}^{\prime}\in\mathbb{Z}^{d}$ and $\xi,\xi^{\prime}\in\{+,-\}$, we have the Töplitz property in the $\bm{k}$-variable:
    $S((\bm{k}^{\prime}+\bm{k},\bm{n},\xi);(\bm{k}^{\prime\prime}+\bm{k},\bm{n}^{\prime},\xi^{\prime}))=S((\bm{k}^{\prime},\bm{n},\xi);(\bm{k}^{\prime\prime},\bm{n}^{\prime},\xi^{\prime})).$
  • We have for some $C_{2}>0$ and $\gamma\in(\frac{1}{2},10),$
    $|S((\bm{k},\bm{n},\xi);(\bm{k}^{\prime},\bm{n}^{\prime},\xi^{\prime}))|\leq C_{2}(1+|\bm{k}-\bm{k}^{\prime}|)^{C_{2}}e^{-\gamma|\bm{k}-\bm{k}^{\prime}|-\gamma|\bm{n}|}\delta_{\bm{n},\bm{n}^{\prime}}.$
The frequency $\bm{\omega}$ satisfies
$\bm{\omega}\in\Omega=\bm{\omega}^{(0)}+[-C\delta,C\delta]^{b},$ (2.2)
where $C>0$ and $\bm{\omega}^{(0)}=\bm{\omega}^{(0)}(\bm{\alpha},\bm{\theta})$ is defined in Theorem 1.1.
2.2. The (generalized) elementary regions on the lattice
For some technical reason, we will introduce the definitions of elementary regions and generalized elementary regions on $\mathbb{Z}^{b+d}$ originated from [BGS02, Bou07]. We refer to [Liu22] (cf. Section 2 in [Liu22]) for some important clarifications and improvements on those concepts.
Denote by $\Lambda_{N}(\bm{x})$ ($\bm{x}\in\mathbb{Z}^{r},\ r=b+d,d$),
$\Lambda_{N}(\bm{x})=\{\bm{y}\in\mathbb{Z}^{r}:\ |\bm{y}-\bm{x}|\leq N\},$
the cube of radius $N$ centered at $\bm{x}$. In particular, we write $\Lambda_{N}=\Lambda_{N}(\bm{0})$. For $\Lambda\subset\mathbb{Z}^{r}$, we define its diameter to be ${\rm diam}\ \Lambda=\sup\limits_{\bm{x},\bm{y}\in\Lambda}|\bm{x}-\bm{y}|$. For $\bm{x}\in\mathbb{Z}^{r}$ and $\bm{w}\in\mathbb{R}_{+}^{r}$, let $R_{\bm{w}}(\bm{x})=\{\bm{y}\in\mathbb{Z}^{r}:\ |y_{\ell}-x_{\ell}|\leq w_{l},\,1\leq\ell\leq r\}$ denote the rectangle.
A generalized elementary region is defined to be a set $\Lambda$ of the form
$\Lambda=R_{\bm{w}}(\bm{x})\setminus(R_{\bm{w}}(\bm{x})+\bm{z}),$
where $\bm{z}\in\mathbb{Z}^{r}$ is arbitrary. The size of a generalized elementary region is simply its diameter. The set of all generalized elementary regions of size at most $M$ will be denoted by $\mathcal{E}_{M}$.
An elementary region $Q_{N}$ of size $N$ and centered at $\bm{0}$ is one of the following regions
$Q_{N}=\Lambda_{N}\ {\rm or}\ Q_{N}=\Lambda_{N}\setminus\{\bm{x}\in\mathbb{Z}^{r}:\ x_{\ell}\square_{\ell}0,\ 1\leq\ell\leq r\},$
where $\square_{\ell}\in\{<,>,\emptyset\}$ and at least two $\square_{\ell}$’s are not $\emptyset.$ Denote by $\mathcal{ER}_{\bm{0}}(N)$ the set of all elementary regions of size $N$ and centered at $\bm{0}$. Let
$\mathcal{ER}(N)=\{\bm{x}+Q_{N}:\ \bm{x}\in\mathbb{Z}^{r},\ Q_{N}\in\mathcal{ER}_{\bm{0}}(N)\}.$
Let $\mathcal{E}_{M}^{L}$ denote the set of all $M$-size generalized elementary regions of width at least $L\leq M$, namely, $\Lambda\in\mathcal{E}_{M}^{L}$ iff $\Lambda\in\mathcal{E}_{M}$ and, for any $\bm{x}\in\Lambda,0<L^{\prime}<L$, there is some $\Lambda^{\prime}\in\mathcal{ER}(L^{\prime})$ so that $\bm{x}\in\Lambda^{\prime}\subset\Lambda$, ${\rm dist}(\bm{x},\Lambda\setminus\Lambda^{\prime})\geq L^{\prime}/2$, where ${\rm dist}(\cdot,\cdot)$ is induced by the supremum norm. So we have $\mathcal{ER}(N)\subset\mathcal{E}_{2N}^{N}$.
With a slight abuse of notation, we also use $\mathcal{ER}_{\bm{0}}(N),\mathcal{ER}(N),\mathcal{E}_{N},\mathcal{E}_{N}^{L},\Lambda_{N}(\bm{x}),\ \Lambda_{N}$ to denote $\mathcal{ER}_{\bm{0}}(N)\times\{+,-\},\mathcal{ER}(N)\times\{+,-\},\mathcal{E}_{N}\times\{+,-\},\mathcal{E}_{N}^{L}\times\{+,-\},\Lambda_{N}(\bm{x})\times\{+,-\},\Lambda_{N}\times\{+,-\}$. Similarly, for any $\Lambda\subset\mathbb{Z}^{r}$, denote by $R_{\Lambda}$ the restriction to $\Lambda\times\{+,-\}$.
2.3. Green’s functions and LDE
We now define Green’s functions on subsets of $\mathbb{Z}_{{\rm pm},*}^{b+d}$. Let $\Lambda\subset\mathbb{Z}_{{\rm pm},*}^{b+d}$ be a non-empty set, $R_{\Lambda}$ the restriction operator to $\Lambda$, and define the $(\#\Lambda)\times(\#\Lambda)$ matrix:
$(R_{\Lambda}{H}(\sigma)R_{\Lambda})((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))={H}(\sigma)((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime})),\ (\bm{x},\xi),(\bm{x}^{\prime},\xi^{\prime})\in\Lambda,$ (2.3)
where $H(\sigma)$ is defined by (2.1). Define the Green’s function to be (if it exists)
$G_{\Lambda}(\sigma)=(R_{\Lambda}{H}(\sigma)R_{\Lambda})^{-1}.$
Before addressing the LDT for Green’s functions, let us first give the definition of large deviation estimates (LDE):
Definition 2.1 (LDE).
Let $\rho>0$, $0<\gamma^{\prime}<\gamma$ and $M\in\mathbb{N}$. We say ${H}(\sigma)$ defined by (2.1) satisfies the $(\rho,\gamma^{\prime},M)$-LDE if there exists a set $\Sigma_{M}\subset\mathbb{R}$ with
${\rm meas}(\Sigma_{M})\leq e^{-M^{\rho}},$
so that for $\sigma\notin\Sigma_{M}$ the following estimates hold: If $\Lambda\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(M)$ satisfies $|\bm{n}|\leq 10M$, then
$\|G_{\Lambda}(\sigma)\|$ $\leq e^{M^{\frac{3}{4}}},$
$|G_{\Lambda}(\sigma)((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))|$ $\leq e^{-\gamma^{\prime}|\bm{x}-\bm{x}^{\prime}|}\ {\rm for}\ |\bm{x}-\bm{x}^{\prime}|\geq M^{\frac{8}{9}}\ {\rm and}\ \xi,\xi^{\prime}\in\{\pm\},$
where $\bm{x}=(\bm{k}^{\prime},\bm{n}^{\prime}),\bm{x}^{\prime}=(\bm{k}^{\prime\prime},\bm{n}^{\prime\prime})$.
We will show in this section that the $(\rho,\frac{\gamma}{2},N)$-LDE hold for some small
$\rho=\rho(b,d)>0$
and all $N\gg 1$, under certain non-resonant conditions on $\bm{\alpha},\bm{\theta}$ and $\bm{\omega}$. This leads to the LDT, Theorem 2.11 below. Since we view $\bm{\omega}$ as an $O(\delta)$ perturbation of $\bm{\omega}^{(0)}$, the properties of $\bm{\omega}^{(0)}$ become essential in the proof of LDT. So, in the following, we first establish non-resonant properties (called Diophantine estimates) of $\bm{\omega}^{(0)}$. Then the proof of LDT will follow from a multi-scale analysis scheme in the spirit of [BGS02, Bou05a, Bou07].
2.4. Diophantine estimates
Recall that
$\bm{\omega}^{(0)}$ $=\bm{\omega}^{(0)}(\bm{\alpha},\bm{\theta})=(V(\bm{n}_{\ell}\bm{\alpha}+\bm{\theta}))_{\ell=1}^{b},$
$\mu_{\bm{n}}$ $=\mu_{\bm{n}}(\bm{\alpha},\bm{\theta})=V(\bm{n}\bm{\alpha}+\bm{\theta}),\ \bm{n}\in\mathbb{Z}^{d}.$
The main purpose of this section is to obtain lower bounds on
$|\bm{k}\cdot\bm{\omega}^{(0)}|\ {\rm for}\ \bm{k}\in\mathbb{Z}^{b}\setminus\{\bm{0}\},$
$|\pm\bm{k}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}}|\ {\rm for}\ (\bm{k},\bm{n})\in\mathbb{Z}^{b+d}\backslash{\mathcal{S}},$
$|\bm{k}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}}-\mu_{\bm{n}^{\prime}}|\ {\rm for}\ \bm{n}\neq\bm{n}^{\prime}\in\mathbb{Z}^{d},$
under certain restrictions on $(\bm{\alpha},\bm{\theta}).$ As mentioned in the Introduction, due to the $d$-dimensional phase space, such Diophantine-type estimates are highly non-trivial. We use a generalized Wronskian approach and work out the details in the Appendix (cf. Theorem A.2 and Corollary A.4). These estimates can be read independently, and are key to the existence of quasi-periodic solutions to the nonlinear equation.
Applying Corollary A.4 then leads to
Theorem 2.2.
There exist some $0<c_{1}=c_{1}(b,d,K,V)<\frac{1}{100b},C_{1}=C_{1}(b,d,K,V)>1$ such that, if $0<\varepsilon+\delta\leq\delta_{0}(b,d,K,V,\max\limits_{1\leq\ell\leq b}|\bm{n}_{\ell}|)\ll 1$, then there is some $\mathcal{M}\subset[0,1]^{2d}$ satisfying
${\rm meas}([0,1]^{2d}\setminus\mathcal{M})\leq\log^{-1}\frac{1}{\varepsilon+\delta},$ (2.4)
so that the following properties hold true for $(\bm{\alpha},\bm{\theta})\in\mathcal{M}$ and $L_{\varepsilon,\delta}:=100(\varepsilon+\delta)^{-c_{1}}$.
  • (1)
    For all $\log\frac{1}{\varepsilon+\delta}\leq L\leq L_{\varepsilon,\delta}$ and all $(\bm{k},\bm{n})\in\Lambda_{L}\setminus\mathcal{S},$ we have
    $\min_{\xi=\pm 1}|\xi\bm{k}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}}|>L^{-C_{1}}.$ (2.5)
  • (2)
    We have for any $\xi=\pm 1,$
    $\sup_{\sigma\in\mathbb{R}}\#\left\{(\bm{k},\bm{n})\in\Lambda_{L_{\varepsilon,\delta}}:\ |\xi(\sigma+\bm{k}\cdot\bm{\omega}^{(0)})+\mu_{\bm{n}}|<\frac{(\varepsilon+\delta)^{\frac{1}{8b}}}{4}\right\}\leq b.$ (2.6)
Remark 2.3.
The estimate in (1) plays an essential role in the initial iteration steps for the nonlinear analysis. The conclusion (2) of this theorem will be only used in the proof of intermediate scales LDE.
Proof.
The proof is similar to that in [SW23]. First, the measure estimate (2.4) is a consequence of Corollary A.4. Next, the inequality (2.5) in (1) follows directly from (iii) of Corollary A.4.
So, it suffices to establish (2.6) of (2). We prove it by the contradiction. Without loss of generality, we consider the case $\xi=+1$. Assume that there are $b+1$ distinct $\{(\bm{k}_{\ell},\bm{n}_{\ell}^{\prime})\}_{1\leq\ell\leq b+1}$ satisfying
$|\sigma+\bm{k}_{\ell}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}_{\ell}^{\prime}}|<\frac{(\varepsilon+\delta)^{\frac{1}{8b}}}{4}.$ (2.7)
We claim that all $\bm{k}_{\ell}$ ($1\leq\ell\leq b+1$) are distinct. Actually, if there are $\bm{k}_{\ell_{1}}=\bm{k}_{\ell_{2}}$ for some $1\leq\ell_{1}\neq\ell_{2}\leq b+1$, then it must be that $\bm{n}_{\ell_{1}}^{\prime}\neq\bm{n}_{\ell_{2}}^{\prime}$. As a result, we obtain using (2.7) that
$|\mu_{\bm{n}_{\ell_{1}}^{\prime}}-\mu_{\bm{n}_{\ell_{2}}^{\prime}}|$ $\leq|\sigma+\bm{k}_{\ell_{1}}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}_{\ell_{1}}^{\prime}}|$
$\ \ \ +|\sigma+\bm{k}_{\ell_{2}}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}_{\ell_{2}}^{\prime}}|$
$\leq\frac{(\varepsilon+\delta)^{\frac{1}{8b}}}{2},$
which contradicts (i) of Corollary A.4. Next, we claim that $\{\bm{n}_{\ell}^{\prime}\}_{1\leq\ell\leq b+1}\subset\{\bm{n}_{\ell}\}_{1\leq\ell\leq b}.$ In fact, if $\bm{n}_{\ell_{1}}^{\prime}\notin\{\bm{n}_{\ell}\}_{1\leq\ell\leq b}$ for some $1\leq\ell_{1}\leq b+1$, we can assume that there is $\ell_{2}\neq\ell_{1}$ so that $\bm{n}_{\ell_{2}}^{\prime}\neq\bm{n}_{\ell_{1}}^{\prime}$. Otherwise, we may have
$|(\bm{k}_{\ell_{1}}-\bm{k}_{\ell_{2}})\cdot\bm{\omega}^{(0)}|\leq\frac{(\varepsilon+\delta)^{\frac{1}{8b}}}{2},$
which contradicts (ii) of Corollary A.4, since $\bm{k}_{\ell_{1}}\neq\bm{k}_{\ell_{2}}$ as claimed above. So we obtain for some $\ell_{2}\neq\ell_{1}$ and $1\leq\ell_{2}\leq b+1$ with $\bm{n}_{\ell_{2}}^{\prime}\neq\bm{n}_{\ell_{1}}^{\prime}$ that
$|(\bm{k}_{\ell_{1}}-\bm{k}_{\ell_{2}})\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}_{\ell_{1}}^{\prime}}-\mu_{\bm{n}_{\ell_{2}}^{\prime}}|\leq\frac{(\varepsilon+\delta)^{\frac{1}{8b}}}{2},$
which contradicts (iv) of Corollary A.4. So it suffices to assume $\{\bm{n}_{\ell}^{\prime}\}_{1\leq\ell\leq b+1}\subset\{\bm{n}_{\ell}\}_{1\leq\ell\leq b}.$ However, in this case, we have by the pigeonhole principle that there are $1\leq\ell_{1}\neq\ell_{2}\leq b+1$ so that $\bm{n}_{\ell_{1}}^{\prime}=\bm{n}_{\ell_{2}}^{\prime}$. Then we get
$|(\bm{k}_{\ell_{1}}-\bm{k}_{\ell_{2}})\cdot\bm{\omega}^{(0)}|\leq\frac{(\varepsilon+\delta)^{\frac{1}{8b}}}{2},$
which contradicts (ii) of Corollary A.4 as shown above. This proves (2.6). ∎
2.5. Bourgain’s geometric lemma
To proceed with the analysis, we will also need Anderson localization-type properties for the linear quasi-periodic Schrödinger operator $\mathcal{H}$ in (1.3), namely,
$\mathcal{H}=\mathcal{H}(\bm{\alpha},\bm{\theta})=\varepsilon\Delta+V(\bm{\theta}+\bm{n}\bm{\alpha})\delta_{\bm{n},\bm{n}^{\prime}}.$
This was proven by Bourgain in the remarkable work [Bou07], where he established Anderson localization for general quasi-periodic Schrödinger operators on $\mathbb{Z}^{d}$ with non-degenerate analytic potentials on $\mathbb{T}^{d}$ (identified with $0,1]^{d}$). This result was later significantly extended by Jitomirskaya-Liu-Shi [JLS20] (cf. also [Shi22, Liu22]) to potentials on $\mathbb{T}^{b}$ for $b\geq d.$
In the present paper, due to the $d$-dimensional phase space, we need to apply Bourgain’s arguments. Using short-range property of $S$ in the $\bm{n}$-variable, Bourgain’s geometric lemma is made available to handle LDE for Green’s functions at large scales. This is a main new aspect not present in [SW23]. The general scheme can be contrast with that in Chapter 19 of [Bou05a], where Bourgain made a great breakthrough and first proved the existence of quasi-periodic solutions for the nonlinear Schrödinger equation on arbitrarily dimensional torus (a KAM theorem was later established by Eliasson-Kuksin in an important paper [EK10]). Among others, Bourgain [Bou05a] employed heavily the separation property of eigenvalues of the Laplace operator on the higher dimensional torus, while the separation property is missing in the present setting (indeed, the spectrum of $\mathcal{H}$ may contain an interval).
We start with introducing the definition of semi-algebraic sets.
Definition 2.4.
A set $X\subset\mathbb{R}^{r}$ is called semi-algebraic if it is a finite union of sets defined by a finite number of polynomial equalities and inequalities. More precisely, let $\{P_{1},\cdots,P_{k}\}\subset\mathbb{R}[x_{1},\cdots,x_{r}]$ be a family of real polynomials whose degrees are bounded by $p$. A (closed) semi-algebraic set X{X} is given by an expression
${X}=\bigcup\limits_{s}\bigcap\limits_{\ell\in\mathcal{K}_{s}}\left\{\bm{x}\in\mathbb{R}^{r}:\ P_{\ell}(\bm{x})\varsigma_{s\ell}0\right\},$ (2.8)
where $\mathcal{K}_{s}\subset\{1,\cdots,k\}$ and $\varsigma_{s\ell}\in\{\geq,\leq,=\}$. Then we say that X{X} has degree at most $kp$. In fact, the degree of X{X} which is denoted by $\deg{X}$, means the smallest $kp$ over all representations as in (2.8).
Denote by $\mathcal{ER}_{\mathbb{Z}^{d}}(M)$ (resp. $\mathcal{ER}_{\mathbb{Z}^{d},\bm{0}}(M)$) the set of all $M$-size elementary regions on $\mathbb{Z}^{d}$ (resp. centered at $\bm{0}$). We have
Lemma 2.5 (cf. the Claim (page 694) in [Bou07] and Theorem 2.7 in [JLS20]).
There are small constants $0<\kappa_{1}<\kappa_{2}<1$ depending only on $d$ so that the following holds true. For $0<\varepsilon\leq\varepsilon_{0}(V,d)\ll 1$ and $N\geq N_{1}:=10^{-\frac{32}{\kappa_{1}^{2}}}\log^{\frac{4}{\kappa_{1}}}(\log\frac{1}{\varepsilon})$, there is a semi-algebraic set $\mathcal{A}_{N}\subset[0,1]^{d}$ satisfying
$\deg\mathcal{A}_{N}\leq N^{4d},\ {\rm meas}(\bigcap_{N\geq N_{1}}\mathcal{A}_{N})=1-O(\log^{-c}\frac{1}{\varepsilon}),\ c=c(d)>0,$
so that the following properties hold true for $\bm{\alpha}\in\mathcal{A}_{N}.$
  • (1)
    For all $E\in\mathbb{R}$ and all $\bm{\theta}\in[0,1]^{d}$, there is $L\in[N^{\kappa_{1}},N^{\kappa_{2}}]$ so that, for all $Q\in\mathcal{ER}_{\mathbb{Z}^{d}}(L_{1})$ satisfying $L_{1}\sim(\log N)^{\frac{4}{\kappa_{1}}}$ and
    $Q\subset[-L,L]^{d}\setminus[-L^{\frac{1}{10d}},L^{\frac{1}{10d}}]^{d},$
    one has
    $\|\mathcal{H}^{-1}_{Q}(E;\bm{\theta})\|$ $\leq e^{\sqrt{L_{1}}},$
    $|\mathcal{H}^{-1}_{Q}(E;\bm{\theta})(\bm{n};\bm{n}^{\prime})|$ $\leq e^{-\frac{1}{2}|\log\varepsilon|\cdot|\bm{n}-\bm{n}^{\prime}|}\ {\rm for}\ |\bm{n}-\bm{n}^{\prime}|\geq L_{1}^{\frac{8}{9}},$
    where
    $\mathcal{H}_{Q}(E;\bm{\theta})$ $=R_{Q}\left(\varepsilon\Delta+V(\bm{\theta}+\bm{n}\bm{\alpha})\delta_{\bm{n},\bm{n}^{\prime}}\right)R_{Q}-E$
    $:=\mathcal{L}_{Q}(\bm{\theta})-E.$
  • (2)
    There is some $\Theta_{N}=\Theta_{N}(\bm{\alpha})\subset[0,1]^{d}$ satisfying
    ${\rm meas}([0,1]^{d}\setminus\Theta_{N})\leq e^{-N^{\frac{\kappa_{1}}{4}}}$
    so that, if $\bm{\theta}\in\Theta_{N}$, then one has for all $\Lambda\in\mathcal{ER}_{\mathbb{Z}^{d},\bm{0}}(N),$
    $\|\mathcal{H}^{-1}_{\Lambda}(\bm{\alpha},\bm{\theta})\|$ $\leq e^{\sqrt{N}},$
    $|\mathcal{H}^{-1}_{\Lambda}(\bm{\alpha},\bm{\theta})(\bm{n};\bm{n}^{\prime})|$ $\leq e^{-\frac{1}{2}|\log\varepsilon|\cdot|\bm{n}-\bm{n}^{\prime}|}\ {\rm for}\ |\bm{n}-\bm{n}^{\prime}|\geq{N^{\frac{8}{9}}}.$
    Denote
    $\mathcal{A}=\bigcap_{N\geq N_{1}}\mathcal{A}_{N},\ \tilde{\Theta}_{N}(\bm{\alpha})=\bigcap_{|\bm{n}|\leq e^{N^{\frac{\kappa_{1}}{5}}}}\left\{\bm{\theta}:\ \bm{\theta}+\bm{n}\bm{\alpha}\in\Theta_{N}\right\},$
    $\mathcal{W}^{\prime}=\left\{(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}:\ \bm{\alpha}\in\mathcal{A},\ \bm{\theta}\in\bigcap_{N\geq N_{1}}\tilde{\Theta}_{N}(\bm{\alpha})\right\}.$ (2.9)
    Then
    ${\rm meas}(\mathcal{W}^{\prime})\geq 1-\log^{-c}\frac{1}{\varepsilon}.$ (2.10)
Remark 2.6.
In this lemma, the constants $\kappa_{1},\kappa_{2}$ correspond to $c_{3},c_{4}$ of [JLS20], respectively. Originally, the off-diagonal exponential decay distance (cf. e.g., decay estimates in (2) of Lemma 2.5) in [Bou07, JLS20] is $|\bm{n}-\bm{n}^{\prime}|\geq\frac{N}{10}$. However, it can be improved to the present sublinear scale $N^{\frac{8}{9}}.$ This issue is important for the nonlinear analysis, and can be resolved with little effort, cf. Remark B.7 and Remark B.4 in the Appendix for more details.
Remark 2.7.
The proof of this lemma is based on analysis of semi-algebraic sets (cf. [Bou07, JLS20]), and applies to general non-degenerate analytic potentials.
Remark 2.8.
We mention that the set $\mathcal{W}$ on which Theorem 1.1 holds satisfies
$\mathcal{W}=\mathcal{W}^{\prime}\cap\mathcal{M},$ (2.11)
where $\mathcal{M}$ is the set on which the Diophantine estimates in Theorem 2.2 hold, and $\mathcal{W}^{\prime}$ is defined by ((2)). So, we have
${\rm meas}(\mathcal{W})\geq 1-\log^{-c}\frac{1}{\varepsilon+\delta},\ c=c(d)>0$
as shown in Theorem 1.1. For $(\bm{\alpha},\bm{\theta})\in\mathcal{W}$, the conclusion in (2) plays an essential role in the nonlinear analysis.
Proof.
It suffices to establish the measure bounds. The proof is a small modification of that of Theorem 4.1 in [JLS20] (cf. pages 475–476). Indeed, the only difference is that we will apply Theorem 3.7 in [JLS20] starting from $N_{1}=10^{-\frac{32}{\kappa_{1}^{2}}}\log^{\frac{4}{\kappa_{1}}}(\log\frac{1}{\varepsilon})$ rather than $\log\log\frac{1}{\varepsilon}$ as in [JLS20]. This change makes sense since the large deviation estimates in [JLS20] hold for the initial scales of $N_{0}(V,d)\leq N\leq N_{2}:=\frac{1}{10}\log^{2}\frac{1}{\varepsilon}$ (via the Neumann series argument), and for $f(x)=e^{x^{\frac{\kappa_{1}}{4}}},$
$N_{1},(f(N_{1}))^{\frac{8}{\kappa_{1}}}]\subset[N_{0},N_{2}].$
This modification then leads to the following measure estimate:
${\rm meas}(\bigcap_{N\geq N_{1}}\mathcal{A}_{N})$ $\geq 1-\sum_{N\geq N_{1}}f(N)^{-\kappa_{1}}$
$\geq 1-\sum_{N\geq N_{1}}e^{-\kappa_{1}N^{\frac{\kappa_{1}}{4}}}$
$\geq 1-C(d)e^{-\frac{\kappa_{1}}{2}N_{1}^{\frac{\kappa_{1}}{4}}}$
$\geq 1-C(d)\log^{-\frac{\kappa_{1}}{2}\times 10^{-\frac{8}{\kappa_{1}}}}\frac{1}{\varepsilon}.$
For the proof of (2.10), it follows directly from applying the Fubini’s theorem and
${\rm meas}(\bigcap_{N\geq N_{1}}\tilde{\Theta}_{N}(\bm{\alpha}))\geq 1-C(d)\log^{-\frac{1}{5}\times 10^{-\frac{8}{\kappa_{1}}}}\frac{1}{\varepsilon}.$
2.6. Large deviation theorem
In this section, we establish that the LDE (cf. Definition 2.1) hold for all sufficiently large scales by a multi-scale analysis. From (2.2), however, the variation of the frequency, $\bm{\omega}$ is only $O(\delta)$ and not $O(1)$, hence as we will see shortly, the proof of LDT will be accomplished in three steps instead of the usual two:
  • The first step deals with the small scales, for which we establish LDE for all $\bm{\omega}\in\Omega$ and scales
    $N_{0}\leq N\leq 10^{-\frac{1}{\rho}}\log^{\frac{1}{\rho}}\frac{1}{\varepsilon+\delta},$
    for some $\rho>0.$ In this step, we only use the Neumann series argument, which leads to the above scales (cf. Lemma 2.15).
  • Next, we establish intermediate scales LDE, i.e., scales in the range
    $10^{-\frac{1}{\rho}}\log^{\frac{1}{\rho}}\frac{1}{\varepsilon+\delta}\leq N\leq(\varepsilon+\delta)^{-c_{1}}$
    for all $\bm{\omega}\in\Omega$, where $c_{1}$ is defined in Theorem 2.2. Note that Theorem 2.2 is available in the above interval of scales. The proof of such intermediate scales LDE is based on preparation type theorem together with clustering properties of the spectrum of the diagonal matrix $D(0)$ (cf. (2.6), Theorem 2.2).
    Remark 2.9.
    The intermediate scales are needed here because for a fixed $(\bm{\alpha},\bm{\theta})$, hence fixed $\bm{\omega}^{(0)}$, $\bm{\omega}$ only varies in an interval of size $O(\delta)$, see (2.2), as mentioned above. Consequently, there is a gap in the range of scales before Bourgain-type analysis becomes applicable. Fortunately, the Diophantine estimates for $(\bm{\alpha},\bm{\theta})\in\mathcal{M}$ can control the resonances directly as in (2.6), leading to LDE for the range of scales in the gap, which we term the intermediate scales.
  • For the large scales (i.e., $N>(\varepsilon+\delta)^{-c_{1}}$) LDE, we will apply matrix-valued Cartan’s lemma and semi-algebraic sets theory of [BGS02, Bou07]. Of particular importance is Bourgain’s geometric lemma (cf. Lemma 2.5). This step requires both the Diophantine condition and the so called weak second Melnikov’s condition on $\bm{\omega}.$
    To be coherent, since the variation in $\bm{\omega}$ is only $O(\delta)$, the excised measure in $\Omega$ needs to be less than $\delta^{b}$. So we make a particular choice of the exponent and impose the following Diophantine condition:
    $\ \ \ {\rm DC}_{\bm{\omega}}(N)$
    $=\left\{\bm{\omega}\in\Omega:\ |\bm{k}\cdot\bm{\omega}|\geq e^{-N^{\rho^{4}}}\ {\rm for}\ \forall\ 0<|\bm{k}|\leq 100N^{2}\right\}.$ (2.12)
    Then we have for $\varepsilon\leq\delta$,
    ${\rm meas}\left(\Omega\setminus\left(\bigcap_{N>(\varepsilon+\delta)^{-c_{1}}}{\rm DC}_{\bm{\omega}}(N)\right)\right)\ll\delta^{b}.$ (2.13)
    The weak second Melnikov’s condition is imposed in order to restrict the time $\bm{k}$-direction where resonances could possibly occur. This greatly reduces the number of resonances, and paves the way for the application of Lemma 2.25.
    Remark 2.10.
    The condition on the size of $\varepsilon$ relative to $\delta$ stems again from the modulation in $\bm{\omega}$ being $O(\delta)$; while the excised measure depends on both $\varepsilon$ and $\delta$, as we will see in the proofs. This also ensures, in addition, that the amplitude-frequency map: $\bm{a}\to\bm{\omega}(\bm{a})$ is non-degenerate in the nonlinear analysis, so that we may indeed vary $\bm{\omega}$ in the linear analysis.
Before stating the LDT, let us summarize various parameters and the relation between $\varepsilon,\delta$.
  • Throughout this paper, we first fix constants
    $\kappa_{1}=\kappa_{1}(d),\kappa_{2}=\kappa_{2}(d)\in(0,1)$
    given in Bourgain’s geometric lemma (cf. Lemma 2.5).
  • The most important parameter $\rho>0$ is then fixed so that it depends only on $b,d$ and satisfies
    $\rho=c_{2}\kappa_{1}\leq\frac{\kappa_{1}}{10^{4}(b+d)^{2}},$ (2.14)
    where $c_{2}=c_{2}(b,d)$ is given by Lemma 2.22.
  • The parameters
    $0<c_{1}=c_{1}(b,d,K,V)<\frac{1}{100b},\ C_{1}=C_{1}(b,d,K,V)>1$
    are fixed, which appear in the Diophantine estimates (cf. Theorem 2.2). Also, the parameter $c_{1}$ is the power-law exponent of the upper bound of intermediate scales.
  • Fix
    $\kappa=\frac{1}{2}\min\left\{\vartheta,\ \frac{1}{10}\right\}\leq\frac{1}{20},$
    where $\vartheta>0$ is an absolute constant appeared in the coupling lemma of [Liu22] (cf. Lemma B.3). Then $\kappa$ (cf. (2.35)) describes the changes of off-diagonal exponential decay rates of Green’s functions along the multi-scale iterations.
  • For convenience, we assume the decay rate of $S$ satisfies $\gamma\in(\frac{1}{2},10)$. It turns out that the $(\rho,\gamma_{\infty},N)$-LDE hold true for all $N\geq\log^{3}\frac{1}{\varepsilon+\delta}$ and for some $\gamma_{\infty}\in[\frac{\gamma}{2},\gamma].$ The constant $C_{2}>0$ is the power-law index in $S.$
  • With these parameters having been fixed, we aim to find
    $\delta_{0}=\delta_{0}(b,d,K,V,\max\limits_{1\leq\ell\leq b}|\bm{n}_{\ell}|,\kappa_{1},\kappa_{2},\rho,c_{1},C_{2})>0,$
    so that, if $0<\varepsilon+\delta\leq\delta_{0}$, then LDT can be established. Since we will use the multi-scale analysis method to prove LDT, various conditions on $\delta_{0}$ are required, and it is unclear a priori if such restrictions are meaningful. Fortunately, we will divide the LDT proof into 3 steps, and at each step, we can find a desired $\delta_{\ell}$ ($\ell=1,2,3$). Finally, it suffices to take $\delta_{0}=\min\{\delta_{1},\delta_{2},\delta_{3}\}.$
  • As we will see later, we further assume $\delta\leq\log^{-1}\frac{1}{\varepsilon}$ when proving large scales LDE.
Let $\rho,\kappa_{1},\kappa_{2},c_{1},\gamma,C_{2}$ be fixed as above. We have
Theorem 2.11 (LDT).
There is some
$\delta_{0}=\delta_{0}(b,d,K,V,\max\limits_{1\leq\ell\leq b}|\bm{n}_{\ell}|,\kappa_{1},\kappa_{2},\rho,c_{1},C_{2})>0$
so that the following holds true for $0<\varepsilon+\delta\leq\delta_{0}.$
  • (1)
    For $\log^{3}\frac{1}{\varepsilon+\delta}<N\leq 10^{-\frac{1}{\rho}}\log^{\frac{1}{\rho}}\frac{1}{\varepsilon+\delta},$ the $(\rho,\gamma_{N},N)$-LDE hold true for all $(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}$ (and thus for all $\bm{\omega}\in\Omega$) with $\gamma_{N}\equiv\gamma_{0}=\gamma-\log^{-2}\frac{1}{\varepsilon+\delta}\in[\frac{\gamma}{2},\gamma].$
  • (2)
    For $10^{-\frac{1}{\rho}}\log^{\frac{1}{\rho}}\frac{1}{\varepsilon+\delta}\leq N\leq(\varepsilon+\delta)^{-c_{1}},$ the $(\rho,\gamma_{N},N)$-LDE hold true for $(\bm{\alpha},\bm{\theta})\in\mathcal{W}$ (cf. (2.11)) and all $\bm{\omega}\in\Omega$ (depending on $(\bm{\alpha},\bm{\theta})$) with $\gamma_{N}=\gamma_{0}-N^{-\vartheta}\in[\frac{\gamma}{2},\gamma].$
  • (3)
    Fix $(\bm{\alpha},\bm{\theta})\in\mathcal{W}$ (cf. (2.11)) and let $\varepsilon\leq\delta\leq\log^{-1}\frac{1}{\varepsilon}$. Let $N>(\varepsilon+\delta)^{-c_{1}}$, $N_{1}\sim N^{\rho^{2}}$ and $N_{2}\sim N^{2\rho}$. Then there are some $\widetilde{\Omega}_{N}\subset\Omega$ (independent of $S$) satisfying ${\rm meas}(\Omega\setminus\widetilde{\Omega}_{N})\leq e^{-\frac{1}{5}N^{\frac{3\kappa_{1}\rho^{2}}{4}}}$ and some $\gamma_{N}\in[\frac{\gamma}{2},\gamma]$ so that the $(\rho,\gamma_{N},N)$-LDE hold true for $\bm{\omega}\in\Omega_{N}:=\widetilde{\Omega}_{N}\cap{\rm DC}_{\bm{\omega}}(N)\cap\Omega_{N_{2}}$ with ${\rm DC}_{\bm{\omega}}(N)$ given by (2.12) and
    $\widetilde{\Omega}_{N}=\bigcap_{1\leq\ell,\ell^{\prime}\leq N^{4(b+d)},\xi=\pm 1,\xi^{\prime}=\pm 1,0<|\bm{k}|\leq 2N^{2}}\widetilde{\Omega}_{\bm{k},\ell,\ell^{\prime},\xi,\xi^{\prime}},$
    where
    $\widetilde{\Omega}_{\bm{k},\ell,\ell^{\prime},\xi,\xi^{\prime}}:=\left\{\bm{\omega}\in\Omega:\ |\bm{k}\cdot\bm{\omega}+\xi{\lambda_{\ell}}-\xi^{\prime}{\lambda_{\ell^{\prime}}}|>10e^{-\frac{1}{4}N^{\frac{3\kappa_{1}\rho^{2}}{4}}}\right\},$
    $\gamma_{N}\geq\gamma_{N_{1}}-N^{-\kappa}$ and $\lambda_{\ell}=\lambda_{\ell}(\bm{\alpha},\bm{\theta})$ (${1\leq\ell\leq N^{4(b+d)}})$ are all eigenvalues of the operators $\mathcal{H}_{Q}(\bm{\alpha},\bm{\theta})$ for all $Q\subset[-100N^{2},100N^{2}]^{d}$ satisfying
[…]
Remark 2.12.
Note that the sets $\Omega_{N}$ are independent of $S$. This will be important for later nonlinear applications. Moreover, we have $\Omega_{N}\subset\Omega_{N-1}\subset\cdots\subset\Omega_{N_{0}}=\Omega$ and $\Omega_{N}$ is a semi-algebraic set of ${\rm deg}\ \Omega_{N}\leq N^{10(b+d)}$ and ${\rm meas}(\Omega_{N^{2\rho}}\setminus\Omega_{N})\leq e^{-\frac{1}{2}N^{\rho^{4}}}.$
Remark 2.13.
One may observe that this LDT is for a fixed operator $H$; while the linearized operator is varying in a Newton scheme. This issue, however, can be remedied by using that the Green’s function estimates in LDE (at scale $N$) are stable under perturbations of order $e^{-N^{2}}$ using Lemma B.1. The sup-exponentially decaying error (cf. (3.19)) in the Newton scheme can ensure this. More precisely, let
$\gamma\geq\gamma_{1}\geq\gamma_{2}\geq\cdots\geq\frac{\gamma}{2},$
$\gamma\geq\gamma_{1}^{\prime}\geq\gamma_{2}^{\prime}\geq\cdots\geq\frac{\gamma}{2},$
$\gamma_{\ell}^{\prime}\leq\min\{\gamma_{[\ell^{\rho^{2}}]+1},\gamma_{[\ell^{\rho^{2}}]+1}^{\prime}\}-\ell^{-\kappa}.$
We can replace $H$ in LDT with a sequence of operators $H^{(\ell)}(\sigma)=D{(\sigma)}+\delta S_{\ell}$ ($\ell\geq 1$) satisfying (we hide the dependence on $\xi,\xi^{\prime}\in\{\pm\}$)
$|S_{\ell}((\bm{k},\bm{n});(\bm{k}^{\prime},\bm{n}^{\prime}))|$ $\leq C_{2}(1+|\bm{k}-\bm{k}^{\prime}|)^{C_{2}}e^{-\gamma_{\ell}|\bm{k}-\bm{k}^{\prime}|-\gamma_{\ell}|\bm{n}|}\delta_{\bm{n},\bm{n}^{\prime}},$
$\|H^{(\ell+1)}-H^{(\ell)}\|$ $\leq e^{-\ell^{\frac{2}{\rho^{2}}}}.$
It is remarkable that in our LDT, $\Omega_{N}$ does not depend on $S$. So with this modification, the conclusion of LDT becomes for $\bm{\omega}\in\Omega_{N},$ $H^{(\ell)}$ has $(\rho,\gamma_{N}^{\prime},N)$-LDE for all $\ell\geq N\geq\log^{3}{\frac{1}{\varepsilon+\delta}}$ ($\Sigma_{N}$ depending only on $S_{\ell}$, $1\leq\ell\leq N$).
Indeed, assume we have established the LDE at scale $N_{1}=N^{\rho^{2}}$ for all $H^{(\ell)}$ with $\ell\geq N_{1}$. Then it suffices to establish $(\rho,\gamma_{N}^{\prime},N)$-LDE for $H^{(N)}$, since for all $\ell>N$,
$\|H^{(\ell)}-H^{(N)}\|\leq\sum\limits_{\ell\geq N}e^{-\ell^{\frac{2}{\rho^{2}}}}<e^{-\frac{N^{\frac{2}{\rho^{2}}}}{2}}\ll e^{-100N},$
and $(\rho,\gamma_{N}^{\prime},N)$-LDE for $H^{(\ell)}$ can be obtained via the perturbation argument (cf. Lemma B.1). Next, we fix $S_{\ell}\equiv S_{N_{1}}$ for $\ell\in[N_{1},N]$. Similar to the proof of Lemma 2.20, we can show using the modified resolvent identities Lemmas B.3, B.5, B.6 (e.g., taking account of the power-law factor in the estimate of $S_{\ell}$) that $H^{(N_{1})}$ satisfies $(\rho,\gamma_{N}^{\prime\prime},N)$-LDE with
$\gamma_{N}^{\prime\prime}\geq\min\{\gamma_{N_{1}},\gamma_{N_{1}}^{\prime}\}-N^{-\kappa}\geq\gamma_{N}^{\prime}.$
Finally, using again the perturbation argument Lemma B.1 and
$\|H^{(N_{1})}-H^{(N)}\|\leq e^{-\frac{N_{1}^{\frac{2}{\rho^{2}}}}{2}}\leq e^{-\frac{N^{2}}{5}}\ll e^{-100N},$
we get that $H^{(N)}$ satisfies $(\rho,\gamma_{N}^{\prime},N)$-LDE by the assumptions on $\gamma_{\ell},\gamma_{\ell}^{\prime}$.
Remark 2.14.
In the large scales case, we assume $\delta\in[\varepsilon,\log^{-1}\frac{1}{\varepsilon}]$. The condition $\delta\geq\varepsilon$ is necessary since (2.13) as mentioned above. The restriction $\delta\leq\log^{-1}\frac{1}{\varepsilon}$, however, is due to the fact that we need to use Bourgain’s geometric lemma in the form of Lemma 2.5 to handle the large $\bm{|}\bm{n}|$ regime estimates. In fact, if we assume that $\mathcal{W}$ (cf. (2.11)) satisfies ${\rm meas}([0,1]^{2d}\setminus\mathcal{W})=o(1)$ (rather than the quantitative version ${\rm meas}([0,1]^{2d}\setminus\mathcal{W})=O(\log^{-c}\frac{1}{\varepsilon+\delta})$ in the present work), we can remove the condition $\delta\leq\log^{-1}\frac{1}{\varepsilon}$ (cf. Remark 2.24 for details).
This LDT will be proved in detail in the following three sections.
2.6.1. The small scales LDE
This section is devoted to the proof of LDE for small scales, i.e.,
$\log^{3}\frac{1}{\varepsilon+\delta}\leq N\leq 10^{-\frac{1}{\rho}}\log^{\frac{1}{\rho}}\frac{1}{\varepsilon+\delta}.$
In this case, we only use the standard Neumann series argument. As a result, we can establish LDE for all $\bm{\omega}\in\Omega.$ We have
Lemma 2.15.
Let
$\ \ \ \Sigma_{N}$
$=\bigcup_{\Lambda\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N),\ |\bm{n}|\leq 10N}\left\{\sigma\in\mathbb{R}:\ \min_{(\bm{k},\bm{n},\xi)\in\Lambda}|\xi(\sigma+\bm{k}\cdot\bm{\omega})+\mu_{\bm{n}}|\leq e^{-2N^{\rho}}\right\}.$
Let $0<\rho<\frac{1}{10^{4}}$. Then there is some $\delta_{1}=\delta_{1}(b,d,C_{2},\rho)>0$ so that if $0<\varepsilon+\delta\leq\delta_{1},$ for
$\log^{3}\frac{1}{\varepsilon+\delta}\leq N\leq 10^{-\frac{1}{\rho}}\log^{\frac{1}{\rho}}\frac{1}{\varepsilon+\delta},$
we have
${\rm meas}(\Sigma_{N})\leq e^{-N^{\rho}}.$ (2.15)
Moreover, if $\sigma\notin\Sigma_{N}$, we have for all $\Lambda\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N)$ with $|\bm{n}|\leq 10N$,
$\|G_{\Lambda}(\sigma)\|$ $\leq e^{N^{\frac{3}{4}}},$ (2.16)
$|G_{\Lambda}(\sigma)((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))|$ $\leq e^{-\gamma_{0}|\bm{x}-\bm{x}^{\prime}|}\ {\rm for}\ |\bm{x}-\bm{x}^{\prime}|\geq N^{\frac{8}{9}},$ (2.17)
where
$\gamma_{0}=\gamma-\log^{-2}\frac{1}{\varepsilon+\delta}\in[\frac{\gamma}{2},\gamma].$ (2.18)
Remark 2.16.
In this lemma, we make no restrictions on $\bm{\omega}.$
Proof.
We first establish the measure bound (2.15). We have
${\rm meas}(\Sigma_{N})$ $\leq C(b,d)N^{C(b,d)}e^{-2N^{\rho}}\leq e^{-N^{\rho}},$
where in the last inequality we use $0<\varepsilon+\delta\leq\delta_{1}^{(1)}(b,d,\rho)\ll 1$ to ensure $N\geq\log^{3}\frac{1}{\varepsilon+\delta}\geq N_{0}(b,d,\rho)\gg 1.$
Next, we let $\sigma\notin\Sigma_{N}$ and fix any $\Lambda\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N)$ with $|\bm{n}|\leq 10N$. Let $A=D_{\Lambda}(\sigma),B=\delta S_{\Lambda}$. We have
$|A^{-1}(\bm{x},\bm{x}^{\prime})|\leq e^{N^{\rho}}\delta_{\bm{x},\bm{x}^{\prime}},\ |B(\bm{x},\bm{x}^{\prime})|\leq C_{2}\delta(1+|\bm{x}-\bm{x}^{\prime}|)^{C_{2}}e^{-\gamma|\bm{x}-\bm{x}^{\prime}|}.$
Note that
$CN^{2(b+d)+C_{2}}\delta e^{N^{\rho}}$ $\leq C(b,d,C_{2})N^{2(b+d)+C_{2}}(\varepsilon+\delta)^{\frac{9}{10}}$
$\leq C(b,d,C_{2})10^{-\frac{2(b+d)+C_{2}}{\rho}}(\varepsilon+\delta)^{\frac{9}{10}}\log^{-\frac{2(b+d)+C_{2}}{\rho}}\frac{1}{\varepsilon+\delta}$
$\leq\frac{1}{2}$
assuming $0<\varepsilon+\delta\leq\delta_{1}^{(2)}(b,d,C_{2},\rho)\ll 1.$ So we can apply Lemma B.1 with $A=D_{\Lambda}(\sigma),B=\delta S_{\Lambda},\epsilon_{2}=C_{2}\delta,\epsilon_{1}=e^{-N^{\rho}},M=0,X=\Lambda$ to obtain (we hide the dependence on $\xi,\xi^{\prime}$)
$\|G_{\Lambda}(\sigma)\|$ $\leq 2e^{N^{\rho}}<e^{N^{\frac{3}{4}}}\ ({\rm since}\ 0<\rho<\frac{1}{10^{4}}),$
$|G_{\Lambda}(\sigma)(\bm{x},\bm{x}^{\prime})|$ $\leq e^{N^{\rho}}e^{-\gamma|\bm{x}-\bm{x}^{\prime}|}.$
Then for $|\bm{x}-\bm{x}^{\prime}|\geq N^{\frac{8}{9}},$
$|G_{\Lambda}(\sigma)(\bm{x},\bm{x}^{\prime})|$ $\leq e^{-|\bm{x}-\bm{x}^{\prime}|(\gamma-\frac{N^{\rho}}{|\bm{x}-\bm{x}^{\prime}|})}$
$\leq e^{-|\bm{x}-\bm{x}^{\prime}|(\gamma-N^{-\frac{7}{8}})}\ \ ({\rm since}\ 0<\rho<\frac{1}{10^{4}})$
$:=e^{-\gamma^{\prime}|\bm{x}-\bm{x}^{\prime}|},$
which implies
$\gamma^{\prime}>\gamma_{0}=\gamma-\log^{-2}\frac{1}{\varepsilon+\delta}.$
Then both (2.16) and (2.17) hold true if we assume that
$0<\varepsilon+\delta\leq\delta_{1}=\min\{\delta_{1}^{(1)}(b,d,C_{2},\rho),\delta_{1}^{(2)}(b,d,C_{2},\rho)\},$
and $\sigma\notin\Sigma_{N}.$
2.6.2. The intermediate scales LDE
This section and the next deal with intermediate scales LDE. We will show that the LDE hold true for all $\bm{\omega}\in\Omega$ and scales in the range
$10^{-\frac{1}{\rho}}\log^{\frac{1}{\rho}}\frac{1}{\varepsilon+\delta}\leq N\leq(\varepsilon+\delta)^{-c_{1}},$
where $c_{1}$ is given by Theorem 2.2. In this case, the analysis in Section 2.4 becomes essential.
Lemma 2.17.
Let $0<\rho<\frac{1}{10^{4}}$. There is some $\delta_{2}=\delta_{2}(b,d,K,V,\rho,c_{1},C_{2})>0$ so that the following holds for $0<\varepsilon+\delta\leq\delta_{2}.$ Assume $(\bm{\alpha},\bm{\theta})\in\mathcal{M}$ (cf. Theorem 2.2) and fix
$10^{-\frac{1}{\rho}}\log^{\frac{1}{\rho}}\frac{1}{\varepsilon+\delta}\leq N\leq(\varepsilon+\delta)^{-c_{1}}.$
Then for each $\bm{\omega}\in\Omega$, there exists a set $\Sigma_{N}\subset\mathbb{R}$ with
${\rm meas}(\Sigma_{N})\leq(\varepsilon+\delta)^{-\frac{9}{10}}e^{-\frac{1}{2b}N^{\frac{3}{8}}}<e^{-N^{\rho}},$
so that if $\sigma\notin\Sigma_{N}$, then for all $\Lambda\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N)$ with $|\bm{n}|\leq 10N$,
$\|G_{\Lambda}(\sigma)\|$ $\leq e^{N^{\frac{3}{4}}},$
$|G_{\Lambda}(\sigma)((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))|$ $\leq e^{-\gamma_{N}|\bm{x}-\bm{x}^{\prime}|}\ {\rm for}\ |\bm{x}-\bm{x}^{\prime}|\geq N^{\frac{8}{9}},$
where
$\gamma_{N}=\gamma_{0}-N^{-\vartheta}\in[\frac{\gamma}{2},\gamma],$
with $\gamma_{0}$ given by (2.18) and $\vartheta>0$ by Lemma B.3.
Remark 2.18.
Note that this lemma also holds for all $\bm{\omega}\in\Omega$.
Proof.
The proof is based on clustering property (cf. (2) of Theorem 2.2), Schur’s completement argument, the Rouché theorem and the iteration of resolvent identities. So we first choose $0<\varepsilon+\delta\leq c(b,d,K,V,\max\limits_{1\leq\ell\leq b}|\bm{n}_{\ell}|)\ll 1$ so that Theorem 2.2 holds true. Then the proof can be decomposed into two steps.
Step 1. We define
$\Sigma_{N}=\bigcup_{\sqrt{N}\leq N^{\prime}\leq N,\ \Lambda\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N^{\prime}),\ |\bm{n}|\leq 10N^{\prime}}\Sigma_{\Lambda},$ (2.19)
where
$\Sigma_{\Lambda}=\left\{\sigma\in\mathbb{R}:\ \|G_{\Lambda}(\sigma)\|\geq e^{({\rm diam}\ \Lambda)^{\frac{3}{4}}}\right\}.$
We first show that
$\Sigma_{\Lambda}\subset\bigcup_{\xi=\pm 1,\bm{i}=(\bm{k},\bm{n})\in\Lambda}I_{\xi,\bm{i}},$ (2.20)
where
$I_{\xi,\bm{i}}=\left\{\sigma\in\mathbb{R}:\ |\xi(\sigma+\bm{k}\cdot\bm{\omega}^{(0)})+\mu_{\bm{n}}|\leq\frac{1}{100}(\varepsilon+\delta)^{\frac{1}{7b}}\right\}.$
In fact, if $\sigma\notin\bigcup\limits_{(\xi,\bm{i})\in\Lambda}I_{\xi,\bm{i}},$ then for each $\bm{\omega}\in\Omega,$
$\ \ \ \ \inf_{(\xi,\bm{i})\in\Lambda}|\xi(\sigma+\bm{k}\cdot\bm{\omega})+\mu_{\bm{n}}|$
$\geq\inf_{(\xi,\bm{i})\in\Lambda}|\xi(\sigma+\bm{k}\cdot\bm{\omega}^{(0)})+\mu_{\bm{n}}|-\sup_{\bm{k}\in\Lambda}|\bm{k}\cdot(\bm{\omega}-\bm{\omega}^{(0)})|$
$\geq\frac{1}{200}(\varepsilon+\delta)^{\frac{1}{7b}}.$
Since $N\geq 10^{-\frac{1}{\rho}}\log^{\frac{1}{\rho}}\frac{1}{\varepsilon+\delta}$, using Neumann series argument implies
$\|G_{\Lambda}(\sigma)\|\leq C(\varepsilon+\delta)^{-\frac{2}{7b}}<e^{({\rm diam}\ \Lambda)^{\frac{3}{4}}}$
when $\sigma\notin\bigcup\limits_{(\xi,\bm{i})\in\Lambda}I_{\xi,\bm{i}}.$ This proves (2.20) and
$\Sigma_{\Lambda}=\bigcup_{(\xi,\bm{i})\in\Lambda}(\Sigma_{\Lambda}\cap I_{\xi,\bm{i}}).$
In the following, we estimate ${\rm meas}(\Sigma_{\Lambda}\cap I_{\xi,\bm{i}}).$ For this purpose, we fix any $\xi^{*}\in\{+,-\}$, $\bm{i}^{*}=(\bm{k}^{*},\bm{n}^{*})$ and denote
$I^{*}=I_{\xi^{*},\bm{i}^{*}}=\left[\sigma^{*}-\frac{1}{100}(\varepsilon+\delta)^{\frac{1}{7b}},\sigma^{*}+\frac{1}{100}(\varepsilon+\delta)^{\frac{1}{7b}}\right],$
where
$\sigma^{*}=-\bm{k}^{*}\cdot\bm{\omega}^{(0)}-{\xi^{*}}\mu_{\bm{n}^{*}}.$ (2.21)
We will study $G_{\Lambda}(\sigma)$ for $\sigma\in I^{*}.$ From (2) of Theorem 2.2, we have for every $\sigma^{*}\in\mathbb{R}$, there exists $B_{*}=B_{*}^{+}\cup B_{*}^{-}\subset\Lambda$ with $B_{*}^{\pm}\leq b$ so that
$\inf_{(\xi,\bm{i})\in\Lambda\setminus B_{*}}|\xi(\sigma^{*}+\bm{k}\cdot\bm{\omega}^{(0)})+\mu_{\bm{n}}|\geq\frac{(\varepsilon+\delta)^{\frac{1}{8b}}}{4}.$
Similar to the proof of (2.20), if (we denote $\mathbb{D}_{r}(z):=\{w\in\mathbb{C}:\ |w-z|\leq r\}$) $z\in\mathbb{D}_{(\varepsilon+\delta)^{\frac{1}{8b}}/10}(\sigma^{*}),$ then
$\inf_{(\xi,\bm{i})\in\Lambda\setminus B_{*}}|\xi(z+\bm{k}\cdot\bm{\omega})+\mu_{\bm{n}}|\geq\frac{(\varepsilon+\delta)^{\frac{1}{8b}}}{8}.$
So we have an analytic extension of ${H}_{\Lambda}(z)$ to $z\in\mathbb{D}_{(\varepsilon+\delta)^{\frac{1}{8b}}/10}(\sigma^{*})$ and
$\sup_{(\bm{x},\xi),(\bm{x}^{\prime},\xi^{\prime})\in\Lambda,\ z\in\mathbb{D}_{(\varepsilon+\delta)^{\frac{1}{8b}}/10}(\sigma^{*})}|{D}_{\Lambda}(z)((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))|\leq C(b)(\varepsilon+\delta)^{-c_{1}}.$
More importantly, if we denote $\Lambda^{c}=\Lambda\setminus B_{*}$, then by the Neumann series argument, we have for all $z\in\mathbb{D}_{(\varepsilon+\delta)^{\frac{1}{8b}}/10}(\sigma^{*}),$
$\|G_{\Lambda^{c}}(z)\|\leq(\varepsilon+\delta)^{-\frac{1}{4b}}.$
In the following, we estimate $G_{\Lambda}(z)$ using the Schur complement reduction. We define
$S(z)=R_{B_{*}}{H}(z)R_{B_{*}}-R_{B_{*}}{H}(z)R_{\Lambda^{c}}G_{\Lambda^{c}}(z)R_{\Lambda^{c}}{H}(z)R_{B_{*}}.$
Obviously, we have
$R_{B_{*}}{H}(z)R_{B_{*}}$ $=R_{B_{*}}{D}(z)R_{B_{*}}+O(\varepsilon+\delta)$ (2.22)
and
$R_{B_{*}}{H}(z)R_{\Lambda^{c}}G_{\Lambda^{c}}(z)R_{\Lambda^{c}}{H}(z)R_{B_{*}}=O((\varepsilon+\delta)^{2-\frac{1}{4b}}).$ (2.23)
Now we consider $\det S(z)$. Combining (2.22) and (2.23) yields
$\ \ \ \det S(z)$
$=\prod_{\bm{i}=(\bm{k},\bm{n})\in B_{*}^{+}}(z+\bm{k}\cdot\bm{\omega}+\mu_{\bm{n}})\prod_{\bm{i}=(\bm{k},\bm{n})\in B_{*}^{-}}(-z-\bm{k}\cdot\bm{\omega}+\mu_{\bm{n}})$
$\ \ \ +O((\varepsilon+\delta)^{3/4})$
$=(-1)^{\#B_{*}^{-}}\prod_{\ell=1}^{\#B_{*}}(z-\sigma_{\ell})+O((\varepsilon+\delta)^{3/4}),$
where $\sigma_{\ell}\in\{-\bm{k}\cdot\bm{\omega}-\mu_{\bm{n}}\}_{(\bm{k},\bm{n})\in B_{*}^{+}}\cup\{-\bm{k}\cdot\bm{\omega}+\mu_{\bm{n}}\}_{(\bm{k},\bm{n})\in B_{*}^{-}}.$ Recalling the definition of $B_{*}$, we must have
$|z-\sigma_{\ell}|\leq\frac{4}{5}(\varepsilon+\delta)^{\frac{1}{8b}}.$
From (2.21), there exists at least one $\sigma_{\ell}$ so that $|\sigma_{\ell}-\sigma^{*}|=O((\varepsilon+\delta)^{1-c_{1}}).$ At this stage, we need a useful result.
Lemma 2.19.
Let $\varphi=\left(\frac{14}{15}\right)^{\frac{1}{5b}}<1$. There is $0\leq\ell_{*}\leq 5b-1$ so that
$\left(\mathbb{A}_{\ell_{*}+1}\setminus\mathbb{A}_{\ell_{*}}\right)\cap\{\sigma_{\ell}\}_{1\leq\ell\leq\#B_{*}}=\emptyset,$
where for $0\leq\ell\leq 5b$,
$\mathbb{A}_{\ell}=\mathbb{D}_{(\varepsilon+\delta)^{\frac{\varphi^{\ell}}{7b}}}(\sigma^{*}).$
Proof of Lemma 2.19.
It suffices to note that
$\mathbb{A}_{5b}\setminus\mathbb{A}_{0}=\bigcup_{\ell=0}^{5b-1}\left(\mathbb{A}_{\ell+1}\setminus\mathbb{A}_{\ell}\right).$
The proof then follows from the pigeonhole principle since $B_{*}\leq 2b$. ∎
Let $r_{*}=(\varepsilon+\delta)^{\frac{\varphi^{\ell_{*}}}{7b}}$, where $\varphi$ and $\ell_{*}$ are defined in Lemma 2.19. From Lemma 2.19, we can restrict our considerations on $\mathbb{A}_{*}$ with
$\mathbb{A}_{*}=\mathbb{D}_{\frac{r_{*}+r_{*}^{\varphi}}{2}}(\sigma^{*}).$
Then
$\mathbb{D}_{(\varepsilon+\delta)^{\frac{1}{7b}}}(\sigma^{*})\subset\mathbb{A}_{\ell_{*}}\subset\mathbb{A}_{*}\subset\mathbb{D}_{(\varepsilon+\delta)^{\frac{1}{7.5b}}}(\sigma^{*}).$
Let
$\mathcal{K}=\{1\leq\ell\leq\#B_{*}:\ \sigma_{\ell}\in\mathbb{D}_{10r_{*}}(\sigma^{*})\}.$
We have $\mathcal{K}\neq\emptyset$ and for $\ell\not\in\mathcal{K}$,
$\sigma_{\ell}\notin\mathbb{A}_{*},\ {\rm dist}(\sigma_{\ell},\partial\mathbb{D}_{20r_{*}}(\sigma^{*}))\geq{\rm dist}(\sigma_{\ell},\partial\mathbb{A}_{*})\geq\frac{r_{*}^{\varphi}}{4}\gg(\varepsilon+\delta)^{\frac{1}{7b}}.$
As a result, we can write for $z\in\mathbb{A}_{*},$
$\frac{\det S(z)}{\prod_{\ell\notin\mathcal{K}}(z-\sigma_{\ell})}=\pm{\prod_{\ell\in\mathcal{K}}(z-\sigma_{\ell})}+O((\varepsilon+\delta)^{\frac{1}{3}}).$ (2.24)
It holds that
${\prod_{\ell\in\mathcal{K}}|z-\sigma_{\ell}|}\big|_{\partial\mathbb{D}_{20r_{*}}(\sigma^{*})}\geq(\varepsilon+\delta)^{\frac{\#\mathcal{K}}{7b}}\geq(\varepsilon+\delta)^{2/7}\gg(\varepsilon+\delta)^{\frac{1}{3}}.$
Applying Rouché theorem shows that the function $\frac{\det S(z)}{\prod_{\ell\notin\mathcal{K}}(z-\sigma_{\ell})}$ has exactly $\mathcal{K}$ zeros (denoted by $z_{\ell},\ \ell\in\mathcal{K}$) in $\mathbb{D}_{20r_{*}}(\sigma^{*}).$ From $\sigma_{\ell}\in\mathbb{D}_{10r_{*}}(\sigma^{*})$, we also get
$|\sigma_{\ell}-z_{\ell}|\leq 30r_{*}.$ (2.25)
From similar analysis, we can obtain that (2.24) also has exactly $\mathcal{K}$ zeros in $\mathbb{A}_{*}$, which implies that $\{z_{\ell}\}_{\ell\in\mathcal{K}}$ are all zeros of (2.24) in $\mathbb{A}_{*}$. So on $\mathbb{A}_{*}$, we have
$\frac{\det S(z)}{\prod_{\ell\notin\mathcal{K}}(z-\sigma_{\ell})}=g(z)\prod_{\ell\in\mathcal{K}}(z-z_{\ell}),$ (2.26)
where $g$ is analytic on $\mathbb{A}_{*}$ with $\inf\limits_{z\in\mathbb{A}_{*}}|g(z)|>0.$ Using (2.25) and (2.26), we have
$g\big|_{\partial\mathbb{A}_{*}}=1+O((\varepsilon+\delta)^{\frac{1}{21}}).$
Using the maximum principle then leads to
$g\big|_{\mathbb{A}_{*}}=1+O((\varepsilon+\delta)^{\frac{1}{21}}).$
We have established on $\mathbb{A}_{*},$
$\det S(z)=g(z)\prod_{\ell\notin\mathcal{K}}(z-\sigma_{\ell})\prod_{\ell\in\mathcal{K}}(z-z_{\ell}).$
In particular, one has for all $z\in\mathbb{A}_{*},$
$|\det S(z)|$ $\geq\frac{4}{5}(\varepsilon+\delta)^{\frac{\#B_{*}-\#\mathcal{K}}{7b}}\prod_{\ell\in\mathcal{K}}|z-z_{\ell}|$
$\geq(\varepsilon+\delta)^{\frac{2}{7}}\prod_{\ell\in\mathcal{K}}|z-z_{\ell}|$
assuming $0<\varepsilon+\delta\leq c(b)\ll 1.$ Thus combining Hadamard’s inequality and Cramer’s rule implies for $z\in\mathbb{A}_{*},$
$\|S^{-1}(z)\|$ $\leq C(b)(\varepsilon+\delta)^{-{2bc_{1}}}(\varepsilon+\delta)^{-\frac{2}{7}}\prod_{\ell\in\mathcal{K}}|z-z_{\ell}|^{-1}$
$\leq(\varepsilon+\delta)^{-\frac{1}{3}}\prod_{\ell\in\mathcal{K}}|z-z_{\ell}|^{-1}\ ({\rm since}\ c_{1}<\frac{1}{100b}).$
Applying the Schur’s complement argument shows that for $z\in\mathbb{A}_{*},$
$\|G_{\Lambda}(z)\|$ $\leq 4(1+\|G_{\Lambda^{c}}(z)\|)^{2}(1+\|S^{-1}(z)\|)$
$\leq(\varepsilon+\delta)^{-\frac{7}{8}}\prod_{\ell\in\mathcal{K}}|z-z_{\ell}|^{-1}.$
Recall that $I^{*}\subset\mathbb{A}_{*}$. So for $\sigma\in I^{*},$ we have
$\|G_{\Lambda}(\sigma)\|$ $\leq(\varepsilon+\delta)^{-\frac{7}{8}}\prod_{\ell\in\mathcal{K}}|\sigma-z_{\ell}|^{-1}$
and
$\Sigma_{\Lambda}\cap I^{*}\subset\bigcup_{\ell\in\mathcal{K}}\left\{\sigma\in I^{*}:\ |\sigma-z_{\ell}|\leq(\varepsilon+\delta)^{-\frac{7}{8}}e^{-\frac{1}{2b}({\rm diam}\ \Lambda)^{\frac{3}{4}}}\right\}.$
This implies
${\rm meas}(\Sigma_{\Lambda}\cap I^{*})\leq C(\varepsilon+\delta)^{-\frac{7}{8}}e^{-\frac{1}{2b}({\rm diam}\ \Lambda)^{\frac{3}{4}}},$
and subsequently
${\rm meas}(\Sigma_{\Lambda})\leq(\varepsilon+\delta)^{-\frac{8}{9}}e^{-\frac{1}{2b}({\rm diam}\ \Lambda)^{\frac{3}{4}}},$
which combined with (2.19) shows
${\rm meas}(\Sigma_{N})\leq(\varepsilon+\delta)^{-\frac{9}{10}}e^{-\frac{1}{2b}N^{\frac{3}{8}}}.$
Step 2. In this step, we prove the off-diagonal exponential decay of $G_{\Lambda}(\sigma)((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))$ assuming $\sigma\not\in\Sigma_{N}$ and $\Lambda\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N)$ with $|\bm{n}|\leq 10N.$ By Theorem 2.2, we know that there exists $B\subset\Lambda$ so that
$\inf_{(\xi,\bm{k},\bm{n})\in\Lambda\setminus B}|\xi(\sigma+\bm{k}\cdot\bm{\omega})+\mu_{\bm{n}}|\geq\frac{(\varepsilon+\delta)^{\frac{1}{8b}}}{8}.$
Using the Neumann series argument (cf. Lemma B.1) yields that for any $\Lambda^{\prime}\subset\Lambda$ with $\Lambda^{\prime}\cap B=\emptyset,$
$\|G_{\Lambda^{\prime}}(\sigma)\|$ $\leq 50(\varepsilon+\delta)^{-\frac{1}{4b}},$ (2.27)
$|G_{\Lambda^{\prime}}(\sigma)((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))|$ $\leq 50(\varepsilon+\delta)^{-\frac{1}{4b}}e^{-\gamma|\bm{x}-\bm{x}^{\prime}|}\ {\rm for}\ \bm{x}\neq\bm{x}^{\prime}.$ (2.28)
We say that $\mathcal{ER}(\sqrt{N})\ni Q,\ Q\subset\Lambda$ is $\sqrt{N}$-regular if both (2.27) and (2.28) hold with $\Lambda^{\prime}=Q.$ Otherwise, $Q$ is called $\sqrt{N}$-singular. Let $\mathcal{F}$ be any family of pairwise disjoint $\sqrt{N}$-singular sets contained in $\Lambda$. We obtain
$\#\mathcal{F}\leq 2b.$
We have established the sublinear bound. Then applying the coupling lemma, i.e., Lemma B.3 of [Liu22] shows
$|G_{\Lambda}(\sigma)((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))|\leq e^{-\gamma_{N}|\bm{x}-\bm{x}^{\prime}|}\ {\rm for}\ |\bm{x}-\bm{x}^{\prime}|>N^{\frac{8}{9}},$
where
$\gamma_{N}$ $\geq\gamma-N^{-\frac{8}{9}}\cdot\log\frac{1}{\varepsilon+\delta}-N^{-\vartheta}$
$\geq\gamma-{\log^{-\frac{8}{9\rho}+1}\frac{1}{\varepsilon+\delta}}-N^{-\vartheta}$
$\geq\gamma_{0}-N^{-\vartheta}\ ({\rm since}\ 0<\rho<\frac{1}{10^{4}}).$
This proves the off-diagonal exponential decay estimates.
We have completed the proof of Lemma 2.17.
2.6.3. The large scales LDE
In this section, we will prove the LDE for
$(\varepsilon+\delta)^{-c_{1}}<N<\infty.$
To perform the multi-scale scheme, we need two small scales $N_{1}<N_{2}$ with
$N_{2}=N_{1}^{\frac{2}{\rho}},\ N=N_{1}^{\frac{1}{\rho^{2}}}.$
We assume the LDE hold for $\bm{\omega}\in\Omega_{N_{2}}\subset\Omega_{N_{1}}$.
Then we have
Lemma 2.20.
There are some $0<c_{2}=c_{2}(b,d)\leq\frac{1}{10^{4}(b+d)^{2}}$ and some
$\delta_{3}=\delta_{3}(b,d,K,V,\max\limits_{1\leq\ell\leq b}|\bm{n}_{\ell}|,\kappa_{1},\kappa_{2},\rho,c_{1},C_{2})>0$
so that the following holds true for $0<\varepsilon+\delta\leq\delta_{3}$, $\varepsilon\leq\delta\leq\log^{-1}\frac{1}{\varepsilon}$ and $0<\rho\leq c_{2}\kappa_{1}$ with $\kappa_{1}$ given by Lemma 2.5. Let $(\bm{\alpha},\bm{\theta})\in\mathcal{W}$ (cf. (2.11)) and NN satisfy
$(\varepsilon+\delta)^{-c_{1}}<N<\infty.$
Then there is a semi-algebraic set $\widetilde{\Omega}_{N}\subset\Omega$ (cf. Lemma 2.26) with ${\rm meas}(\Omega\setminus\widetilde{\Omega}_{N})\leq e^{-\frac{1}{5}N^{\frac{3\kappa_{1}\rho^{2}}{4}}}$ and $\deg\ \widetilde{\Omega}_{N}\leq N^{5(b+d)}$ so that, for $\bm{\omega}\in{\rm DC}_{\bm{\omega}}(N)\cap\widetilde{\Omega}_{N}\cap{\Omega_{N_{2}}}$ (with ${\rm DC}_{\bm{\omega}}(N)$ given by (2.12)), there is some $\Sigma_{N}\subset\mathbb{R}$ with ${\rm meas}(\Sigma_{N})\leq e^{-N^{\rho}}$, so that, if $\sigma\notin\Sigma_{N}$, then for all $\Lambda\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N)$ with $|\bm{n}|\leq 10N,$
$\|G_{\Lambda}(\sigma)\|$ $\leq e^{N^{\frac{3}{4}}},$
$|G_{\Lambda}(\sigma)((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))|$ $\leq e^{-\gamma_{N}|\bm{x}-\bm{x}^{\prime}|}\ {\rm for}\ |\bm{x}-\bm{x}^{\prime}|\geq N^{\frac{8}{9}},$
where $\frac{\gamma}{2},\gamma]\ni\gamma_{N}\geq\gamma_{N_{1}}-N^{-\kappa}$ (cf. (2.35)).
We begin with a useful definition.
Definition 2.21 ($\sigma$-good).
For $\Lambda\in\mathcal{ER}(L)$, we say $\Lambda$ is $\sigma$-good if
$\|G_{\Lambda}(\sigma)\|$ $\leq e^{L^{\frac{3}{4}}},$
$|G_{\Lambda}(\sigma)(\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))|$ $\leq e^{-\gamma_{L}|\bm{x}-\bm{x}^{\prime}|}\ {\rm for}\ |\bm{x}-\bm{x}^{\prime}|\geq L^{\frac{8}{9}},$
where $\gamma_{L}\in[\frac{\gamma}{2},\gamma]$, and $\sigma$-bad, if one of the inequalities is violated.
The proof of Lemma 2.20 is based on matrix-valued Cartan’s lemma (Lemma 2.25) and iterations of resolvent identities. To apply Cartan’s lemma, one needs to establish the sublinear bound on $\sigma$-bad $\Lambda\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ contained in $\Lambda_{N}$: If $|\bm{n}|\leq 10N_{1}$, the sublinear bound can be obtained via the Töplitz property of $H$ in the $\bm{k}$-direction together with the semi-algebraic sets theory (cf. Lemma 2.22); for $|\bm{n}|>10N_{1}$, we will use the short-range property of $S$ and Bourgain’s geometric lemma (Lemma 2.5) to prove the sublinear bound. Once the sublinear bound is derived, one can use the Cartan’s lemma and resolvent identities to finish the proof of large scales LDE.
We need the following ingredients to prove Lemma 2.20.
Lemma 2.22.
There is some $0<c_{2}=c_{2}(b,d)\leq\frac{1}{10^{4}(b+d)^{2}}$ so that the following holds true. Assume that LDE hold true at scale $N_{1}$ with
$N_{1}\geq N_{0}(b,d,\rho,K,V).$
If $0<\rho\leq c_{2}$, then for all $\sigma\in\mathbb{R}$, the number of $\sigma$-bad $\Lambda\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ satisfying $|\bm{k}|\leq N^{2}$ and $|\bm{n}|\leq 10N_{1}$ is at most $N^{\frac{1}{10^{3}(b+d)^{2}}}.$
Proof.
First, we define $\widetilde{\Sigma}_{N_{1}}$ to be the set of $\sigma\in\mathbb{R}$ so that, there is $\Lambda\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ satifying $|\bm{n}|\leq 10N_{1}$, which is $\sigma$-bad. We then obtain using LDE at scale $N_{1},$
${\rm meas}(\widetilde{\Sigma}_{N_{1}})\leq C(b,d)N_{1}^{C(b,d)}e^{-N_{1}^{\rho}}<e^{-\frac{1}{2}N_{1}^{\rho}}$
for $N_{1}\geq N_{0}(b,d,\rho).$ Using the Hilbert-Schmidt norm and Cramer’s rule, we can identify $\widetilde{\Sigma}_{N_{1}}$ with a semi-algebraic set of $\deg\widetilde{\Sigma}_{N_{1}}\leq N_{1}^{C(b,d)}$. Then using Basu-Pollack-Roy Theorem [BPR96] on Betti numbers of a semi-algebraic set (cf. also Proposition 9.2 of [Bou05a]), we have a decomposition of
$\widetilde{\Sigma}_{N_{1}}\subset\bigcup_{1\leq\ell\leq N_{1}^{C(b,d)}}I_{\ell},$
where each $I_{\ell}$ is an interval of length $|I_{\ell}|\leq e^{-\frac{1}{2}N_{1}^{\rho}}.$ Now assume that $\Lambda\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ and $\Lambda^{\prime}\in(\bm{k}^{\prime},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$, which are all $\sigma$-bad for some $|\bm{n}|\leq 10N_{1},\bm{k}\neq\bm{k}^{\prime}$ and $|\bm{k}|,|\bm{k}^{\prime}|\leq N^{2}$. Then from the Töplitz property of ${H}(\sigma)$ in the $\bm{k}$-direction, we obtain
$\sigma+\bm{k}\cdot\bm{\omega},\ \sigma+\bm{k}^{\prime}\cdot\bm{\omega}\in\widetilde{\Sigma}_{N_{1}}.$
As a result, by $\bm{k}\neq\bm{k}^{\prime}$ with $|\bm{k}|,|\bm{k}^{\prime}|\leq N^{2}$, $\bm{\omega}\in{\rm DC}_{\bm{\omega}}(N)$ (cf. (2.12)) and $N\sim N_{1}^{\frac{1}{\rho^{2}}}$, we have for $N_{1}\geq N_{0}(b,d,\rho,K,V),$
${e^{-\frac{1}{2}N_{1}^{\rho}}}=e^{-\frac{1}{2}N^{\rho^{3}}}<e^{-N^{\rho^{4}}}\leq|(\bm{k}-\bm{k}^{\prime})\cdot\bm{\omega}|,$
which implies $\bm{k}$ and $\bm{k}^{\prime}$ can not stay in the same interval $I_{\ell}$ for $1\leq\ell\leq N_{1}^{C(b,d)}.$ Thus we have shown that the number of $\sigma$-bad $\Lambda\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ with $|\bm{n}|\leq 10N_{1}$ and $|\bm{k}|\leq N^{2}$ is at most
$N_{1}^{C(b,d)}\leq N^{\rho^{2}\cdot C(b,d)}\leq N^{\frac{1}{10^{3}(b+d)^{2}}},$
assuming $0<\rho\leq c_{2}(b,d)\leq\frac{1}{10^{4}(b+d)^{2}}.$ This proves Lemma 2.22. ∎
Next, we will deal with $|\bm{n}|>10N_{1}$ by combining short-range property of $S$ and Bourgain’s geometric lemma (cf. Lemma 2.5). For $\Lambda\subset\mathbb{Z}^{b+d}$, denote by $\Pi_{b}\Lambda$ (resp. $\Pi_{d}\Lambda$) the projection of $\Lambda$ on $\mathbb{Z}^{b}$ (resp. $\mathbb{Z}^{d}$). For each $\bm{k}\in\Pi_{b}\Lambda,$ define
$\Lambda(\bm{k})=\{\bm{n}\in\mathbb{Z}^{d}:\ (\bm{k},\bm{n})\in\Lambda\}.$
We have
Lemma 2.23.
Assume that $(\bm{\alpha},\bm{\theta})\in\mathcal{\mathcal{W}^{\prime}}$ (cf. Lemma 2.5) and $L\geq 10^{-\frac{32}{\kappa_{1}^{2}}}\log^{\frac{4}{\kappa_{1}}}(\log\frac{1}{\varepsilon})$. We have
  • (1)
    Fix $\Lambda\in(\bm{k}^{\prime},\bm{n}^{\prime})+\mathcal{ER}_{\bm{0}}(L)$ with $|\bm{n}^{\prime}|>10L$. There is a collection of $\{\lambda_{\ell}=\lambda_{\ell}(\bm{\alpha},\bm{\theta})\}_{1\leq\ell\leq L^{b+2d}}$ depending only on $\Lambda,\bm{\alpha},\bm{\theta},V$ so that, if
    $\min_{\bm{k}\in\Pi_{b}\Lambda,1\leq\ell\leq L^{b+2d},\xi=\pm 1}|\sigma+\bm{k}\cdot\bm{\omega}+\xi{\lambda_{\ell}}|>e^{-\frac{1}{4}L^{\frac{3\kappa_{1}}{4}}},$
    then $\Lambda$ is $\sigma$-good with $\gamma_{L}=\gamma$, where $\kappa_{1}$ is given in Lemma 2.5.
  • (2)
    Let $L\leq N\leq e^{\log^{2}L}$. There exists some $\tilde{\Sigma}_{L}\subset\mathbb{R}$ with ${\rm meas}(\tilde{\Sigma}_{L})\leq e^{-\frac{1}{5}L^{\frac{3\kappa_{1}}{4}}}$ so that, if $\sigma\notin\tilde{\Sigma}_{L}$, then each $\Lambda\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(L)$ satisfying $|(\bm{k},\bm{n})|\leq 100N^{2}$ and $|\bm{n}|>10L$ is $\sigma$-good with $\gamma_{L}=\gamma$.
Proof.
(1) Let $\bm{k}\in\Pi_{b}\Lambda$. It is important that either $\Lambda({\bm{k}})\in\mathcal{ER}_{\mathbb{Z}^{d}}(L),$ or $\Lambda({\bm{k}})$ is a rectangle of width at least $L$ (cf. Lemma B.8). In each of the two cases, we have $\Lambda({\bm{k}})\in\mathcal{E}_{2L}^{L}$.
For $Q\subset\Lambda({\bm{k}}),$ write
$A_{\bm{k},Q}=A_{\bm{k},Q}(\sigma)$ $=R_{Q}({D}(\sigma)+\varepsilon(\Delta\oplus\Delta))R_{Q}$
$=(-\sigma-\bm{k}\cdot\bm{\omega}+\mathcal{L}_{Q}(\bm{\theta}))\oplus(\sigma+\bm{k}\cdot\bm{\omega}+\mathcal{L}_{Q}(\bm{\theta})),$
where $\mathcal{L}_{Q}(\bm{\theta}):=\mathcal{H}_{Q}(\bm{\alpha},\bm{\theta})$ is defined in Lemma 2.5. Then the set of all eigenvalues of $A_{\bm{k},Q}$ is given by
$\{\pm(\sigma+\bm{k}\cdot\bm{\omega})+\lambda_{\ell,Q}(\bm{\alpha},\bm{\theta})\}_{\ell=1}^{\#Q},$
where $\{\lambda_{\ell}:=\lambda_{\ell,Q}(\bm{\alpha},\bm{\theta})\}_{\ell=1}^{\#Q}$ denotes the spectrum of $\mathcal{L}_{Q}(\bm{\theta}).$ So we have
$\|A_{\bm{k},Q}^{-1}\|$ $\leq\max_{\xi=\pm 1,1\leq\ell\leq\#Q}|\xi(\sigma+\bm{k}\cdot\bm{\omega})+\lambda_{\ell}|^{-1}.$
[…]
$Q_{\bm{m}}\subset\bm{m}+[-\tilde{L},\tilde{L}]^{d},\ \bm{m}+[-\tilde{L}^{\frac{1}{10d}},\tilde{L}^{\frac{1}{10d}}]^{d}\subset\tilde{Q}_{\bm{m}},$
$\bm{m}\in Q_{\bm{m}}\subset\Lambda(\bm{k}),{\rm dist}(\bm{m},\Lambda({\bm{k}})\setminus Q_{\bm{m}})\geq\frac{\tilde{L}}{2},\ {\rm diam}\ \tilde{Q}_{\bm{m}}\leq 4\tilde{L}^{\frac{1}{10d}}.$
Also for any $\bm{n}\in Q_{\bm{m}}\setminus\tilde{Q}_{\bm{m}}$, there exist $L_{1}\sim(\log L)^{\frac{4}{\kappa_{1}}}$ and some $\mathcal{ER}_{\mathbb{Z}^{d}}(L_{1})\ni W\subset Q_{\bm{m}}\setminus\tilde{Q}_{\bm{m}}$ such that
$\bm{n}\in W,\ {\rm dist}(\bm{n},Q_{\bm{m}}\setminus\tilde{Q}_{\bm{m}}\setminus W)\geq\frac{L_{1}}{2},$
and
$\|\mathcal{H}^{-1}_{W}(E_{\bm{k}};\bm{\theta})\|$ $\leq e^{\sqrt{L_{1}}},$
$|\mathcal{H}^{-1}_{W}(E_{\bm{k}};\bm{\theta})(\bm{n};\bm{n}^{\prime})|$ $\leq e^{-\frac{1}{2}|\log\varepsilon|\cdot|\bm{n}-\bm{n}^{\prime}|}\ {\rm for}\ |\bm{n}-\bm{n}^{\prime}|\geq{L_{1}^{\frac{8}{9}}}.$
Next, assume that (set $Q=Q_{\bm{m}}$)
$\min_{\bm{m}\in\Lambda(\bm{k}),\xi=\pm 1,1\leq\ell\leq\#Q_{\bm{m}}}|\sigma+\bm{k}\cdot\bm{\omega}+\xi{\lambda_{\ell}}|>e^{-\frac{1}{4}\tilde{L}^{\frac{3}{4}}}.$
Then applying Lemma B.6 of [JLS20] yields for
$A_{\bm{k}}:=A_{\bm{k}}(\sigma)=A_{\bm{k},\Lambda(\bm{k})}(\sigma),$
the following estimates
$\|A^{-1}_{\bm{k}}\|$ $\leq e^{\frac{1}{2}L^{\frac{3}{4}}},$
$|A_{\bm{k}}^{-1}(\bm{n};\bm{n}^{\prime})|$ $\leq e^{-\frac{1}{4}|\log\varepsilon|\cdot|\bm{n}-\bm{n}^{\prime}|}\ {\rm for}\ |\bm{n}-\bm{n}^{\prime}|\geq{L}^{\frac{8}{9}}.$
Note that
$\tilde{G}_{\Lambda}(\sigma)=\bigoplus_{\bm{k}\in\Pi_{b}\Lambda}A_{\bm{k}}^{-1},$
where
$\tilde{G}_{\Lambda}(\sigma)=\left(R_{\Lambda}({D}(\sigma)+\varepsilon(\Delta\oplus\Delta))R_{\Lambda}\right)^{-1}.$
By taking account of all $\bm{k}\in\Pi_{b}\Lambda,$ we obtain that if
$\min_{\bm{k}\in\Pi_{b}\Lambda,\bm{m}\in\Lambda(\bm{k}),\xi=\pm 1,1\leq\ell\leq\#Q_{\bm{m}}}|\sigma+\bm{k}\cdot\bm{\omega}+\xi{\lambda_{\ell}}|>e^{-\frac{1}{4}L^{\frac{3\kappa_{1}}{4}}},$ (2.29)
then
$\|\tilde{G}_{\Lambda}(\sigma)\|$ $\leq e^{\frac{1}{2}L^{\frac{3}{4}}},$
$|\tilde{G}_{\Lambda}(\sigma)((\bm{k},\bm{n},\xi);(\bm{k}^{\prime},\bm{n}^{\prime},\xi^{\prime}))|$ $\leq\delta_{\bm{k},\bm{k}^{\prime}}e^{-\frac{1}{4}|\log\varepsilon|\cdot|\bm{n}-\bm{n}^{\prime}|}\ {\rm for}\ |\bm{n}-\bm{n}^{\prime}|\geq L^{\frac{8}{9}}.$
Note that the total number of $\lambda_{\ell}$ in (2.29) is at most
$C(b,d)L^{b}L^{d}L^{\kappa_{2}}L^{\kappa_{2}d}\ll L^{b+2d}.$
Finally, since $|\bm{n}|>10L$, we know that $S$ has the decay estimate
$|S((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))|\leq C_{2}(1+|\bm{x}-\bm{x}^{\prime}|)^{C_{2}}e^{-6\gamma L}e^{-2\gamma|\bm{x}-\bm{x}^{\prime}|}.$
Then using the perturbation Lemma B.1 implies that $\Lambda$ is $\sigma$-good assuming (2.29) holds true. In fact, we can apply Lemma B.1 (we hide the dependence on $\xi,\xi^{\prime}$) with $B=\delta S_{\Lambda}(\bm{x},\bm{x}^{\prime})$ , $A^{-1}(\bm{x},\bm{x}^{\prime})=\tilde{G}_{\Lambda}(\bm{x},\bm{x}^{\prime})$ and $c=2\gamma,\epsilon_{1}=e^{-\frac{1}{2}L^{\frac{3}{4}}},M=L^{\frac{8}{9}},\epsilon_{2}=C\delta e^{-6\gamma L}$. Since $\gamma\in(\frac{1}{2},10)$, we obtain for $L\geq N_{0}(b,d,C_{2})\gg 1,$
$C(b,d,C_{2})L^{2(b+d)+C_{2}}\delta e^{-6\gamma L}e^{2\gamma L^{\frac{8}{9}}}e^{\frac{1}{2}L^{\frac{3}{4}}}\leq\frac{1}{2}.$
This verifies all assumptions of Lemma B.1. So using Lemma B.1, we get
$\|G_{\Lambda}(\sigma)\|$ $\leq 2e^{\frac{1}{2}L^{\frac{3}{4}}}<e^{L^{\frac{3}{4}}},$
and for $|\bm{x}-\bm{x}^{\prime}|\geq L^{\frac{8}{9}}$ (since $\frac{1}{4}|\log\varepsilon|>2\gamma>1$),
$|G_{\Lambda}(\sigma)((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))|$ $\leq|\tilde{G}_{\Lambda}(\sigma)((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))|+e^{\frac{1}{2}L^{\frac{3}{4}}}e^{-2\gamma|\bm{x}-\bm{x}^{\prime}|}$
$\leq e^{L^{\frac{3}{4}}-L^{\frac{8}{9}}}e^{-\gamma|\bm{x}-\bm{x}^{\prime}|}$
$\leq e^{-\gamma|\bm{x}-\bm{x}^{\prime}|}.$
This proves (1) of Lemma 2.23.
Remark 2.24.
  • Let $L\leq N\leq e^{(\log L)^{2}}$. We would also like to remark that if we take into account of all these $\Lambda\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(L)$ with $|(\bm{k},\bm{n})|\leq 100N^{2}$ and $|\bm{n}|>10L$ (the total number of these $\Lambda$ is at most $C(b,d)N^{2(b+d)}$), then we get a sequence of real numbers $\{{\lambda_{\ell}(\bm{\alpha},\bm{\theta})}\}_{1\leq\ell\leq N^{4(b+d)}}$. As a result, there is a set $\tilde{\Sigma}_{L}\subset\mathbb{R}$ satisfying
    ${\rm meas}(\tilde{\Sigma}_{L})\leq e^{C(b,d)\log^{2}L}e^{-\frac{1}{4}L^{\frac{3}{4}\kappa_{1}}}\leq e^{-\frac{1}{5}L^{\frac{3}{4}\kappa_{1}}}$
    so that, for $\sigma\not\in\tilde{\Sigma}_{L}$, all $\Lambda\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(L)$ satisfying $|(\bm{k},\bm{n})|\leq 100N^{2}$ and $|\bm{n}|>10L$ are $\sigma$-good.
  • In the application, we typically choose $L=N_{1}\geq 10^{-\frac{32}{\kappa_{1}^{2}}}\log^{\frac{4}{\kappa_{1}}}(\log\frac{1}{\varepsilon})$. It then follows from $N_{1}\sim N^{\rho^{2}}\geq(\varepsilon+\delta)^{-c_{1}\rho^{2}}$ that a restriction like $\delta\leq\log^{-1}\frac{1}{\varepsilon}$ is needed.
(2) The conclusion is just that in Remark 2.24.
This completes the proof of Lemma 2.23.∎
By combining Lemma 2.22 and Lemma 2.23, we are ready to prove LDE at the scale $N$ (indeed all scales in $N,N^{2}]$). We first recall an important lemma.
Lemma 2.25 (Matrix-valued Cartan’s lemma, cf. Proposition 14.1 in [Bou05a]).
Let $T(\sigma)$ be a self-adjoint $N\times N$ matrix-valued function of a parameter $\sigma\in[-\beta,\beta]$ satisfying the following conditions:
  • (i)
    $T(\sigma)$ is real analytic in $\sigma$ and has a holomorphic extension to
    $\mathbb{D}=\left\{z\in\mathbb{C}:\ |\Re z|\leq\beta,\ |\Im z|\leq\beta_{1}\right\}$
    satisfying $\sup\limits_{z\in\mathbb{D}}\|T(z)\|\leq K_{1},\ K_{1}\geq 1.$
  • (ii)
    For each $\sigma\in[-\beta,\beta]$, there is a subset $Y\subset[1,N]$ with $Y\leq M$ such that
    $\|(R_{[1,N]\setminus Y}T(\sigma)R_{[1,N]\setminus Y})^{-1}\|\leq K_{2},\ K_{2}\geq 1.$
  • (iii)
    Assume
    $\mathrm{meas}\left(\{\sigma\in[-{\beta},{\beta}]:\ \|T^{-1}(\sigma)\|\geq K_{3}\}\right)\leq 10^{-3}\beta_{1}(1+K_{1})^{-1}(1+K_{2})^{-1}.$
Let $0<\epsilon\leq(1+K_{1}+K_{2})^{-10M}.$ Then we have
$\mathrm{meas}\left(\left\{\sigma\in\left[-{\beta}/{2},{\beta}/{2}\right]:\ \|T^{-1}(\sigma)\|\geq\frac{1}{\epsilon}\right\}\right)\leq Ce^{\frac{-c\log\frac{1}{\epsilon}}{M\log(M+K_{1}+K_{2}+K_{3})}},$ (2.30)
where $C,c>0$ are some absolute constants.
Now, we can prove Lemma 2.20.
Proof of Lemma 2.20.
Recalling
$N_{2}=N_{1}^{\frac{2}{\rho}},\ N=N_{1}^{\frac{1}{\rho^{2}}},$
we assume that for $L=N_{1},N_{2}$ and $\Lambda\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(L)$ with $|\bm{n}|\leq 10L$, $G_{\Lambda}(\sigma)$ satisfies the LDE.
Without loss of generality, we only establish LDE for the Green’s function on $\Lambda_{{\rm pm},N}=\{(\bm{x},\xi)\in\mathbb{Z}_{{\rm pm},*}^{b+d}:\ |\bm{x}|\leq N\}=(\Lambda_{N}\times\{+,-\})\cap\mathbb{Z}_{{\rm pm},*}^{b+d}$, the other cases (i.e., $\Lambda\in(\bm{0},\bm{n})+{\mathcal{ER}_{\bm{0}}(N_{3})}$ for $|\bm{n}|\leq 10N_{3},N_{3}\in[N,N^{2}]$) can be dealt with similarly. We define three intermediate scales between $N_{2},N]$:
$L_{1}=N^{\frac{1}{100(b+d)^{2}}},\ L_{2}=N^{\frac{1}{50(b+d)^{2}}},\ L_{3}=N^{\frac{1}{10(b+d)}}.$ (2.31)
Note that from $0<\rho\leq c_{2}(b,d)\leq{\frac{1}{10^{4}(b+d)^{2}}}$ (cf. Lemma 2.22), we have
$\log N_{1}\ll\log N_{2}\ll\log L_{1}=\frac{1}{100(b+d)^{2}}\log N.$
We first cover the $\Lambda_{{\rm pm},N}$ with two overlapped subregions, $\Lambda_{{\rm pm},N}=\mathcal{R}_{1}\cup\mathcal{R}_{2}$ satisfying
$\mathcal{R}_{1}=\{(\bm{k},\bm{n},\xi)\in\Lambda_{{\rm pm},N}:\ |\bm{n}|\leq L_{2}\},$
$\mathcal{R}_{2}=\{(\bm{k},\bm{n},\xi)\in\Lambda_{{\rm pm},N}:\ |\bm{n}|\geq L_{1}\}.$
We will apply matrix-valued Cartan’s lemma (cf. Lemma 2.25) and Bourgain’s geometric lemma [Bou07] (cf. Lemma 2.5) to obtain the sub-exponential growth estimates on Green’s functions. Then we apply the coupling lemmas of [BGS02, JLS20, Liu22] (cf. e.g., Lemmas B.3, B.5, B.6) based on the iteration of resolvent identities to prove the off-diagonal exponential decay of $G_{\Lambda_{{\rm pm},N}}(\sigma)$ if desired $\sigma$ are removed.
Before we enter the details, let us first explain why we need to divide $\Lambda_{{\rm pm},N}$ into the above two regions. This is quite different from that in [SW23]. Based on Lemma 2.23, we can impose the so called weak second Melnikov’s condition on $\bm{\omega}$ so that all $\sigma$-bad (cf. Definition 2.21) $\Lambda\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})\subset\Lambda_{{\rm pm},N}$ can stay only in a narrow strip region (cf. Figure 1), namely, there is some $\bm{k}_{*}\in[-N,N]^{b}$ so that
$|\bm{k}-\bm{k}_{*}|\leq 10N_{1}\leq N^{0+}.$
This gives the desired control on $\sigma$-bad $\Lambda$ in the $\bm{k}$-direction. To establish the sublinear bound, it remains to handle the $\bm{n}$-direction. In [SW23], the $\bm{n}$-direction estimates can be obtained directly using the fact that $\theta\in\mathbb{R}$, Diophantine $\bm{\omega}$ and LDT for the corresponding quasi-periodic Schrödinger operators. Such arguments inevitably fail for $\bm{\theta}\in\mathbb{R}^{d}$: There is simply no one-dimensional interval decomposition of $d$-dimensional semi-algebraic sets, and more essentially, the Diophantine property of $\bm{\omega}$ cannot ensure the sublinear estimate in the $\bm{n}$-direction. This motivates us to restrict $|\bm{n}|\leq L_{2}$ (i.e., the region $\mathcal{R}_{1}$). We can use the Cartan’s lemma and resolvent identities to deal with $\mathcal{R}_{1}$. For $\mathcal{R}_{2}$, we take advantage of Bourgain’s geometric lemma and the short-range property of $H$ using Lemma 2.23. Finally, it suffices to cover $\Lambda_{{\rm pm},N}$ with four types of regions of sizes $N_{1},L_{1}^{\frac{9}{10}},L_{1},L_{3}$ (will be specified below) contained in $\mathcal{R}_{1}$, $\mathcal{R}_{2}$, and apply the resolvent identities again. See Figure 1.
We first deal with $\mathcal{R}_{2}$, and then $\mathcal{R}_{1}$.
Analysis in $\mathcal{R}_{2}$
In this region, we can apply directly the conclusion of Lemma 2.23. In this case, we make no restriction on $\bm{\omega}.$ We let
$L_{0}=L_{1}^{\frac{9}{10}}=N^{\frac{9}{1000(b+d)^{2}}}.$
For any $\Lambda\subset\mathcal{R}_{2}$ satisfying $\Lambda=(\bm{k},\bm{n})+\mathcal{ER}(L_{0})$, we have $|\bm{n}|>L_{1}>10L_{0}$. Then applying Lemma 2.23 with $L=L_{0}$ and since $0<\rho<c_{2}(b,d)\kappa_{1}\leq\frac{\kappa_{1}}{10^{4}(b+d)^{2}},$ we get a set $\tilde{\Sigma}_{N,1}\subset\mathbb{R}$ satisfying
$\mathrm{meas}(\tilde{\Sigma}_{N,1})\leq e^{-\frac{1}{5}L_{0}^{\frac{3\kappa_{1}}{4}}}\leq e^{-\frac{1}{5}N^{\frac{27\kappa_{1}}{4000(b+d)^{2}}}}<e^{-N^{3\rho}},$
so that, if $\sigma\notin\tilde{\Sigma}_{N,1}$, all $\Lambda\in\mathcal{ER}(L_{0})$ contained in $\mathcal{R}_{2}$ is $\sigma$-good (cf. Definition 2.21).
We remark that in this case, we do not use Cartan’s lemma.
Analysis in $\mathcal{R}_{1}$
In this region, we need to make further restriction on $\bm{\omega}$ called the weak second Melnikov’s condition. More precisely, we have
Lemma 2.26.
There is a sequence of real numbers $\{\lambda_{\ell}=\lambda_{\ell}(\bm{\alpha},\bm{\theta})\}_{1\leq\ell\leq N^{4(b+d)}}$ depending only on $V,\bm{\alpha},\bm{\theta}$ so that the following holds true. If $\bm{\omega}\in\widetilde{\Omega}_{N}$ with
$\ \ \ \widetilde{\Omega}_{N}$
$:=\bigcap_{1\leq\ell,\ell^{\prime}\leq N^{4(b+d)},\xi=\pm 1,\xi^{\prime}=\pm 1,0<|\bm{k}|\leq 2N^{2}}\left\{\bm{\omega}\in\Omega:\ |\bm{k}\cdot\bm{\omega}+\xi{\lambda_{\ell}}-\xi^{\prime}{\lambda_{\ell^{\prime}}}|>10e^{-\frac{1}{4}N^{\frac{3\kappa_{1}\rho^{2}}{4}}}\right\},$ (2.32)
then for any $\sigma\in\mathbb{R},$ there is some $\bm{k}_{*}\in[-N^{2},N^{2}]^{b}$ so that, all $\sigma$-bad $\Lambda\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ satisfying $|\bm{k}|\leq N^{2}$ and $10N_{1}\leq|\bm{n}|\leq 100N^{2}$ must obey
$\Pi_{b}\Lambda\subset[\bm{k}_{*}-10N_{1},\bm{k}_{*}+10N_{1}]^{b}.$ (2.33)
In particular, we have
${\rm meas}(\Omega\setminus\widetilde{\Omega}_{N})\leq e^{-\frac{1}{5}N^{\frac{3\kappa_{1}\rho^{2}}{4}}}.$ (2.34)
Remark 2.27.
Indeed, in this lemma, $\lambda_{\ell}(\bm{\alpha},\bm{\theta})$ (${1\leq\ell\leq N^{4(b+d)}})$ are eigenvalues of the operator $\mathcal{H}_{Q}(\bm{\alpha},\bm{\theta})$ for all $Q\subset[-100N^{2},100N^{2}]^{d}$ satisfying $Q\in\mathcal{ER}_{\mathbb{Z}^{d}}(\tilde{N}_{1})$ for some $\tilde{N}_{1}\in[\frac{1}{4}N_{1}^{\kappa_{1}},N_{1}^{\kappa_{2}}]$.
Proof of Lemma 2.26.
The proof follows from Lemma 2.23. Applying (1) of Lemma 2.23 (cf. also Remark 2.24) with $L=N_{1}$ and taking into account of all $\Lambda\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ with $|(\bm{k},\bm{n})|\leq 100N^{2}$ and $|\bm{n}|>10N_{1}$ yield a sequence $\{{\lambda_{\ell}}\}_{\leq\ell\leq N^{4(b+d)}}\subset\mathbb{R}$ and each $\lambda_{\ell}$ depending only on $\bm{\alpha},\bm{\theta}$ (but not on $\sigma,\bm{\omega}$), so that if $\Lambda=(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ satisfies $|(\bm{k},\bm{n})|\leq 100N^{2}$, $|\bm{n}|>10N_{1}$ and
$\min_{\bm{k}\in\Pi_{b}\Lambda,\xi=\pm 1,1\leq\ell\leq N^{4(b+d)}}|\sigma+\bm{k}\cdot\bm{\omega}+\xi{\lambda_{\ell}}|>e^{-\frac{1}{4}N_{1}^{\frac{3\kappa_{1}}{4}}}=e^{-\frac{1}{4}N^{\frac{3\kappa_{1}\rho^{2}}{4}}},$
then $\Lambda$ is $\sigma$-good. So recalling (2.32), we have
${\rm meas}(\Omega\setminus\widetilde{\Omega}_{N})\leq N^{C(b,d)}e^{-\frac{1}{4}N^{\frac{3\kappa_{1}\rho^{2}}{4}}}\leq e^{-\frac{1}{5}N^{\frac{3\kappa_{1}\rho^{2}}{4}}},$
which implies (2.34).
In the following, we assume $\bm{\omega}\in\widetilde{\Omega}_{N}.$
Suppose now $\Lambda^{\prime}$ with $\Lambda^{\prime}\in(\bm{k}^{\prime},\bm{n}^{\prime})+\mathcal{ER}_{\bm{0}}(N_{1})$ satisfying $|\bm{k}^{\prime}|\leq N^{2}$ and $10N_{1}<|\bm{n}^{\prime}|\leq 100N^{2}$ is $\sigma$-bad. Then there are some $\bm{k}_{*}\in\Pi_{b}\Lambda^{\prime}\subset[-N^{2},N^{2}]^{b}$, $\xi_{*}=\pm 1$ and some $\ell_{*}\in[1,N^{4(b+d)}]$ so that
$|\sigma+\bm{k}_{*}\cdot\bm{\omega}+\xi_{*}{\lambda_{\ell_{*}}}|\leq e^{-\frac{1}{4}N_{1}^{\frac{3\kappa_{1}}{4}}}.$
Now let $\Lambda^{\prime\prime}$ (satisfying $\Lambda^{\prime\prime}\in(\bm{k}^{\prime\prime},\bm{n}^{\prime\prime})+\mathcal{ER}_{\bm{0}}(N_{1})$ with $|\bm{k}^{\prime\prime}|\leq N^{2}$ and $10N_{1}<|\bm{n}^{\prime\prime}|\leq 100N^{2}$) be another $\sigma$-bad region. From $\bm{\omega}\in\widetilde{\Omega}_{N},$ we must have
$\bm{k}_{*}\in\Pi_{b}\Lambda^{\prime\prime}.$
Otherwise, there must be some $\bm{k}^{\prime\prime\prime}\in\Pi_{b}\Lambda^{\prime\prime},\xi^{\prime}\in\{\pm 1\},\ell^{\prime}$ so that, $\bm{k}_{*}\neq\bm{k}^{\prime\prime\prime}$ and
$|\sigma+\bm{k}^{\prime\prime\prime}\cdot\bm{\omega}+\xi^{\prime}{\lambda_{\ell^{\prime}}}|\leq e^{-\frac{1}{4}N_{1}^{\frac{3\kappa_{1}}{4}}},$
which contradicts $\bm{\omega}\in\widetilde{\Omega}_{N}$. We have established that all $\sigma$-bad elementary $N_{1}$-regions $\Lambda$ with centers $(\bm{k},\bm{n})$ satisfying $|\bm{k}|\leq N^{2}$ and $10N_{1}<|\bm{n}|\leq 100N^{2}$ must satisfy (2.33).
In the following, we always assume $\bm{\omega}\in\widetilde{\Omega}_{N}$. Under this condition, we first give a geometric construction of the a region $\Lambda^{\dagger}$, on which the worst resonances appear. It turns out that this region has the annulus structure similar to [Bou07, JLS20].
Lemma 2.28.
Let $\bm{\omega}\in\widetilde{\Omega}_{N}$ and $\sigma\in\mathbb{R}$. Let $\bm{k}_{*}\in[-N,N]^{b}$ (if it exists) be defined by Lemma 2.26. Then there are some $\Lambda^{\dagger}_{1}\subset\Lambda^{\dagger}\subset\mathcal{R}_{1}$ satisfying ${\rm diam}\ \Lambda_{1}^{\dagger}\leq 4L_{1},\bm{k}_{*}\in\Lambda_{1}^{\dagger}$ and
$\Lambda^{\dagger}$ $=((\bm{k}_{*}+[-L_{3}+O(L_{1}),L_{3}+O(L_{1})]^{b})\times[-L_{2},L_{2}]^{d})\cap\mathcal{R}_{1},$
so that the following holds true (recalling $L_{1}\ll L_{2}\ll L_{3}$ given by (2.31)). For any $\bm{x}\in\Lambda^{\dagger}\setminus\Lambda^{\dagger}_{1}$, there are some $L\in\{N_{1},L_{1},L_{0}\}$ and some $\mathcal{ER}(L)\ni W\subset\Lambda^{\dagger}\setminus\Lambda^{\dagger}_{1}$ so that
$\bm{x}\in W,\ {\rm dist}(\bm{x},\Lambda^{\dagger}\setminus\Lambda^{\dagger}_{1}\setminus W)\geq\frac{L}{2}.$
Proof.
The proof is similar to that of Theorem 3.7 of [JLS20] (cf. pages 471–472). We initially set $Q^{\dagger}=(\bm{k}_{*}+[-L_{3},L_{3}]^{b})\times[-L_{2},L_{2}]^{d}$ and $Q_{1}=(\bm{k}_{*}+[-L_{1},L_{1}]^{b})\times[-L_{1},L_{1}]^{d}$. Then we slightly change $Q^{\dagger},Q_{1}^{\dagger}$ to $\Lambda^{\dagger},\Lambda_{1}^{\dagger}$ with the diameter modulations of length $2L_{1}$ if the boundary of $Q^{\dagger},Q_{1}^{\dagger}$ approaches to the boundary of $\mathcal{R}_{1}$ with the distance of $0<L\leq 2L_{1}$.
Next, let $\bm{x}=(\bm{k},\bm{n})\in\Lambda^{\dagger}\setminus\Lambda_{1}^{\dagger}$. For $|\bm{n}|\leq 10N_{1}$, we use $L_{1}$ size regions to cover $\bm{x}$; for $10N_{1}<|\bm{n}|\leq L_{1}$, we use the $N_{1}$ size regions to cover $\bm{x}$; for $L_{1}<|\bm{n}|\leq L_{2}$, we use the $L_{0}$ size regions to cover $\bm{x}.$
Next, we show that all $Q\in(\bm{k},\bm{0})+\mathcal{ER}_{\bm{0}}(L_{1})$ satisfying $|\bm{k}|\leq N$ and $|\bm{k}-\bm{k}_{*}|>10N_{1}$ are $\sigma$-good for “most” $\sigma$, by using Cartan’s lemma and coupling lemma of [Liu22]. More precisely, we have
Lemma 2.29.
Let $\bm{\omega}\in\widetilde{\Omega}_{N}\cap\Omega_{N_{2}}$. There is some $\tilde{\Sigma}_{N,2}\subset\mathbb{R}$ satisfying
${\rm meas}(\tilde{\Sigma}_{N,2})\leq e^{-N^{3\rho}},$
so that for $\sigma\notin\tilde{\Sigma}_{N,2}$, the following holds true. If $Q\subset\mathcal{R}_{1}$ satisfies
$Q\in\mathcal{Q}:=\bigcup_{|\bm{k}-\bm{k}_{*}|>10N_{1},|\bm{k}|\leq N}\left\{(\bm{k},\bm{0})+Q:\ Q\in\mathcal{ER}_{\bm{0}}(L_{1})\right\},$
then $Q$ is $\sigma$-good with $\gamma_{L_{1}}=\gamma_{N_{1}}-N_{1}^{-\vartheta}$, where $\vartheta>0$ is an absolute constant defined in Lemma B.3.
Proof of Lemma 2.29.
We choose any $Q\in\mathcal{ER}(L_{1})$ with $Q\in\mathcal{Q}$. Then we take any $Q^{\prime}\in\mathcal{R}_{L_{*}}^{{\sqrt{L_{1}}}}$ satisfying $Q^{\prime}\subset Q$ and $\sqrt{{L_{1}}}\leq L_{*}\leq 2L_{1}$. From Lemma 2.22 and $\bm{\omega}\in\widetilde{\Omega}_{N}$, we have for any $\sigma\in\mathbb{R},$
$\{\Lambda\in\mathcal{ER}(N_{1}):\ \Lambda\ {\rm is}\ \sigma{-\rm bad},\ \Lambda\subset Q^{\prime}\}\leq N^{\frac{1}{10^{3}(b+d)^{2}}}.$
Denote by $Y_{1}(\sigma):=\bigcup\limits_{1\leq\ell\leq k}Q_{\bm{x}_{\ell}},k\leq N^{\frac{1}{10^{3}(b+d)^{2}}}$ all those $\sigma$-bad $N_{1}$ size elementary regions. We claim that, there is some $Y=Y(\sigma)\subset Q^{\prime}$ satisfying $Y\leq C(b,d)N_{1}^{b+d}N^{\frac{1}{10^{3}(b+d)^{2}}}$, so that if $\bm{x}\in Q^{\prime}\setminus Y$, then there is some $\sigma$-good $W\in\mathcal{ER}(N_{1}),W\subset Q^{\prime}\setminus Y$ satisfying
$\bm{x}\in W,\ {\rm dist}(\bm{x},Q^{\prime}\setminus Y\setminus W)\geq\frac{N_{1}}{2}.$
Indeed, we pave $Q^{\prime}$ with $\Lambda_{\alpha}$ for $\Lambda_{\alpha}=Q^{\prime}\cap\Lambda_{10N_{1}}(\bm{x}),\bm{x}\in 10N_{1}\mathbb{Z}^{b+d}$. Then we can enlarge $\Lambda_{\alpha}$ to $\tilde{\Lambda}_{\alpha}\subset Q^{\prime}$ so that $\tilde{\Lambda}_{\alpha}$ has width at least $N_{1}$ and diameter at most $C(b,d)N_{1}$, and $\bigcup\limits_{\alpha}\tilde{\Lambda}_{\alpha}$ remains a tiling of $Q^{\prime}$. It suffices to take
$Y(\sigma)=\bigcup\limits_{\alpha:\ \tilde{\Lambda}_{\alpha}\cap Y_{1}(\sigma)\neq\emptyset}\tilde{\Lambda}_{\alpha}.$
Since each $Q_{\bm{x}_{\ell}}$ can only intersect with at most $C(b,d)$ many $\tilde{\Lambda}_{\alpha}$, we have
$Y(\sigma)\leq C(b,d)N_{1}^{b+d}N^{\frac{1}{10^{3}(b+d)^{2}}}.$
We are ready to apply Cartan’s lemma (cf. Lemma 2.25). From Lemma B.1, any $\sigma$-good $N_{1}$ region remains essentially $\sigma^{\prime}$-good if $|\sigma^{\prime}-\sigma|\leq e^{-10\gamma N_{1}}$. So on an interval $I$ of length at most $e^{-10\gamma N_{1}}$, we can fix $Y$ to associate with the midpoint of $I$. In addition, we can assume $|\sigma|\leq C(b,d)N$ since if $|\sigma|>CN$, then $\|D_{Q^{\prime}}^{-1}(\sigma)\|\leq 1$. In this case, $Q^{\prime}$ becomes $\sigma$-good via the Neumann series argument. We apply Lemma 2.25 with
$T(\sigma)={{H}}_{Q^{\prime}}(\sigma),\ \beta=\beta_{1}=e^{-10\gamma N_{1}}.$
It remains to verify the assumptions of Lemma 2.25. First, one has by the definition of the operator $H(\sigma),$
$K_{1}=CN,C=C(V,b,d,C_{2})>0.$
We let $Y$ be defined in the above claim. Then using sublinear bound conclusion and Lemma B.5, we have
$M=\#Y\leq C(b,d)N^{\rho^{2}(b+d)}N^{\frac{1}{10^{3}(b+d)^{2}}}<N^{\frac{1}{999(b+d)^{2}}}\leq L_{*}^{\frac{1}{4}}$
$({\rm since}\ \rho\leq c_{2}\leq\frac{1}{10^{4}(b+d)^{2}},\ L_{*}\geq N^{\frac{1}{200(b+d)^{2}}}),$
$K_{2}=e^{2{N_{1}}^{\frac{3}{4}}}.$
To apply the matrix-valued Cartan’s lemma, we need the scale
$N_{2}=N_{1}^{\frac{2}{\rho}}.$
Recall that the LDE hold at scale $N_{2}$. Applying the resolvent identity Lemma B.5 and using (2) of Lemma 2.23 with $L=N_{2}$ yield
$\|T^{-1}(\sigma)\|\leq e^{2{N_{2}}^{\frac{3}{4}}}=K_{3}$
for $\sigma$ away from a set of measure at most (since $0<\rho<c_{2}\kappa_{1}$)
$CN^{C(b,d)}e^{-{N_{2}^{\rho}}}+e^{-\frac{1}{5}{N_{2}^{\frac{3\kappa_{1}}{4}}}}\leq e^{-{N_{2}^{\rho}}/{2}}.$
It follows from $N_{1}=N_{2}^{\frac{\rho}{2}}$ that
$10^{-3}\beta_{1}(1+K_{1})^{-1}(1+K_{2})^{-1}$ $\geq e^{-10\gamma N_{1}}N^{-2}e^{-2N_{1}^{\frac{3}{4}}}$
$>e^{-{N_{2}^{\rho}}/{2}}.$
This verifies (iii) of Lemma 2.25. If $\epsilon=e^{-{L_{*}}^{\frac{1}{2}}}$, then one has $\epsilon<(1+K_{1}+K_{2})^{-10M}.$ Cover [CN,CN][-CN,CN] with disjoint intervals of length $e^{-10\gamma N_{1}}$. Define
$\tilde{\Sigma}_{Q^{\prime}}=\{\sigma\in\mathbb{R}:\ \|G_{Q^{\prime}}(\sigma)\|\geq e^{L_{*}^{\frac{3}{4}}}\}.$
Then by (2.30) of Lemma 2.25, one obtains
$\mathrm{meas}(\tilde{\Sigma}_{Q^{\prime}})\leq CN^{C}e^{10\gamma N_{1}}e^{-\frac{c{{L_{*}}^{\frac{3}{4}}}}{CN_{1}N_{2}L_{*}^{\frac{1}{4}}\log N}}\leq e^{-{L_{*}}^{\frac{1}{3}}}\ll e^{-N^{3\rho}}.$
Note that the number of $Q^{\prime}\subset Q$ with $Q^{\prime}\in\mathcal{R}_{L_{*}}^{\sqrt{L_{1}}},\sqrt{{L_{1}}}\leq L_{*}\leq 2L_{1}$ is at most $L^{C(b,d)}$ and the number of $Q\in\mathcal{Q}$ is at most $N^{(b,d)}$. We take the union on all these $Q^{\prime}\subset Q,Q\in\mathcal{Q}$ leading to the desired $\tilde{\Sigma}_{N,2}\subset\mathbb{R}$ with (since $0<\rho<c_{2}\leq\frac{1}{10^{4}(b+d)^{2}}$)
${\rm meas}(\tilde{\Sigma}_{N,2})\leq N^{C(b,d)}e^{-L_{1}^{\frac{1}{6}}}<e^{-N^{\frac{1}{600(b+d)^{2}}}}\ll e^{-N^{3\rho}}.$
To finish the proof, it suffices to prove the off-diagonal exponential decay of $G_{Q}(\sigma)$ for $Q\in\mathcal{Q}$ and $\sigma\notin\tilde{\Sigma}_{N,2}$. This follows from using the coupling lemma of [Liu22]. Indeed, fix $Q\in\mathcal{Q}$ (then $Q\in\mathcal{ER}(L_{1})$). Let $\mathcal{F}$ be any family of pairwise disjoint $\sqrt{L_{1}}$ size elementary regions in $Q$. Then using Lemma 2.22 and $\bm{\omega}\in\widetilde{\Omega}_{N},$ we get
$\{\Lambda\in\mathcal{F}:\ \Lambda\ {\rm is}\ \sigma{-\rm bad}\}\leq N^{\frac{1}{10^{3}(b+d)^{2}}}\leq L_{1}^{\frac{1}{9}}.$
So applying Lemma B.3 leads to the off-diagonal exponential decay of $G_{Q}(\sigma)$ with the decay rate
$\gamma_{L_{1}}=\gamma_{N_{1}}-N_{1}^{-\vartheta}.$
We then deal with $G_{\Lambda^{\dagger}}(\sigma)$ with $\Lambda^{\dagger}$ given by Lemma 2.28. We have
Lemma 2.30.
Let $\bm{\omega}\in\widetilde{\Omega}_{N}\cap\Omega_{N_{2}}$. Then there is some $\widetilde{\Sigma}_{N,3}\subset\mathbb{R}$ satisfying
${\rm meas}(\tilde{\Sigma}_{N,3})<e^{-N^{3\rho}}$
so that, if $\sigma\notin\bigcup\limits_{\ell=1,2,3}\tilde{\Sigma}_{N,\ell},$ then $G_{\Lambda^{\dagger}}(\sigma)$ is $\sigma$-good in the sense of Definition 2.21 (replacing $\Lambda$ with $\Lambda^{\dagger}$) with $\gamma_{L_{2}}=\gamma_{L_{1}}-N_{1}^{-\frac{1}{10}}$.
Proof of Lemma 2.30.
The proof is similar to that of Lemma 2.29, but is more in the spirit of [JLS20] when dealing with off-diagonal exponential decay estimate, as $\Lambda^{\dagger}$ has the special annulus structure.
To get the sub-exponential growth estimate of $\|G_{\Lambda^{\dagger}}(\sigma)\|$, the key point is to apply Cartan’s lemma only involving $N_{1},N_{2}$ sizes regions similar to the proof of Lemma 2.29. This means precisely that we do not choose $Y=\Lambda_{1}^{\dagger}$ when applying Lemma 2.25.
From $\bm{\omega}\in\widetilde{\Omega}_{N}$, the definition of $\Lambda^{\dagger}$ (cf. Lemma 2.28) and Lemma 2.22, we have for any $\sigma\in\mathbb{R},$
$\ \ \ \#\{\Lambda\in\mathcal{ER}(N_{1}):\ \Lambda\ {\rm is}\ \sigma{-\rm bad},\ \Lambda\subset\Lambda^{\dagger}\}$
$\leq N^{\frac{1}{10^{3}(b+d)^{2}}}+C(b,d)N_{1}^{b}L_{2}^{d}$
$\leq N^{\frac{1}{10^{3}(b+d)^{2}}}+C(b,d)N^{\rho^{2}b+\frac{1}{50(b+d)}}$
$\leq 2N^{\rho^{2}b+\frac{1}{50(b+d)}}.$
Similar to the proof of Lemma 2.29, we can find some $Y=Y(\sigma)\subset\Lambda^{\dagger}$ satisfying $Y\leq C(b,d)N^{2\rho^{2}(b+d)}N^{\frac{1}{50(b+d)}}<N^{\frac{1}{49(b+d)}}$ so that, if $\bm{x}\in\Lambda^{\dagger}\setminus Y$, then there is some $\sigma$-good $W\in\mathcal{ER}(N_{1}),W\subset\Lambda^{\dagger}\setminus Y$ satisfying
$\bm{x}\in W,\ {\rm dist}(\bm{x},\Lambda^{\dagger}\setminus Y\setminus W)\geq\frac{N_{1}}{2}.$
Let $\bm{\omega}\in\widetilde{\Omega}_{N}$. Again, both the site $\bm{k}_{*}\in[-N,N]^{b}$ (given in Lemma 2.26) and $Y(\sigma)$ depend on $\sigma$. From Lemma B.1, any $\sigma$-good $N_{1}$ size region remains essentially $\sigma^{\prime}$-good if $|\sigma^{\prime}-\sigma|\leq e^{-10\gamma N_{1}}$. So on an interval $I$ of length at most $e^{-10\gamma N_{1}}$, we can choose $Y,\bm{k}_{*}$ to associate with the midpoint of $I$. Also, we can assume $|\sigma|\leq C(b,d)N$ since if $|\sigma|>CN$, then $\|D_{\Lambda^{\dagger}}^{-1}(\sigma)\|\leq 1$, and $\Lambda^{\dagger}$ becomes $\sigma$-good via the Neumann series argument.
We now apply Lemma 2.25 with
$T(\sigma)={{H}}_{\Lambda^{\dagger}}(\sigma),\ \beta=\beta_{1}=e^{-10\gamma N_{1}}.$
We have $K_{1}=CN.$ Using the sublinear bound conclusion and the resolvent identity Lemma B.5, we have
$M=\#Y$ $\leq N^{\frac{1}{49(b+d)}}<L_{3}^{\frac{1}{4}},\ K_{2}=e^{2{N_{1}}^{\frac{3}{4}}}.$
We again use the scale $N_{2}=N_{1}^{\frac{2}{\rho}}$, at which the LDE hold. Similar to the proof of Lemma 2.29, we can verify all assumptions of Lemma 2.25. Define
$\tilde{\Sigma}_{N,3}=\{\sigma\in\mathbb{R}:\ \|G_{\Lambda^{\dagger}}(\sigma)\|\geq e^{L_{3}^{\frac{3}{4}}}\}.$
Then by (2.30) of Lemma 2.25, one obtains
$\mathrm{meas}(\tilde{\Sigma}_{N,3})\leq CN^{C}e^{10\gamma N_{1}}e^{-\frac{c{{L_{3}}^{\frac{3}{4}}}}{CN_{1}N_{2}L_{3}^{\frac{1}{4}}\log N}}\leq e^{-{L_{3}}^{\frac{1}{3}}}\ll e^{-N^{3\rho}}.$
Finally, it suffices to prove the off-diagonal exponential decay of $G_{\Lambda^{\dagger}}(\sigma).$ This follows directly from the resolvent identity of [JLS20], i.e., Lemma B.6. In fact, by the analysis in $\mathcal{R}_{2}$, Lemma 2.29 and Lemma 2.28, all elementary regions contained in $\Lambda^{\dagger}\setminus\Lambda_{1}^{\dagger}$ with sizes $N_{1},L_{0},L_{1}$ are $\sigma$-good. Then applying Lemma B.6 with $\Lambda=\Lambda^{\dagger},\Lambda_{1}=\Lambda_{1}^{\dagger}$ (since ${\rm diam}\ \Lambda^{\dagger}\sim L_{3}>100^{2(b+d)}L_{1}^{2(b+d)}\geq({\rm diam}\ \Lambda_{1}^{\dagger})^{2(b+d)}$) leads to the desired off-diagonal exponential decay estimate of $G_{\Lambda^{\dagger}}(\sigma)$ with
$\gamma_{L_{2}}=\gamma_{L_{1}}-N_{1}^{-\frac{1}{10}}=\gamma_{N_{1}}-N_{1}^{-\vartheta}-N_{1}^{\frac{1}{10}}.$
Application of the resolvent identity again
$10N_{1}$$L_{1}\sim N^{\frac{1}{100(b+d)^{2}}}$$L_{2}\sim N^{\frac{1}{50(b+d)^{2}}}$$L_{3}\sim N^{\frac{1}{10(b+d)}}$$L_{0}\sim L_{1}^{\frac{9}{10}}$$\bm{n}$$\bm{k}$$\log N_{1}\ll\log N$
Figure 1. The distribution of resonant blocks
Finally, we define $\Sigma_{N}=\bigcup\limits_{\ell=1,2,3}\tilde{\Sigma}_{N,\ell}$ with $\tilde{\Sigma}_{N,1},\tilde{\Sigma}_{N,2},\tilde{\Sigma}_{N,3}$ given by the analysis in $\mathcal{R}_{2}$, Lemma 2.29, Lemma 2.30, respectively. Then
${\rm meas}(\Sigma_{N})\leq 3e^{-2N^{3\rho}}\ll e^{-N^{2\rho}},$
and for $\sigma\notin\Sigma_{N}$, we can cover $\Lambda_{{\rm pm},N}$ with $\sigma$-good regions of sizes $N_{1},L_{0},L_{1},L_{3}$. More precisely, we have (cf. Figure 1)
  • For $\bm{x}\in\mathcal{R}_{2}$, cover $\bm{x}$ with $\Lambda\in\mathcal{ER}(L_{0})$ with $L_{0}\sim N^{\frac{9}{1000(b+d)^{2}}}$.
  • For $\bm{x}\in\Lambda_{1}^{\dagger}$, cover $\bm{x}$ with $\Lambda^{\dagger}$.
  • For $\bm{x}\in\mathcal{R}_{1}\setminus\Lambda_{1}^{\dagger}$, cover $\bm{x}$ with regions in
    $\bigcup_{L=N_{1},L_{0},L_{1}}\mathcal{ER}(L).$
As a result, for $\sigma\notin\Sigma_{N}$, we can apply Lemma B.5 and Lemma B.6 to show that $\Lambda_{{\rm pm},N}$ is $\sigma$-good. Indeed, we first apply Lemma B.5 with
$\Lambda=\Lambda_{{\rm pm},N},W\in\bigcup_{L=N_{1},L_{0},L_{1}}(\mathcal{ER}(L)\cup\{\Lambda^{\dagger}\})$
to obtain
$\|G_{\Lambda_{{\rm pm},N}}(\sigma)\|\leq e^{N^{\frac{3}{4}}}.$
Next, for the off-diagonal exponential decay, we use Lemma B.6 with $\Lambda=\Lambda_{{\rm pm},N},\ \Lambda_{1}=\emptyset$, $M_{0}=N_{1}$ to obtain (recalling the decay rate estimates in Lemmas 2.29, 2.30) for $|\bm{x}-\bm{x}^{\prime}|>N^{\frac{8}{9}},$
$|G_{\Lambda_{{\rm pm},N}}(\sigma)((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))|$ $\leq e^{-(\gamma_{L_{2}}-N_{1}^{-\frac{1}{10}})|\bm{x}-\bm{x}^{\prime}|}$
$\leq e^{-(\gamma_{N_{1}}-N_{1}^{-\vartheta}-2N_{1}^{-\frac{1}{10}})|\bm{x}-\bm{x}^{\prime}|}$
$=e^{-\gamma_{N}|\bm{x}-\bm{x}^{\prime}|},$
where
$\gamma_{N}\geq\gamma_{N_{1}}-N^{-\kappa},\ \kappa={\frac{1}{2}}\min\{\vartheta,\frac{1}{10}\}\leq\frac{1}{20}.$ (2.35)
Since $N\sim N_{1}^{\frac{1}{\rho^{2}}}$, we choose $\ell_{*}$ so that $N^{\rho^{2\ell_{*}}}:=N_{*}\sim\log^{\frac{1}{\kappa}}\frac{1}{\varepsilon+\delta}$. It follows from applying Lemma 2.15 (i.e., $\gamma_{N^{\rho^{2\ell_{*}}}}=\gamma_{0}=\gamma-\log^{-2}\frac{1}{\varepsilon+\delta}$) and iterating (2.35) that
$\gamma_{N}$ $\geq\gamma_{N_{*}}-\sum_{\ell=1}^{\ell_{*}}N_{*}^{-\frac{1}{\rho^{2\ell}}\kappa}$
$\geq\gamma_{N^{\rho^{2\ell_{*}}}}-N_{*}^{-\frac{\kappa}{\rho^{2}}}-\sum_{\ell\geq 2}N_{*}^{-\frac{1}{\rho^{2\ell}}\kappa}$
$\geq\gamma-\log^{-2}{\frac{1}{\varepsilon+\delta}}-2N_{*}^{-\frac{\kappa}{\rho^{2}}}$
$\geq\gamma-3\log^{-2}{\frac{1}{\varepsilon+\delta}}\geq\frac{\gamma}{2}$ (2.36)
provided $0<\varepsilon+\delta\leq c(\rho,c_{1},\kappa).$
The above arguments hold indeed for all scales in $N,N^{2}]$. Then we can propagate the LDE from small scales interval $N_{1},N_{2}]$ to large scales one $N,N^{2}]$ similar to the proof of Theorem 4.1 in [JLS20]. This completes the proof of Lemma 2.20.
Proof of Theorem 2.11.
The proof is just a combination of Lemmas 2.152.17  and 2.20 by taking
$0<\varepsilon+\delta\leq\delta_{0}=\min\{\delta_{1},\delta_{2},\delta_{3}\}.$
3. Nonlinear Analysis
In this section, we focus on nonlinear analysis. We work directly with the nonlinear equation (1.1). Recall that
$\mu_{\bm{n}}=V(\bm{n}\bm{\alpha}+\bm{\theta}),\ \bm{\omega}^{(0)}=\left(\omega^{(0)}_{\ell}=\mu_{\bm{n}_{\ell}}\right)_{\ell=1}^{b}.$
We aim to construct solutions to (1.1) of the form
$u(t,\bm{n})=\sum_{(\bm{k},\bm{n})\in\mathbb{Z}^{b}\times\mathbb{Z}^{d}}\hat{u}(\bm{k},\bm{n})e^{i\bm{k}\cdot\bm{\omega}t}.$
Using the Ansatz, we get the nonlinear lattice equation
${\rm diag}_{(\bm{k},\bm{n})}(-\bm{k}\cdot\bm{\omega}+\mu_{\bm{n}})\hat{u}+\varepsilon\Delta\hat{u}+\delta(\hat{u}*\hat{v})^{*p}*\hat{u}=0,$ (3.1)
where $\hat{v}(\bm{k},\bm{n})=\overline{\hat{u}(-\bm{k},\bm{n})}$, the convolution is only in the $\bm{k}$-variable:
$\hat{u}*\hat{v}(\bm{k},\bm{n})=\sum_{\bm{k}_{1}+\bm{k}_{2}=\bm{k}}\hat{u}(\bm{k}_{1},\bm{n})\hat{v}(\bm{k}_{2},\bm{n})$
and
$(\hat{u}*\hat{v})^{*p}=\underbrace{(\hat{u}*\hat{v})*\cdots*(\hat{u}*\hat{v})}_{p\ {\rm times}}.$
Note that $\Delta$ acts only on the $\bm{n}$-variable. To solve for $\hat{u},\hat{v}$, we also need a conjugate equation of (3.1):
${\rm diag}_{(\bm{k},\bm{n})}(\bm{k}\cdot\bm{\omega}+\mu_{\bm{n}})\hat{v}+\varepsilon\Delta\hat{v}+\delta(\hat{u}*\hat{v})^{*p}*\hat{v}=0.$
As a result, we obtain the following vector-valued nonlinear lattice equation:
$F(\vec{u})=0,\ \vec{u}=\binom{\hat{u}}{\hat{v}}$, (3.2)
where
$F(\vec{u})=H_{0}\vec{u}+\delta P(\vec{u})$
with
$H_{0}$ $={\rm diag}_{(\bm{k},\bm{n})\in\mathbb{Z}^{b+d}}\left(\begin{array}[]{cc}{-\bm{k}\cdot\bm{\omega}+\mu_{\bm{n}}+\varepsilon\Delta}&{0}\\ {0}&{\bm{k}\cdot\bm{\omega}+\mu_{\bm{n}}+\varepsilon\Delta}\\ \end{array}\right)$ (3.5)
$:=D(0)+\varepsilon(\Delta\oplus\Delta)$ (3.6)
and
$P(\vec{u})=\binom{(\hat{u}*\hat{v})^{*p}*\hat{u}}{(\hat{u}*\hat{v})^{*p}*\hat{v}}$.
The linearized operator of $P$ at $\vec{u}$ is then given by
${T}_{\vec{u}}=\left(\begin{array}[]{cc}{(p+1)(\hat{u}*\hat{v})^{*p}*}&{p(\hat{u}*\hat{v})^{*(p-1)}*\hat{u}*\hat{u}*}\\ {p(\hat{u}*\hat{v})^{*(p-1)}*\hat{v}*\hat{v}*}&{(p+1)(\hat{u}*\hat{v})^{*p}*}\\ \end{array}\right).$ (3.9)
3.1. The Lyapunov-Schmidt decomposition
We look for quasi-periodic solutions to (1.1) near
$u^{(0)}=u^{(0)}(t,\bm{n})=\sum_{\ell=1}^{b}a_{\ell}e^{i\omega_{\ell}^{(0)}t}\delta_{\bm{n}_{\ell},\bm{n}},$
using the Lyapunov-Schmidt decomposition. The Fourier support of $u^{(0)}$ is $\mathcal{S}_{+}=\{(\bm{e}_{\ell},\bm{n}_{\ell})_{\ell=1}^{b}\}\subset\mathbb{Z}^{b+d}$ (we identify $\mathcal{S}_{+}$ with $\{(\bm{e}_{\ell},\bm{n}_{\ell})_{\ell=1}^{b}\}\times\{+\}\subset\mathbb{Z}^{b+d}_{{\rm pm}}$). Similarly, let $\mathcal{S}_{-}=\{(-\bm{e}_{\ell},\bm{n}_{\ell})_{\ell=1}^{b}\}\subset\mathbb{Z}^{b+d}$ be the Fourier support of $v^{(0)}:=\overline{u^{(0)}}$, which is then identified with $\{(-\bm{e}_{\ell},\bm{n}_{\ell})_{\ell=1}^{b}\}\times\{-\}\subset\mathbb{Z}^{b+d}_{{\rm pm}}$. Let $\mathcal{S}=\mathcal{S}_{+}\cup\mathcal{S}_{-}$ and write $\mathbb{Z}^{b+d}_{{\rm pm},*}=\mathbb{Z}^{b+d}_{{\rm pm}}\setminus\mathcal{S}:=\mathcal{S}^{c}.$
We divide the equation (3.2) into the $P$-equations
${F}\left(\vec{u}\right)\big|_{\mathcal{S}^{c}}=0,\$(3.10)
and the $Q$-equations
${F}(\vec{u})\big|_{\mathcal{S}}=0.$ (3.11)
We iteratively solve the $Q$-equations and then the $P$-equations.
3.2. The $Q$-equations and extraction of parameters
The $Q$-equations are of $2b$-dimensional.
On $\mathcal{S}$, $\vec{u}$ is held fixed, and the $Q$-equations (3.11) are viewed instead as equations for $\bm{\omega}$, and will be solved using the implicit function theorem. The $Q$-equations are used to relate the frequencies $\bm{\omega}$ to the amplitudes $\bm{a}$, permitting amplitude-frequency modulation to solve the $P$-equations.
Note that
$u^{(0)}(t,\bm{n})=\sum_{(\bm{k},\bm{n})\in\mathcal{S}_{+}}\hat{u}^{(0)}(\bm{k},\bm{n})e^{i\bm{k}\cdot\bm{\omega}^{(0)}t},$ (3.12)
where
$\hat{u}^{(0)}(\bm{e}_{\ell},\bm{n}_{\ell})=\hat{v}^{(0)}(-\bm{e}_{\ell},\bm{n}_{\ell})=a_{\ell}\ {\rm for}\ 1\leq\ell\leq b.$
The solutions $\vec{u}=(\hat{u},\hat{v})^{t}$ (with $(\cdot)^{t}$ denoting the transpose) on $\mathcal{S}$ are held fixed: $\vec{u}\equiv(\hat{u}^{(0)},\hat{v}^{(0)})^{t}=\vec{u}^{(0)}$ on $\mathcal{S}$, the $Q$-equations are used instead to solve for the frequencies. Due to symmetry, the $Q$-equations for $\hat{v}$ are the same as that for $\hat{u}$. So we only need to consider that for the $\hat{u}$. Solving the equations at the initial step
${F}(\vec{u}^{(0)})\big|_{\mathcal{S_{+}}}=0$
leads to, for $1\leq\ell\leq b$, the initial modulated frequencies
$\omega_{\ell}=\omega_{\ell}^{(1)}$ $=\omega_{\ell}^{(0)}+\delta a_{\ell}^{2p}$ (3.13)
satisfying
$\det\left(\frac{\partial\bm{\omega}^{(1)}}{\partial\bm{a}}\right)=\delta^{b}\prod_{\ell=1}^{b}(2pa_{\ell}^{2p-1}).$
Along the way, we will show that $\bm{\omega}^{(r)}=\bm{\omega}^{(0)}+O(\delta)$, for all $r=0,1,2,\cdots$. So $\bm{\omega}^{(r)}\in\Omega$, the frequency set in Theorem 2.11, hence Theorem 2.11 is at our disposal to solve the $P$-equations. To construct approximate solutions to the $P$-equations, it is convenient to view $\bm{\omega}\in\Omega$ as an independent parameter, and work in the $(\bm{\omega},\bm{a})$-variable, $(\bm{\omega},\bm{a})\in\Omega\times[1,2]^{b}:=(\bm{\omega}^{(0)}+[\delta,2^{2p}\delta]^{b})\times[1,2]^{b}$. The approximate solutions to the nonlinear matrix equation (3.1) will then be obtained by taking into account the solutions to the $Q$-equations as well, which restricts $(\bm{\omega},\bm{a})$ to the $b$-dimensional hypersurface $\bm{\omega}=\bm{\omega}^{(r)}(\bm{a})$, at the $r$-th iteration.
3.3. The $P$-equations
The $P$-equations are of infinitely dimensional.
We use (3.10) to solve for $\vec{u}|_{\mathcal{S}^{c}}$. Solving the $P$-equations requires analyzing the invertibility of the linearized operator $H=H_{0}+\delta T_{\vec{u}}$ restricted to $\mathbb{Z}^{b+d}_{{\rm pm},*}$, which we do in the previous section. Then the resolution of the $P$-equations combines the Newton scheme and the linear analysis in Section 2. The large deviation estimates in $\sigma$ in Theorem 2.11 will be converted into estimates in the amplitudes $\bm{a}$ by using a semi-algebraic projection lemma (cf. Lemma 3.7 below), and amplitude-frequency modulation, $\bm{\omega}=\bm{\omega}(\bm{a})$. This is an established scheme, which has recently been extensively elaborated on with detailed proofs in sects. IV-VI, [KLW24] and [LW24]. They will serve as the basic reference point for the present section.
Recall first the formal Newton scheme. Let $\vec{u}$ be an approximate solution to the $P$-equations (3.10). For the next approximation $\vec{u}$$^{\prime}$, write
${\vec{u}}^{\prime}=\vec{u}+\Delta_{\rm cor}\vec{u}^{\prime}.$
Then $\Delta_{\rm cor}\vec{u}^{\prime}$ is set to be
$\Delta_{\rm cor}\vec{u}^{\prime}=-(H(\vec{u}))^{-1}F(\vec{u}),$
where the linearized operator $H(\vec{u})=H_{0}+\delta T_{\vec{u}}(\bm{\omega},\bm{a})$ (cf. (3.5) and (3.9)). This is, however, only indicative since we assume that $H(\vec{u})$ is invertible. Due to the small divisor difficulties in the inversion process, we need to regularize the linearized operator $H(\vec{u})$ and do a multi-scale Newton iteration instead.
Toward that purpose, let $M$ be a large integer and consider the geometric sequence of scales $M^{\ell}$, $\ell=1,2,\cdots$. Denote by $\vec{u}^{(r)}$, the $r$-th approximate solution to the $P$-equations (3.10). For the $(r+1)$-th approximation, write
$\vec{u}^{(r+1)}=\vec{u}^{(r)}+\Delta_{\rm cor}\vec{u}^{(r+1)}.$
We define (if it exsits)
$\Delta_{\rm cor}\vec{u}^{(r+1)}=-(R_{N}{{H(\vec{u}^{(r)})}}R_{N})^{-1}F(\vec{u}^{(r)}),$
where $N=M^{r+1}$ and $R_{N}$ is the restriction operator to the cube $\Lambda_{N}\cap\mathbb{Z}^{b+d}_{\rm pm,*}.$ This leads to estimate the inverse (i.e., the Green’s function),
${{G}}_{N}(\vec{u}):=\left(R_{{N}}{{H(\vec{u}^{(r)})}}R_{N}\right)^{-1},\ r=0,1,2,\cdots.$
The following estimates are crucial for the convergence of the approximate solutions with exponential decay:
$\|{{G}}_{N}(\vec{u}^{(r)})\|$ $\leq e^{(\log N)^{C}},$
$|{{G}}_{N}(\vec{u}^{(r)})((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))|$ $\leq e^{-c|\bm{x}-\bm{x}^{\prime}|}\ {\rm for}\ |\bm{x}-\bm{x}^{\prime}|\geq(\log N)^{C},$
for $\bm{x}=(\bm{k},\bm{n})$ and some $c,C>0$, see sect. V, [KLW24].
The LDT Theorem 2.11 in the previous section plays an essential role in the above Green’s function estimates. The operator $T_{\vec{u}}$ plays the role of $S$ in (2.1) (cf. Remark 2.13 for details). The following lemma implies that $T_{\vec{u}}$ satisfies the decay assumption of $S$ (cf. also Remark 2.13). This will then enable us to use Theorem 2.11.
Lemma 3.1.
  • (1)
    For all $\bm{k},\bm{k}^{\prime},\bm{k}^{\prime\prime}\in\mathbb{Z}^{b}$, $\bm{n},\bm{n}^{\prime}\in\mathbb{Z}^{d}$ and $\xi,\xi^{\prime}\in\{+,-\}$, we have the Töplitz property in the $\bm{k}$-variable:
    $T_{\vec{u}}((\bm{k}^{\prime}+\bm{k},\bm{n},\xi);(\bm{k}^{\prime\prime}+\bm{k},\bm{n}^{\prime},\xi^{\prime}))=T_{\vec{u}}((\bm{k}^{\prime},\bm{n},\xi);(\bm{k}^{\prime\prime},\bm{n}^{\prime},\xi^{\prime})).$
  • (2)
    Assume that $|\vec{u}(\bm{k},\bm{n})|\leq e^{-c(|\bm{k}|+|\bm{n}|)},c>0$. Then we have for some $C=C(b,p)>0,$
    $|T_{\vec{u}}((\bm{k},\bm{n},\xi);(\bm{k}^{\prime},\bm{n}^{\prime},\xi^{\prime}))|\leq C(1+|\bm{k}-\bm{k}^{\prime}|)^{C}e^{-c|\bm{k}-\bm{k}^{\prime}|-c|\bm{n}|}\delta_{\bm{n},\bm{n}^{\prime}}.$
Proof.
The proof is similar to that of Lemma 5.1 in [LW24] and Proposition 4.1 in [LSZ25] (our case is easier since the convolution here involves only the $\bm{k}$-variable). We omit the details. ∎
3.4. The induction hypothesis and the proof
Let $M$ be a large integer and recall that $\Lambda_{R}\subset\mathbb{Z}^{b+d}$ denotes the $\ell^{\infty}$-norm induced cube of radius $R$ and centered at the origin.
In the following, $C>0$ is a large constant and $c>0$ is a small one.
We begin with the analysis of initial induction steps, which relies on the Neumann series argument and Diophantine estimates (2.5) of Theorem 2.2.
3.4.1. The initial $2r_{\star}-r_{0}$ steps
We choose $1<r_{0}<r_{\star}$ such that
$M^{r_{0}}\leq\log\frac{1}{\varepsilon+\delta}\leq M^{r_{0}+1}$
and
$M^{2r_{\star}}\leq(\varepsilon+\delta)^{-c_{1}}<M^{2r_{\star}+1},$ (3.14)
where $c_{1}>0$ is given in Theorem 2.11.
Let $\vec{u}^{(0)}(\bm{\omega},\bm{a})=\vec{u}^{(0)}(\bm{\omega}^{(0)},\bm{a})$ with $\bm{\omega}=\bm{\omega}^{(1)}$ given by (3.13).
We start by constructing $\vec{u}_{\rm in}^{(1)}(\bm{\omega},\bm{a})$. Obviously,
${F}(\vec{u}^{(0)})=O(\varepsilon+\delta)$
and (the support) ${\rm supp}\ {F}(\vec{u}^{(0)})\subset\Lambda_{M_{0}}$ for some $M_{0}=M_{0}(p,\mathcal{S})>0.$ Let $L_{1}=M^{r_{0}+1}>M_{0}.$ It suffices to estimate
$H_{L_{1}}^{-1}=(R_{\Lambda_{L_{1}}}({D}(0)+\varepsilon(\Delta\oplus\Delta)+\delta{T}_{\vec{u}^{(0)}})R_{\Lambda_{L_{1}}})^{-1}.$
It follows from (2.5) of Theorem 2.2, $\bm{\omega}=\bm{\omega}^{(0)}+O(\delta)$ and the Neumann series argument (cf. Lemma B.1) that
$\|H_{L_{1}}^{-1}\|$ $\leq 2L_{1}^{C_{1}},$
$|H_{L_{1}}^{-1}((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))|$ $\leq L_{1}^{C_{1}}e^{-\frac{\log\frac{1}{\varepsilon+\delta}}{2}|\bm{x}-\bm{x}^{\prime}|}\ {\rm for}\ \bm{x}\neq\bm{x}^{\prime}\in\Lambda_{L_{1}},$
since $L_{1}=M^{r_{0}+1}\ll(\varepsilon+\delta)^{-c_{1}}$. Then we have
$\Delta_{\rm cor}\vec{u}_{\rm in}^{(1)}=-H_{L_{1}}^{-1}{F}(\vec{u}^{(0)}),$
and consequently
$\vec{u}_{\rm in}^{(1)}=\vec{u}^{(0)}+\Delta_{\rm cor}\vec{u}_{\rm in}^{(1)}=\vec{u}^{(0)}-H_{L_{1}}^{-1}{F}(\vec{u}^{(0)}).$
Substituting $\vec{u}_{\rm in}^{(1)}$ into the $Q$-equations using
$\hat{u}_{\rm in}^{(1)}\big|_{\mathcal{S}+}=\hat{v}_{\rm in}^{(1)}\big|_{\mathcal{S}-}=\bm{a},$
we obtain
$\omega_{\ell}$ $=\omega_{\ell}^{(0)}+\varepsilon\left(\frac{1}{a_{\ell}}(\Delta\hat{u}_{\rm in}^{(1)})(\bm{e}_{\ell},\bm{n}_{\ell})\right)(\bm{\omega},\bm{a})$
$\ \ +\delta\left(\frac{1}{a_{\ell}}(\hat{u}_{\rm in}^{(1)}*\hat{v}_{\rm in}^{(1)})^{*p}*\hat{u}_{\rm in}^{(1)}(\bm{e}_{\ell},\bm{n}_{\ell})\right)(\bm{\omega},\bm{a}),\,\ell=1,2,\cdots,b.$
Since the second and the third terms are smooth in $\bm{\omega}$ and $\bm{a}$, using the implicit function theorem yields $\bm{\omega}_{\rm in}^{(2)}=\bm{\omega}_{\rm in}^{(2)}(\bm{a})=(\omega_{{\rm in},\ell}^{(2)}(\bm{a}))_{\ell=1}^{b}$, written in the form:
$\omega_{{\rm in},\ell}^{(2)}(\bm{a})=\omega_{\ell}^{(0)}+(\varepsilon+\delta)\varphi_{{\rm in},\ell}^{(2)}(\bm{a}),\,\ell=1,2,\cdots,b,$
with a smooth function $\varphi_{{\rm in},\ell}^{(2)}$ (comparing with (3.13)).
The above constructions can be inductively performed for $2r_{\star}-r_{0}$ steps, with $r_{\star}$ satisfying (3.14). So we have obtained $\vec{u}_{\rm in}^{(r)}=\vec{u}_{\rm in}^{(r)}(\bm{\omega},\bm{a})$ ($1\leq r\leq 2r_{\star}-r_{0}$) for all $(\bm{\omega},\bm{a})\in\Omega\times[1,2]^{b}$.
3.4.2. The Inductive Theorem
Recall that
$\Omega=\bm{\omega}^{(0)}+[\delta,2^{2p}\delta]^{b}\subset\mathbb{R}^{b}.$
We first state the induction hypothesis, which can be verified if $0<\varepsilon\leq\delta\leq\log^{-1}\frac{1}{\varepsilon}\leq\delta_{0}\ll 1$ and $(\bm{\alpha},\bm{\theta})\in\mathcal{W}$ (cf. (2.11)). Given its significance, we refer to it as the induction theorem.
Theorem 3.2.
For $r\geq r_{\star},$ we have
  • (Hi)
    ${\rm supp}\ \vec{u}^{(r)}\subset\Lambda_{M^{\tilde{r}}}$, $\tilde{r}=r+r_{\star}.$
  • (Hii)
    $\|\Delta_{\rm cor}\vec{u}^{(r)}\|<\delta_{r},\ \|\partial\Delta_{\rm cor}\vec{u}^{(r)}\|<{\bar{\delta}}_{r}$, where $\partial$ refers to derivation in $\bm{\omega}$ or $\bm{a}$,
    $\vec{u}^{(r)}=\vec{u}^{(r-1)}+\Delta_{\rm cor}\vec{u}^{(r)},$
    and $\|\cdot\|=\sup_{\bm{\omega},\bm{a}}\|\cdot\|_{\ell^{2}(\mathbb{Z}_{\rm pm}^{b+d})}.$
    Remark 3.3.
    The size of $\delta_{r},\bar{\delta}_{r}$ will satisfy $\log\log\frac{1}{\delta_{r}+\bar{\delta}_{r}}\sim r.$
  • (Hiii)
    $|\vec{u}^{(r)}(\bm{k},\bm{n})|\leq e^{-c(|\bm{k}|+|\bm{n}|)}$ for some $c>0.$
    Remark 3.4.
    The constant $c>0$ will decrease slightly along the iterations but remain bounded away from 0. This will become clear in the proof, cf. sect. V, G. Step 4, [KLW24]. We also remark that $\vec{u}^{(r)}$ can be defined as a $C^{1}$ function on the entire parameter space $(\bm{\omega},\bm{a})\in\Omega\times[1,2]^{b}$ by using an extension argument, cf., [Bou98, BW08] and sect. V, L. Step 9, [KLW24]. So one can apply the implicit function theorem to solve the $Q$-equations leading to
    $\omega_{\ell}^{(r)}(\bm{a})=\omega_{\ell}^{(0)}+(\varepsilon+\delta)\varphi_{\ell}^{(r)}(\bm{a})\ (1\leq\ell\leq b),$ (3.15)
    where $\|\partial\bm{\varphi}^{(r)}\|=\sup_{1\leq\ell\leq b}\|\partial\varphi_{\ell}^{(r)}\|\lesssim 1$ and $\bm{\varphi}^{(r)}=(\varphi^{(r)}_{\ell})_{\ell=1}^{b}.$ By (Hii), we have
    $|\bm{\varphi}^{(r)}-\bm{\varphi}^{(r-1)}|\lesssim(\varepsilon+\delta)\|\vec{u}^{(r)}-\vec{u}^{(r-1)}\|\lesssim{(\varepsilon+\delta)}\delta_{r}.$
    Denote by $\Gamma_{r}$ the graph of $\bm{\omega}^{(r)}=\bm{\omega}^{(r)}(\bm{a})$. We have $\|\Gamma_{r}-\Gamma_{r-1}\|\lesssim{(\varepsilon+\delta)}\delta_{r}.$ Recall that $\bm{\omega}$ is given by (3.13) and $\bm{\varphi}^{(0)}=\bm{0}$. Thus we have a diffeomorphism from $\bm{\omega}^{(r)}\in\Omega$ to $\bm{a}\in[1,2]^{b}$.
  • (Hiv)
    There is a collection $\mathcal{I}_{r}$ of intervals $I\subset\Omega\times[1,2]^{b}$ of size $M^{-{\tilde{r}}^{10C}},$ so that
    • (a)
      On each $I\in\mathcal{I}_{r},$ both $\hat{u}^{(r)}(\bm{\omega},\bm{a})$ and $\hat{v}^{(r)}(\bm{\omega},\bm{a})$ are given by rational functions in $(\bm{\omega},\bm{a})$ of degree at most $M^{{\tilde{r}}^{3}}$.
    • (b)
      For $(\bm{\omega},\bm{a})\in\bigcup_{I\in\mathcal{I}_{r}}I$,
      $\|F(\vec{u}^{(r)})\|\leq\kappa_{r},\ \|\partial F(\vec{u}^{(r)})\|\leq\bar{\kappa}_{r},$
      where $\partial$ refers to derivation in $\bm{\omega}$ or $\bm{a}$, and $\log\log\frac{1}{\kappa_{r}+\bar{\kappa}_{r}}\sim r$.
    • (c)
      For $(\bm{\omega},\bm{a})\in\bigcup_{I\in\mathcal{I}_{r}}I$ and $H=H(\vec{u}^{(r-1)})$, one has
      $\|H_{M^{\tilde{r}}}^{-1}\|$ $\leq M^{{\tilde{r}}^{C}},$ (3.16)
      $|H_{M^{\tilde{r}}}^{-1}((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))|$ $\leq e^{-c|\bm{x}-\bm{x}^{\prime}|}\ {\rm for}\ |\bm{x}-\bm{x}^{\prime}|>{\tilde{r}}^{C},$ (3.17)
      where $\bm{x}=(\bm{k},\bm{n}),\bm{x}^{\prime}=(\bm{k}^{\prime},\bm{n}^{\prime})$ and $H_{M^{\tilde{r}}}$ refers to the restriction of $H$ to $\Lambda_{M^{\tilde{r}}}$.
    • (d)
      Each $I\in\mathcal{I}_{r}$ is contained in some $I^{\prime}\in\mathcal{I}_{r-1}$ and
      ${\rm meas}\left(\Pi_{\bm{a}}(\Gamma_{r-1}\cap(\bigcup_{I^{\prime}\in\mathcal{I}_{r-1}}I^{\prime}\setminus\bigcup_{I\in\mathcal{I}_{r}}I))\right)\leq M^{-\frac{\tilde{r}}{C(b)}},$
      where $C(b)>0$ depends only on $b$, and $\Pi_{\bm{a}}$ denotes the projection of the set on the $\bm{a}$-variable.
  • (Hv)
    The following precise relations hold true
    $\delta_{r}={(\varepsilon+\delta)}^{\frac{1}{2}}M^{-(\frac{4}{3})^{r}},\,\bar{\delta}_{r}={(\varepsilon+\delta)}^{\frac{1}{8}}M^{-\frac{1}{2}(\frac{4}{3})^{r}};$
    $\kappa_{r}={(\varepsilon+\delta)}^{\frac{3}{4}}M^{-(\frac{4}{3})^{r+2}},\,\bar{\kappa}_{r}={(\varepsilon+\delta)}^{\frac{3}{8}}M^{-\frac{1}{2}(\frac{4}{3})^{r+2}}.$
3.5. Proof of the Inductive Theorem
From the analysis in Section 3.4.1, we have constructed $\vec{u}_{\rm in}^{(r)}$ for $1\leq r\leq 2r_{\star}-r_{0}$. In particular, we obtain that the Inductive Theorem holds for $r=r_{\star}$ as we can set $\vec{u}^{(r_{\star})}=\vec{u}_{\rm in}^{(2r_{\star}-r_{0})}$.
Assume now the Inductive Theorem holds for $r>r_{\star}$. We will prove it for $\ell=r+1.$
We first check the degree bound on $\vec{u}^{(r+1)}$ (i.e., (Hiv, a) for $r+1$). In fact, the Newton scheme gives
$\Delta_{\rm cor}\vec{u}^{(r+1)}=-H^{-1}_{N}(\vec{u}^{(r)})F(\vec{u}^{(r)}),\ N=M^{\tilde{r}+1},\ \tilde{r}=r+r_{\star},$ (3.18)
and by (Hiv, a), $\vec{u}^{(r)}(\bm{\omega},\bm{a})$ is a rational function of degree at most $M^{\tilde{r}^{3}}$ (in $(\bm{\omega},\bm{a})$). By the convolution structure of $T_{\vec{u}^{(r)}}$, we have
${\rm deg}\ F(\vec{u}^{(r)})\leq(2p+1)M^{\tilde{r}^{3}},\ {\rm deg}\ H_{N}(\vec{u}^{(r)})\leq 2pM^{\tilde{r}^{3}}.$
Then by Cramer’s rule, we have
$H_{N}^{-1}(\vec{u}^{(r)})=\frac{A}{\det H_{N}(\vec{u}^{(r)})},$
where $A$ is the adjacent matrix of $H_{N}(\vec{u}^{(r)})$. Hence,
${\rm deg}\ H^{-1}_{N}(\vec{u}^{(r)})\leq(4p+2)M^{\tilde{r}+1}M^{\tilde{r}^{3}}.$
Thus, using (3.18) shows
${\rm deg}\ \Delta_{\rm cor}\vec{u}^{(r+1)}\leq(4p+2)M^{\tilde{r}+1}M^{\tilde{r}^{3}}+(2p+1)M^{\tilde{r}^{3}}\leq M^{(\tilde{r}+1)^{3}}.$
This proves the (Hiv, a) for $r+1$.
The other inductive assumptions can be verified in the usual way, cf. [LW24] and sect. V, [KLW24], e.g., the verification of (Hiv, b) can be found in I. Step 6 of [KLW24]; the verification of the constant relations (Hv) can be found in Appendix E of [LW24].
It remains to establish (Hiv, c) and (Hiv, d) for $r+1$. For this purpose, we set (recalling $\tilde{r}=r+r_{\star}$)
$N=M^{\tilde{r}+1}$
and a smaller scale,
$N_{1}=(\log N)^{\frac{4}{\rho^{4}}},$
where $0<\rho\ll 1$ is given in Theorem 2.11.
To establish (Hiv, c), we first make approximations on $T_{\vec{u}^{(\ell)}}$ for different $\vec{u}^{(\ell)}$. For the set
$W=\left(([-N,N]^{b+d}\setminus[-N/2,N/2]^{b+d})\times\{+,-\}\right)\cap\mathbb{Z}_{{\rm pm},*}^{b+d},$
we use $N_{1}$ size regions to do estimates, in order to lower the degree of the associated semi-algebraic set, which will be essential for the upcoming semi-algebraic projection arguments. We set
$r_{1}=2\left[\frac{\log N_{1}}{\log\frac{4}{3}}\right]+1.$
Then
$\|T_{\vec{u}^{(r)}}-T_{\vec{u}^{(r_{1})}}\|\lesssim\delta_{r_{1}}\lesssim\sqrt{\varepsilon+\delta}e^{-(\frac{4}{3})^{r_{1}}}\ll e^{-N_{1}^{2}}.$ (3.19)
So we can estimate $H_{W}^{-1}(\vec{u}^{(r)})$ using $H_{Q}^{-1}(\vec{u}^{(r_{1})})$ for $Q\subset W$. More precisely, we want to show that the following estimates hold for every $Q\subset W$ with $Q\in\mathcal{ER}(N_{1})$:
$\|H_{Q}^{-1}(\vec{u}^{(r_{1})})\|$ $\leq e^{N_{1}^{\frac{3}{4}}},$ (3.20)
$|H_{Q}^{-1}(\vec{u}^{(r_{1})})((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))|$ $\leq e^{-c|\bm{x}-\bm{x}^{\prime}|}\ {\rm for}\ |\bm{x}-\bm{x}^{\prime}|>N_{1}^{\frac{8}{9}}.$ (3.21)
This requires additional restrictions on $(\bm{\omega},\bm{a}).$ For this, we divide into the following cases:
  • Case 1.
    Assume $Q\in(\bm{k},\bm{n})+\mathcal{ER}_{0}(N_{1})$ with $|\bm{n}|>10N_{1}$ and $Q\subset W.$ Denote by $\mathcal{C}_{1}$ the set of all these $Q$. To estimate $H_{Q}^{-1}(\vec{u}^{(r_{1})})$, we can impose as in the proof of Lemma 2.23 (with $\sigma=0$) the following condition
    $\min_{\bm{k}\in\Pi_{b}Q,\xi=\pm 1,1\leq\ell\leq N_{1}^{b+2d}}|\bm{k}\cdot\bm{\omega}+\xi{\lambda_{\ell}}|>e^{-\frac{1}{4}N_{1}^{\frac{3\kappa_{1}}{4}}},$
    where $\lambda_{\ell}$ is given by Lemma 2.23. For $\bm{k}\neq\bm{0}$, we can remove $\bm{\omega}$ directly (i.e., a set in $\bm{\omega}$ of measure at most $N^{C}e^{-\frac{1}{4}N_{1}^{\frac{3\kappa_{1}}{4}}}$) to control $A_{\bm{k}}^{-1}(0)$ (cf. the proof of Lemma 2.23 for this notation). For the case $\bm{k}=\bm{0}$ of which we cannot remove $\bm{\omega}$, we employ Green’s function estimates in (2) of Lemma 2.5. Indeed, similar to the proof of Lemma 2.23 (1), the estimate of $A_{\bm{0}}^{-1}(0)=\mathcal{L}^{-1}_{Q(\bm{0})}(\bm{\theta})\oplus\mathcal{L}^{-1}_{Q(\bm{0})}(\bm{\theta})$ follows directly from (2) of Lemma 2.5 since we assume $(\bm{\alpha},\bm{\theta})\in\mathcal{W}$ (cf. (2.11)). So by taking account of all $\bm{k}\in\Pi_{b}Q$, all $Q\in\mathcal{C}_{1}$ and by Fubini’s theorem, we can find $I_{1}\subset\Omega\times[1,2]^{b}$ with (recalling $\rho\leq c_{2}\kappa_{1}\leq\frac{\kappa_{1}}{10^{4}(b+d)^{2}}$)
    ${\rm meas}(I_{1})\leq N^{C}e^{-\frac{1}{4}(\log N)^{\frac{3\kappa_{1}}{\rho^{4}}}}\leq e^{-(\log N)^{10}}\ll(\varepsilon+\delta)^{b+1}M^{-{\tilde{r}}}$,
    so that, for all $(\bm{\omega},\bm{a})\notin I_{1}$ and all $Q\in\mathcal{C}_{1}$, $H_{Q}^{-1}(\vec{u}^{(r_{1})})$ satisfies (3.20) and (3.21).
  • Case 2.
    Assume $Q\in(\bm{k},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ with $|\bm{n}|\leq 10N_{1}$ and $Q\subset W.$ Denote by $\mathcal{C}_{2}$ the set of all these $Q$. In this case, it must be that $|\bm{k}|\geq N/2.$ We will use the projection lemma and LDT to remove $\bm{\omega}.$ For any $Q\in\mathcal{C}_{2}$, we can write $Q=(\bm{k},\bm{n})+Q_{0}$ with $Q_{0}\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ for $|\bm{n}|\leq 10N_{1},$ and $N/2\leq|\bm{k}|\leq N.$ This motivates us to consider
    $H(\sigma)=H(\vec{u}^{(r_{1})};\sigma)=D(\sigma)+\varepsilon(\Delta\oplus\Delta)+\delta T_{\vec{u}^{(r_{1})}},$
    which has been investigated in Section 2. Recall that $G_{Q}(\sigma)$ denotes the Green’s function of $H(\sigma)$ restricted to $Q$. Then the Toeplitz property in the $\bm{k}$-direction of $H(\sigma)$ implies
    $H^{-1}_{Q}(\vec{u}^{(r_{1})})=G_{Q}(0)=G_{Q_{0}}(\bm{k}\cdot\bm{\omega}).$
    Note that $\Omega_{N_{1}}$ given by Theorem 2.11 is a semi-algebraic set, independent of $\bm{a}$, and of degree at most $N_{1}^{10(b+d)}$. Without loss of generality, we may assume $\Pi_{\bm{\omega}}(I\cap\Gamma_{r_{1}})\subset\Omega_{N_{1}}$ for each $I\in\mathcal{I}_{r_{1}}$. For otherwise, we can replace $\Pi_{\bm{\omega}}(I\cap\Gamma_{r_{1}})$ with $(\Pi_{\bm{\omega}}(I\cap\Gamma_{r_{1}}))\cap(\Omega_{N_{1}^{\prime}}\setminus\Omega_{N_{1}})$, where $N_{1}^{\prime}=(\log M^{\tilde{r}})^{\frac{4}{\rho^{4}}}\ll N_{1}$ since by Remark 2.12,
    ${\rm meas}(\Omega_{N_{1}^{\prime}}\setminus\Omega_{N_{1}})\leq e^{-\frac{1}{2}N_{1}^{\rho^{4}}}\leq e^{-\frac{1}{2}(\log N)^{4}}\ll(\varepsilon+\delta)^{b+1}N^{-10}.$
    Then we apply the LDT at scale $N_{1}$ on each $I\in\mathcal{I}_{r_{1}}$ (cf. Remark 2.13) to obtain that for all $Q\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ with $|\bm{n}|\leq 10N_{1},$ and for $\sigma$ outside a set of measure at most $e^{-N_{1}^{\rho}}$, the following estimates hold true:
    $\|G_{Q}(\sigma)\|$ $\leq e^{N_{1}^{\frac{3}{4}}},$ (3.22)
    $|G_{Q}(\sigma)((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))|$ $\leq e^{-c|\bm{x}-\bm{x}^{\prime}|}\ {\rm for}\ |\bm{x}-\bm{x}^{\prime}|>N_{1}^{\frac{8}{9}}.$ (3.23)
    For the usage of projection lemma in the present setting (i.e., ${\rm meas}(\Omega)\sim\delta^{b}$), we need to make more precise descriptions of the admitted $\sigma$ first. We say $\sigma$ is $Q$-bad if either (3.22) or (3.23) fails.
    Lemma 3.5.
    Fix $Q\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ with $|\bm{n}|\leq 10N_{1}.$ Denote by $\Sigma_{Q}$ the set of $Q$-bad $\sigma\in\mathbb{R}$. Then
    $\Sigma_{Q}\subset\bigcup_{1\leq\ell\leq N_{1}^{C}}J_{\ell},$
    where each $J_{\ell}$ is an interval of length $\sim(\varepsilon+\delta).$
    Proof.
    Note that $N=M^{\tilde{r}+1}\geq M^{2r_{\star}+1}>(\varepsilon+\delta)^{-c_{1}}$ since $r>r_{\star}$. So we have
    $N_{1}=(\log N)^{\frac{4}{\rho^{4}}}>c_{1}\log^{\frac{4}{\rho^{4}}}\frac{1}{\varepsilon+\delta}>\log^{\frac{2}{\rho^{4}}}\frac{1}{\varepsilon+\delta},$
    which implies
    $(\varepsilon+\delta)^{-1}<e^{N_{1}^{\rho^{4}/2}}.$ (3.24)
    So we can define for each $\bm{x}=(\bm{k},\bm{n})\in Q$ and $\xi=\pm 1$ the set
    $J_{\bm{x},\xi}=\{\sigma\in\mathbb{R}:\ |\sigma+\bm{k}\cdot\bm{\omega}+\xi\mu_{\bm{n}}|\leq C(\varepsilon+\delta)\},$
    where $C>1$ depends only on $\|\Delta\|,\|T_{\vec{u}^{(r_{1})}}\|.$ From (3.24) and the Neumann series argument (cf. Lemma B.1), we have
    $\Sigma_{Q}\subset\bigcup_{\bm{x}\in Q,\ \xi=\pm 1}J_{\bm{x},\xi}.$
    This proves Lemma 3.5.
    Now fix $I_{0}\in\mathcal{I}_{r_{1}}.$ Solving the $Q$-equations at $r=r_{1}$ leads to the graph $\Gamma_{r_{1}}$. Then $\tilde{I}=\Pi_{\bm{\omega}}(\Gamma_{r_{1}}\cap I_{0})$ is an interval of size at most $\varepsilon+\delta.$ For $\bm{\omega}\in\tilde{I}$, let $\tilde{\Sigma}$ be the set of $\sigma\in\mathbb{R}$ so that either (3.22) or (3.23) fails for some $Q\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ with $|\bm{n}|\leq 10N_{1}$. Then by Lemma 3.5, $\tilde{\Sigma}$ can be covered by $N_{1}^{C}$ intervals of size $\sim(\varepsilon+\delta).$ Pick $J$ to be one of such intervals and consider $\mathcal{K}:=\tilde{I}\times(\tilde{\Sigma}\cap J)\subset\tilde{I}\times J\subset\mathbb{R}^{b}\times\mathbb{R}.$ We will show that $\mathcal{K}$ is a semi-algebraic set of degree $\deg\mathcal{K}\leq N_{1}^{C}M^{C{\tilde{r}}_{1}^{3}}$ and measure ${\rm meas}(\mathcal{K})\leq C(\delta+\varepsilon)^{b}e^{-N_{1}^{\rho}}$. This measure bound follows directly from the LDT at scale $N_{1}$ and the Fubini’s theorem. For the semi-algebraic description of $\mathcal{K}$, we need the following lemma.
    Lemma 3.6 (Tarski-Seidenberg Principle, cf. e.g., [Bou07]).
    Denote by $(\bm{x},\bm{y})\in\mathbb{R}^{d_{1}+d_{2}}$ the product variable. If $X\subset\mathbb{R}^{d_{1}+d_{2}}$ is semi-algebraic of degree $B$, then its projections $\Pi_{\bm{x}}X\subset\mathbb{R}^{d_{1}}$ and $\Pi_{\bm{y}}X\subset\mathbb{R}^{d_{2}}$ are semi-algebraic of degree at most $B^{C}$, where $C=C(d_{1},d_{2})>0$.
    We define $X\subset I_{0}\times J$ to be the set of all $(\bm{\omega},\bm{a},\sigma)$ so that either (3.22) or (3.23) fails for some $Q\in(\bm{0},\bm{n})+\mathcal{ER}_{\bm{0}}(N_{1})$ with $|\bm{n}|\leq 10N_{1}$. Similar to the proof of Lemma 2.22, we can regard $X$ as a semi-algebraic set of degree $N_{1}^{C}$. From (Hiv) and solving the $Q$-equations for $r=r_{1}$, we know that $\Gamma_{r_{1}}\cap I_{0}$ is given by an algebraic equation in $\bm{\omega},\bm{a}$ (cf. (3.15)) of degree $M^{C{\tilde{r}}_{1}^{3}}$. Note also that
    $\mathcal{K}=\Pi_{\bm{\omega},\sigma}\left(X\cap((\Gamma_{r_{1}}\cap I_{0})\times\mathbb{R})\right),$
    which together with Lemma 3.6 implies $\deg\mathcal{K}\leq N_{1}^{C}M^{C{\tilde{r}}_{1}^{3}}.$
    Our next aim is to estimate
    ${\rm meas}\left(\bigcup_{N/2\leq|\bm{k}|\leq N}\left\{\bm{\omega}:\ (\bm{\omega},\bm{k}\cdot\bm{\omega})\in\mathcal{K}\right\}\right).$ (3.25)
    At this stage, we need an important projection lemma.
    Lemma 3.7 (cf. e.g., [Bou07]).
    Let $Y\subset[0,1]^{d=d_{1}+d_{2}}$ be a semi-algebraic set of degree $\deg Y=B$ and $\mathrm{meas}(Y)\leq\eta$, where $\log B\ll\log\frac{1}{\eta}.$ Denote by $(\bm{x},\bm{y})\in[0,1]^{d_{1}}\times[0,1]^{d_{2}}$ the product variable. Suppose $\eta^{\frac{1}{d}}\leq\epsilon.$ Then there is a decomposition of $Y$ as
    $Y=Y_{1}\cup Y_{2}$
    with the following properties: The projection of $Y_{1}$ on $0,1]^{d_{1}}$ has small measure
    $\mathrm{meas}(\Pi_{\bm{x}}Y_{1})\leq{B}^{C(d)}\epsilon,$
    and $Y_{2}$ has the transversality property
    $\mathrm{meas}(\mathbb{H}\cap Y_{2})\leq B^{C(d)}\epsilon^{-1}\eta^{\frac{1}{d}},$
    where $\mathbb{H}$ is any $d_{2}$-dimensional hyperplane in $\mathbb{R}^{d},$ s.t., $\max\limits_{1\leq j\leq d_{1}}|\Pi_{\mathbb{H}}(\bm{e}_{j})|<{\epsilon}$ (we denote by $\bm{e}_{1},\cdots,{\bm{e}}_{d_{1}}$ the $\bm{x}$-coordinate standard basis).
    Remark 3.8.
    This lemma permits scaling. It is based on the Yomdin-Gromov triangulation theorem [Gro87]; for a complete proof, see [BN19].
    Taking $\eta=C(\varepsilon+\delta)^{b}e^{-N_{1}^{\rho}},\epsilon=10/N$, we obtain
    $C(\varepsilon+\delta)^{b}e^{-\frac{1}{b+1}N_{1}^{\rho}}\leq C(\varepsilon+\delta)^{b}e^{-\frac{1}{b+1}(\log N)^{4}}\ll\epsilon.$
    Using Lemma 3.7, we have a decomposition
    $\mathcal{K}=\mathcal{K}_{1}\cup\mathcal{K}_{2},$
    where ${\rm meas}\ (\Pi_{\bm{\omega}}\mathcal{K}_{1})\leq C(\varepsilon+\delta)^{b+1}N_{1}^{C}M^{C{\tilde{r}}_{1}^{3}}M^{-{\tilde{r}}}\leq(\varepsilon+\delta)^{b+1}M^{-\frac{\tilde{r}}{2}}$ and the factor $(\varepsilon+\delta)^{b+1}$ comes from scaling when applying Lemma 3.7. Note however that $|\bm{k}|\geq N/2$ cannot ensure $\min_{1\leq\ell\leq b}|k_{\ell}|\geq N/2$. So the hyperplane $\{(\bm{\omega},\bm{k}\cdot\bm{\omega})\}$ does not satisfy the steepness condition. To address this issue, we will need to take into account of all possible directions and apply Lemma 3.7 for $b$ times, as done by Bourgain (cf. (3.26) of [Bou07] and also (142) of [LW24]). This then leads to an upper bound $C(\varepsilon+\delta)^{b+1}M^{-\frac{{\tilde{r}}+1}{C(b)}}$ on (3.25), where $C(b)>0$ only depends on $b$.
    Taking into account of all $J$, we have shown the existence of $\tilde{I}_{1}\subset\tilde{I}$ with ${\rm meas}_{b}(\tilde{I}_{1})\leq C(\varepsilon+\delta)^{b+1}N_{1}^{C}M^{-\frac{{\tilde{r}}+1}{C(b)}}\leq C(\varepsilon+\delta)^{b+1}M^{-\frac{{\tilde{r}}+1}{C(b)}}$ so that for $\bm{\omega}\in\tilde{I}\setminus\tilde{I}_{1}$, the estimates (3.20) and (3.21) hold for all $Q\in\mathcal{C}_{2}$. Let $I_{0}$ range over $\mathcal{I}_{r_{1}}$. The total measure removed from $\Pi_{\bm{\omega}}\Gamma_{r_{1}}$ is at most $C(\varepsilon+\delta)^{b+1}M^{{\tilde{r}}_{1}^{C}}M^{-\frac{{\tilde{r}}+1}{C(b)}}\leq(\varepsilon+\delta)^{b+1}M^{-\frac{{\tilde{r}}+1}{C(b)}}$. Since (3.20)–(3.21) allow $O(e^{-N^{2}_{1}})$ perturbation of $(\bm{\omega},\bm{a})$ (again using Lemma B.1) and $\|\Gamma_{r}-\Gamma_{r_{1}}\|\lesssim\delta_{r_{1}}\ll e^{-N^{2}_{1}}$, we obtain a subset $\Gamma_{r}^{\prime}\subset\Gamma_{r}$ with ${\rm meas}(\Pi_{\bm{\omega}}\Gamma_{r}^{\prime})\leq(\varepsilon+\delta)^{b+1}M^{-\frac{{\tilde{r}}+1}{C(b)}}$ so that (3.20) and (3.21) hold on
    $\bigcup_{I\in\mathcal{I}_{r_{1}}}(I\cap(\Gamma_{r}\setminus\Gamma_{r}^{\prime})),$
    and hence on
    $\bigcup_{I\in\mathcal{I}_{r}}(I\cap(\Gamma_{r}\setminus\Gamma_{r}^{\prime})).$
Next, by the induction hypothesis at step $r$, we have that (3.16) and (3.17) hold for $r$. Also, by (Hii) at step $r$, if $|\bm{x}-\bm{x}^{\prime}|\leq(\tilde{r}+1)^{5C}\log M,$ then (we hide the dependence on $\xi,\xi^{\prime},\bm{\omega},\bm{a}$)
$\ \ \ |{T}_{\vec{u}^{(r)}}(\bm{x},\bm{x}^{\prime})-{T}_{\vec{u}^{(r-1)}}(\bm{x},\bm{x}^{\prime})|$
$\leq C\|\Delta_{\rm cor}\vec{u}^{(r)}\|$
$\leq C\delta_{r}\leq CM^{-(\tilde{r}+1)^{10C}}$
$\leq M^{-10(\tilde{r}+1)^{C}}e^{-c|\bm{x}-\bm{x}^{\prime}|}.$
If $|\bm{x}-\bm{x}^{\prime}|>(\tilde{r}+1)^{5C}\log M$, by (Hiii) at step $r$ and Lemma 3.1, we have
$\ \ \ |{T}_{\vec{u}^{(r)}}(\bm{x},\bm{x}^{\prime})-{T}_{\vec{u}^{(r-1)}}(\bm{x},\bm{x}^{\prime})|$
$\leq C(1+|\bm{x}-\bm{x}^{\prime}|)^{C}e^{-c|\bm{x}-\bm{x}^{\prime}|}$
$\leq M^{-10(\tilde{r}+1)^{C}}e^{-(c-(\tilde{r}+1)^{-3C})|\bm{x}-\bm{x}^{\prime}|}.$
The above estimates imply that for all $\bm{x},\bm{x}^{\prime},$
$|{T}_{\vec{u}^{(r)}}(\bm{x},\bm{x}^{\prime})-{T}_{\vec{u}^{(r-1)}}(\bm{x},\bm{x}^{\prime})|\leq M^{-10(\tilde{r}+1)^{C}}e^{-(c-(\tilde{r}+1)^{-3C})|\bm{x}-\bm{x}^{\prime}|}.$
Hence, combining (3.16)–(3.17) (at step $r$) and Lemma B.1, it follows that for any $(\bm{\omega},\bm{a})\in\cup_{I\in\mathcal{I}_{r}}I$, we can replace $\vec{u}^{(r-1)}$ in (3.16)–(3.17) with $\vec{u}^{(r)}$. That is to say,
$\|H_{M^{\tilde{r}}}^{-1}(\vec{u}^{(r)})\|\leq 2M^{{\tilde{r}}^{C}},$ (3.26)
and for $|\bm{x}-\bm{x}^{\prime}|>{\tilde{r}}^{C},$
$|H_{M^{\tilde{r}}}^{-1}(\vec{u}^{(r)})(\bm{x},\bm{x}^{\prime})|\leq e^{-(c-(\tilde{r}+1)^{-3C})|\bm{x}-\bm{x}^{\prime}|}.$ (3.27)
Assume $|(\bm{\omega},\bm{a})-(\bm{\omega}^{\prime},\bm{a}^{\prime})|\leq M^{-({\tilde{r}}+1)^{10C}}$. Similar to the above proof, we have for all $\bm{x},\bm{x}^{\prime},$
$|{H}(\vec{u}^{(r)};(\bm{\omega},\bm{a}))(\bm{x},\bm{x}^{\prime})-{H}(\vec{u}^{(r)};(\bm{\omega}^{\prime},\bm{a}^{\prime}))(\bm{x},\bm{x}^{\prime})|$
$\leq$ $M^{-10(\tilde{r}+1)^{C}}e^{-(c-(\tilde{r}+1)^{-3C})|\bm{x}-\bm{x}^{\prime}|}.$
Using Lemma B.1 again, we have that the estimates (3.26)–(3.27) and (3.20)–(3.21) (replacing $\vec{u}^{(r_{1})}$ with $\vec{u}^{(r)}$) allow $O(M^{-({\tilde{r}}+1)^{10C}})$ perturbation of $(\bm{\omega},\bm{a})$. Combining conclusions in the above Case 1Case 2 and the resolvent identity of [HSSY20] (cf. Lemma 3.6) gives a collection $\mathcal{I}_{r+1}$ of intervals of size $M^{-({\tilde{r}}+1)^{10C}}$ so that for $I\in\mathcal{I}_{r+1}$,
$\|H_{N}^{-1}(\vec{u}^{(r)})\|$ $\leq M^{({\tilde{r}}+1)^{C}},$
$|H_{N}^{-1}(\vec{u}^{(r)})((\bm{x},\xi);(\bm{x}^{\prime},\xi^{\prime}))|$ $\leq e^{-c_{r+1}|\bm{x}-\bm{x}^{\prime}|}\ {\rm for}\ |\bm{x}-\bm{x}^{\prime}|>({\tilde{r}}+1)^{C},$
which concludes (Hiv, c) at scale $r+1.$ We remark that in the above off-diagonal exponential decay estimates, $c_{r+1}\geq c_{r}-(\log M)^{C}({\tilde{r}}+1)^{-C}$, which implies $\inf_{r}c_{r}>0$. This explains why we need the sublinear distant off-diagonal decay in LDT.
Finally, we have
$\ \ \ {\rm meas}\left(\Pi_{\bm{\omega}}(\Gamma_{r}\cap(\bigcup_{I^{\prime}\in\mathcal{I}_{r}}I^{\prime}\setminus\bigcup_{I\in\mathcal{I}_{r+1}}I))\right)$
$\leq(\varepsilon+\delta)^{b+1}M^{-\frac{{\tilde{r}}+1}{C(b)}}+e^{-\frac{1}{2}N_{1}^{\rho^{4}}}<(\varepsilon+\delta)^{b+1}M^{-\frac{{\tilde{r}}+1}{C(b)}}.$
Recalling that $\bm{\omega}\to\bm{a}$ is a $C^{1}$ diffeomorphism and $\det(\frac{\partial\bm{\omega}}{\partial\bm{a}})\sim\delta^{b}$, we obtain since $\varepsilon\leq\delta$ that
${\rm meas}\left(\Pi_{\bm{a}}(\Gamma_{r}\cap(\bigcup_{I^{\prime}\in\mathcal{I}_{r}}I^{\prime}\setminus\bigcup_{I\in\mathcal{I}_{r+1}}I))\right)\leq M^{-\frac{{\tilde{r}}+1}{C(b)}},$
which yields (Hiv, d) at step $r+1.$ $\square$
Proof of Theorem 1.1.
The proof of Theorem 1.1 is a direct corollary of the Inductive Theorem. We refer to [BW08, LW24] and sects. V and VI of [KLW24] for further details. ∎
Acknowledgements
Y. Shi was supported by the National Key R&D Program (2021YFA1001600) and the NSFC (12522110). W.-M. Wang acknowledges support from the CY Initiative of Excellence, “Investissements d’Avenir” Grant No. ANR-16-IDEX-0008. The authors would like to thank the handling editor and the anonymous reviewers for valuable suggestions.
Appendix A Diophantine estimates
Below we provide Diophantine estimates when $V(\bm{\theta})$ are (arbitrary) trigonometric polynomials. These estimates hold on a large set in $(\bm{\alpha},\bm{\theta})$. The main idea is to utilize an appropriate generalized Wronskian approach. To our knowledge, these estimates did not appear to be known in the literature before. Given the relation between analytic functions (hence trigonometric polynomials) and generalized Wronskians, this approach also seems natural in hindsight.
To our knowledge, using transversality properties to make Diophantine approximations on manifolds was first developed by Pyartli [Pya69]. This method was later introduced into the study of KAM theory [XYQ97, Eli02, Bam03]. It turns out that this idea also plays a key role in the Craig-Wayne-Bourgain type argument used in the present paper. In addition, we can handle Diophantine estimates on general quasi-periodic polynomials on $0,1]^{d}$. The main difference between [XYQ97, Eli02, Bam03] and the present work is that we have no prior transversality estimate, which requires much additional effort.
Recall that
$V(\bm{\theta})=\sum_{\bm{\ell}\in\Gamma_{K}}v_{\bm{\ell}}\cos 2\pi(\bm{\ell}\cdot\bm{\theta})\ {\rm with}\ \bm{\theta}\in[0,1]^{d},$
where $\Gamma_{K}\subset[-K,K]^{d}\setminus\{\bm{0}\}$ ($K\in\mathbb{N}$) is maximal so that
$\text{if }\bm{\ell}\in\Gamma_{K},{\rm then},\ell_{s}\neq 0,\forall 1\leq s\leq d;$ (A.1)
$\text{and if }\ \bm{\ell},\bm{\ell}^{\prime}\in\Gamma_{K},\ {\rm then}\ \bm{\ell}+\bm{\ell}^{\prime}\neq\bm{0},$
and
$\bm{v}=(v_{\bm{\ell}})_{\bm{\ell}\in\Gamma_{K}}\in\mathbb{R}^{\#\Gamma_{K}}\setminus\{\bm{0}\}.$
We also assume $V(\bm{\theta})$ is non-degenerate.
To handle $\mu_{\bm{n}}=V(\bm{n}\bm{\alpha}+\bm{\theta})$, we need the following Łojasiewicz type Lemma.
Lemma A.1 (cf. e.g., Lemma 4.2 in [JLS20]).
There exist some constants $C_{V}>0,0<c_{V}<1$ depending only on $V$ so that, for any $\eta>0,$ one has
${\rm meas}\left(\left\{\bm{\theta}\in[0,1]^{d}:\ |V(\bm{\theta})|\leq\eta\right\}\right)\leq C_{V}\eta^{c_{V}}.$
Next, we fix
$\bm{n}_{1}^{\prime},\cdots,\bm{n}_{D}^{\prime}\in\mathbb{Z}^{d}$
to be $D$ ($D\geq 2$) distinct lattice points. We are mainly concerned with Diophantine estimates for
$\bm{\omega}^{\prime}=(\omega_{1}^{\prime},\cdots,\omega_{D}^{\prime}),\ \omega_{s}^{\prime}=V({\bm{\theta}+\bm{n}_{s}^{\prime}\bm{\alpha}})\ (1\leq s\leq D).$
Throughout this Appendix, all constants $C,c>0$ are assumed to (depend only on $d,K,D,V$) be independent on $\bm{n}_{s}^{\prime}$ ($1\leq s\leq D$).
Our main result on the Diophantine estimates is
Theorem A.2.
Let $0<\eta<1$. We have for $R=R_{D}=D\cdot(\#\Gamma_{K})=D\frac{(2K+1)^{d}-1}{2}$,
${\rm meas}({\rm DC}_{\rm nls}(\eta))\geq 1-C\cdot(\max_{1\leq s\leq D}|\bm{n}_{s}^{\prime}|)^{5dR^{2}}\eta^{\frac{1}{10dR^{2}}},$
where $C=C(d,K,D,V)>0$ and
$\ \ \ {\rm DC}_{\rm nls}(\eta)$
$=\left\{(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}:\ |\bm{k}^{\prime}\cdot\bm{\omega}^{\prime}|\geq\frac{\eta}{|\bm{k}^{\prime}|^{4R^{2}}}\ {\rm for}\ \forall\ \bm{k}^{\prime}\in\mathbb{Z}^{D}\setminus\{\bm{0}\}\right\}.$
We also have for any $\bm{k}^{\prime}\in\mathbb{Z}^{D}\setminus\{\bm{0}\},$
${\rm meas}(\{(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}:\ |\bm{k}^{\prime}\cdot\bm{\omega}^{\prime}|\leq\eta\})\leq C\cdot(\max_{1\leq s\leq D}|\bm{n}_{s}^{\prime}|)^{5dR^{2}}\eta^{\frac{1}{10dR^{2}}}.$
Remark A.3.
In contrast, if $\bm{\omega}^{\prime}$ were a free parameter varying in $0,1]^{D}$, we would have
${\rm meas}\left(\left\{\bm{\omega}^{\prime}\in[0,1]^{D}:\ |\bm{k}^{\prime}\cdot\bm{\omega}^{\prime}|\geq\frac{\eta}{|\bm{k}^{\prime}|^{4R^{2}}}\ {\rm for}\ \forall\ \bm{k}^{\prime}\in\mathbb{Z}^{D}\setminus\{\bm{0}\}\right\}\right)=1-O(\eta).$
So the absence of independence of the coordinates of $\bm{\omega}^{\prime}$ may lead to the corresponding measure bound (from $1-O(\eta)$ to) $1-O(\eta^{\frac{1}{10dR^{2}}}).$
Recall that we have fixed in Theorem 1.1,
$\bm{n}_{1},\cdots,\bm{n}_{b}\in\mathbb{Z}^{d},\ \bm{\omega}^{(0)}=(V(\bm{n}_{1}\bm{\alpha}+\bm{\theta}),\cdots,V(\bm{n}_{b}\bm{\alpha}+\bm{\theta})),$
and for arbitrary $\bm{n}$, we have $\mu_{\bm{n}}=V(\bm{n}\bm{\alpha}+\bm{\theta}).$
As a consequence of Lemma A.1 and Theorem A.2, we obtain
Corollary A.4.
There exist some $0<c_{1}=c_{1}(b,d,K,V)<\frac{1}{100b},C_{1}=C_{1}(b,d,K,V)>1$ such that, if $0<\varepsilon+\delta\leq\delta_{0}(b,d,K,V,\max\limits_{1\leq\ell\leq b}|\bm{n}_{\ell}|)\ll 1$, then there is some $\mathcal{M}\subset[0,1]^{2d}$ satisfying
${\rm meas}([0,1]^{2d}\setminus\mathcal{M})\leq\log^{-1}\frac{1}{\varepsilon+\delta}$
so that the following properties hold true for $(\bm{\alpha},\bm{\theta})\in\mathcal{M}$ and $L_{\varepsilon,\delta}=100(\varepsilon+\delta)^{-c_{1}}$.
  • (i).
    For any $\bm{n}\neq\bm{n}^{\prime}$ satisfying $|\bm{n}|,|\bm{n}^{\prime}|\leq L_{\varepsilon,\delta}$,
    $|\mu_{\bm{n}}-\mu_{\bm{n}^{\prime}}|\geq(\varepsilon+\delta)^{\frac{1}{8b}}.$
  • (ii).
    For any $\bm{k}\in\mathbb{Z}^{b}$ satisfying $0<|\bm{k}|\leq 2L_{\varepsilon,\delta},$
    $|\bm{k}\cdot\bm{\omega}^{(0)}|\geq(\varepsilon+\delta)^{\frac{1}{8b}}.$
  • (iii).
    For any $\log\frac{1}{\varepsilon+\delta}\leq L\leq L_{\varepsilon,\delta}$ and all $(\bm{k},\bm{n})\in\Lambda_{L}\setminus\mathcal{S},$
    $\min_{\xi=\pm 1}|\xi\bm{k}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}}|\geq L^{-C_{1}}.$
  • (iv).
    For any $(\bm{k},\bm{n},\bm{n}^{\prime})\in\mathbb{Z}^{b}\times\mathbb{Z}^{d}\times\mathbb{Z}^{d}$ satisfying $|\bm{k}|\leq 2L_{\varepsilon,\delta},|(\bm{n},\bm{n}^{\prime})|\leq L_{\varepsilon,\delta}$ and $\bm{k}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}}-\mu_{\bm{n}^{\prime}}\not\equiv 0,$
    $|\bm{k}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}}-\mu_{\bm{n}^{\prime}}|\geq(\varepsilon+\delta)^{\frac{1}{8b}}.$
Proof.
It suffices to apply Lemma A.1 (i.e., $D=1$) and Theorem A.2 with $\eta=(\varepsilon+\delta)^{\frac{1}{8b}},L^{-C_{1}}$ and $D=2,b+1,b+2$. Without loss of generality, we only deal with (iii) and (iv) since (i) and (iv) can be handled similarly.
(iii). Denote by $\mathcal{M}_{{\rm iii}}$ the set of $(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}$ so that the conclusion of (iii) holds true.
For the case of $\bm{k}=\bm{0}$, applying Lemma A.1 yields for all $\bm{\alpha}\in[0,1]^{d},$
$\ \ \ {\rm meas}\left(\bigcup_{\log\frac{1}{\varepsilon+\delta}\leq L\leq L_{\varepsilon,\delta}}\bigcup_{|\bm{n}|\leq L}\left\{\bm{\theta}\in[0,1]^{d}:\ |\mu_{\bm{n}}=V(\bm{n}\bm{\alpha}+\bm{\theta})|\leq L^{-C_{1}}\right\}\right)$
$\leq\sum_{\log\frac{1}{\varepsilon+\delta}\leq L\leq L_{\varepsilon,\delta}}\sum_{|\bm{n}|\leq L}C_{V}L^{-C_{1}c_{V}}$
$\leq\sum_{\log\frac{1}{\varepsilon+\delta}\leq L\leq L_{\varepsilon,\delta}}C(V,d)L^{-C_{1}c_{V}+d}$
$\leq\log^{-3}\frac{1}{\varepsilon+\delta}\ ({\rm choosing}\ C_{1}\ {\rm\ so\ that}\ C_{1}c_{V}-d>10),$
where in the last inequality we assume $0<\varepsilon+\delta\leq\delta_{0}(V,d)\ll 1.$
For the case of $\bm{k}\neq\bm{0}$ and $\bm{n}\notin\{\bm{n}_{1},\cdots,\bm{n}_{b}\}$, we apply Theorem A.2 with $D=b+1,\eta=L^{-C_{1}}$ and $\bm{k}^{\prime}=(\pm\bm{k},1)$ to get for
$\mathcal{M}_{\rm iii}^{(1)}(L):=\bigcup_{|(\bm{k},\bm{n})|\leq L,\bm{n}\notin\{\bm{n}_{1},\cdots,\bm{n}_{b}\}}\left\{(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}:\ \min_{\xi=\pm 1}|\xi\bm{k}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}}|\leq L^{-C_{1}}\right\},$
the following
$\ \ \ {\rm meas}\left(\bigcup_{\log\frac{1}{\varepsilon+\delta}\leq L\leq L_{\varepsilon,\delta}}\mathcal{M}_{\rm iii}^{(1)}(L)\right)$
$\leq\sum_{\log\frac{1}{\varepsilon+\delta}\leq L\leq L_{\varepsilon,\delta}}C(b,d,K,V)L^{2d}L^{5dR^{2}_{b+1}}L^{-\frac{C_{1}}{10dR^{2}_{b+1}}}\ ({\rm assume}\ L>\max\limits_{1\leq\ell\leq b}|\bm{n}_{\ell}|)$
$\leq\sum_{\log\frac{1}{\varepsilon+\delta}\leq L\leq L_{\varepsilon,\delta}}C(b,d,K,V)L^{-10}\ ({\rm choosing}\ C_{1}\geq C({b,d,K})\gg 1)$
$\leq\log^{-3}\frac{1}{\varepsilon+\delta},$
where in the last inequality we assume $0<\varepsilon+\delta\leq\delta_{0}(b,d,K,V,\max\limits_{1\leq\ell\leq b}|\bm{n}_{\ell}|)\ll 1.$ The case of $\bm{k}\neq\bm{0}$ and $\bm{n}\in\{\bm{n}_{1},\cdots,\bm{n}_{b}\}$ can be handled similarly and we omit the details. Since there are only 3 cases to deal with, $C_{1}=C_{1}(b,d,K,V)>0$ can be well-defined.
By combining all the above estimates, we obtain the desired measure bound on $\mathcal{M}_{{\rm iii}}$, namely,
${\rm meas}([0,1]^{2d}\setminus\mathcal{M}_{\rm iii})\leq\log^{-2}\frac{1}{\varepsilon+\delta}.$
(iv). If $x,y\in\mathbb{R}$, we denote $x\wedge y=\min\{x,y\}$ and $x\vee y=\max\{x,y\}$. Denote by $\mathcal{M}_{{\rm iv}}$ the set of $(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}$ so that the conclusion of (iv) holds true. First we consider the case of $(\bm{k},\bm{n},\bm{n}^{\prime})\in\mathcal{T}_{1}$ with $\mathcal{T}_{1}$ the set of all $(\bm{k},\bm{n},\bm{n}^{\prime})$ satisfying: $0<|\bm{k}|\leq 2L_{\varepsilon,\delta},$ $|(\bm{n},\bm{n}^{\prime})|\leq L_{\varepsilon,\delta}$, $\bm{n}\neq\bm{n}^{\prime}$, ${\rm both}\ \bm{n}\ {\rm and}\ \bm{n}^{\prime}\ \notin\{\bm{n}_{1},\cdots,\bm{n}_{b}\}.$ We let
$\ \ \ \mathcal{M}_{\rm iv}^{(1)}:=\bigcup_{(\bm{k},\bm{n},\bm{n}^{\prime})\in\mathcal{T}_{1}}\left\{(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}:\ \min_{\xi=\pm 1}|\bm{k}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}}-\mu_{\bm{n}^{\prime}}|\leq(\varepsilon+\delta)^{\frac{1}{8b}}\right\}.$
So we can apply Theorem A.2 with $D=b+2,\eta=(\varepsilon+\delta)^{\frac{1}{8b}}$ and $\bm{k}^{\prime}=(\bm{k},1,-1)$ to get
${\rm meas}\left(\mathcal{M}_{\rm iv}^{(1)}\right)$ $\leq\sum_{(\bm{k},\bm{n},\bm{n}^{\prime})\in\mathcal{T}_{1}}C(b,d,K,V)((\max\limits_{1\leq\ell\leq b}|\bm{n}_{\ell}|)\vee|\bm{n}|\vee|\bm{n}^{\prime}|)^{5dR^{2}_{b+2}}(\varepsilon+\delta)^{\frac{1}{80bdR^{2}_{b+2}}}$
$\leq C(b,d,K,V)L_{\varepsilon,\delta}^{b+2d+5dR^{2}_{b+1}}(\varepsilon+\delta)^{\frac{1}{80bdR^{2}_{b+2}}}\ ({\rm assume}\ L_{\varepsilon,\delta}>\max\limits_{1\leq\ell\leq b}|\bm{n}_{\ell}|)$
$\leq C(b,d,K,V){(\varepsilon+\delta)}^{-c_{1}(b+2d+5dR^{2}_{b+1})+\frac{1}{80bdR^{2}_{b+2}}}$
$\leq{(\varepsilon+\delta)}^{\frac{1}{100bdR^{2}_{b+2}}}$
$({\rm choosing}\ 0<c_{1}\leq\frac{1}{(b+2d+5dR^{2}_{b+1})720bdR^{2}_{b+2}}),$
where in the last inequality we assume $0<\varepsilon+\delta\leq\delta_{0}(b,d,K,V,\max\limits_{1\leq\ell\leq b}|\bm{n}_{\ell}|)\ll 1.$
Next, we consider the case of $(\bm{k},\bm{n},\bm{n}^{\prime})$ with $\bm{k}\cdot\bm{\omega}^{(0)}+\mu_{\bm{n}}\equiv 0$. In this case, applying Lemma A.1 yields for all $\bm{\alpha}\in[0,1]^{d},$
$\ \ \ {\rm meas}\left(\bigcup_{|\bm{n}^{\prime}|\leq L_{\varepsilon,\delta}}\left\{\bm{\theta}\in[0,1]^{d}:\ |\mu_{\bm{n}^{\prime}}=V(\bm{n}^{\prime}\bm{\alpha}+\bm{\theta})|\leq(\varepsilon+\delta)^{\frac{1}{8b}}\right\}\right)$
$\leq\sum_{|\bm{n}^{\prime}|\leq L_{\varepsilon,\delta}}C_{V}(\varepsilon+\delta)^{\frac{c_{V}}{8b}}$
$\leq C(V,d)(\varepsilon+\delta)^{-c_{1}d+\frac{c_{V}}{8b}}$
$\leq(\varepsilon+\delta)^{\frac{c_{V}}{10b}}\ ({\rm choosing}\ 0<c_{1}\leq\frac{c_{V}}{72db}),$
where in the last inequality we assume $0<\varepsilon+\delta\leq\delta_{0}(b,d,V)\ll 1.$
Other cases are easier to handle, and we omit the details. By taking account of all the above estimates, we have
${\rm meas}([0,1]^{2d}\setminus\mathcal{M}_{\rm iv})\leq(\varepsilon+\delta)^{\frac{c_{V}}{1000bdR^{2}_{b+2}}}.$
Note that since there are only finitely many (say, at most 10) cases to deal with, $c_{1}=c_{1}(b,d,K,V)>0$ can be well-defined. ∎
A.1. Some useful lemmas
In this section, we will introduce some useful lemmas.
Lemma A.5 ([KM98]).
Let $I\subset\mathbb{R}$ be an interval of finite length (i.e., $0<|I|<\infty$) and $k\geq 1.$ If $f\in C^{k}(I;\mathbb{R})$ satisfies
$\inf_{x\in I}|\frac{d^{k}}{dx^{k}}f(x)|\geq A>0,$ (A.2)
then for all $\varepsilon>0,$
${\rm meas}(\{x\in I:\ |f(x)|\leq\varepsilon\})\leq\zeta_{k}(\frac{\varepsilon}{A})^{\frac{1}{k}},$ (A.3)
where $\zeta_{k}=k(k+1)((k+1)!)^{\frac{1}{k}}.$
Proof.
The following proof is taken from Kleinbock-Margulis [KM98].
From (A.2) and the Rolle’s theorem, the equation $\frac{d}{dx}f(x)=0$ has at most $k$ zeros on $I$. This then implies that the set $\{x\in I:\ |f(x)|\leq\varepsilon\}$ consists of at most $k+1$ intervals. Denote by $I_{1}$ one of those subintervals having maximal length. So we get
${\rm meas}(\{x\in I:\ |f(x)|\leq\varepsilon\})\leq(k+1)|I_{1}|.$
To prove (A.3), it remains to estimate $|I_{1}|>0$. We divide $I_{1}$ into $k$ equal parts by points $x_{1},\cdots,x_{k+1}.$ Let $P(x)$ be the Lagrange polynomial of $f$ on points $x_{1},\cdots,x_{k+1},$ namely,
$P(x)=\sum_{i=1}^{k+1}f(x_{i})\frac{\prod_{1\leq j\neq i\leq k+1}(x-x_{i})}{\prod_{1\leq j\neq i\leq k+1}(x_{j}-x_{i})}.$
By applying the Rolle’s theorem $k$ times and since $f(x_{i})=P(x_{i})$ ($1\leq i\leq k+1$), we can find $x_{0}\in I_{1}$ so that
$\frac{d^{k}}{dx^{k}}f(x_{0})=\frac{d^{k}}{dx^{k}}P(x_{0})=k!\sum_{i=1}^{k+1}\frac{f(x_{i})}{\prod_{1\leq j\neq i\leq k+1}(x_{j}-x_{i})},$
which together with (A.2) yields
$A\leq k!\sum_{i=1}^{k+1}\frac{\varepsilon}{|I_{1}|^{k}k^{-k}}.$
This implies
$|I_{1}|\leq k(k+1)!)^{\frac{1}{k}}(\frac{\varepsilon}{A})^{\frac{1}{k}}.$
We have finished the proof. ∎
In the following, we turn to the analysis of multi-variable functions. We first introduce the notations.
  • For any ${\bm{x}}=(x_{1},\cdots,x_{d})$ and any $d\geq 1$, let
    $|{\bm{x}}|_{2}=\sqrt{\sum_{i=1}^{d}|x_{i}|^{2}},\ |\bm{x}|_{1}=\sum_{i=1}^{d}|x_{i}|.$
  • Let $\bm{a}=(a_{1},\cdots,a_{d}),\bm{b}=(b_{1},\cdots,b_{d})$. Define the $d$-dimensional interval to be
    $I=I_{\bm{a},\bm{b}}=\prod_{i=1}^{d}[a_{i},b_{i}]\subset\mathbb{R}^{d}.$ (A.4)
    Denote
    $|\bm{a}\vee\bm{b}|=\sum_{i=1}^{d}\max(|a_{i}|,|b_{i}|).$
  • For ${\bm{\beta}}=(\beta_{1},\cdots,\beta_{d})\in\mathbb{R}^{d}\setminus\{\bm{0}\}$, define
    $d_{\bm{\beta}}:=\sum_{i=1}^{d}\beta_{i}{\partial}_{i},\ \partial_{i}:=\frac{\partial}{\partial x_{i}}.$
    We also write for $\ell\geq 1,$
    $d_{\bm{\beta}}^{\ell}=d_{\bm{\beta}}\cdots d_{\bm{\beta}}\ (\ell-{\rm times}).$
    Denote
    $\nabla=(\partial_{1},\cdots,\partial_{d}).$
  • For a multi-index ${\bm{\gamma}}=(\gamma_{1},\cdots,\gamma_{d})\in\mathbb{N}^{d}$, denote
    $|\bm{\gamma}|=\sum_{i=1}^{d}\gamma_{i},\ \partial^{\bm{\gamma}}=\partial^{\gamma_{1}}_{1}\cdots\partial^{\gamma_{d}}_{d}.$
We have
Lemma A.6.
Fix $k\in\mathbb{N}$. Let $I=I_{\bm{a},\bm{b}}\subset\mathbb{R}^{d}$ be defined by (A.4). Assume that the function $f\in C^{k+1}(I;\mathbb{R})$ satisfies for some $A>0,$
$\inf_{\bm{x}\in I}\sup_{1\leq\ell\leq k}\left|{d_{\bm{\beta}}^{\ell}}f(\bm{x})\right|\geq A.$ (A.5)
Let
$\|f\|_{k+1}=\sup_{\bm{x}\in I}\sup_{1\leq|\bm{\gamma}|\leq k+1}\left|{\partial^{\bm{\gamma}}}f(\bm{x})\right|<\infty.$
Then for $0<\varepsilon<1,$
$\ \ \ {\rm meas}(\{\bm{x}\in I:\ |f(\bm{x})|\leq\varepsilon\})$
$\leq C(\bm{\beta},k,d)(\|f\|_{k+1}|\bm{b}-\bm{a}|_{2}+1)^{d}{|\bm{a}\vee\bm{b}|}^{d-1}A^{-d}(A\wedge 1)^{-1}{\varepsilon}^{\frac{1}{k}},$ (A.6)
where $C=C(\bm{\beta},k,d)>0$ depends only on $\bm{\beta},k,d$ (but not on $f$).
Remark A.7.
If $f$ is a polynomial on $\mathbb{R}^{d}$, [CW01] proved some strong distributional inequalities, which may directly imply (A.6). However, in the present case, we have a weak transversality condition (A.5). So we would like to give an elementary proof applying for more general functions (cf. [LSZ25] for an application), which also produces an upper bound in the estimate (A.6) with the explicit dependence on $\|f\|_{k+1}$.
Proof.
The proof can be divided into two steps:
Step 1. Assume that for some $1\leq k_{1}\leq k$ and interval $I_{1}=\prod_{i=1}^{d}[h_{i},m_{i}],$
$\inf_{\bm{x}\in I_{1}}\left|{d_{\bm{\beta}}^{k_{1}}}f(\bm{x})\right|\geq A.$ (A.7)
We will combine Lemma A.5 and Fubini’s theorem to deal with the measure estimate. Let $\bm{e}_{1}=\frac{\bm{\beta}}{|\bm{\beta}|_{2}}$ and choose a normalized orthogonal basis $\bm{e}_{2},\cdots,\bm{e}_{d}\in\mathbb{R}^{d}$ of
$\bm{\beta}^{\perp}=\{\bm{x}\in\mathbb{R}^{d}:\ \bm{x}\cdot\bm{\beta}=0\}.$
We can then define the coordinate transform:
$\bm{x}=\bm{\varphi}(s,\bm{y}),(s,\bm{y})\in\mathbb{R}\times\mathbb{R}^{d-1},$
via
$\bm{\varphi}(s,\bm{y})=s\bm{e}_{1}+\sum_{i=2}^{d}y_{i}\bm{e}_{i}.$
So the Jacobian satisfies
$|\det\frac{\partial\bm{x}}{\partial(s,\bm{y})}|=1.$
Define
$S_{\varepsilon}=\{\bm{x}\in I_{1}:\ |f(\bm{x})|\leq\varepsilon\},$
and denote by $\chi_{(\cdot)}$ the indicator function. We have
${\rm meas}(S_{\varepsilon})$ $=\int_{I_{1}}\chi_{S_{\varepsilon}}(\bm{x})d\bm{x}$
$=\int_{\varphi^{-1}(I_{1})}\chi_{S_{\varepsilon}}(\bm{\varphi}(s,\bm{y}))dsd\bm{y}.$ (A.8)
We proceed to control $\bm{\varphi}^{-1}(I_{1})$. From $\bm{x}=s\bm{e}_{1}+\sum_{i=2}^{d}y_{i}\bm{e}_{i}\in I_{1}$, we obtain
$s=\bm{x}\cdot\bm{e}_{1}\in[-|\bm{h}\vee\bm{m}|,|\bm{h}\vee\bm{m}|].$
Similarly, we also have $y_{i}\in[-|\bm{h}\vee\bm{m}|,|\bm{h}\vee\bm{m}|]$ for $2\leq i\leq d,$ which yields $\bm{\varphi}^{-1}(I_{1})\subset[-|\bm{h}\vee\bm{m}|,|\bm{h}\vee\bm{m}|]^{d}$. We define
$\Pi_{1}\bm{\varphi}^{-1}(I_{1})=\{\bm{y}\in\mathbb{R}^{d-1}:\ \exists\ s\in\mathbb{R}\ s.t.,\ (s,\bm{y})\in\bm{\varphi}^{-1}(I_{1})\}$
and, define for $\bm{y}\in\Pi_{1}\bm{\varphi}^{-1}(I_{1}),$ the interval $I_{\bm{y}}=\{s\in\mathbb{R}:\ (s,\bm{y})\in\bm{\varphi}^{-1}(I_{1})\}$ of finite length. Applying the Fubini’s theorem together with (A.8) implies that
${\rm meas}(S_{\varepsilon})$ $=\int_{\Pi_{1}\bm{\varphi}^{-1}(I_{1})}d\bm{y}\int_{I_{\bm{y}}}\chi_{S_{\varepsilon}}(\bm{\varphi}(s,\bm{y}))ds$
$\leq{\rm meas}_{d-1}(\Pi_{1}\bm{\varphi}^{-1}(I_{1}))\sup_{\bm{y}\in\Pi_{1}\bm{\varphi}^{-1}(I_{1})}\int_{I_{\bm{y}}}\chi_{S_{\varepsilon}}(\bm{\varphi}(s,\bm{y}))ds$
$\leq 2^{d-1}|\bm{h}\vee\bm{m}|^{d-1}\sup_{\bm{y}\in\Pi_{1}\bm{\varphi}^{-1}(I_{1})}\int_{I_{\bm{y}}}\chi_{S_{\varepsilon}}(\bm{\varphi}(s,\bm{y}))ds$
$\ \ \ ({\rm since}\ \Pi_{1}\bm{\varphi}^{-1}(I_{1})\subset[-|\bm{h}\vee\bm{m}|,|\bm{h}\vee\bm{m}|]^{d-1}).$
For fixed $\bm{y}\in\Pi_{1}\bm{\varphi}^{-1}(I_{1})$, we get
$\ \ \ \int_{I_{\bm{y}}}\chi_{S_{\varepsilon}}(\bm{\varphi}(s,\bm{y}))ds$
$={\rm meas}(\{s\in I_{\bm{y}}:\ |f(s\bm{e}_{1}+\sum_{i=2}^{d}y_{i}\bm{e}_{i})|\leq\varepsilon\}).$
Denote $g(s)=f(s\bm{e}_{1}+\sum_{i=2}^{d}y_{i}\bm{e}_{i})$. From (A.7), we have for all $s\in I_{\bm{y}},$
$|\frac{d^{k_{1}}}{ds^{k_{1}}}g(s)|=|\frac{1}{|\bm{\beta}|_{2}^{k_{1}}}d^{k_{1}}_{\bm{\beta}}f(s\bm{e}_{1}+\sum_{i=2}^{d}y_{i}\bm{e}_{i})|\geq\frac{A}{|\bm{\beta}|_{2}^{k_{1}}}.$
So applying Lemma A.5 yields
${\rm meas}(S_{\varepsilon})\leq\zeta_{k_{1}}2^{d-1}|\bm{h}\vee\bm{m}|^{d-1}|\bm{\beta}|_{2}(\frac{\varepsilon}{A})^{\frac{1}{k_{1}}}.$ (A.9)
Step 2. We deal with the general case in this step. So we first divide each one dimensional interval $a_{i},b_{i}]$ into $N$ (will be specified below) equal subintervals $I_{i,j_{i}}$ ($1\leq j_{i}\leq N$). This then induces a decomposition of $I$, namely,
$I=\bigcup_{1\leq j_{1},\cdots,j_{d}\leq N}\prod_{i=1}^{d}I_{i,j_{i}}=\bigcup_{\bm{J}\in([1,N]\cap\mathbb{N})^{d}}I_{\bm{J}}.$
We will apply the argument proved in Step 1 on each $I_{\bm{J}}$. For this purpose, we fix $I_{\bm{J}}$. For any $\bm{x},\bm{y}\in I_{\bm{J}}$, we have
$|\bm{x}-\bm{y}|_{2}$ $\leq\sqrt{\sum_{i=1}^{d}\frac{|b_{i}-a_{i}|^{2}}{N^{2}}}$
$\leq\frac{|\bm{b}-\bm{a}|_{2}}{N}.$
Fixing any $\bm{x}_{0}\in I_{\bm{J}}$, we get from the assumption (A.5) that there exists $1\leq k_{1}\leq k$ with
$|d_{\bm{\beta}}^{k_{1}}f(\bm{x}_{0})|\geq A>0.$
As a result, for any $\bm{x}\in I_{\bm{J}}$, we obtain
$|d_{\bm{\beta}}^{k_{1}}f(\bm{x})|$ $\geq|d_{\bm{\beta}}^{k_{1}}f(\bm{x}_{0})|-|d_{\bm{\beta}}^{k_{1}}f(\bm{x})-d_{\bm{\beta}}^{k_{1}}f(\bm{x}_{0})|$
$\geq A-\sup_{\bm{x}\in I_{\bm{J}}}|\nabla d_{\bm{\beta}}^{k_{1}}f(\bm{x})|_{2}\cdot|\bm{x}-\bm{x}_{0}|_{2}$
$\geq A-\frac{\|f\|_{k_{1}+1}L_{\bm{\beta},d}|\bm{b}-\bm{a}|_{2}}{N}$
$\geq A-\frac{\|f\|_{k+1}L_{\bm{\beta},d}|\bm{b}-\bm{a}|_{2}}{N},$ (A.10)
where
$L_{\bm{\beta},d}=\sqrt{d}\sum_{1\leq l_{1},\cdots,l_{d}\leq d}|\beta_{l_{1}}\cdots\beta_{l_{d}}|.$
In fact, $L_{\bm{\beta},d}$ is derived from
$|\nabla d_{\bm{\beta}}^{k_{1}}f(\bm{x})|_{2}$ $=\sqrt{\sum_{j=1}^{d}|\sum_{1\leq l_{1},\cdots,l_{d}\leq d}\beta_{l_{1}}\cdots\beta_{l_{d}}\partial_{jl_{1}\cdots l_{d}}^{k_{1}+1}f(\bm{x})|^{2}}$
$\leq\|f\|_{k_{1}+1}\sqrt{\sum_{j=1}^{d}(\sum_{1\leq l_{1},\cdots,l_{d}\leq d}|\beta_{l_{1}}\cdots\beta_{l_{d}}|)^{2}}$
$:=\|f\|_{k_{1}+1}L_{\bm{\beta},d}.$
From (A.10), we can set
$N=[2\|f\|_{k+1}L_{\bm{\beta},d}|\bm{b}-\bm{a}|_{2}A^{-1}]+1$
so that
$\inf_{\bm{x}\in I_{\bm{J}}}|d_{\bm{\beta}}^{k_{1}}f(\bm{x})|\geq A/2>0,$
where $[x]$ denotes the integer part of $x\in\mathbb{R}.$ Thus applying (A.9) yields (since $I_{J}\subset I,0<\varepsilon<1$)
${\rm meas}(\{\bm{x}\in I_{\bm{J}}:\ |f(\bm{x})|\leq\varepsilon\})$ $\leq\zeta_{k_{1}}2^{d-1}|\bm{a}\vee\bm{b}|^{d-1}2^{\frac{1}{k_{1}}}|\bm{\beta}|_{2}(\frac{\varepsilon}{A})^{\frac{1}{k_{1}}}$
$\leq C(k,d)|\bm{a}\vee\bm{b}|^{d-1}|\bm{\beta}|_{2}(A\wedge 1)^{-1}\varepsilon^{\frac{1}{k}}.$
Collecting all $I_{\bm{J}}$ leads to,
$\ \ \ {\rm meas}(\{\bm{x}\in I:\ |f(\bm{x})|\leq\varepsilon\})$
$\leq C(k,d)N^{d}|\bm{a}\vee\bm{b}|^{d-1}|\bm{\beta}|_{2}(A\wedge 1)^{-1}\varepsilon^{\frac{1}{k}}$
$\leq C(\bm{\beta},k,d)(\|f\|_{k+1}|\bm{b}-\bm{a}|_{2}A^{-1}+1)^{d}|\bm{a}\vee\bm{b}|^{d-1}(A\wedge 1)^{-1}\varepsilon^{\frac{1}{k}}$
$\leq C(\bm{\beta},k,d)(\|f\|_{k+1}|\bm{b}-\bm{a}|_{2}+1)^{d}|\bm{a}\vee\bm{b}|^{d-1}A^{-d}(A\wedge 1)^{-1}{\varepsilon}^{\frac{1}{k}}.$
We have proven (A.6). ∎
Finally, we introduce a key lemma (cf. Proposition in Appendix B, [BGG85] for a more precise form) on determinant estimates.
Lemma A.8.
Let $\bm{v}^{(1)},\cdots,\bm{v}^{(r)}\in\mathbb{R}^{r}$ be $r$ linearly independent vectors with $|\bm{v}^{(\ell)}|_{1}\leq M$ for $1\leq\ell\leq r.$ Then for any $\bm{w}\in\mathbb{R}^{r}$, we have
$\max_{1\leq\ell\leq r}|\bm{w}\cdot\bm{v}^{(\ell)}|\geq r^{-3/2}M^{1-r}|\bm{w}|_{2}\cdot|\det\left[\bm{v}^{(\ell)}\right]_{1\leq\ell\leq r}|.$
Proof.
For completeness, we give a proof based on Cramer’s rule and the Hadamard’s inequality. We assume $\bm{v}^{(\ell)}$ is the row vector in $\mathbb{R}^{r}$ ($1\leq\ell\leq r$). Denote by $X=[\bm{v}^{(1)},\cdots,\bm{v}^{(r)}]$ the $r\times r$ matrix and consider the equation
$X\bm{w}=\bm{\xi},\ \bm{\xi}=(\xi_{1},\cdots,\xi_{r})\ {\rm with}\ \xi_{\ell}=\bm{w}\cdot\bm{v}^{(\ell)}.$
By the Cramer’s rule, we obtain
$\bm{w}=X^{-1}\bm{\xi}=\frac{X^{*}}{\det X}\bm{\xi},$
where $X^{*}=[\tilde{v}_{ij}]_{1\leq i,j\leq r}$ denotes the adjugate matrix of $X$. As a result, we get
$|\bm{w}|_{2}\cdot|\det X|$ $=|X^{*}\bm{\xi}|_{2}$
$=\sqrt{\sum_{i=1}^{r}|\sum_{j=1}^{r}\tilde{v}_{ij}\xi_{j}|^{2}}$
$\leq\max_{1\leq j\leq r}|\xi_{j}|\sqrt{\sum_{i=1}^{r}(\sum_{j=1}^{r}|\tilde{v}_{ij}|)^{2}}$
$\leq r^{3/2}\max_{1\leq j\leq r}|\xi_{j}|\max_{1\leq i,j\leq r}|\tilde{v}_{ij}|.$
It remains to bound $\max_{1\leq i,j\leq r}|\tilde{v}_{ij}|,$ which will be completed by using the Hadamard’s inequality. In fact, we have
$|\tilde{v}_{ij}|$ $\leq\prod_{1\leq\ell\neq i\leq r}\sqrt{\sum_{1\leq m\neq j\leq r}|v^{(\ell)}_{m}|^{2}}$
$\leq\prod_{1\leq\ell\neq i\leq r}{\sum_{1\leq m\neq j\leq r}|v^{(\ell)}_{m}|}$
$\leq M^{r-1}\ ({\rm since}\ \max_{1\leq\ell\leq r}|\bm{v}^{(\ell)}|_{1}\leq M).$
This finishes the proof. ∎
A.2. Proof of Theorem A.2
In this section, we prove Theorem A.2.
Proof of Theorem A.2.
Recall that
$\bm{k}^{\prime}\cdot\bm{\omega}^{\prime}$ $=\sum_{s=1}^{D}k_{s}^{\prime}V(\bm{\theta}+\bm{n}_{s}^{\prime}\bm{\alpha})$
$=\sum_{s=1}^{D}\sum_{\bm{\ell}\in\Gamma_{K}}k_{s}^{\prime}v_{\bm{\ell}}\cos 2\pi\bm{\ell}\cdot(\bm{\theta}+\bm{n}_{s}^{\prime}\bm{\alpha})$
$=(\bm{k}^{\prime}\otimes\bm{v})\cdot\bm{V},$
where
$\bm{k}^{\prime}\otimes\bm{v}=(k_{s}^{\prime}v_{\bm{\ell}})_{1\leq s\leq D,\bm{\ell}\in\Gamma_{K}},$
$\bm{V}=(V_{(s,\bm{\ell})})_{1\leq s\leq D,\bm{\ell}\in\Gamma_{K}}:=(\cos 2\pi\bm{\ell}\cdot(\bm{\theta}+\bm{n}_{s}^{\prime}\bm{\alpha}))_{1\leq s\leq D,\bm{\ell}\in\Gamma_{K}}.$
Denote $\bm{\xi}=(\bm{\beta},\bm{q})$ for $\bm{\beta},\bm{q}\in[0,1]^{d}$. Write also
$d_{\bm{\xi}}=\sum_{i=1}^{d}(\beta_{i}\frac{\partial}{\partial\alpha_{i}}+q_{i}\frac{\partial}{\partial\theta_{i}})$
and accordingly $d_{\bm{\xi}}^{\ell}$ for $\ell\geq 1.$
A key observation is that for all $j\geq 1$ and $\bm{\ell}\in\Gamma_{K}$,
$\ \ \ d_{\bm{\xi}}^{2j}\cos 2\pi\bm{\ell}\cdot(\bm{\theta}+\bm{n}_{s}^{\prime}\bm{\alpha})$
$=(-(2\pi(\bm{\ell}\bm{n}_{s}^{\prime})\cdot\bm{\beta}+2\pi\bm{q}\cdot\bm{\ell})^{2})^{j}\cos 2\pi\bm{\ell}\cdot(\bm{\theta}+\bm{n}_{s}^{\prime}\bm{\alpha}).$
This motivates us to consider the Wronskian
$W=[d_{\bm{\xi}}^{2j}{V}_{(s,\bm{\ell})}]_{s,\bm{\ell}}^{1\leq j\leq R}\ {\rm with}\ R=D\cdot(\#\Gamma_{K}),$
which is a $R\times R$ real matrix. Direct computations show
$\ \ \ |\det W|$
$=\prod_{s,\bm{\ell}}|\cos 2\pi{\bm{\ell}}\cdot(\bm{\theta}+\bm{n}_{s}^{\prime}\bm{\alpha})|\times\prod_{s,\bm{\ell}}(2\pi)^{2}((\bm{\ell}\bm{n}_{s}^{\prime})\cdot\bm{\beta}+\bm{q}\cdot\bm{\ell})^{2}\times|\det W_{1}|,$ (A.11)
where $W_{1}$ is a Vandermonde matrix with
$|\det W_{1}|$ $=\sqrt{\prod_{(s,\bm{\ell})\neq(s^{\prime},\bm{\ell}^{\prime})}(2\pi)^{2}|((\bm{\ell}\bm{n}_{s}^{\prime})\cdot\bm{\beta}+\bm{q}\cdot\bm{\ell})^{2}-((\bm{\ell}^{\prime}\bm{n}_{s^{\prime}}^{\prime})\cdot\bm{\beta}+\bm{q}\cdot\bm{\ell}^{\prime})^{2}|}.$ (A.12)
Then the proof can be divided into three steps:
Step 1. Prior conditions on $(\bm{\alpha},\bm{\theta})$
We aim to get a lower bound for (A.11), which requires the prior restrictions on $\bm{\alpha},\bm{\theta}.$ Let $\|x\|_{\mathbb{T}/2}={\rm dist}(x,\mathbb{Z}/2)$. From
$4\|x-\frac{1}{4}\|_{\mathbb{T}/2}\leq|\cos 2\pi x|\leq 2\pi\|x-\frac{1}{4}\|_{\mathbb{T}/2},$
we can define for $\delta>0$ the resonant set (of $\bm{\alpha},\bm{\theta}$)
$\ \ \ S_{\delta}$
$=\bigcup_{1\leq s\leq D,\bm{\ell}\in\Gamma_{K}}\left\{(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}:\ \|\bm{\ell}\cdot(\bm{\theta}+\bm{n}_{s}^{\prime}\bm{\alpha})-\frac{1}{4}\|_{\mathbb{T}/2}\leq\delta\right\}.$ (A.13)
Denote $L=\max\limits_{1\leq s\leq D}K|\bm{n}_{s}^{\prime}|$. We have
Lemma A.9.
Let $0<\delta<1$. Then $S_{\delta}$ can be covered by $CL^{2d-1}\delta^{-2d+1}$ intervals of side length $\leq\delta,$ where $C=C(d,R)>0.$
Proof.
It suffices to cover for each $\bm{n}_{s}^{\prime},\bm{\ell}$ the set
$S_{\bm{n}_{s}^{\prime},\bm{\ell}}=\{(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}:\ \|\bm{\ell}\cdot(\bm{\theta}+\bm{n}_{s}^{\prime}\bm{\alpha})-\frac{1}{4}\|_{\mathbb{T}/2}\leq\delta\}$
with $C(d)L^{2d-1}\delta^{-2d+1}$ intervals of side length $\leq\delta$ since the total number of all $\bm{n}_{s}^{\prime},\bm{\ell}$ is at most $R.$ Since $|\bm{\ell}|>0,$ we may assume $\ell_{1}\neq 0$ without loss of generality. For any $\bm{x}=(x_{1},\cdots,x_{d}),$ write $\bm{x}=(x_{1},\bm{x}^{c}_{1})$. So we can divide $0,1]^{d}\times[0,1]^{d-1}\ni(\bm{\alpha},\bm{\theta}_{1}^{c})$ into $C_{1}=[100L\delta^{-1}]^{2d-1}$ cubes $\bm{C}_{i}$ ($1\leq i\leq C_{1}$) with disjoint exteriors of the same side length $\leq\frac{\delta}{L}$. Denote by $(\bm{\alpha}_{i},\bm{\theta}^{c}(i))\in[0,1]^{d}\times[0,1]^{d-1}$ ($1\leq i\leq C_{1}$) the centers of those cubes. We further define
$\ \ \ I_{\bm{n}_{s}^{\prime},\bm{\ell}}$
$=\bigcup_{1\leq i\leq C_{1}}\{(\bm{\alpha},\bm{\theta}_{1}^{c},\theta_{1})\in\bm{C}_{i}\times[0,1]:\ \|\ell_{1}\theta_{1}+\bm{\ell}_{1}^{c}\cdot\bm{\theta}^{c}(i)+(\bm{n}_{s}^{\prime}\bm{\ell})\cdot\bm{\alpha}_{i}-\frac{1}{4}\|_{\mathbb{T}/2}\leq 2\delta\}$
$:=\bigcup_{1\leq i\leq C_{1}}J_{i}.$
We first claim that $S_{\bm{n}_{s}^{\prime},\bm{\ell}}\subset I_{\bm{n}_{s}^{\prime},\bm{\ell}}.$ Let $(\bm{\alpha},\bm{\theta})\in S_{\bm{n}_{s}^{\prime},\bm{\ell}}$. Then by the definition (of $S_{\bm{n}_{s}^{\prime},\bm{\ell}}$), we have
$\|\ell_{1}\theta_{1}+\bm{\ell}_{1}^{c}\cdot\bm{\theta}_{1}^{c}+(\bm{\ell}\bm{n}_{s}^{\prime})\cdot\bm{\alpha}-\frac{1}{4}\|_{\mathbb{T}/2}\leq\delta.$
Since $\bm{C}_{i}$ ($1\leq i\leq C_{1}$) covers $0,1]^{2d-1}$, there is $1\leq i_{0}\leq C_{1}$ so that $(\bm{\alpha},\bm{\theta}_{1}^{c})\in\bm{C}_{i_{0}}$, namely, $|\bm{\alpha}-\bm{\alpha}_{i_{0}}|+|\bm{\theta}_{1}^{c}-\bm{\theta}^{c}(i_{0})|\leq\frac{\delta}{2L}.$ This implies that
$\ \ \ \|\ell_{1}\theta_{1}+\bm{\ell}_{1}^{c}\cdot\bm{\theta}^{c}(i_{0})+(\bm{n}_{s}^{\prime}\bm{\ell})\cdot\bm{\alpha}_{i_{0}}-\frac{1}{4}\|_{\mathbb{T}/2}$
$\leq\|\bm{\ell}\cdot(\bm{\theta}+\bm{n}_{s}^{\prime}\bm{\alpha})-\frac{1}{4}\|_{\mathbb{T}/2}+|\bm{\ell}_{1}^{c}\cdot(\bm{\theta}^{c}(i_{0})-\bm{\theta}_{1}^{c})|$
$\ \ +|(\bm{\ell}\bm{n}_{s}^{\prime})\cdot(\bm{\alpha}-\bm{\alpha}_{i_{0}})|$
$\leq 2\delta.$
The claim is proved. So it remains to cover each $J_{i}$ with intervals of side length $\leq\delta$. This can be accomplished by the fact that the set $\{\theta_{1}\in[0,1]:\ \|\ell_{1}\theta_{1}+x\|_{\mathbb{T}/2}\leq 2\delta\}$ ($x\in\mathbb{R}$) consists of $r\sim|\ell_{1}|$ one dimensional intervals of length $2\delta/|\ell_{1}|$.
We have finished the proof. ∎
Given $0<\delta<1/10,$ we divide $[0,1]$ into $N=[\delta^{-2}]+1$ intervals of equal length $\frac{1}{N}\sim\delta^{2}$. This then induces a partition of $0,1]^{2d}$ into cubes of equal side length $1/N$. We denote by $\Lambda_{1/N}(\bm{x}_{i})$ ($1\leq i\leq N^{2d}$) all these cubes. We have
Lemma A.10.
Let $0<\delta<1/10$ and let $S_{\delta}$ be defined by (A.13). Then we have for $N=[\delta^{-2}]+1$,
$\#\{1\leq i\leq N^{2d}:\ \Lambda_{1/N}(\bm{x}_{i})\cap S_{\delta}\neq\emptyset\}\leq C(d,R)L^{2d-1}\delta^{-4d+1}.$ (A.14)
In particular, we have
${\rm meas}\left(\bigcup_{1\leq i\leq N^{2d}:\ \Lambda_{1/N}(\bm{x}_{i})\cap S_{\delta}\neq\emptyset}\Lambda_{1/N}(\bm{x}_{i})\right)\leq C(d,R)L^{2d-1}\delta.$ (A.15)
Proof.
Let $\bm{C}_{i}$ be given as in the proof of Lemma A.9. Then $S_{\delta}\subset\cup_{1\leq i\leq C(d,R)L^{2d-1}\delta^{-2d+1}}\bm{C}_{i}$. So it suffices to bound
$L_{1}=\#\{1\leq i\leq N^{2d}:\ (\cup_{1\leq j\leq CL^{2d-1}\delta^{-2d+1}}\bm{C}_{j})\cap\Lambda_{1/N}(\bm{x}_{i})\neq\emptyset\}.$
Note that if $\Lambda_{1/N}(\bm{x}_{i})\cap\bm{C}_{j}\neq\emptyset$, then
$\Lambda_{1/N}(\bm{x}_{i})\subset\bm{C}_{j}^{\prime},$
where $\bm{C}_{j}^{\prime}$ is also an interval containing $\bm{C}_{j}$ and satisfying
${\rm dist}(\partial\bm{C}_{j},\partial\bm{C}_{j}^{\prime})\leq 10/N,$
with $\partial$ the boundary. Regarding the disjointness (of the exterior) of $\Lambda_{1/N}(\bm{x}_{i})$ for different $i$, we must have
$L_{1}\leq CL^{2d-1}\delta^{-2d+1}(\delta+10/N)^{2d}\delta^{-4d}\leq CL^{2d-1}\delta^{-4d+1}.$
This implies the bound (A.14). Then the measure bound follows directly from (A.15). ∎
From this lemma, we get immediately that
Lemma A.11.
Let $0<\delta<1/10$ and define $N=[\delta^{-2}]+1$. Then there is a collection of cubes $\Lambda_{1/N}(\bm{x}_{i})\subset[0,1]^{d+d}$ ($1\leq i\leq N_{1}$) with disjoint exteriors so that
${\rm meas}\left([0,1]^{2d}\setminus\cup_{1\leq i\leq N_{1}}\Lambda_{1/N}(\bm{x}_{i})\right)\leq CL^{2d-1}\delta,$
and if $(\bm{\alpha},\bm{\theta})\in\cup_{1\leq i\leq N_{1}}\Lambda_{1/N}(\bm{x}_{i}),$ then
$\prod_{1\leq s\leq D,\bm{\ell}\in\Gamma_{K}}|\cos 2\pi\bm{\ell}\cdot(\bm{\theta}+\bm{n}_{s}^{\prime}\bm{\alpha})|\geq c\delta^{R},\ c=c(d,R)>0.$
Step 2. Conditions on $(\bm{\beta},\bm{q})$
Next we impose conditions on $(\bm{\beta},\bm{q})$.
Lemma A.12.
Let $\epsilon\in(0,1)$ and $\bm{B}\subset[0,1]^{d}\times[0,1]^{d}$ be the set of $\bm{\xi}=(\bm{\beta},\bm{q})$ satisfying
$\min_{s,\bm{\ell}}((\bm{\ell}\bm{n}_{s}^{\prime})\cdot\bm{\beta}+\bm{q}\cdot\bm{\ell})^{2}\geq\epsilon>0,$
$\min_{(s,\bm{\ell})\neq(s^{\prime},\bm{\ell}^{\prime})}|(\bm{\ell}\bm{n}_{s}^{\prime}-\bm{\ell}^{\prime}\bm{n}_{s^{\prime}}^{\prime})\cdot\bm{\beta}+\bm{q}\cdot(\bm{\ell}-\bm{\ell}^{\prime})|\geq\epsilon>0,$
$\min_{(s,\bm{\ell})\neq(s^{\prime},\bm{\ell}^{\prime})}|(\bm{\ell}\bm{n}_{s}^{\prime}+\bm{\ell}^{\prime}\bm{n}_{s^{\prime}}^{\prime})\cdot\bm{\beta}+\bm{q}\cdot(\bm{\ell}+\bm{\ell}^{\prime})|\geq\epsilon>0.$
Then
${\rm meas}\left([0,1]^{2d}\setminus\bm{B}\right)\leq 3R^{2}\sqrt{\epsilon}.$
Moreover, for $\bm{\xi}\in\bm{B},$ we have
$\ \ \ \prod_{s,\bm{\ell}}((\bm{\ell}\bm{n}_{s}^{\prime})\cdot\bm{\beta}+\bm{q}\cdot\bm{\ell})^{2}\times|\det W_{1}|$
$=\prod_{s,\bm{\ell}}((\bm{\ell}\bm{n}_{s}^{\prime})\cdot\bm{\beta}+\bm{q}\cdot\bm{\ell})^{2}$
$\ \ \times\prod_{(s,\bm{\ell})\neq(s^{\prime},\bm{\ell}^{\prime})}|((\bm{\ell}\bm{n}_{s}^{\prime})\cdot\bm{\beta}+\bm{q}\cdot\bm{\ell})^{2}-((\bm{\ell}^{\prime}\bm{n}_{s^{\prime}}^{\prime})\cdot\bm{\beta}+\bm{q}\cdot\bm{\ell}^{\prime})^{2}|$
$\geq\epsilon^{R^{3}}.$
Proof.
The proof is based on the Fubini’s theorem, (A.1) and the assumption on $\bm{n}_{s}^{\prime}$ ($1\leq s\leq D$). Actually, in all cases we have
$(\bm{\ell}\bm{n}_{s}^{\prime},\bm{\ell})\neq\bm{0},$
$(\bm{\ell}\bm{n}_{s}^{\prime}-\bm{\ell}^{\prime}\bm{n}_{s^{\prime}}^{\prime},\bm{\ell}-\bm{\ell}^{\prime})\neq\bm{0},$
$(\bm{\ell}\bm{n}_{s}+\bm{\ell}^{\prime}\bm{n}_{s^{\prime}}^{\prime},\bm{\ell}+\bm{\ell}^{\prime})\neq\bm{0},$
which combined with $s\leq D$ and $\bm{\ell}\in\Gamma_{K}$ leads to the excision of $(\bm{\beta},\bm{q})$ in a set of measure $O(\sqrt{\epsilon})$. ∎
Step 3. Usage of the Lemma A.8
Now combining Lemma A.11 and Lemma A.12 (we set $\epsilon=10^{-2}R^{-4}$) leads to
Lemma A.13.
Let $0<\delta<1/10$ and define $N=[\delta^{-2}]+1$. Then there are a collection of cubes $\Lambda_{1/N}(\bm{x}_{i})\subset[0,1]^{2d}$ ($1\leq i\leq N_{1}$ for some $N_{1}>1$) with disjoint exteriors, and a set $\bm{B}\subset[0,1]^{2d}$ so that
${\rm meas}\left([0,1]^{2d}\setminus\cup_{1\leq i\leq N_{1}}\Lambda_{1/N}(\bm{x}_{i})\right)$ $\leq CL^{2d-1}\delta,$
${\rm meas}([0,1]^{2d}\setminus\bm{B})$ $<1/10.$
Moreover, if $(\bm{\alpha},\bm{\theta})\in\cup_{1\leq i\leq N_{1}}\Lambda_{1/N}(\bm{x}_{i})$ and $\bm{\xi}\in\bm{B},$ then
$|\det W|\geq c\delta^{R},\ c=c(d,R)>0.$
We are ready to conclude the proof of Theorem A.2. Fix $\bm{\xi}\in\bm{B}$ and a $1/N\sim\delta^{2}$-cube $\Lambda_{i}=\Lambda_{1/N}(\bm{x}_{i})\subset[0,1]^{2d}$ in Lemma A.13. We will deal with $(\bm{\alpha},\bm{\theta})\in\Lambda_{i}.$ Denote for $\bm{k}^{\prime}\in\mathbb{Z}^{D}\setminus\{\bm{0}\}$,
$f(\bm{\alpha},\bm{\theta})=\bm{k}^{\prime}\cdot\bm{\omega}^{\prime}=(\bm{k}^{\prime}\otimes\bm{v})\cdot\bm{V}.$
Direct computations show that for $\bm{\xi}=(\bm{\beta},\bm{q}),$
$\ \ \ \sup_{(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}}\sup_{1\leq j\leq R}|\partial^{2j}_{\bm{\xi}}\bm{V}|_{1}$
$=\sup_{(\bm{\alpha},\bm{\theta})\in[0,1]^{2d}}\sup_{1\leq j\leq R}\sum_{1\leq s\leq D,\bm{\ell}\in\Gamma_{K}}|(2\pi(\bm{\ell}\bm{n}_{s}^{\prime})\cdot\bm{\beta}+2\pi\bm{q}\cdot\bm{\ell})^{2j}\cos 2\pi\bm{\ell}\cdot(\bm{\theta}+\bm{n}_{s}^{\prime}\bm{\alpha})|$
$\leq CR(\sup_{s}|\bm{n}_{s}^{\prime}|+dK)^{2R}:=M.$
We have by Lemma A.13 and Lemma A.8 that
$\inf_{(\bm{\alpha},\bm{\theta})\in\Lambda_{i}}\max_{1\leq j\leq R}|d^{2j}_{\bm{\xi}}f(\bm{\alpha},\bm{\theta})|\geq c\delta^{R}M^{1-R}|\bm{k}^{\prime}|\cdot|\bm{v}|,$ (A.16)
where $c=c(d,R,V)>0.$
It suffices to apply Lemma A.6 by letting $k=2R$. The transversality condition has been verified by (A.16). For $f(\bm{\alpha},\bm{\theta})=\bm{k}^{\prime}\cdot\bm{\omega}^{\prime}$, we have
$\|f\|_{2R+1}=\sup_{|\bm{\gamma}|\leq 2R+1,(\bm{\alpha},\bm{\theta})\in\Lambda_{i}}|\partial^{\bm{\gamma}}_{\bm{\xi}}f(\bm{\alpha},\bm{\theta})|\leq C(d,R,V)\cdot(\sup_{1\leq s\leq D}|\bm{n}_{s}^{\prime}|)^{2R+1}|\bm{k}^{\prime}|.$
So using Lemma A.6 by setting $\varepsilon=\frac{\eta}{|\bm{k}^{\prime}|^{4R^{2}}}$ leads to
$\ \ \ {\rm meas}\left(\left\{(\bm{\alpha},\bm{\theta})\in\Lambda_{i}:\ |\bm{k}^{\prime}\cdot\bm{\omega}^{\prime}|\leq\frac{\eta}{|\bm{k}^{\prime}|^{4R^{2}}}\right\}\right)$
$\leq C(d,R,V)(\max_{1\leq s\leq D}|\bm{n}_{s}^{\prime}|)^{3Rd}M^{2Rd}\delta^{-R(d+1)}\eta^{\frac{1}{2R}}|\bm{k}^{\prime}|^{-{2R}}$
$\leq C(d,R,V)(\max_{1\leq s\leq D}|\bm{n}_{s}^{\prime}|)^{5dR^{2}}\frac{\delta^{3dR}}{|\bm{k}^{\prime}|^{2D}}\ ({\rm choose}\ \delta=\frac{1}{10}\eta^{\frac{1}{10dR^{2}}}).$
Together with Lemma A.13 this yields
$\ \ \ {\rm meas}([0,1]^{2d}\setminus{\rm DC}_{\rm nls}(\eta))$
$\leq C(d,R,V)L^{2d-1}\delta+C(d,R,V)(\max_{1\leq s\leq D}|\bm{n}_{s}^{\prime}|)^{5dR^{2}}\delta^{-4d}\sum_{\bm{k}^{\prime}\in\mathbb{Z}^{D}\setminus\{\bm{0}\}}\frac{\delta^{3dR}}{|\bm{k}^{\prime}|^{2D}}$
$\leq C(d,K,D,V)(\max_{1\leq s\leq D}|\bm{n}_{s}^{\prime}|)^{5dR^{2}}\delta\ ({\rm since}\ R\geq 2\ {\rm and}\ L=K\max_{1\leq s\leq D}|\bm{n}_{s}^{\prime}|)$
$\leq C(d,K,D,V)(\max_{1\leq s\leq D}|\bm{n}_{s}^{\prime}|)^{5dR^{2}}\eta^{\frac{1}{10dR^{2}}}.$
This proves Theorem A.2. ∎
Appendix B Some important lemmas
Let $M_{2}(\mathbb{C})$ denote the Banach algebra of all $(2\times 2)$-complex matrices equipped with the standard operator norm. Since the index $\xi\in\{+,-\}$, we can identify the matrix from $\mathbb{Z}^{b+d}_{\rm pm}$ to $\mathbb{C}$ with that from $\mathbb{Z}^{b+d}$ to $M_{2}(\mathbb{C})$. So, in the following analysis, we only focus on matrices with their entries indexes $\bm{x}=(\bm{k},\bm{n})\in\mathbb{Z}^{b+d}$.
We first introduce an important perturbation lemma.
Lemma B.1 (cf. Lemma 4.2 of [LW24] and Lemma A.1 of [Shi22]).
Let $X\subset\mathbb{Z}^{b+d}$ with |X|=#X|X|=\#X. Assume that $A$ and $B$ are two matrices with entries $A(\bm{x},\bm{x}^{\prime})$ and $B(\bm{x},\bm{x}^{\prime})$, where $\bm{x},\bm{x}^{\prime}\in X$. Assume that for some positive constants $\epsilon_{1},\epsilon_{2}\in(0,1)$, $C_{2}\geq 0$, $c>0$ and $0\leq M<{\rm diam}\ X$: (1). for all $\bm{x}$ and $\bm{x}^{\prime}$, $|B(\bm{x},\bm{x}^{\prime})|\leq\epsilon_{2}(1+|\bm{x}-\bm{x}^{\prime}|)^{C_{2}}e^{-c|\bm{x}-\bm{x}^{\prime}|}$; (2). for $|\bm{x}-\bm{x}^{\prime}|>M$, $|A^{-1}(\bm{x},\bm{x}^{\prime})|\leq e^{-c|\bm{x}-\bm{x}^{\prime}|}$. Suppose that $\|A^{-1}\|\leq\epsilon_{1}^{-1}$ and
$|X|^{2}e^{2cM}(1+{\rm diam}\ X)^{C_{2}}\cdot\epsilon_{2}\epsilon_{1}^{-1}\sum_{\bm{x}\in\mathbb{Z}^{b+d}}e^{-c|\bm{x}|}\leq\frac{1}{2}.$
Then
$\|(A+B)^{-1}\|$ $\leq 2\epsilon_{1}^{-1},$
$|(A+B)^{-1}(\bm{x},\bm{x}^{\prime})-A^{-1}(\bm{x},\bm{x}^{\prime})|$ $\leq\epsilon_{1}^{-1}e^{-c|\bm{x}-\bm{x}^{\prime}|}.$
Proof.
The proof is based on the Neumann series argument similar to [Shi22, LW24]. For completeness, we give a proof.
First, we note that $|A^{-1}(\bm{x},\bm{x}^{\prime})|\leq\epsilon_{1}^{-1}e^{cM}e^{-c|\bm{x}-\bm{x}^{\prime}|}$ for all $\bm{x},\bm{x}^{\prime}\in X$. For $\bm{x}^{0}=\bm{x},\bm{x}^{s}=\bm{x}^{\prime}\in X$ and $s\geq 1$, we can obtain by direct computations that
$(BA^{-1})^{s}(\bm{x},\bm{x}^{\prime})=\sum_{\bm{x}^{1},\cdots,\bm{x}^{s-1},\bm{y}^{1},\cdots,\bm{y}^{s}\in X}\prod_{t=1}^{s}B(\bm{x}^{t-1},\bm{y}^{t})A^{-1}(\bm{y}^{t},\bm{x}^{t}).$
Thus for $s\geq 1$, one has
$\ \ \ |(BA^{-1})^{s}(\bm{x},\bm{x}^{\prime})|$
$\leq\sum_{\bm{x}^{1},\cdots,\bm{x}^{s-1},\bm{y}^{1},\cdots,\bm{y}^{s}\in X}e^{scM}(1+{\rm diam}\ X)^{sC_{2}}\cdot\epsilon_{2}^{s}\epsilon_{1}^{-s}e^{-c|\bm{x}-\bm{x}^{\prime}|}$
$\leq|X|^{2s-1}e^{scM}(1+{\rm diam}\ X)^{sC_{2}}\cdot\epsilon_{2}^{s}\epsilon_{1}^{-s}e^{-c|\bm{x}-\bm{x}^{\prime}|},$
and for $\bm{x}\neq\bm{x}^{\prime},$
$\ \ \ \left|\left(\sum_{s\geq 1}(-BA^{-1})^{s}\right)(\bm{x},\bm{x}^{\prime})\right|$
$\leq\sum_{s\geq 1}|X|^{2s-1}e^{scM}(1+{\rm diam}\ X)^{sC_{2}}\cdot\epsilon_{2}^{s}\epsilon_{1}^{-s}e^{-c|\bm{x}-\bm{x}^{\prime}|}$
$=|X|e^{cM}(1+{\rm diam}\ X)^{C_{2}}\cdot\epsilon_{2}\epsilon_{1}^{-1}\cdot\frac{1}{1-|X|^{2}e^{cM}(1+{\rm diam}\ X)^{C_{2}}\cdot\epsilon_{2}\epsilon_{1}^{-1}}e^{-c|\bm{x}-\bm{x}^{\prime}|}$
$\leq 2|X|e^{cM}(1+{\rm diam}\ X)^{C_{2}}\cdot\epsilon_{2}\epsilon_{1}^{-1}e^{-c|\bm{x}-\bm{x}^{\prime}|}.$
Hence, from the Schur’s test, we have
$\|BA^{-1}\|\leq|X|e^{cM}(1+{\rm diam}\ X)^{C_{2}}\cdot\epsilon_{2}\epsilon_{1}^{-1}\cdot\sum_{\bm{x}\in\mathbb{Z}^{b+d}}e^{-c|\bm{x}|}\leq\frac{1}{2}.$
Using the Neumann series argument, we get
$(A+B)^{-1}=A^{-1}\sum_{s\geq 0}(-BA^{-1})^{s}.$
Thus
$\|(A+B)^{-1}\|\leq\|A^{-1}\|\frac{1}{1-\|BA^{-1}\|}\leq 2\epsilon_{1}^{-1},$
and for $\forall\ \bm{x},\bm{x}^{\prime}\in X$,
$\ \ \ |(A+B)^{-1}(\bm{x},\bm{x}^{\prime})-A^{-1}(\bm{x},\bm{x}^{\prime})|$
$\leq\sum_{\bm{y}\in X}|A^{-1}(\bm{x},\bm{y})|\cdot\left|\left(\sum_{s\geq 1}(-BA^{-1})^{s}\right)(\bm{y},\bm{x}^{\prime})\right|$
$\leq\sum_{\bm{y}\in X}\epsilon_{1}^{-1}e^{cM}e^{-c|\bm{x}-\bm{y}|}\cdot 2|X|e^{cM}(1+{\rm diam}\ X)^{C_{2}}\cdot\epsilon_{2}\epsilon_{1}^{-1}e^{-c|\bm{y}-\bm{x}^{\prime}|}$
$\leq 2|X|^{2}e^{2cM}(1+{\rm diam}\ X)^{C_{2}}\cdot\epsilon_{2}\epsilon_{1}^{-2}e^{-c|\bm{x}-\bm{x}^{\prime}|}$
$\leq\epsilon_{1}^{-1}e^{-c|\bm{x}-\bm{x}^{\prime}|}.$
We complete the proof. ∎
In the following, let $A$ be a matrix on $\mathbb{Z}^{b+d}$ satisfying
$|A(\bm{x},\bm{x}^{\prime})|\leq C_{2}(1+|\bm{x}-\bm{x}^{\prime}|)^{C_{2}}e^{-b_{1}|\bm{x}-\bm{x}^{\prime}|},\ C_{2},b_{1}>0.$
We introduce two types of coupling lemmas by iterating resolvent identities, which are repeatedly used to derive off-diagonal exponential decay estimates for Green’s functions.
Remark B.2.
In (linear) spectral problems (cf. [Bou07, JLS20]), one typically deals with operators such as
$|A(\bm{x},\bm{x}^{\prime})|\leq C_{2}e^{-b_{1}|\bm{x}-\bm{x}^{\prime}|}$
without the polynomial factor $(|\bm{x}-\bm{x}^{\prime}|+1)^{C_{2}}$. However, the extra polynomial factor does not significantly affect the proofs in [JLS20] (cf. Lemma 3.2 and Theorem 3.3) and [Liu22] (cf. Theorem 2.1), as one can replace, at each step, the estimate
$\sum_{\bm{x}_{1}\in W\atop\bm{x}_{2}\in\Lambda\setminus W}e^{-b_{1}|\bm{x}-\bm{x}^{\prime}|}|G_{\Lambda}(\bm{x}_{2},\bm{x}^{\prime})|\leq(2N+1)^{2d}\sup_{\bm{x}_{2}\in\Lambda\setminus W}|G_{\Lambda}(\bm{x}_{2},\bm{x}^{\prime})|$
with
$\sum_{\bm{x}_{1}\in W\atop\bm{x}_{2}\in\Lambda\setminus W}(|\bm{x}-\bm{x}^{\prime}|+1)^{C_{2}}e^{-b_{1}|\bm{x}-\bm{x}^{\prime}|}|G_{\Lambda}(\bm{x}_{2},\bm{x}^{\prime})|\leq(2N+1)^{2d+C_{2}}\sup_{\bm{x}_{2}\in\Lambda\setminus W}|G_{\Lambda}(\bm{x}_{2},\bm{x}^{\prime})|.$
We say that an elementary region $\Lambda\in\mathcal{ER}(L)$ is in the class $\mathbb{G}$ (good) if ($G_{\Lambda}=(R_{\Lambda}AR_{\Lambda})^{-1}$)
$|G_{\Lambda}(\bm{x},\bm{x}^{\prime})|\leq e^{-b_{2}|\bm{x}-\bm{x}^{\prime}|}\ {\rm for}\ |\bm{x}-\bm{x}^{\prime}|\geq L^{\frac{8}{9}},$ (B.1)
where $\frac{1}{2}b_{1}<b_{2}\leq b_{1}$.
We have a revised version of the coupling lemma of [Liu22] (cf. also [BGS02]).
Lemma B.3 (cf. Theorem 2.1, [Liu22]).
Let $\tilde{\Lambda}_{0}\in\mathcal{ER}(N)$ be an elementary region with the property that for all $\Lambda\subset\tilde{\Lambda}_{0},\Lambda\in\mathcal{R}_{L}^{\sqrt{N}}$ with $\sqrt{N}\leq L\leq N$, the Green’s function $(R_{\Lambda}AR_{\Lambda})^{-1}$ satisfies
$\|(R_{\Lambda}AR_{\Lambda})^{-1}\|\leq e^{L^{\frac{3}{4}}}.$
Assume that for any family $\mathcal{F}$ of pairwise disjoint elementary regions of size $M=[N^{\xi}]$ contained in $\tilde{\Lambda}_{0},$
$\{\Lambda\in\mathcal{F}:\ \Lambda\ {\rm is\ not\ in\ class}\ \mathbb{G}\}\leq{N^{\frac{1}{4}}}.$
Then for large $N$ (depending on $C_{2},b_{1}$), we have
$|(R_{\tilde{\Lambda}_{0}}AR_{\tilde{\Lambda}_{0}})^{-1}(\bm{x},\bm{x}^{\prime})|\leq e^{-(b_{2}-N^{-\vartheta})|\bm{x}-\bm{x}^{\prime}|}\ {\rm for}\ |\bm{x}-\bm{x}^{\prime}|\geq N^{\frac{8}{9}},$ (B.2)
where $\vartheta=\vartheta\in(0,1)$ is an absolute constant.
Remark B.4.
There is an important difference between the original form of Lemma B.3 in [Liu22] and the present one. In [Liu22], (B.1)–(B.2) hold under the condition of $|\bm{x}-\bm{x}^{\prime}|\geq\frac{N}{10}$, $0<b_{2}\leq\frac{4}{5}b_{1}$. However, the relation $0<b_{2}\leq\frac{4}{5}b_{1}$ leads to the deterioration of decay rates in the Newton iteration when solving the $P$-equations. To address this issue, we impose a stronger restriction $|\bm{x}-\bm{x}^{\prime}|>N^{\frac{8}{9}}$ (this idea was first introduced in [HSSY20]), which ensures that $0<b_{2}\leq b_{1}$. So Lemma B.3 is proved via the argument of [Liu22] combined with the argument of [HSSY20] (we can take $b=\frac{3}{4},\tau=\frac{1}{2},\theta=\frac{8}{9}$ in Lemma 4.2 of [HSSY20]). We omit the details here.
We also need the following lemmas of [JLS20] (with again minor modifications).
Lemma B.5 (cf. Lemma 3.2, [JLS20]).
Let $M_{0}\geq(\log N)^{\frac{4}{3}},\ \frac{b_{1}}{2}<b_{2}\leq b_{1}$ and $M_{1}\leq N$. Suppose that $\Lambda\subset\mathbb{Z}^{b+d}$ is connected and ${\rm diam}(\Lambda)\leq 2N+1$. Suppose that for any $\bm{y}\in\Lambda,$ there exists some $W=W(\bm{y})\in\mathcal{E}_{M}$ with $M_{0}\leq M\leq M_{1}$ such that $\bm{y}\in W\subset\Lambda,{\rm dist}(\bm{y},\Lambda\setminus W)\geq\frac{M}{2}$ and
$\|(R_{W}AR_{W})^{-1}\|\leq e^{M^{\frac{3}{4}}},$
$|(R_{W}AR_{W})^{-1}(\bm{x},\bm{x}^{\prime})|\leq e^{-b_{2}|\bm{x}-\bm{x}^{\prime}|}\ {\rm for}\ |\bm{x}-\bm{x}^{\prime}|\geq{M}^{\frac{8}{9}}.$
Assume further that $N$ is large enough such that
$\sup_{M_{0}\leq M\leq M_{1}}e^{M^{\frac{3}{4}}}(2M+1)^{b+d+C_{2}}e^{b_{2}\sqrt{N}}\sum_{j=0}^{\infty}(M+2j+1)^{b+d}e^{-b_{2}(j+M/2)}\leq\frac{1}{2}.$
Then we have
$\|(R_{\Lambda}AR_{\Lambda})^{-1}\|\leq 4(2M_{1}+1)^{b+d}e^{M_{1}^{\frac{3}{4}}}.$
Lemma B.6 (cf. Theorem 3.3, [JLS20]).
Assume $\Lambda\subset\mathbb{Z}^{b+d}$ is connected and ${\rm diam}\ \Lambda\leq 2N+1$. Assume ${\rm diam}\ \Lambda_{1}\leq N^{\frac{1}{2(b+d)}}$. Let $M_{0}\geq(\log N)^{\frac{4}{3}},\ \frac{b_{1}}{2}<b_{2}\leq b_{1}$. Suppose that for any $\bm{y}\in\Lambda,$ there exists some $W=W(\bm{y})\in\mathcal{ER}(M)$ with $M_{0}\leq M\leq N^{\frac{3}{4}}$ such that $\bm{y}\in W\subset\Lambda,{\rm dist}(\bm{y},\Lambda\setminus\Lambda_{1}\setminus W)\geq\frac{M}{2}$ and
$\|(R_{W}AR_{W})^{-1}\|$ $\leq e^{M^{\frac{3}{4}}},$
$|(R_{W}AR_{W})^{-1}(\bm{x},\bm{x}^{\prime})|$ $\leq e^{-b_{2}|\bm{x}-\bm{x}^{\prime}|}\ {\rm for}\ |\bm{x}-\bm{x}^{\prime}|\geq{M}^{\frac{8}{9}}.$
Suppose that
$\|R_{\Lambda}AR_{\Lambda}\|\leq e^{N^{\frac{3}{4}}}.$
Then
$|(R_{\Lambda}AR_{\Lambda})^{-1}(\bm{x},\bm{x}^{\prime})|$ $\leq e^{-(b_{2}-\frac{C(b,d)}{M_{0}^{\frac{1}{9}}})|\bm{x}-\bm{x}^{\prime}|}\ {\rm for}\ |\bm{x}-\bm{x}^{\prime}|\geq{M}^{\frac{8}{9}}.$
Remark B.7.
The similar issue in Lemma B.3 also appears Lemma B.5 and Lemma B.6. Again, we can improve the original estimates of [JLS20] from $|\bm{x}-\bm{x}^{\prime}|\geq\frac{N}{10}$ to $|\bm{x}-\bm{x}^{\prime}|\geq N^{\frac{8}{9}}.$
The following lemma deals with projection property of the elementary regions.
Lemma B.8.
Let $\Lambda\in\mathcal{ER}_{\mathbb{Z}^{b+d}}(N)$ and denote by $\Pi_{b}:\ \mathbb{Z}^{b+d}\to\mathbb{Z}^{b}$ the canonical projection map. Then for any $\bm{k}\in\Pi_{b}\Lambda$, the section
$\Lambda(\bm{k})=\{\bm{n}\in\mathbb{Z}^{d}:(\bm{k},\bm{n})\in\Lambda\}$
either belongs to $\mathcal{ER}_{\mathbb{Z}^{d}}(N)$ or is a rectangle of width at least $N$.
Proof.
By translation invariance, it suffices to consider the regions centered at the origin, i.e., $\Lambda\in\mathcal{ER}_{\bm{0},\mathbb{Z}^{b+d}}(N)$. We proceed by analyzing the two possible forms of $\Lambda$.
Case 1: $\Lambda=\Lambda_{N,\mathbb{Z}^{b+d}}:=[-N,N]^{b+d}\cap\mathbb{Z}^{b+d}$.
In this case, the section at any $\bm{k}\in\Pi_{b}\Lambda$ is simply $\Lambda(\bm{k})=[-N,N]^{d}\cap\mathbb{Z}^{d}$, which belongs to $\mathcal{ER}_{\mathbb{Z}^{d}}(N)$.
Case 2: $\Lambda=\Lambda_{N,\mathbb{Z}^{b+d}}^{\bm{\iota}}$ for some constraint vector $\bm{\iota}=(\iota_{i})_{1\leq i\leq b+d}\in\{>,<,\emptyset\}^{b+d}$ with at least two non-empty components.
By definition, $\Lambda$ can be expressed as:
$\Lambda$ $=\Lambda_{N,\mathbb{Z}^{b+d}}\setminus\left\{\bm{x}=(x_{i})_{1\leq i\leq b+d}:\ \ x_{i}\ \iota_{i}\ 0,\ \forall\ 1\leq i\leq b+d\right\}$
$=\Lambda_{N,\mathbb{Z}^{b+d}}\setminus\left(\bigcap_{1\leq i\leq b+d}\ \left\{\bm{x}:\ \ x_{i}\ \iota_{i}\ 0\right\}\right).$ (B.3)
Using De Morgan’s laws, we can rewrite (B.3) as:
$\Lambda=\Lambda_{N,\mathbb{Z}^{b+d}}\bigcap\left(\bigcup_{1\leq i\leq b+d}\ \left\{\bm{x}:\ \ x_{i}\ \iota_{i}\ 0\right\}^{c}\right),$ (B.4)
where the complement sets are given by
$\left\{\bm{x}\ \big|\ x_{i}\ \iota_{i}\ 0\right\}^{c}=\begin{cases}\emptyset,&\text{if }\iota_{i}=\emptyset,\\ \left\{\bm{x}:\ \ x_{i}\ \overline{\iota_{i}}\ 0\right\},&\text{otherwise},\end{cases}$
with the convention $\overline{>}:=<$ and $\overline{<}:=>$.
Note that the sectioning operation $A\mapsto A(\bm{k})=\{\bm{n}:\ (\bm{k},\bm{n})\in A\}$ has the following properties. Given $A,B\in\mathbb{Z}^{b+d}$:
  • For $\bm{k}\in\prod_{b}(A\cup B)$, one has $(A\cup B)(\bm{k})=A(\bm{k})\cup B(\bm{k})$ since
    $\{\bm{n}:\ (\bm{k},\bm{n})\in(A\cup B)\}=\{\bm{n}:\ (\bm{k},\bm{n})\in A\}\cup\{\bm{n}:\ (\bm{k},\bm{n})\in B\}.$
  • For $\bm{k}\in\prod_{b}(A\cap B)$, one has $(A\cap B)(\bm{k})=A(\bm{k})\cap B(\bm{k})$ since
    $\{\bm{n}:\ (\bm{k},\bm{n})\in(A\cap B)\}=\{\bm{n}:\ (\bm{k},\bm{n})\in A\}\cap\{\bm{n}:\ (\bm{k},\bm{n})\in B\}.$
Therefore, for a fixed $\bm{k}\in\Pi_{b}\Lambda$, taking the section of (B.4) yields:
$\Lambda(\bm{k})$ $=\Lambda_{N,\mathbb{Z}^{d}}\bigcap\left(\bigcup_{1\leq i\leq b+d}\ \left\{\bm{x}:\ \ x_{i}\ \iota_{i}\ 0\right\}^{c}(\bm{k})\right).$ (B.5)
We now analyze the sections of the complements for $1\leq i\leq b+d$. If $\iota_{i}=\emptyset$, the section is trivially empty. Thus it is sufficient to consider the case $\iota_{i}\neq\emptyset$. We decouple the Cartesian product $\bm{x}\in\mathbb{Z}^{b+d}$ as $(\bm{k},\bm{n})\in\mathbb{Z}^{b}\times\mathbb{Z}^{d}$:
[…]
Combining (B.5) and (B.6), we observe that if there exists any $1\leq i\leq b$ such that $k_{i}\ \overline{\iota_{i}}\ 0$ ($\iota_{i}\neq\emptyset$), the union in (B.5) evaluates to $\mathbb{Z}^{d}$. In this scenario, $\Lambda(\bm{k})=\Lambda_{N,\mathbb{Z}^{d}}$, which belongs to $\mathcal{ER}_{\mathbb{Z}^{d}}(N)$.
Otherwise, the first $b$ constraints do not contribute to the union, and the geometric shape is entirely determined by the remaining $d$ spatial constraints. Thus, there exists a projected constraint vector $\bm{\iota}^{\prime}\in\{>,<,\emptyset\}^{d}$ such that
$\bigcup_{1\leq i\leq b+d}\left\{\bm{x}:\ \ x_{i}\ \iota_{i}\ 0\right\}^{c}(\bm{k})$ $=\bigcup_{1\leq j\leq d}\ \left\{\bm{n}\in\mathbb{Z}^{d}:\ \ n_{j}\ \iota_{j}^{\prime}\ 0\right\}^{c}.$
Substituting this back into (B.5) and applying De Morgan’s laws once more, we obtain
$\Lambda(\bm{k})$ $=\Lambda_{N,\mathbb{Z}^{d}}\setminus\left(\bigcap_{1\leq j\leq d}\ \left\{\bm{n}\in\mathbb{Z}^{d}:\ \ n_{j}\ \iota_{j}^{\prime}\ 0\right\}\right).$
Let $|\bm{\iota}^{\prime}|$ denote the number of non-empty components (i.e., $\iota_{j}^{\prime}\neq\emptyset$) in $\bm{\iota}^{\prime}$. We classify $\Lambda(\bm{k})$ as follows:
  • If $|\bm{\iota}^{\prime}|=0$, then $\Lambda(\bm{k})=\Lambda_{N,\mathbb{Z}^{d}}$, which belongs to $\mathcal{ER}_{\mathbb{Z}^{d}}(N)$.
  • If $|\bm{\iota}^{\prime}|=1$, $\Lambda(\bm{k})$ is the full box $\Lambda_{N,\mathbb{Z}^{d}}$ minus a single half-space. This leaves a rectangle where (d1)(d-1) sides have length $2N$ and one side has length $N$. Hence, it is a rectangle of width at least $N$.
  • If $|\bm{\iota}^{\prime}|\geq 2$, by the definition, $\Lambda(\bm{k})=\Lambda_{N,\mathbb{Z}^{d}}^{\bm{\iota}^{\prime}}$, which implies $\Lambda(\bm{k})\in\mathcal{ER}_{\mathbb{Z}^{d}}(N)$.
This completes the proof for the origin centered case. ∎